### v3

# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

Abstract.
The hybrid quantum group was firstly introduced by Gaitsgory, whose category $\mathcal{O}$ can be viewed as a quantum analogue of BGG category $\mathcal{O}$. We give a coherent model for its principal block at roots of unity, using the non-commutative Springer resolution defined by Bezrukavnikov–Mirković. In particular, the principal block is derived equivalent to the affine Hecke category. As an application, we endow the principal block with a canonical grading, and show that the graded multiplicity of simple module in Verma module is given by the generic Kazhdan–Lusztig polynomial.
1. Introduction
1.1. Hybrid quantum group and its category ${\mathcal{O}}$
Let $G$ be a complex connected and simply-connected semisimple algebraic group, with a Borel subgroup $B$ and a Cartan subgroup $T\subset B$. Let $U_{q}$ be Lusztig’s quantum group and let ${\mathfrak{U}}_{q}$ be De Concini–Kac’s quantum group, associated to $G$. The hybrid (or mixed) quantum group $U^{\mathrm{hb}}_{q}$ is an algebra admitting a triangular decomposition  [^1] In the present paper we will use the definition $U^{\mathrm{hb}}_{q}={\mathfrak{U}}^{-}_{q}\otimes{\mathfrak{U}}^{0}_{q}\otimes U^{+}_{q}$, for the convenience of deforming category ${\mathcal{O}}$. In fact these two definitions lead to the same category ${\mathcal{O}}$ (see [9, Lem 3.1]).
$U^{\mathrm{hb}}_{q}={\mathfrak{U}}^{-}_{q}\otimes U^{0}_{q}\otimes U^{+}_{q},$
which was firstly introduced by Gaitsgory [22] with the perspective of generalizing the Kazhdan–Lusztig equivalence. It also appears naturally when one considers the quantum analogue of the BGG category ${\mathcal{O}}$. Recall that the BGG category ${\mathcal{O}}$ (with integral weights) for the Lie algebra ${\mathfrak{g}}=\mathrm{Lie}(G)$ can be interpreted as the category
${\mathfrak{g}}\text{-}\mathrm{mod}^{B}$
of finitely generated ${\mathfrak{g}}$-modules whose restriction on ${\mathfrak{b}}=\mathrm{Lie}(B)$ is integrable. Therefore, one can define the quantum category ${\mathcal{O}}$ to be the category
${\mathcal{O}}_{q}={\mathfrak{U}}_{q}\text{-}\mathrm{mod}^{U^{\geq}_{q}}$
of finitely generated ${\mathfrak{U}}_{q}$-modules equipped with an integrable $U^{\geq}_{q}$-module structure that is compatible through ${\mathfrak{U}}^{\geq}_{q}\rightarrow U^{\geq}_{q}$. It is exactly the category ${\mathcal{O}}$ associated to $U^{\mathrm{hb}}_{q}$ with its triangular decomposition.
1.2. Main result
Let ${\mathcal{O}}^{0}_{q}$ be the principal block for ${\mathcal{O}}_{q}$. From now on, we specialize $q$ to a root of unity $\zeta$ whose order satisfies some reasonable conditions specified in §2.1 and §3.2.2 (see also Remark 5.14). Our main goal is to give a coherent model for ${\mathcal{O}}^{0}_{\zeta}$, using the non-commutative Springer (Grothendieck) resolution defined in [12].
Let ${\mathrm{A}}$ be the ($G$-equivariant) ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$-algebra defined in loc. cit. as the non-commutative counterpart of the Grothendieck resolution $\tilde{{\mathfrak{g}}}=G\times^{B}{\mathfrak{b}}\rightarrow{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}$ (where ${\mathfrak{t}}=\mathrm{Lie}(T)$ and $W$ is the Weyl group). Let $\pi_{{}_{\tilde{{\mathcal{N}}}}}:\tilde{{\mathcal{N}}}=G\times^{B}{\mathfrak{n}}\rightarrow{\mathfrak{g}}$ be the Springer resolution (where ${\mathfrak{n}}$ is the nilpotent radical of ${\mathfrak{b}}$). Our main result is
Theorem A (=Corollaries 5.15, 6.8).
There is an equivalence of abelian categories
(1.1) ${\mathcal{O}}^{0}_{\zeta}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})$
intertwining the reflection functors on both sides.
Note that a similar result for modular category ${\mathcal{O}}$ was obtained by Losev [26] recently. The category $\mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})$ firstly appeared in [10], which was shown to be derived equivalent to a version of the affine Hecke category and coincide with the heart of the “new t-structure” defined by Frenkel–Gaitsgory [21]. It would be interesting to explore the relationship between (1.1) and a conjectural equivalence of Gaitsgory [22, Conj. 9.2.2] recently proved by Chen–Fu [16].
After this paper is written, Ivan Losev [27] informed us that he has a different approach to prove the equivalence (1.1) independently.
Let us explain the ingredients of the proof of Theorem A.
1.3. Equivalence of Steinberg block and functor ${\mathbb{V}}$
A key ingredient is an equivalence of the Steinberg block ${\mathcal{O}}^{-\rho}_{\zeta}$ (the block containing the Steinberg representation) established in our previous work [36].
Theorem 1.1 ([36]).
There is an equivalence of abelian categories
(1.2) ${\mathcal{O}}^{-\rho}_{\zeta}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\tilde{{\mathcal{N}}}).$
Based on the equivalence above, we establish functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ on both sides of (1.1) that are fully-faithful on the objects admitting Verma flags (in the deformed setting, see §1.5 below).
The functors ${\mathbb{V}}$ are constructed as follows. Let $C={\mathbb{C}}[0\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ be the coinvariant ring. There is an algebra homomorphism (see [9])
$C\rightarrow Z({\mathcal{O}}^{0}_{\zeta}).$
We consider the category ${\mathcal{O}}^{-\rho}_{\zeta,C}$ of $U^{\mathrm{hb}}_{\zeta}\otimes C$-modules that are contained in ${\mathcal{O}}^{-\rho}_{\zeta}$ when viewed as $U^{\mathrm{hb}}_{\zeta}$-modules. We define a functor
${\mathbb{V}}_{r}:\ {\mathcal{O}}^{0}_{\zeta}\rightarrow{\mathcal{O}}^{-\rho}_{\zeta,C}$
by the translation functor from ${\mathcal{O}}^{0}_{\zeta}$ to ${\mathcal{O}}^{-\rho}_{\zeta}$ and remembering at the same time the action of $C$ on ${\mathcal{O}}^{0}_{\zeta}$. If the parameter $q$ is generic, then the Steinberg block ${\mathcal{O}}^{-\rho}_{q}$ is trivial, so the functor ${\mathbb{V}}_{r}$ is given by
${\mathbb{V}}_{r}:\ {\mathcal{O}}^{0}_{q}\rightarrow{\mathcal{O}}^{-\rho}_{q,C}\simeq C\text{-}\mathrm{mod}.$
In this case, it coincides with the Soergel’s functor ${\mathbb{V}}$ [38] for the classical BGG category ${\mathcal{O}}$. When $q=\zeta$, the category ${\mathcal{O}}^{-\rho}_{\zeta}$ is non-trivial. Hence we may view ${\mathbb{V}}_{r}$ as a “relative analogue” of Soergel’s functor.
On the coherent side, we define a functor
${\mathbb{V}}_{c}:\ \mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})\rightarrow\mathrm{Coh}^{G}(\tilde{{\mathcal{N}}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$
by base change along $\pi_{{}_{\tilde{{\mathcal{N}}}}}:\tilde{{\mathcal{N}}}\rightarrow{\mathfrak{g}}$ of the non-commutative resolution ${\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$. To relate the images of two functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$, we use a $C$-linear extension of the equivalence (1.2), i.e.
${\mathcal{O}}^{-\rho}_{\zeta,C}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\tilde{{\mathcal{N}}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}).$
Then we compare the images of the projective objects under the functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$. To that end, we firstly establish a “de-equvariantization” of Theorem A.
1.4. A de-equvariantization
More precisely, let ${\mathfrak{U}}^{b}_{\zeta}$ be the image of ${\mathfrak{U}}_{\zeta}\rightarrow U^{\mathrm{hb}}_{\zeta}$, and consider the category ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T}$ of $X^{*}(T)$-graded representations of ${\mathfrak{U}}^{b}_{\zeta}$. Let ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T,0}$ be its principal block. Note that the restriction of $\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}}$ on the fiber ${\mathfrak{n}}\hookrightarrow\tilde{{\mathcal{N}}}$ is ${\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]$. Hence we can identify
$\mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})=\mathrm{Coh}^{B}({\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}])$
by restriction to ${\mathfrak{n}}$. We have the following $T$-equivariant version of Theorem A.
Proposition 1.2 (=Proposition 4.1).
There is an equivalence of abelian categories
(1.3) ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T,0}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{T}({\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]),$
intertwining the reflection functors on both sides.
We observe that a generating family of projective objects of the categories in (1.1) can be obtained from the inductions of the projective objects of categories in (1.3). Therefore, we construct similar functors ${\mathbb{V}}$ for the categories in (1.3), and show their compatibility with the induction functors. So that we can identify the images of projective objects under the functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$.
The proof of Proposition 1.2 is based on results of Tanisaki [42], where the localization theorem for ${\mathfrak{U}}_{\zeta}$ was established and it further related ${\mathfrak{U}}_{\zeta}$-modules to ${\mathrm{A}}$-modules by some splitting bundles for the sheaf of differential operators on the quantized flag variety. Note that it is not clear from their constructions whether the splitting bundles are $U_{\zeta}$-equivariant or not. So à priori, one can not derive (1.1) from (1.3), by adding the $U^{\geq}_{\zeta}$-equivariant and $B$-equivariant conditions on both sides. However, after proving our main result, we can show that (1.3) dose respect the $B$-actions on both categories as a consequence (see Corollary 6.7).
Finally, we have to show the fully-faithfulness for ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ on the objects admitting Verma flags, which is true only in the deformed settings (see Remark 4.15). Hence we need deformations of category ${\mathcal{O}}$, which we mainly work with in this paper.
1.5. Deformations
For the classical category ${\mathcal{O}}$, the deformation technique played an important role on studying basic structures, e.g. [38], [40] and [20]. Deformation of ${\mathcal{O}}_{\zeta}$ was studied in [37]. Let $S={\mathbb{C}}[{\mathfrak{t}}]_{\hat{0}}$ be the completion of ${\mathbb{C}}[{\mathfrak{t}}]$ at $0\in{\mathfrak{t}}$. The deformation category for ${\mathcal{O}}_{\zeta}$ is an $S$-linear category ${\mathcal{O}}_{\zeta,S}$ whose specialization to the residue field is ${\mathcal{O}}_{\zeta}$. Let ${\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ be the completion of ${\mathrm{A}}$ at $0\in{\mathfrak{t}}/W$. We have a deformed version of Theorem A
Theorem B (=Theorem 5.13).
There is an equivalence of $S$-linear abelian categories
(1.4) ${\mathcal{O}}^{0}_{\zeta,S}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}),$
where ${\mathcal{O}}^{0}_{\zeta,S}$ is the principal block of ${\mathcal{O}}_{\zeta,S}$ and $\pi:\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ is the natural projection.
We have similar constructions as in §1.3 and §1.4 in the deformed setting, which together with the fact that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are fully-faithful on objects admitting Verma flags (see Proposition 5.11) imply Theorem B. Then the equivalence (1.1) follows from (1.4) by $S$-linearity.
1.6. Kazhdan–Lusztig polynomial
Consider the ${\mathbb{C}}^{\times}$-action on ${\mathfrak{g}}$ by $t.x=t^{-2}x$, for any $t\in{\mathbb{C}}^{\times}$, $x\in{\mathfrak{g}}$. The ${\mathbb{C}}[{\mathfrak{g}}]$-algebra ${\mathrm{A}}$ can be equipped with a ${\mathbb{C}}^{\times}$-equivariant structure. By the equivalence (1.1), we can define a grading on ${\mathcal{O}}^{0}_{\zeta}$ via
${\mathcal{O}}^{0,{\mathrm{gr}}}_{\zeta}:=\mathrm{Coh}^{G\times{\mathbb{C}}^{\times}}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})\ \rightarrow\ {\mathcal{O}}^{0}_{\zeta}.$
We compute the Kazhdan–Lusztig polynomials in ${\mathcal{O}}^{0,{\mathrm{gr}}}_{\zeta}$:
Theorem C (=Theorem 7.4).
The Verma modules and simple modules in ${\mathcal{O}}^{0}_{\zeta}$ admit liftings in ${\mathcal{O}}^{0,{\mathrm{gr}}}_{\zeta}$. Moreover, the graded multiplicities of simple module in Verma module are given by the generic Kazhdan–Lusztig polynomials.
1.7. Organization of the paper
In Section 2 we introduce notations for quantum groups and recall some basic properties of the category ${\mathcal{O}}$ for $U^{\mathrm{hb}}_{\zeta}$ from [37]. In Section 3 we give some reminders on the non-commutative Springer resolutions following [12] and the localization theorem for ${\mathfrak{U}}_{\zeta}$ following [42].
In Section 4 we prove Proposition 1.2, and construct the functors ${\mathbb{V}}_{r}^{b}$, ${\mathbb{V}}^{b}_{c}$ in §4.2, whose fully-faithfulness will be proved in §4.4. In §4.2 we also study the induction functors from the categories in (1.3) to the categories in (1.1). In §4.3 we investigate the image of Verma modules under the equivalence (1.3).
In Section 5, we define the functors ${\mathbb{V}}_{r}$, ${\mathbb{V}}_{c}$, and show their compatibility with reflection functors. §5.2 is a slightly technical section. The projective objects in ${\mathcal{O}}_{\zeta}$ only exist after truncations, hence we have to define corresponding truncations on the coherent side and study projective objects in them, which will be done in §5.2. In §5.3 we show the fully-faithfulness of ${\mathbb{V}}_{r}$, ${\mathbb{V}}_{c}$ and then prove the main theorem.
In Section 6, we firstly recall the formalism of a rational group action on a category, then equip a $B$-action on LHS of (1.3) using coinduction functor. We apply our main theorem to show that the equivalence (1.3) respects $B$-actions on both categories, and the equivalence (1.1) can be obtained as the $B$-equivariantization of (1.3).
In Section 7, we firstly recall the periodic and generic polynomials. Then we discuss the liftings of the simple, Verma and projective objects in the graded principal block and compute the related graded multiplicities.
1.8. Conventions
Let $A$ be an algebra over a commutative Noetherian ring $R$. For any closed point $z\in\mathrm{Spec}R$, we denote by $A_{\hat{z}}$ the completion of $A$ at $z$. Suppose $f:X\rightarrow\mathrm{Spec}R$ is a morphism of schemes. By formal neighborhood of $Y=f^{-1}(z)$ in $X$, we mean a scheme
$\mathrm{FN}_{X}(Y):=X_{\hat{z}}:=X\times_{\mathrm{Spec}R}\mathrm{Spec}(R_{\hat{z}}).$
Let ${\mathcal{A}}$ be a quasi-coherent sheaf of algebras on a scheme $X$. We will denote by ${\mathcal{A}}\text{-}\mathrm{mod}$ (resp. ${\mathcal{A}}\text{-}\mathrm{Mod}$) the category of coherent (resp. quasi-coherent) ${\mathcal{A}}$-modules. For a closed subscheme $Y$ in $X$, we denote by ${\mathcal{A}}\text{-}\mathrm{mod}_{Y}$ (resp. ${\mathcal{A}}\text{-}\mathrm{Mod}_{Y}$) the category of sheaves in ${\mathcal{A}}\text{-}\mathrm{mod}$ (resp. ${\mathcal{A}}\text{-}\mathrm{Mod}$) that are set-theoretically supported on $Y$. Abbreviate $\mathrm{Coh}_{Y}(X)={\mathcal{A}}\text{-}\mathrm{mod}_{Y}$ if ${\mathcal{A}}={\mathcal{O}}_{X}$. We will view any algebra $A$ as a sheaf on $\mathrm{Spec}Z(A)$.
For an algebraic group $K$, we will denote by $\mathrm{Rep}(K)$ the category of rational representations of $K$, and by $\mathrm{rep}(K)$ the full subcategory of finite dimensional representations in $\mathrm{Rep}(K)$.
For a Hopf algebra $H$, we usually denote by $\Delta$ its comultiplication and by $\mathrm{S}$ its antipode. We will use the Sweedler’s notation $\Delta(u)=u^{(1)}\otimes u^{(2)}$, for any $u\in H$.
For a category ${\mathcal{C}}$, we denote by $Z({\mathcal{C}})=\mathrm{End}(1_{\mathcal{C}})$ its center. We will write $\mathrm{Hom}=\mathrm{Hom}_{\mathcal{C}}$ if without confusions.
1.9. Acknowledgements
The author sincerely thanks his supervisor Peng Shan for helpful discussions. The author is deeply grateful to Simon Riche for his invitation of visiting Université Clermont Auvergne and for the enlightening discussions. The author is deeply grateful to Eric Vasserot for his invitation of visiting IMJ-PRG, Université Paris Cité (where the major part of this work was done), and for the constant discussions and enlightening suggestions. These visits are supported by Tsinghua Scholarship for Overseas Graduate Studies. The author also thanks Lin Chen, Gurbir Dhillon, Yuchen Fu, Ivan Losev, Cris Negron and Ruotao Yang for stimulating discussions. The author is supported by NSFC Grant No. 12225108.
2. Quantum groups and their category ${\mathcal{O}}$
2.1. Root data
Let $G$ be a connected and simply-connected semisimple algebraic group over ${\mathbb{C}}$, with a Borel subgroup $B$ and a maximal torus $T$ contained in $B$. Denote their Lie algebras by
${\mathfrak{g}}=\mathrm{Lie}(G),\quad{\mathfrak{b}}=\mathrm{Lie}(B),\quad{\mathfrak{n}}=\mathrm{Lie}(N),\quad{\mathfrak{t}}=\mathrm{Lie}(T).$
We identify $T_{\hat{1}}={\mathfrak{t}}_{\hat{0}}$ via the exponential map ${\mathfrak{t}}\rightarrow T$. Denote by ${\mathcal{B}}=G/B$ the flag variety. Let $W=N_{G}(T)/T$ be the Weyl group for $G$. Let ${\mathrm{I}}\subset W$ be the subset of simple reflections. Let $w_{0}\in W$ be the longest element, and set $B^{-}=w_{0}Bw^{-1}_{0}$, $N^{-}=w_{0}Nw^{-1}_{0}$.
Let $(X^{*}(T),X_{*}(T),\Phi,\check{\Phi})$ be the root datum associated with $G$. Let $\langle$-,$\rangle$ - , - ⟩ be the natural pairing of $\Lambda:=X^{*}(T)$ and $X_{*}(T)$. Let $\Phi^{+}\subset\Phi$ be the subsets of positive roots. Denote by $\alpha_{s}\in\Phi^{+}$ the simple root associated to $s\in{\mathrm{I}}$. Let ${\mathrm{Q}}\subset\Lambda$ be the root lattices. There is a partial order $\leq$ on $\Lambda$ given by $\mu\leq\lambda$ if $\lambda-\mu\in\sum_{s\in{\mathrm{I}}}{\mathbb{N}}\alpha_{s}$. Let $(d_{s})_{s\in{\mathrm{I}}}\in{\mathbb{N}}^{\mathrm{I}}$ be the unique tuple of relatively prime positive integers such that the matrix $\big{(}d_{s}\langle\check{\alpha}_{s},\alpha_{t}\rangle\big{)}_{s,t\in{\mathrm{I}}}$ is symmetry and positive definite. It defines a pairing $(-,-):{\mathrm{Q}}\times{\mathrm{Q}}\rightarrow{\mathbb{Z}}$ by $(\alpha_{s},\alpha_{t}):=d_{s}\langle\check{\alpha}_{s},\alpha_{t}\rangle$ and extents to
$(-,-):\Lambda\times\Lambda\rightarrow\frac{1}{{\mathrm{e}}}{\mathbb{Z}},\quad{\mathrm{e}}:=|\Lambda/{\mathrm{Q}}|.$
Let $l$ be an odd positive integer larger than the Coxeter number of $G$ which is prime to ${\mathrm{e}}$, and to $3$ if $G$ contains a component of type $G_{2}$. Let $\zeta_{\mathrm{e}}\in{\mathbb{C}}$ be a primitive $l$-th root of unity, and let $\zeta=(\zeta_{\mathrm{e}})^{\mathrm{e}}$.
2.1.1. Affine Weyl group
Let $W_{\mathrm{af}}=W\ltimes{\mathrm{Q}}$ be the affine Weyl group. Let $W_{\mathrm{ex}}=W\ltimes\Lambda$ be the extended affine Weyl group. We have $W_{\mathrm{ex}}\simeq W_{\mathrm{af}}\rtimes\Lambda/{\mathrm{Q}}$, where $\Lambda/{\mathrm{Q}}$-acts by automorphisms of the affine root system. Let ${\mathrm{I}}_{\mathrm{af}}$ be the subset of simple reflections in $W_{\mathrm{af}}$. Denote by $t(\lambda)$ the translation by $\lambda\in\Lambda$ in $W_{\mathrm{ex}}$. For any $n\in{\mathbb{Z}}$, there is an $n$-dilated $(-\rho)$-shifted action of $W_{\mathrm{ex}}$ on $\Lambda$ given by
$wt(\nu)\bullet_{n}\lambda=w(\lambda+n\nu+\rho)-\rho,\quad\forall w\in W,\ \forall\nu,\lambda\in\Lambda,$
where $\rho=\frac{1}{2}\sum_{\alpha\in\Phi^{+}}\alpha$. We abbreviate $\bullet=\bullet_{n}$ if $n=1$.
We set $\Xi=\Lambda/(W_{\mathrm{ex}},\bullet_{l})$. For any $\omega\in\Xi$, we will fix a representative (still denoted by $\omega$) in the fundamental alcove $\Xi_{\mathrm{sc}}$ of $W_{\mathrm{af}}$-action on $\Lambda$, where
[…]
2.2. Quantum groups
Let $q$ be a formal variable. The quantum group $\mathscr{U}_{q}$ associated with $G$ is the ${\mathbb{C}}(q^{\frac{1}{{\mathrm{e}}}})$-algebra generated by the standard generators $E_{i},F_{i},K_{\lambda}\ (i\in{\mathrm{I}},\lambda\in\Lambda)$ modulo the quantum Chevalley–Serre relations. It is a Hopf algebra with comultiplication, antipode and counit given by
$\Delta(E_{i})=E_{i}\otimes K_{\alpha_{i}}+1\otimes E_{i},\quad\Delta(F_{i})=F_{i}\otimes 1+K_{\alpha_{i}}^{-1}\otimes F_{i},\quad\Delta(K_{\lambda})=K_{\lambda}\otimes K_{\lambda},$
$\mathrm{S}(E_{i})=-E_{i}K_{\alpha_{i}}^{-1},\quad\mathrm{S}(F_{i})=-K_{\alpha_{i}}F_{i},\quad\mathrm{S}(K_{\lambda})=K_{\lambda}^{-1},$
$\varepsilon(E_{i})=\varepsilon(F_{i})=0,\quad\varepsilon(K_{\lambda})=1.$
The Lusztig’s quantum group $U_{q}$ is a ${\mathbb{C}}[q^{\pm\frac{1}{{\mathrm{e}}}}]$-subalgebra of $\mathscr{U}_{q}$ generated by $E_{i}^{(n)},F_{i}^{(n)},K_{\lambda}$; the De Concini–Kac’s quantum group ${\mathfrak{U}}_{q}$ is generated by $E_{i},F_{i},K_{\lambda}$. The hybrid quantum group $U^{\mathrm{hb}}_{q}$ is the ${\mathbb{C}}[q^{\pm\frac{1}{{\mathrm{e}}}}]$-subalgebra generated by $E_{i}^{(n)},F_{i},K_{\lambda}$. There are all Hopf subalgebras with inclusions
${\mathfrak{U}}_{q}\subset U_{q}^{\mathrm{hb}}\subset U_{q}.$
We denote the ${\mathbb{C}}(q^{\frac{1}{{\mathrm{e}}}})$-subalgebras $\mathscr{U}^{+}_{q}:=\langle E_{i}\rangle_{i\in{\mathrm{I}}}$, $\mathscr{U}^{-}_{q}:=\langle F_{i}\rangle_{i\in{\mathrm{I}}}$ and $\mathscr{U}^{0}_{q}:=\langle K_{\lambda}\rangle_{\lambda\in\Lambda}$, and denote by ${\mathfrak{U}}_{q}^{+},{\mathfrak{U}}_{q}^{-},{\mathfrak{U}}^{0}_{q}$ and $U_{q}^{+},U_{q}^{-},U^{0}_{q}$ their intersections with ${\mathfrak{U}}_{q}$ and $U_{q}$. There is a triangular decomposition
$U_{q}^{\mathrm{hb}}={\mathfrak{U}}_{q}^{-}\otimes{\mathfrak{U}}_{q}^{0}\otimes U_{q}^{+}.$
We will identify $\mathscr{U}^{0}_{q}={\mathbb{C}}(q^{\frac{1}{{\mathrm{e}}}})[\Lambda]={\mathbb{C}}(q^{\frac{1}{{\mathrm{e}}}})[T]$.
For any integral form $A_{q}$ above, we let the ${\mathbb{C}}$-algebra $A_{\zeta}:=A_{q}\otimes_{{\mathbb{C}}[q^{\pm\frac{1}{{\mathrm{e}}}}]}{\mathbb{C}}$ be the specialization at $q^{\frac{1}{{\mathrm{e}}}}=\zeta_{\mathrm{e}}$. The specialization yields a chain of maps
${\mathfrak{U}}_{\zeta}\rightarrow U_{\zeta}^{\mathrm{hb}}\rightarrow U_{\zeta}.$
Let $u_{\zeta}$ be the small quantum group in $U_{\zeta}$, which coincides with the image of ${\mathfrak{U}}_{\zeta}\rightarrow U_{\zeta}$. Denote by ${\mathfrak{U}}^{b}_{\zeta}$ the image of ${\mathfrak{U}}_{\zeta}\rightarrow U_{\zeta}^{\mathrm{hb}}$. Then there are triangular decompositions
$u_{\zeta}=u_{\zeta}^{-}\otimes u_{\zeta}^{0}\otimes u_{\zeta}^{+},\quad{\mathfrak{U}}^{b}_{\zeta}={\mathfrak{U}}^{-}_{\zeta}\otimes{\mathfrak{U}}^{0}_{\zeta}\otimes u_{\zeta}^{+}.$
We set ${\mathfrak{U}}^{t}_{\zeta}$ be the algebra with triangular decomposition ${\mathfrak{U}}^{t}_{\zeta}=u_{\zeta}^{-}\otimes{\mathfrak{U}}_{\zeta}^{0}\otimes u_{\zeta}^{+}$.
2.2.1. Harish-Chandra center
We set $\mathscr{U}_{q}^{0,ev}:={\mathbb{C}}(q^{\frac{1}{{\mathrm{e}}}})\langle K_{2\lambda}\rangle_{\lambda\in\Lambda}$. There is an algebra isomorphism
$Z(\mathscr{U}_{q}){\ \xrightarrow{\sim}\ }(\mathscr{U}_{q}^{0,ev})^{(W,\bullet)}$
given by projecting $Z(\mathscr{U}_{q})$ to $\mathscr{U}_{q}^{0}$ under the triangular decomposition, where $(W,\bullet)$ is the shifted action by $w\bullet K_{\lambda}=q^{(w\lambda-\lambda,\rho)}K_{w\lambda}$, for any $w\in W$, $\lambda\in\Lambda$. The natural identification
$(\mathscr{U}_{q}^{0,ev})^{(W,\bullet)}={\mathbb{C}}(q^{\frac{1}{{\mathrm{e}}}})[T]^{(W,\bullet)},\quad f(K_{2\lambda})\mapsto f(K_{\lambda}),$
gives an isomorphism
$Z(\mathscr{U}_{q}){\ \xrightarrow{\sim}\ }{\mathbb{C}}(q^{\frac{1}{{\mathrm{e}}}})[T/(W,\bullet)].$
The center of ${\mathfrak{U}}_{q}$ is by $Z({\mathfrak{U}}_{q})=Z(\mathscr{U}_{q})\cap{\mathfrak{U}}_{q}$. We obtain an isomorphism
$Z_{\mathrm{HC}}:=Z({\mathfrak{U}}_{q})/(q^{\frac{1}{{\mathrm{e}}}}-\zeta_{\mathrm{e}})Z({\mathfrak{U}}_{q}){\ \xrightarrow{\sim}\ }{\mathbb{C}}[T/(W,\bullet)].$
Here $Z_{\mathrm{HC}}$ is called the Harish-Chandra center of ${\mathfrak{U}}_{\zeta}$. For any $\lambda\in\Lambda$, we denote by $\chi_{\lambda}$ the character of $Z_{\mathrm{HC}}$ corresponding to the image of $\zeta^{\lambda}$ under the map $T\rightarrow T/(W,\bullet)$, by $t\mapsto W\bullet t^{2}$ for any $t\in T$.
2.2.2. Frobenius center
The Frobenius center of ${\mathfrak{U}}_{\zeta}$ is the ${\mathbb{C}}$-subalgebra
$Z_{\mathrm{Fr}}:=\langle K^{l}_{\lambda},F^{l}_{\beta},E^{l}_{\beta}\rangle_{\lambda\in\Lambda,\beta\in\Phi^{+}},$
where $E_{\beta}\in{\mathfrak{U}}^{+}_{\zeta}$ and $F_{\beta}\in{\mathfrak{U}}^{-}_{\zeta}$ are the root vectors. We abbreviate $Z_{\mathrm{Fr}}^{\flat}:=Z_{\mathrm{Fr}}\cap{\mathfrak{U}}_{\zeta}^{\flat}$ for $\flat=-,+,0,\leq$ and $\geq$. By [18] (see also [44]) there are isomorphisms of ${\mathbb{C}}$-algebras
(2.1) $Z_{\mathrm{Fr}}^{-}{\ \xrightarrow{\sim}\ }{\mathbb{C}}[N],\quad Z_{\mathrm{Fr}}^{+}{\ \xrightarrow{\sim}\ }{\mathbb{C}}[N^{-}],\quad\text{and}\quad Z_{\mathrm{Fr}}^{0}{\ \xrightarrow{\sim}\ }{\mathbb{C}}[T],$
which give an isomorphism
$Z_{\mathrm{Fr}}{\ \xrightarrow{\sim}\ }{\mathbb{C}}[G^{*}],$
where $G^{*}=N^{-}\times T\times N$ is the Poisson dual group of $G$. There is a unramified covering to the big open cell $\kappa:G^{*}\rightarrow G$, sending $(n_{1},t,n_{2})\in N^{-}\times T\times N$ to $n_{1}t^{2}n_{2}^{-1}$. We have the following isomorphisms by [19],
(2.2) $Z({\mathfrak{U}}_{\zeta})=Z_{\mathrm{Fr}}\otimes_{Z_{\mathrm{Fr}}\cap Z_{\mathrm{HC}}}Z_{\mathrm{HC}}={\mathbb{C}}[G^{*}\times_{T/W}T/(W,\bullet)],$
where $G^{*}\rightarrow T/W$ is the composition of $\kappa$ with $G\rightarrow G/\!/G=T/W$, and $T/(W,\bullet)\rightarrow T/W$ is induced by taking $l$-th power.
2.3. Module categories
There is a $\Lambda$-action on $\mathscr{U}_{q}^{0}$ such that any $\mu\in\Lambda$ corresponds to the ${\mathbb{C}}(q^{\frac{1}{{\mathrm{e}}}})$-algebra automorphism
$\tau_{\mu}:K_{\lambda}\mapsto q^{(\mu,\lambda)}K_{\lambda},\quad\forall\lambda\in\Lambda.$
The $\Lambda$-action on $\mathscr{U}_{q}^{0}$ preserves the integral forms ${\mathfrak{U}}^{0}_{q}$ and $U^{0}_{q}$, and specializes to an action on ${\mathfrak{U}}^{0}_{\zeta}$ and $U^{0}_{\zeta}$. Let $U^{0}$ be a Noetherian sub-quotient algebra of ${\mathfrak{U}}_{\zeta}^{0}$ or $U_{\zeta}^{0}$ such that the $\Lambda$-action induces an action on $U^{0}$. Let $U=\bigoplus_{\lambda\in{\mathrm{Q}}}U_{\lambda}$ be a ${\mathrm{Q}}$-graded ${\mathbb{C}}$-algebra with triangular decomposition $U=U^{-}\otimes U^{0}\otimes U^{+}$, such that
$fm=m\tau_{\lambda}(f),\quad\forall f\in U^{0},\ \forall m\in U_{\lambda},$
and satisfies further conditions as in [37, §2.3]. We abbreviate $U^{\leq}=U^{-}U^{0}$ and $U^{\geq}=U^{+}U^{0}$. A deformation ring $R$ for $U$ is a commutative Noetherian $U^{0}$-algebra. Let $\iota:U^{0}\rightarrow R$ be the structure map. If $U^{0}$ is an augmented algebra, then we will view ${\mathbb{C}}$ as a deformation ring via the counit map $U^{0}\rightarrow{\mathbb{C}}$.
We define $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ to be the category consisting of $U\otimes R$-modules $M$ endowed with a decomposition $M=\bigoplus\limits_{\mu\in\Lambda}M_{\mu}$ of $R$-modules (called the weight spaces), such that $M_{\mu}$ is killed by the elements in $U\otimes R$ of the form
$f\otimes 1-1\otimes\iota({\tau_{\mu}}(f)),\quad f\in U^{0}.$
Let $U\text{-}\mathrm{mod}_{R}^{\Lambda}$ be the full subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ consisting of finitely generated $U\otimes R$-modules whose weight spaces are finitely generated $R$-modules. Define the category ${\mathcal{O}}$ for $U$ to be the full subcategory ${\mathcal{O}}^{U}_{R}$ of $U\text{-}\mathrm{mod}_{R}^{\Lambda}$ of modules that are locally unipotent for the action of $U^{+}$. It is an abelian subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$.
Define the Verma module
$M^{U}(\lambda)_{R}:=U\otimes_{U^{\geq}}R_{\lambda}\ \in{\mathcal{O}}^{U}_{R}$
where $R_{\lambda}$ is a $U^{\geq}$-module via $U^{\geq}\twoheadrightarrow U^{0}\xrightarrow{\iota\circ\tau_{\lambda}}R_{\lambda}$. We have an isomorphism $M^{U}(\lambda)_{R}=U^{-}\otimes R_{\lambda}$ as $U^{\leq}$-modules. A Verma flag of an object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ is a (decreasing, separated, ${\mathbb{Z}}_{\leq 0}$-) filtration with composition factors by Verma modules. The composition factors of such an object are called Verma factors, which are independent on the choice of Verma flags.
If $R={\mathbb{F}}$ is a field, then $M^{U}(\lambda)_{\mathbb{F}}$ admits a unique simple quotient $L^{U}(\lambda)_{\mathbb{F}}$, and $\{L^{U}(\lambda)_{\mathbb{F}}\}_{\lambda\in\Lambda}$ provides a complete family of pairwise non-isomorphic simple objects in ${\mathcal{O}}^{U}_{R}$.
2.3.1. Truncated categories
For any $\nu\in\Lambda$, there is a truncated subcategory $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ consists of the module $M$ in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ such that $M_{\mu}=0$ unless $\mu\leq\nu$. Note that $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ is a Serre subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$. The category ${\mathcal{O}}^{U,\leq\nu}_{R}:={\mathcal{O}}^{U}_{R}\cap U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ always admits enough projective objects, in contrast to ${\mathcal{O}}^{U}_{R}$. And any object in ${\mathcal{O}}^{U}_{R}$ admits a finite decomposition with each direct factor contained in ${\mathcal{O}}^{U,\leq\nu}_{R}$ for some $\nu\in\Lambda$. Any projective module in ${\mathcal{O}}^{U,\leq\nu}_{R}$ admits Verma flags. If $R$ is a complete local domain with residue field ${\mathbb{F}}$, we write $Q^{U}(\lambda)_{R}^{\leq\nu}$ as the projective cover of $L^{U}(\lambda)_{\mathbb{F}}$ in ${\mathcal{O}}^{U,\leq\nu}_{R}$.
The truncation functor
$U\text{-}\mathrm{Mod}_{R}^{\Lambda}\rightarrow U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu},\quad M\mapsto\ M^{\leq\nu}:=M\big{/}U.\big{(}\bigoplus_{\mu\nleq\nu}M_{\mu}\big{)}$
by taking the maximal quotient in $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ is left adjoint to the natural inclusion $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}\hookrightarrow U\text{-}\mathrm{Mod}_{R}^{\Lambda}$.
Lemma 2.1.
  1. (1)
    We have
    $\mathrm{Hom}(M^{U}(\lambda)_{R},M^{U}(\mu)_{R})\neq 0\quad\text{only if}\ \lambda\leq\mu,$
    and
    $\mathrm{Ext}^{1}(M^{U}(\lambda)_{R},M^{U}(\mu)_{R})\neq 0\quad\text{only if}\ \lambda<\mu.$
  2. (2)
    For any object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ admitting Verma flags, its truncation in $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ is the quotient by the submodule extended by its Verma factors in $\{M^{U}(\lambda)_{R}\}_{\lambda\nleq\nu}$.
2.3.2. Projectives and simples in ${\mathcal{O}}_{R}$
We will drop the superscripts $U$ from the notations above when $U=U^{\mathrm{hb}}_{\zeta}$. We set
$S={\mathbb{C}}[T]_{\hat{1}}={\mathbb{C}}[{\mathfrak{t}}]_{\hat{0}}.$
Then $S$ is naturally an algebra of ${\mathfrak{U}}^{0}_{\zeta}={\mathbb{C}}\langle K_{\lambda}\rangle_{\lambda\in\Lambda}={\mathbb{C}}[T]$. Note that $S$ is a deformation ring for both ${\mathfrak{U}}^{b}_{\zeta}$ and $U^{\mathrm{hb}}_{\zeta}$.
Lemma 2.2 ([37, Lem 3.1]).
The $U_{\zeta}^{\mathrm{hb}}$-module $L(\lambda)_{\mathbb{C}}$ remains simple as a ${\mathfrak{U}}^{b}_{\zeta}$-module, and the ${\mathfrak{U}}^{b}_{\zeta}$-action on $L(\lambda)_{\mathbb{C}}$ factors through the quotient ${\mathfrak{U}}^{b}_{\zeta}\twoheadrightarrow u_{\zeta}$. Hence $L(\lambda)_{\mathbb{C}}$ is a simple $u_{\zeta}$-module.
Let $R$ be a complete local domain over ${\mathfrak{U}}^{0}_{\zeta}$. Let $P^{b}(\lambda)_{R}$ be the projective cover of the simple module of highest weight $\lambda$ in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{R}$. Consider the induced module
$Q(\lambda)_{R}=U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}P^{b}(\lambda)_{R}$
in $U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{R}$.
Lemma 2.3 ([9, Lem 3.7] and [37, Lem 3.2]).
  1. (1)
    The functor $\mathrm{Hom}_{U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{R}}(Q(\lambda)_{R},-)$ is exact on ${\mathcal{O}}_{R}$;
  2. (2)
    The projective cover for $L(\lambda)_{\mathbb{C}}$ in ${\mathcal{O}}^{\leq\nu}_{S}$ (resp. ${\mathcal{O}}^{\leq\nu}_{{\mathbb{C}}}$) is $Q(\lambda)_{S}^{\leq\nu}=\tau^{\leq\nu}Q(\lambda)_{S}$ (resp. $Q(\lambda)_{\mathbb{C}}^{\leq\nu}=\tau^{\leq\nu}Q(\lambda)_{\mathbb{C}}$).
2.3.3. Block decomposition
Till the end of this section, we let $U$ be either ${\mathfrak{U}}^{t}_{\zeta}$, ${\mathfrak{U}}^{t}_{\zeta}U^{+}_{\zeta}$, ${\mathfrak{U}}^{b}_{\zeta}$, ${\mathfrak{U}}_{\zeta}$ or $U^{\mathrm{hb}}_{\zeta}$. Then there are algebra homomorphisms $Z_{\mathrm{HC}}\hookrightarrow{\mathfrak{U}}_{\zeta}\rightarrow U$. By the Harish-Chandra isomorphism, we have an algebra homomorphism
${\mathbb{C}}[T/(W,\bullet)]\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda}_{S}).$
By the $S$-linearity, it extends to a homomorphism
(2.3) ${\mathbb{C}}[T_{\hat{1}}\times_{T/W}T/(W,\bullet)]\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda}_{S}).$
There are algebra isomorphisms (see a specialized version in [15, Thm 4.5])
(2.4) ${\mathbb{C}}[T_{\hat{1}}\times_{T/W}T/(W,\bullet)]$ $=\bigoplus_{\omega\in\Xi}{\mathbb{C}}[T]_{\hat{1}}\otimes_{{\mathbb{C}}[T/W]_{\hat{1}}}{\mathbb{C}}[T/(W,\bullet)]_{\widehat{\chi_{\omega}}}$
$\simeq\bigoplus_{\omega\in\Xi}{\mathbb{C}}[{\mathfrak{t}}]_{\hat{0}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]_{\hat{0}}}{\mathbb{C}}[{\mathfrak{t}}/W_{\omega}]_{\hat{0}}$
$=\bigoplus_{\omega\in\Xi}{\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}/W_{\omega}]_{\hat{0}},$
where $W_{\omega}$ is the stabilizer of the $(W,\bullet)$-action on $\zeta^{\omega}\in T$. The idempotents on the RHS of (2.4) yield a block decomposition
$U\text{-}\mathrm{Mod}^{\Lambda}_{S}=\bigoplus_{\omega\in\Xi}U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S},$
such that the Verma module $M^{U}(\lambda)_{S}$ is contained in $U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S}$ if and only if $\omega=W_{\mathrm{ex}}\bullet_{l}\lambda$. We have algebra homomorphisms
${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}/W_{\omega}]_{\hat{0}}\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S}),\quad\forall\omega\in\Xi.$
The block decomposition induces a decomposition for the subcategory ${\mathcal{O}}^{U}_{S}$, and extends to a decomposition
${\mathcal{O}}^{U}_{R}=\bigoplus_{\omega\in\Xi}{\mathcal{O}}^{U,\omega}_{R},$
for any commutative Noetherian $S$-algebra $R$. Note that when $U={\mathfrak{U}}^{t}_{\zeta}$ or ${\mathfrak{U}}^{b}_{\zeta}$, the category ${\mathcal{O}}$ for $U$ is exactly $U\text{-}\mathrm{mod}^{\Lambda}_{R}$.
Remark 2.4.
Since $U$ is ${\mathrm{Q}}$-graded and $U\text{-}\mathrm{Mod}^{\Lambda}_{S}$ consists of $\Lambda$-graded modules, the blocks above labelled by $\omega\in\Xi$ further decompose into direct sums of ${\mathrm{e}}$ equivalent sub-blocks.
Since $u_{\zeta}$, $u_{\zeta}U^{+}_{\zeta}$ are quotients of ${\mathfrak{U}}_{\zeta}^{t}$, ${\mathfrak{U}}^{t}_{\zeta}U^{+}_{\zeta}$, we have block decompositions
$u_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}=\bigoplus_{\omega\in\Xi}u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,\omega}_{\mathbb{C}}\quad\text{and}\quad{\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta}}_{\mathbb{C}}=\bigoplus_{\omega\in\Xi}{\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta},\omega}_{\mathbb{C}}.$
2.3.4. Translation functors
For any $\omega_{1},\omega_{2}\in\Xi$, there is a unique dominant element $\nu$ in the $W$-orbit of $\omega_{2}-\omega_{1}$. Denote by $V(\nu)$ the Weyl module for $U_{\zeta}$ of highest weight $\nu$. We view $V(\nu)$ as a $U$-module via the natural map $U\rightarrow U_{\zeta}$. We define the translation functors by
$T_{\omega_{1}}^{\omega_{2}}:U\text{-}\mathrm{Mod}^{\Lambda,\omega_{1}}_{S}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda,\omega_{2}}_{S},\quad M\mapsto\mathrm{pr}_{\omega_{2}}(M\otimes V(\nu)),$
$T^{\omega_{1}}_{\omega_{2}}:U\text{-}\mathrm{Mod}^{\Lambda,\omega_{2}}_{S}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda,\omega_{1}}_{S},\quad M\mapsto\mathrm{pr}_{\omega_{1}}(M\otimes V(\nu)^{*}),$
where $\mathrm{pr}_{\omega_{i}}$ is the natural projection to the block $U\text{-}\mathrm{Mod}^{\Lambda,\omega_{i}}_{S}$. Note that $T_{\omega_{1}}^{\omega_{2}}$ and $T^{\omega_{1}}_{\omega_{2}}$ are exact and biadjoint to each other. The translation functors preserve the subcategory ${\mathcal{O}}^{U}_{S}$ and induce translation functors between blocks in ${\mathcal{O}}^{U}_{R}$, for any commutative Noetherian $S$-algebra $R$.
For any simple reflection $s\in{\mathrm{I}}_{\mathrm{af}}$, we fix an element $\omega_{s}\in\Xi_{\mathrm{sc}}$ whose stabilizer under the $W_{\mathrm{af}}$-action is $\{1,s\}$. We abbreviate $T^{0}_{s}=T^{0}_{\omega_{s}}$ and $T_{0}^{s}=T_{0}^{\omega_{s}}$. Define the reflection functor on $U\text{-}\mathrm{Mod}^{\Lambda,0}_{S}$ by
$\Theta^{r}_{s}:=T^{0}_{s}\circ T_{0}^{s}.$
[…]
We also obtain translations and reflections functors for the blocks of $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$ and ${\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta}}_{\mathbb{C}}$.
We similarly define translation functors and reflection functors on the categories ${\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{\chi_{\omega}}$ with $\omega\in\Xi$.
2.3.5. $\Lambda$-translations
Let $U^{\prime}$ be $U$ as in §2.3.3, or be $u_{\zeta}$, $u_{\zeta}U^{+}_{\zeta}$. Let $R$ be a deformation ring for $U^{\prime}$. For any $\nu\in\Lambda$, there is a trivial $U^{\prime}$-module ${\mathbb{C}}_{l\nu}$ supported on the weight $l\nu$. It gives auto-equivalences $-\otimes{\mathbb{C}}_{l\nu}$, $\nu\in\Lambda$ on $U^{\prime}\text{-}\mathrm{Mod}^{\Lambda}_{R}$ preserving each blocks, called the $\Lambda$-translations.
3. Non-commutative Springer resolutions and localization theorem
In this section, we give reminders on the non-commutative Springer resolutions developed in [12] and on the localization theorem for ${\mathfrak{U}}_{\zeta}$ following [42].
3.1. Non-commutative Springer resolutions
For any standard parabolic subgroup $P$ in $G$, we set $\tilde{{\mathfrak{g}}}_{P}=G\times^{P}\mathrm{Lie}(P)$ and abbreviate $\tilde{{\mathfrak{g}}}=\tilde{{\mathfrak{g}}}_{B}$. There is a natural map $\pi_{P}:\tilde{{\mathfrak{g}}}\rightarrow\tilde{{\mathfrak{g}}}_{P}$. The Grothendieck resolution is $\pi:\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$. If $P=P_{s}$ is the minimal parabolic subgroup associated to $s\in{\mathrm{I}}$, we will abbreviate $\tilde{{\mathfrak{g}}}_{s}=\tilde{{\mathfrak{g}}}_{P_{s}}$ and $\pi_{s}=\pi_{P_{s}}$. Denote by ${\mathcal{O}}_{\mathcal{B}}(\lambda)$ the line bundle $G\times^{B}{\mathbb{C}}_{\lambda}\rightarrow{\mathcal{B}}$, and denote by ${\mathcal{O}}_{Y}(\lambda)$ its pullback along any $Y\rightarrow{\mathcal{B}}$. Consider the ${\mathbb{C}}^{\times}$-action on ${\mathfrak{g}}$ via $t.x=t^{-2}x$, for any $t\in{\mathbb{C}}^{\times}$, $x\in{\mathfrak{g}}$. It induces ${\mathbb{C}}^{\times}$-actions on $\tilde{{\mathfrak{g}}}_{P}$ for any $P$.
3.1.1. Affine braid group actions
Let $X$ be a ${\mathfrak{g}}$-scheme satisfying
${{\mathcal{T}}or}_{i}^{{\mathcal{O}}_{\mathfrak{g}}}({\mathcal{O}}_{\tilde{{\mathfrak{g}}}},{\mathcal{O}}_{X})=0,\quad\forall i>0.$
We set $\tilde{{\mathfrak{g}}}_{P,X}:=\tilde{{\mathfrak{g}}}_{P}\times_{\mathfrak{g}}X$ and write $\pi_{P}=\pi_{P}\times_{\mathfrak{g}}X$ by abuse of notations. There is a functor $\Theta_{P}^{c}:=\pi_{P}^{*}\pi_{P*}$ on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$. We abbreviate $\Theta^{c}_{s}=\Theta^{c}_{P_{s}}$ for any $s\in{\mathrm{I}}$, which are called the reflection functors.
Theorem 3.1 ([14]).
For any $s\in{\mathrm{I}}$, there is a self-equivalence ${\mathbb{T}}^{c}_{s}$ on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$ with natural transformations
[…]
which is a distinguished triangle when applying on any object in $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$. The assignments
$\tilde{s}\mapsto({\mathbb{T}}^{c}_{s})^{-1}\quad\text{and}\quad\theta_{\lambda}\mapsto-\otimes^{L}_{{\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{X}}}{\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{X}}(\lambda),\qquad\forall s\in{\mathrm{I}},\ \forall\lambda\in\Lambda,$
define a (weak, right) $B_{\mathrm{ex}}$-action on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$. Moreover, for any closed subgroup $K\subset G\times{\mathbb{C}}^{\times}$ and $K$-equivariant morphism $X\rightarrow{\mathfrak{g}}$, the constructions above are naturally $K$-equivariant.
[…]
3.1.2. Non-commutative Springer resolutions
Let $B^{+}_{\mathrm{ex}}$ be the semi-group of $B_{\mathrm{ex}}$ generated by the canonical lifts $\tilde{s}$ with $s\in{\mathrm{I}}_{\mathrm{af}}$.
Theorem 3.2 ([12]).
  1. (1)
    There is a unique t-structure (called the exotic t-structure) on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$ defined by
    $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})^{\leq}=\{{\mathcal{F}}|\ \pi_{*}(b{\mathcal{F}})\in D^{\leq}\mathrm{Coh}X,\ \forall b\in B^{+}_{\mathrm{ex}}\};$
    $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})^{\geq}=\{{\mathcal{F}}|\ \pi_{*}(b^{-1}{\mathcal{F}})\in D^{\geq}\mathrm{Coh}X,\ \forall b\in B^{+}_{\mathrm{ex}}\}.$
  2. (2)
    There exists a $G\times{\mathbb{C}}^{\times}$-equivariant tilting bundle ${\mathcal{E}}$ on $\tilde{{\mathfrak{g}}}$ such that
    $R\mathcal{H}om_{\tilde{{\mathfrak{g}}}_{X}}({\mathcal{E}}_{X},-):\ D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})\rightarrow D^{b}({\mathrm{A}}_{X}\text{-}\mathrm{mod})$
    is an equivalence, where ${\mathcal{E}}_{X}:={\mathcal{E}}\times_{\mathfrak{g}}X$, ${\mathrm{A}}_{X}={\mathrm{A}}\otimes_{{\mathcal{O}}_{\mathfrak{g}}}{\mathcal{O}}_{X}$ and ${\mathrm{A}}=\mathcal{E}nd_{\tilde{{\mathfrak{g}}}}^{\mathrm{op}}({\mathcal{E}})$, and such that this equivalence is t-exact with respect to the exotic t-structure on the LHS and the trivial t-structure on the RHS.
  3. (3)
    Moreover, for any closed subgroup $K\subset G\times{\mathbb{C}}^{\times}$ and $K$-equivariant morphism $X\rightarrow{\mathfrak{g}}$, we have similar statements for $D^{b}\mathrm{Coh}^{K}(\tilde{{\mathfrak{g}}}_{X})$.
By definition, the functor $\pi_{*}$ is t-exact for the exotic t-structure on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$. It yields a functor $\pi_{*}:{\mathrm{A}}_{X}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}X$. By [12], the functors $\Theta^{c}_{P}$, $\Theta^{c}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$) are also exact for the exotic t-structure. We thus view them as functors on ${\mathrm{A}}_{X}\text{-}\mathrm{mod}$.
Consider the completion $X={\mathfrak{g}}_{\hat{0}}$, then there is an equivalence
$D^{b}({\mathrm{A}}_{{\mathfrak{g}}_{\hat{0}}}\text{-}\mathrm{mod})\simeq D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{\hat{0}}),$
whose restriction on the full subcategories of objects set-theoretically supported at $0\in{\mathfrak{g}}$ induces an equivalence
(3.1) $D^{b}({\mathrm{A}}\text{-}\mathrm{mod}_{0})\simeq D^{b}\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}).$
3.2. Localization theorem for ${\mathfrak{U}}_{\zeta}$
The localization theorem for ${\mathfrak{U}}_{\zeta}$ was established by Backelin–Kremnizer [5] and Tanisaki [44], [45], [42]. In this subsection, we give a reminder on the localization theorem mainly following [42].
3.2.1. Quantum differential operators
Lunts—Rosenberg [28] introduced the quantized flag variety ${\mathcal{B}}_{\zeta}$ as a non-commutative scheme over ${\mathcal{B}}$. Let ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$ be the pushforward of the structure sheave of ${\mathcal{B}}_{\zeta}$ on ${\mathcal{B}}$, which is a coherent sheaf of (in general non-commutative) algebras on ${\mathcal{B}}$. In [43], Tanisaki defined the ring of enhanced differential operators on ${\mathcal{B}}_{\zeta}$ (note that in loc. cit. the convention for flag variety is $B^{-}\backslash G$, which is different from us). Let $\widetilde{{\mathcal{D}}}_{\zeta}$ be its localization on ${\mathcal{B}}$. There is an algebra homomorphism
(3.2) ${\mathfrak{U}}_{\zeta}\otimes{\mathbb{C}}[T]\rightarrow\Gamma(\widetilde{{\mathcal{D}}}_{\zeta}).$
Consider the scheme
$\widetilde{G^{*}}=\{(gB,\chi)\in{\mathcal{B}}\times G^{*}\ |\ g^{-1}\kappa(\chi)g\in B\},$
and set
${\mathcal{V}}=\widetilde{G^{*}}\times_{T}T=\{(gB,\chi,t)\in{\mathcal{B}}\times G^{*}\times T\ |\ g^{-1}\kappa(\chi)g\in t^{l}N\},$
where $T\rightarrow T$ is by taking $l$-th power, and $\widetilde{G^{*}}\rightarrow T$ is induced by $B=N\times T\rightarrow T$, $nt\mapsto t^{2}$ for any $n\in N$ and $t\in T$. Note that ${\mathcal{V}}$ is affine over ${\mathcal{B}}$. By [44, Thm 5.2], there is an isomorphism from the central subsheaf of $\widetilde{{\mathcal{D}}}_{\zeta}$ to the structure sheaf ${\mathcal{O}}_{{\mathcal{V}}}$ of ${\mathcal{V}}$. We thus also view $\widetilde{{\mathcal{D}}}_{\zeta}$ as a (coherent) sheaf of algebras on ${\mathcal{V}}$. The homomorphism ${\mathfrak{U}}_{\zeta}\rightarrow\widetilde{{\mathcal{D}}}_{\zeta}$ restricting on their centers is compatible with the natural map
$\varpi:\ {\mathcal{V}}\rightarrow{G^{*}}\times_{T/W}T\rightarrow{G^{*}}\times_{T/W}T/(W,\bullet).$
3.2.2. Localization theorem
From now on, we assume moreover that $l$ is a power of prime, following [42].
The homomorphism ${\mathfrak{U}}_{\zeta}\rightarrow\Gamma(\widetilde{{\mathcal{D}}}_{\zeta})$ yields a functor
$R\Gamma:D^{b}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod})\rightarrow D^{b}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}).$
Theorem 3.3 ([42]).
For any $(\chi,\zeta^{2\lambda})\in{G^{*}}\times_{T/W}T$ (where $\lambda\in\Lambda$), if $\lambda$ is regular (i.e. the stabilizer of $(W_{\mathrm{ex}},\bullet_{l})$-action on $\lambda$ is trivial), then there is an equivalence
$R\Gamma:D^{b}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod}_{(\chi,\zeta^{2\lambda})}){\ \xrightarrow{\sim}\ }D^{b}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(\chi,\chi_{\lambda})}).$
3.2.3. Splitting Azumaya algebras
For any $(\chi,t)\in{G^{*}}\times_{T/W}T$, we denote by ${{\mathcal{V}}}_{(\chi,t)}$ the fiber of $(\chi,t)$ in ${\mathcal{V}}$. Although the following result holds for any ${{\mathcal{V}}}_{(\chi,t)}$ with more assumptions on $l$, we only consider the case when $t=\zeta^{\lambda}$ for some $\lambda\in\Lambda$.
Theorem 3.4 ([42]).
The sheaf $\widetilde{{\mathcal{D}}}_{\zeta}$ is an Azumaya algebra on ${\mathcal{V}}$, whose restriction on the formal neighborhood of ${{\mathcal{V}}}_{(\chi,\zeta^{\lambda})}$ splits, for any $(\chi,\zeta^{\lambda})\in{G^{*}}\times_{T/W}T$. Therefore, for any splitting bundle ${\mathfrak{M}}^{\chi}_{\zeta^{\lambda}}$ of $\widetilde{{\mathcal{D}}}_{\zeta}$ on $\mathrm{FN}({\mathcal{V}}_{(\chi,\zeta^{\lambda})})$, there is an equivalence
(3.3) ${\mathfrak{M}}^{\chi}_{\zeta^{\lambda}}\otimes_{{\mathcal{O}}_{\mathrm{FN}({{\mathcal{V}}}_{(\chi,\zeta^{\lambda})})}}-:\mathrm{Coh}_{{{\mathcal{V}}}_{(\chi,\zeta^{\lambda})}}({\mathcal{V}}){\ \xrightarrow{\sim}\ }\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod}_{(\chi,\zeta^{\lambda})}.$
Let $(\chi,t)=(1,\zeta^{\lambda})\in{G^{*}}\times_{T/W}T$. Then we have
${\mathcal{V}}_{(1,\zeta^{\lambda})}={\mathcal{B}}\times\{(1,\zeta^{\lambda})\}={\mathcal{B}}\quad\text{in}\quad{\mathcal{V}}.$
We abbreviate ${\mathfrak{M}}_{\lambda}={\mathfrak{M}}^{1}_{\zeta^{2\lambda}}$.
The equivalence (3.3) above depends on the choice of splitting bundles. We will fix and recall the construction of splitting bundles ${\mathfrak{M}}_{\lambda}$ given in [42, §5.4] as follows. Firstly, we consider the completion of ${\mathfrak{U}}_{\zeta}$ at $(1,\chi_{-\rho})\in G^{*}\times_{T/W}T/(W,\bullet)$. By [15, Thms 2.5 and 4.9], the algebra $({\mathfrak{U}}_{\zeta})_{\widehat{(1,\chi_{-\rho})}}$ is a split Azumaya algebra over its center
${\mathbb{C}}[G^{*}\times_{T/W}T/(W,\bullet)]_{\widehat{(1,\chi_{-\rho})}}={\mathbb{C}}[G^{*}]_{\hat{1}}.$
Fix a splitting module $M_{\chi_{-\rho}}$ of $({\mathfrak{U}}_{\zeta})_{\widehat{(1,\chi_{-\rho})}}$, then we have an equivalence
(3.4) $M_{\chi_{-\rho}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}}-:\ \mathrm{Coh}_{1}(G^{*}){\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{-\rho})}.$
On the other hand, by [44] the natural map $\varpi^{*}{\mathfrak{U}}_{\zeta}\rightarrow\widetilde{{\mathcal{D}}}_{\zeta}$ yields an isomorphism
(3.5) $\varpi^{*}{\mathfrak{U}}_{\zeta}|_{\mathrm{FN}(1,\chi_{-\rho})}{\ \xrightarrow{\sim}\ }\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({{\mathcal{V}}}_{(1,\zeta^{-2\rho})})}.$
Hence ${\mathfrak{M}}_{{-\rho}}=\varpi^{*}M_{\chi_{-\rho}}$ is a splitting bundle for $\widetilde{{\mathcal{D}}}_{\zeta}$ on $\mathrm{FN}({{\mathcal{V}}}_{(1,\zeta^{-2\rho})})$. In general, in [43] the author constructed a locally free ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$-module ${\mathcal{O}}^{\lambda}_{{\mathcal{B}}_{\zeta}}$ of rank 1 for any $\lambda\in\Lambda$. By [42, Thm 5.3], one can set ${\mathfrak{M}}_{\lambda}={\mathcal{O}}_{{\mathcal{B}}_{\zeta}}^{\lambda+\rho}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathfrak{M}}_{{-\rho}}$.
3.2.4. Formal neighborhood of ${\mathcal{B}}$
We set $\widetilde{G}=G\times^{B}B$ where $B$ acts on itself by conjugation. There is a natural map $\widetilde{G}\rightarrow G$. Note that $\widetilde{G^{*}}$ fits into the Cartesian diagram
[…]
Since $\kappa$ is étale, $\tilde{\kappa}$ is étale. It induces isomorphisms of formal neighborhoods
$\mathrm{FN}_{{\mathcal{V}}}({\mathcal{B}}\times\{(1,\zeta^{\lambda})\})=\mathrm{FN}_{\widetilde{G^{*}}}({\mathcal{B}})=\mathrm{FN}_{\widetilde{G}}({\mathcal{B}}),\quad\forall\lambda\in\Lambda.$
The exponential map $\exp:{\mathfrak{g}}\rightarrow G$ induces isomorphisms of schemes
${\mathfrak{g}}_{\hat{0}}{\ \xrightarrow{\sim}\ }G_{\hat{1}}\quad\text{and}\quad\tilde{{\mathfrak{g}}}_{\hat{0}}{\ \xrightarrow{\sim}\ }\widetilde{G}_{\hat{1}}.$
Hence we can identify $\mathrm{FN}_{\widetilde{G}}({\mathcal{B}})=\mathrm{FN}_{\tilde{{\mathfrak{g}}}}({\mathcal{B}})$, and it follows that
$\mathrm{Coh}_{{\mathcal{B}}\times\{(1,\zeta^{2\lambda})\}}({\mathcal{V}})=\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}),\quad\forall\lambda\in\Lambda.$
Therefore, by Theorems 3.3 and 3.4, we have an equivalence
(3.6) $D^{b}\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}){\ \xrightarrow{\sim}\ }D^{b}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{0})}).$
We set ${\mathrm{A}}_{0}$ be the specialization of ${\mathrm{A}}$ at $0\in{\mathfrak{g}}$.
Proposition 3.5 ([42]).
The equivalence (3.6) is exact for the exotic t-structure on the LHS and the trivial t-structure on the RHS. Therefore, there are equivalences
(3.7) ${\mathrm{A}}\text{-}\mathrm{mod}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{0})}$
Moreover, this equivalence intertwines the reflection functors $\Theta^{c}_{s}$ and $\Theta^{r}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$), and there is a commutative diagram
[…]
By construction, the equivalence (3.7) is linear with respect to
(3.8) ${\mathbb{C}}[{\mathfrak{g}}_{\hat{0}}\times_{({\mathfrak{t}}/W)_{\hat{0}}}{\mathfrak{t}}_{\hat{0}}]{\ \xrightarrow{\sim}\ }{\mathbb{C}}[G^{*}_{\hat{1}}\times_{(T/W)_{\hat{1}}}\big{(}T/(W,\bullet)\big{)}_{\widehat{\chi_{0}}}].$
Therefore, (3.7) can be specialized to an equivalence
${\mathrm{A}}_{0}\text{-}\mathrm{mod}{\ \xrightarrow{\sim}\ }u_{\zeta}\text{-}\mathrm{mod}_{\chi_{0}}.$
3.2.5. Graded version
The constructions above can be made $T$-equivariantly as follows. Define ${\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{\lambda})}$ to be the category of $\Lambda$-graded ${\mathfrak{U}}_{\zeta}$-modules $M$ such that $M\in{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{\lambda})}$, and that for any $K_{\mu}$ and $m_{\nu}\in M_{\nu}$ there exists $n>0$ satisfying
$(K_{\mu}-\zeta^{(\mu,\nu)})^{n}.m_{\nu}=0.$
By [42, §7.4] there exists a $\Lambda$-grading (in completion sense) on $M_{\chi_{-\rho}}$. It yields an equivalence
$M_{\chi_{-\rho}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}}-:\ \mathrm{Coh}_{0}^{T}({\mathfrak{g}}){\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{-\rho})}.$
Moreover, the splitting bundles ${\mathfrak{M}}_{\lambda}$, $\lambda\in\Lambda$ are endowed with $T$-equivariant structures (in completion sense) by their constructions in §3.2.3.
We have the following graded version of Proposition 3.5.
Proposition 3.6 ([42]).
The $T$-equivariant structure on ${\mathfrak{M}}_{{0}}$ induces equivalences
(3.9) ${\mathrm{A}}\text{-}\mathrm{mod}^{T}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})},$
intertwining the reflection functors $\Theta^{c}_{s}$ and $\Theta^{r}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$). There is a commutative diagram
[…]
We can identify $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{{\mathbb{C}}}$ with the subcategory of modules in ${\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{\lambda})}$ on which $Z_{\mathrm{Fr}}$ acts trivially. Hence (3.9) can be specialized to an equivalence
${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{{\mathbb{C}}}.$
4. Equivalence ${\mathfrak{F}}^{b}$
We define the following scheme and algebras
${\mathfrak{b}}_{S}={\mathfrak{b}}\times_{\mathfrak{t}}{\mathfrak{t}}_{\hat{0}},\quad{\mathrm{A}}^{b}={\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{b}}],\quad\text{and}\quad{\mathrm{A}}^{b}_{S}={\mathrm{A}}^{b}\otimes_{{\mathbb{C}}[{\mathfrak{b}}]}{\mathbb{C}}[{\mathfrak{b}}_{S}].$
The exponential map induces isomorphisms of schemes ${\mathfrak{n}}{\ \xrightarrow{\sim}\ }N$ and ${\mathfrak{b}}_{S}{\ \xrightarrow{\sim}\ }B\times_{T}T_{\hat{1}}$. Since the equivalence (3.9) is ${\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}$-linear, it specializes to an equivalence
(4.1) ${\mathfrak{F}}^{\prime}:\ {\mathrm{A}}^{b}\text{-}\mathrm{mod}^{T}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}.$
In this section, we firstly extend (4.1) to an equivalence
${\mathfrak{F}}^{b}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$
and then investigate some properties of this equivalence.
4.1. Equivalence
Note that ${\mathrm{A}}^{b}\text{-}\mathrm{mod}^{T}_{0}$ is the full subcategory of ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ of the modules set-theoretically supported at $0\in{\mathfrak{b}}_{S}$. We also view ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}$ as a full subcategory of ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ as follows. Let $M$ be in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}$, then by definition the operator $\zeta^{-(\mu,\lambda)}K_{\mu}$ on $M_{\lambda}$ is unipotent for any $\mu,\lambda\in\Lambda$. Hence the assignment
$K_{\mu}\mapsto\bigoplus\limits_{\lambda}\zeta^{-(\mu,\lambda)}K_{\mu},\quad\forall\mu\in\Lambda$
induces an $S$-module structure on $M=\bigoplus_{\lambda}M_{\lambda}$, making $M$ as a module in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$.
Since the reflection functors $\Theta^{c}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$) on ${\mathrm{A}}\text{-}\mathrm{mod}$ are $G$-equivariant and ${\mathbb{C}}[{\mathfrak{g}}]$-linear, they induce reflection functors (still denoted by $\Theta^{c}_{s}$) on ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{K}$ by base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$, for any subgroup $K$ in $B$.
Proposition 4.1.
The equivalence ${\mathfrak{F}}^{\prime}$ in (4.1) extends to an $S$-linear equivalence
(4.2) ${\mathfrak{F}}^{b}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S},$
compatible with the reflection functors $\Theta^{c}_{s}$, $\Theta^{r}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$) and $\Lambda$-translations on both sides.
Proof.
Let ${\mathfrak{m}}_{0}$ (resp. ${\mathfrak{m}}_{1}$) be the maximal ideal of ${\mathbb{C}}[{\mathfrak{b}}]$ (resp. of $Z^{\leq}_{\mathrm{Fr}}={\mathbb{C}}[B]$) corresponding to $0\in{\mathfrak{b}}$ (resp. $1\in B$). Consider the specializations
$S_{n}=S/({\mathfrak{m}}_{1}\cap{\mathbb{C}}[T])^{n+1}S,\qquad{\mathrm{A}}^{b}_{n}={\mathrm{A}}^{b}_{S}/{\mathfrak{m}}_{0}^{n+1}{\mathrm{A}}^{b}_{S}\quad\text{and}\quad{\mathfrak{U}}^{b}_{\zeta,n}={\mathfrak{U}}^{b}_{\zeta}/{\mathfrak{m}}_{1}^{n+1}{\mathfrak{U}}^{b}_{\zeta}.$
Since ${\mathfrak{F}}^{\prime}$ is ${\mathbb{C}}[{\mathfrak{b}}]_{\hat{0}}$-linear, it yields equivalences
${\mathfrak{F}}^{\prime}:\ {\mathrm{A}}^{b}_{n}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}},\quad\forall n\geq 0.$
There are natural functors by specializations
$\tau_{n}:{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}\rightarrow{\mathrm{A}}^{b}_{n}\text{-}\mathrm{mod}^{T},\quad\tau_{n}:{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}\rightarrow{\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}},$
and ${\mathfrak{F}}^{\prime}$ is clearly compatible with $\tau_{n}$. Note that ${\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{n})$ is projective in ${\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}}$, the family $\{{\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{n})\}_{n}$ admits a lifting $P^{b}$ in $\mathrm{Proj}({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S})$. We choose a family of isomorphisms ${\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{n})\cong\tau_{n}(P^{b})$, $n\geq 0$ that are compatible with specializations. To proceed, we need the following lemmas.
Lemma 4.2.
Let $R=\bigoplus_{i\geq 0}R_{i}$ be a commutative graded ring that is finitely generated over a Noetherian complete local ring $R_{0}$. Let $I$ be a graded ideal of $R$ such that $I_{0}$ is the maximal ideal of $R_{0}$. Then for any finitely generated graded $R$-module $M$ that is complete over $R_{0}$, the natural maps
[…]
are isomorphisms.
Proof of Lemma 4.2.
Let $n_{0}$ be the minimal integer such that $M_{n_{0}}\neq 0$. By degree consideration, we have $(I^{n}M)_{i}\subseteq I_{0}^{n+n_{0}-i}M_{i}$ for any $n$ large enough. Hence
[…]
Lemma 4.3.
  1. (1)
    The natural maps
[…]
    are isomorphisms.
  2. (2)
    Let $P,M\in{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{S}$. Suppose that $P$ is projective and $M$ admits Verma flags, then the natural map
[…]
    is an isomorphism.
Proof of Lemma 4.3.
(1) Since $\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ is a projective morphism to an affine scheme, ${\mathrm{A}}=\mathrm{End}^{\mathrm{op}}_{\tilde{{\mathfrak{g}}}}({\mathcal{E}})$ is a finitely generated ${\mathbb{C}}[{\mathfrak{g}}]$-module by Serre’s Theorem. Hence ${\mathrm{A}}^{b}_{S}$ is a finitely generated ${\mathbb{C}}[{\mathfrak{b}}_{S}]$-module. Note that ${\mathrm{A}}^{b}_{S}$ is moreover $\Lambda$-graded and complete over $S$. We view ${\mathbb{C}}[{\mathfrak{b}}_{S}]$ as a ${\mathbb{Z}}_{\geq 0}$-graded ring via ${\mathbb{C}}[{\mathfrak{b}}_{S}]_{i}=\bigoplus_{\mathrm{ht}(\eta)=i}{\mathbb{C}}[{\mathfrak{b}}_{S}]_{-\eta}$, and consider the graded ideal ${\mathfrak{m}}_{0}{\mathbb{C}}[{\mathfrak{b}}_{S}]$, then our assertion follows from Lemma 4.2.
[…]
There are isomorphisms of $\Lambda$-graded algebras
[…]
where the first and the second last equalities are by Lemma 4.3. Note that the simple modules of ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ are contained in $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}={\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{0}\text{-}\mathrm{mod}^{T})$, and therefore they admit a surjection from $\bigoplus_{\nu}{\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{1}\otimes{\mathbb{C}}_{\nu})=\bigoplus_{\nu}\tau_{1}(P^{b}\otimes{\mathbb{C}}_{l\nu})$. It implies that $\{P^{b}\otimes{\mathbb{C}}_{l\nu}\}_{\nu}$ forms a generating family of projective modules in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$. Hence there is a Morita equivalence from ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ to the category of $\Lambda$-graded modules of the algebra $\bigoplus_{\nu}\mathrm{Hom}\big{(}P^{b},P^{b}\otimes{\mathbb{C}}_{l\nu})$. Now our desired equivalence (4.2) follows from the isomorphism (4.4).
By Proposition 3.6, the equivalence (4.1) is compatible with reflection functors and $\Lambda$-translations. The compatibility for (4.2) follows. ∎
We set ${\mathrm{A}}_{\mathfrak{n}}:={\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]$.
Corollary 4.4.
There is an equivalence of abelian categories
(4.5) ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}.$
4.2. Steinberg block and functor ${\mathbb{V}}^{b}$
4.2.1. Equivalence for Steinberg block
Recall the equivalence (3.4). By the ${\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}$-linearity it specializes to an equivalence
$M^{b}_{\chi_{-\rho}}\otimes_{{\mathbb{C}}[{\mathfrak{b}}]_{\hat{0}}}-:\ \mathrm{Coh}_{0}^{T}({\mathfrak{b}}){\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{-\rho})},$
where $M^{b}_{\chi_{-\rho}}=M_{\chi_{-\rho}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}}{\mathbb{C}}[{\mathfrak{b}}]_{\hat{0}}$. As in Proposition 4.1, we can extend it to an equivalence
(4.6) ${\mathfrak{G}}^{b}:\ \mathrm{Coh}^{T}({\mathfrak{b}}_{S}){\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S}.$
Let us describe ${\mathfrak{G}}^{b}$ more precisely. Note that $M^{b}_{\chi_{-\rho}}$ and $M(-\rho)_{S}\otimes_{{\mathbb{C}}[{\mathfrak{b}}_{S}]}{\mathbb{C}}[{\mathfrak{b}}]_{\hat{0}}$ are both (the unique) indecomposable projective modules of $({\mathfrak{U}}^{b}_{\zeta})_{\widehat{(1,\chi_{-\rho})}}$. Hence there is an isomorphism of $({\mathfrak{U}}^{b}_{\zeta})_{\widehat{(1,\chi_{-\rho})}}$-modules
(4.7) $M^{b}_{\chi_{-\rho}}\cong M(-\rho)_{S}\otimes_{{\mathbb{C}}[{\mathfrak{b}}_{S}]}{\mathbb{C}}[{\mathfrak{b}}]_{\hat{0}}.$
We can adjust the $\Lambda$-grading on $M_{\chi_{-\rho}}$ such that the isomorphism (4.7) respects the $\Lambda$-gradings on both sides. Therefore, the equivalence (4.6) is given by
${\mathfrak{G}}^{b}=M(-\rho)_{S}\otimes_{{\mathbb{C}}[{\mathfrak{b}}_{S}]}-.$
For any commutative Noetherian $S$-algebra $R$, the equivalence ${\mathfrak{G}}^{b}$ induces an equivalence
${\mathfrak{G}}^{b}:\ \mathrm{Coh}^{T}({\mathfrak{b}}_{R}){\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{R},$
where ${\mathfrak{b}}_{R}:={\mathfrak{b}}_{S}\times_{\mathrm{Spec}S}\mathrm{Spec}R$.
Since the functor $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}})$ is ${\mathbb{C}}[{\mathfrak{g}}]$-linear and $G$-equivariant, it induces a functor $\pi_{*}:{\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}\rightarrow\mathrm{Coh}^{T}({\mathfrak{b}}_{S})$. From Proposition 3.6 we deduce the following
Proposition 4.5.
There is a commutative diagram
[…]
Recall that by [36, §3.1.2] there is a subalgebra ${\mathfrak{U}}^{\mathrm{hb}}_{\zeta}=Z_{\mathrm{Fr}}^{-}\otimes{\mathfrak{U}}^{0}_{\zeta}\otimes U^{+}_{\zeta}$ of $U^{\mathrm{hb}}_{\zeta}$, equipped with a Frobenius map
$\mathrm{Fr}:\ {\mathfrak{U}}^{\mathrm{hb}}_{\zeta}\rightarrow{\mathbb{C}}[B\times_{T}T]\rtimes U{\mathfrak{n}}$
compatible with the quantum Frobenius map $U^{+}_{\zeta}\rightarrow U{\mathfrak{n}}$, and the isomorphism $Z^{-}_{\mathrm{Fr}}\otimes{\mathfrak{U}}^{0}_{\zeta}\rightarrow{\mathbb{C}}[B\times_{T}T]$ by (2.1), where $T\rightarrow T$ is by the $l$-th power. The map $\mathrm{Fr}$ is a homomorphism of $\Lambda$-graded algebras, where we view ${\mathbb{C}}[B\times_{T}T]\rtimes U{\mathfrak{n}}$ as an $l\Lambda$-graded ring. In loc. cit. we defined a functor for any deformation ring $R$ of $U^{\mathrm{hb}}_{\zeta}$,
${\mathrm{I}}^{\mathrm{hb}}_{1}:\ ({\mathbb{C}}[B\times_{T}T]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{\Lambda}_{R}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{R},$
$\text{by}\qquad V\mapsto U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{\mathrm{hb}}_{\zeta}}\mathrm{Fr}^{*}(V\otimes{\mathbb{C}}_{-\rho}),$
where ${\mathbb{C}}_{-\rho}$ is a one dimensional representation of ${\mathbb{C}}[B\times_{T}T]\rtimes U{\mathfrak{n}}$ whose restriction on ${\mathbb{C}}[T]={\mathbb{C}}\langle K_{\lambda}\rangle_{\lambda}$ is by the character $\zeta^{-\rho}\in T$.
When $R$ is a commutative Noetherian $S$-algebra, we will identify
$({\mathbb{C}}[B\times_{T}T]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{l\Lambda}_{R}=({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}.$
Define an induction functor
$\mathrm{ind}_{c}:=({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\otimes_{{\mathbb{C}}[{\mathfrak{b}}_{R}]}-:\ \mathrm{Coh}^{T}({\mathfrak{b}}_{R})\rightarrow({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}.$
Note that $\mathrm{ind}_{c}M=U{\mathfrak{n}}\otimes M$ as vector spaces for any ${\mathbb{C}}[{\mathfrak{b}}_{R}]$-module $M$. We also consider the induction
$\mathrm{ind}_{r}:=U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}-:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{R}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{R}.$
Note that $\mathrm{ind}_{r}$ preserves the block decomposition.
Proposition 4.6.
There is a commutative diagram
[…]
Proof.
On one hand, since $U{\mathfrak{n}}=U^{+}_{\zeta}\otimes_{u^{+}_{\zeta}}{\mathbb{C}}$ as $U^{+}_{\zeta}$-modules, we have natural isomorphisms
${\mathrm{I}}^{\mathrm{hb}}_{1}\circ\mathrm{ind}_{c}$ $=U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{\mathrm{hb}}_{\zeta}}\mathrm{Fr}^{*}\big{(}U{\mathfrak{n}}\otimes-\otimes{\mathbb{C}}_{-\rho}\big{)}$
$=U^{\mathrm{hb}}_{\zeta}\otimes_{Z^{-}_{\mathrm{Fr}}\otimes{\mathfrak{U}}^{0}_{\zeta}\otimes u^{+}_{\zeta}}({\mathbb{C}}_{-\rho}\otimes-).$
On the other hand, since $M(-\rho)_{R}={\mathfrak{U}}^{b}_{\zeta}\otimes_{{\mathfrak{U}}^{0}_{\zeta}\otimes u^{+}_{\zeta}}R_{-\rho}$ as ${\mathfrak{U}}^{b}_{\zeta}$-modules, we have natural isomorphisms
$\mathrm{ind}_{r}\circ{\mathfrak{G}}^{b}$ $=U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}(M(-\rho)_{R}\otimes_{{\mathbb{C}}[{\mathfrak{b}}_{R}]}-)$
$=U^{\mathrm{hb}}_{\zeta}\otimes_{Z_{\mathrm{Fr}}^{-}\otimes{\mathfrak{U}}^{0}_{\zeta}\otimes u_{\zeta}^{+}}(R_{-\rho}\otimes_{R}-).$
We thus obtain a natural isomorphism ${\mathrm{I}}^{\mathrm{hb}}_{1}\circ\mathrm{ind}_{c}=\mathrm{ind}_{r}\circ{\mathfrak{G}}^{b}$. ∎
4.2.2. Functors ${\mathbb{V}}^{b}_{r}$ and ${\mathbb{V}}^{b}_{c}$
Recall that ${\mathrm{A}}=\mathrm{End}^{\mathrm{op}}_{\tilde{{\mathfrak{g}}}}({\mathcal{E}})$ for some tilting bundle ${\mathcal{E}}$ on $\tilde{{\mathfrak{g}}}$. By [12, Lem 2.5.3], ${\mathcal{E}}$ has ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}}$ as a direct summand. It follows that ${\mathrm{A}}$ is an algebra over $\Gamma({\mathcal{O}}_{\tilde{{\mathfrak{g}}}})={\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$, and that ${\mathrm{A}}$ has ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ as a direct summand as a ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$-module. Moreover, the subalgebra ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ is central in ${\mathrm{A}}$. Hence ${\mathrm{A}}_{S}^{b}$ is an algebra over ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ and ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\subset Z({\mathrm{A}}_{S}^{b})$. We have ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]^{T}={\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}$. Therefore, we have an algebra homomorphism
(4.8) ${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}\rightarrow Z({\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}).$
Note that the functor $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}})$ is represented by the $({\mathbb{C}}[{\mathfrak{g}}],{\mathrm{A}})$-bimodule $\Gamma(\tilde{{\mathfrak{g}}},{\mathcal{E}})$. Hence $\pi_{*}$ can be enhanced to $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$ and induces a functor
${\mathbb{V}}^{b}_{c}:\ {\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}\rightarrow\mathrm{Coh}^{T}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$
by base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$.
We identify the rings
${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}=S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]=S\otimes_{S^{W}}S.$
Recall that by §2.3.3 there is an algebra homomorphism
(4.9) ${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}\rightarrow Z({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}).$
Hence any module in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ is naturally a module in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$. Consider the following composition
${\mathbb{V}}^{b}_{r}:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}\rightarrow{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}\xrightarrow{T^{-\rho}_{0}}{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
By the identification (3.8) and the compatibility of ${\mathfrak{F}}^{\prime}$ and ${\mathfrak{F}}^{b}$, the two maps (4.8) and (4.9) coincide under the equivalence ${\mathfrak{F}}^{b}$. Therefore, we obtain the following enhancement of Proposition 4.5.
Corollary 4.7.
There is a commutative diagram
[…]
4.3. Verma modules
In this section, we investigate the objects in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ corresponding to the Verma modules under the equivalence ${\mathfrak{F}}^{b}$, and then show that they admit natural $B$-equivariant structures.
We define an algebra structure on $(U^{\leq}_{\zeta})^{*}$ by the comultiplication of $U^{\leq}_{\zeta}$, and define an action of $U^{\leq}_{\zeta}$ on $(U^{\leq}_{\zeta})^{*}$ by the right multiplication of $U^{\leq}_{\zeta}$ on itself. Consider the quantum coordinate ring of $N^{-}$,
${\mathbb{C}}_{\zeta}[N^{-}]:=\bigoplus_{\eta\leq 0}(U^{-}_{\zeta,\eta})^{*}.$
We embed ${\mathbb{C}}_{\zeta}[N^{-}]$ into $(U^{\leq}_{\zeta})^{*}$ by $\varphi(ku)=\varepsilon(k)\varphi(u)$, for any $u\in U^{-}_{\zeta}$ and $k\in U^{0}_{\zeta}$ (here $\varepsilon$ is the counit of $U_{q}$). Then ${\mathbb{C}}_{\zeta}[N^{-}]$ is an algebra of $U^{\leq}_{\zeta}$-modules.
Set $e=B/B\in{\mathcal{B}}$ and let $U_{e}=N^{-}B/B$ be the big open cell. By [44, Proof of Prop 3.12], there is an algebra isomorphism
$\Gamma(U_{e},{\mathcal{O}}_{{\mathcal{B}}_{\zeta}})\simeq{\mathbb{C}}_{\zeta}[N^{-}],$
which is compatible via the quantum Frobenius map with the identification
$\Gamma(U_{e},{\mathcal{O}}_{\mathcal{B}})={\mathbb{C}}[N^{-}]=\bigoplus_{\eta\leq 0}\big{(}(U{\mathfrak{n}}^{-})_{\eta}\big{)}^{*}.$
The counit $\epsilon:{\mathbb{C}}_{\zeta}[N^{-}]\rightarrow{\mathbb{C}}$ by evaluating at $1\in U^{\leq}_{\zeta}$ defines an ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$-module ${\mathcal{O}}_{\epsilon}$ supported on $e$. As ${\mathcal{O}}_{{\mathcal{B}}}$-modules, ${\mathcal{O}}_{\epsilon}={\mathcal{O}}_{e}$.
Let $({\mathfrak{U}}_{\zeta}^{0})_{-2\rho}$ be the rank one ${\mathfrak{U}}_{\zeta}^{0}$-module given by the automorphism $\tau_{-2\rho}:{\mathfrak{U}}_{\zeta}^{0}\rightarrow{\mathfrak{U}}_{\zeta}^{0}$, and view it as a ${\mathfrak{U}}^{\geq}_{\zeta}$-module via the projection ${\mathfrak{U}}^{\geq}_{\zeta}\rightarrow{\mathfrak{U}}_{\zeta}^{0}$. Define the universal Verma module ${\mathbb{M}}={\mathfrak{U}}_{\zeta}\otimes_{{\mathfrak{U}}^{\geq}_{\zeta}}({\mathfrak{U}}_{\zeta}^{0})_{-2\rho}$.
Lemma 4.8.
$R\Gamma(\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathcal{O}}_{\epsilon})={\mathbb{M}}$ as ${\mathfrak{U}}_{\zeta}$-modules.
Proof.
Note that $\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathcal{O}}_{\epsilon}$ is supported on $e\in{\mathcal{B}}$. Since $U_{e}$ is affine, we have
$R\Gamma({\mathcal{B}},\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathcal{O}}_{\epsilon})=\Gamma(U_{e},\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathcal{O}}_{\epsilon})=\Gamma(U_{e},\widetilde{{\mathcal{D}}}_{\zeta})\otimes_{{\mathbb{C}}_{\zeta}[N^{-}]}{\mathbb{C}}.$
We define an algebra structure on ${\mathbb{C}}_{\zeta}[N^{-}]\otimes{\mathfrak{U}}^{\leq}_{\zeta}$ by
$u\otimes\varphi=u^{(1)}.\varphi\otimes u^{(2)},\quad\forall u\in{\mathfrak{U}}^{\leq}_{\zeta},\ \forall\varphi\in{\mathbb{C}}_{\zeta}[N^{-}].$
There is an algebra homomorphism
(4.10) $\big{(}{\mathbb{C}}_{\zeta}[N^{-}]\otimes{\mathfrak{U}}^{\leq}_{\zeta}\big{)}\otimes_{{\mathbb{C}}[N^{-}\times B]}\Gamma(U_{e},{\mathcal{O}}_{\mathcal{V}})\rightarrow\Gamma(U_{e},\widetilde{{\mathcal{D}}}_{\zeta}).$
Under the projection ${\mathcal{V}}\rightarrow G^{*}\times_{T}T$, the fiber of $e$ in ${\mathcal{V}}$ is isomorphic to $B\times_{T}T$, where $B\rightarrow T$ is given by $nt\mapsto t^{2}$, $\forall n\in N$, $\forall t\in T$, and $T\rightarrow T$ is by taking $l$-th power. We write
${\mathbb{C}}[B\times_{T}T]=Z^{\leq}_{\mathrm{Fr}}\otimes_{{\mathbb{C}}[T]}{\mathbb{C}}[T],$
where ${\mathbb{C}}[T]\rightarrow Z^{\leq}_{\mathrm{Fr}}$ is given by $k_{\mu}\mapsto K^{2l}_{\mu}$, and ${\mathbb{C}}[T]\rightarrow{\mathbb{C}}[T]$ is given by $k_{\mu}\mapsto k^{l}_{\mu}$ (here we use $\{k_{\mu}\}_{\mu\in\Lambda}$ as the canonical basis of ${\mathbb{C}}[T]$). By [44, Lem 6.3], the map (4.10) induces an isomorphism of ${\mathfrak{U}}^{-}_{\zeta}\otimes Z^{0}_{\mathrm{Fr}}$-modules
(4.11) ${\mathfrak{U}}^{-}_{\zeta}\otimes Z^{0}_{\mathrm{Fr}}\otimes_{{\mathbb{C}}[T]}{\mathbb{C}}[T]={\mathfrak{U}}^{-}_{\zeta}\otimes_{{\mathbb{C}}[N]}{\mathbb{C}}[B\times_{T}T]{\ \xrightarrow{\sim}\ }\Gamma(U_{e},\widetilde{{\mathcal{D}}}_{\zeta})\otimes_{{\mathbb{C}}_{\zeta}[N^{-}]}{\mathbb{C}}.$
We study the action of ${\mathfrak{U}}^{0}_{\zeta}$ on the LHS of (4.11). Fix basis $\{x_{r}\}_{r}$ in $U^{+}_{\zeta}$ and $\{y_{r}\}_{r}$ in ${\mathfrak{U}}^{-}_{\zeta}$ which are dual to each other under the canonical pairing between $U^{+}_{\zeta}$ and ${\mathfrak{U}}^{-}_{\zeta}$. By [44, Prop 4.13]  [^2] Because of our different conventions on the flag variety and the comultiplications of the quantum group, the formulas in loc. cit. are adapted after exchanging $E_{i}\leftrightarrow F_{i}$ and $K_{\lambda}\leftrightarrow K^{-1}_{\lambda}$. , on the LHS of (4.10), there is an equality for any dominant $\lambda\in\Lambda$,
$K^{-2}_{\lambda}=\sum_{r}\varphi_{r}\otimes y_{r}\otimes k_{\lambda},\quad\varphi_{r}:\ u\in U^{-}_{\zeta}\mapsto\langle 1^{*}_{\lambda},\mathrm{S}(ux_{r}).1_{\lambda}\rangle,$
where $1_{\lambda}$ is a highest vector of ${\mathrm{W}}(\lambda)$ (the coWeyl module of $U_{\zeta}$ of highest weight $\lambda$), and $1^{*}_{\lambda}\in{\mathrm{W}}(\lambda)^{*}$ is given by $\langle 1^{*}_{\lambda},1_{\lambda}\rangle=1$ and $\langle 1^{*}_{\lambda},{\mathrm{W}}(\lambda)_{\mu}\rangle=0$ if $\mu\neq\lambda$. Modulo the augmentation ideal of ${\mathbb{C}}_{\zeta}[N^{-}]$ from the right, we have
(4.12) $K^{-2}_{\lambda}$ $=\sum_{r}y_{r}^{(2)}\cdot\mathrm{S}^{-1}(y_{r}^{(1)}).\varphi_{r}\otimes k_{\lambda}$
$\equiv\sum_{r}\langle 1^{*}_{\lambda},\mathrm{S}(x_{r})y_{r}^{(1)}.1_{\lambda}\rangle\cdot y_{r}^{(2)}\cdot k_{\lambda}$
$=\sum_{r}\langle 1^{*}_{\lambda},\mathrm{S}(x_{r})y_{r}.1_{\lambda}\rangle\cdot k_{\lambda}$
$=\zeta^{4(\rho,\lambda)}k_{\lambda}\ ,$
where the third equality is because that $\Delta(y_{r})=y_{r}\otimes 1+\sum_{\text{deg}(y_{r}^{(2)})<0}y_{r}^{(1)}\otimes y_{r}^{(2)}$, and the last equality is because that the quantum Casimir element $\sum_{r}\mathrm{S}(x_{r})y_{r}$ acts on $1_{\lambda}\in{\mathrm{W}}(\lambda)$ by scalar $\zeta^{4(\rho,\lambda)}$. Finally, we identify
$Z^{0}_{\mathrm{Fr}}\otimes_{{\mathbb{C}}[T]}{\mathbb{C}}[T]={\mathfrak{U}}^{0}_{\zeta},\quad\text{by}\quad K^{l}_{\mu}\otimes k_{\lambda}\mapsto K_{l\mu+2\lambda},\quad\forall\mu,\lambda\in\Lambda.$
Then we have ${\mathbb{M}}=\Gamma(U_{e},\widetilde{{\mathcal{D}}}_{\zeta})\otimes_{{\mathbb{C}}_{\zeta}[N^{-}]}{\mathbb{C}}$ as ${\mathfrak{U}}^{\leq}_{\zeta}$-modules. It is in fact an isomorphism of ${\mathfrak{U}}_{\zeta}$-modules, since these two modules are $\Lambda$-graded. ∎
Recall the tilting bundle ${\mathcal{E}}$ in $\tilde{{\mathfrak{g}}}$. Let ${\mathcal{E}}^{\vee,b}$ be the restriction of the dual bundle ${\mathcal{E}}^{\vee}$ on $\{e\}\times{\mathfrak{b}}$ in $\tilde{{\mathfrak{g}}}$, and set ${\mathcal{E}}^{\vee,b}_{S}={\mathcal{E}}^{\vee,b}\otimes_{{\mathbb{C}}[{\mathfrak{t}}]}S$, which is an ${\mathrm{A}}^{b}_{S}$-module.
Proposition 4.9.
We have ${\mathfrak{F}}^{b}({\mathcal{E}}^{\vee,b}_{S})=M(-2\rho)_{S}$.
Proof.
As a sheave on ${\mathcal{V}}$, $\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathcal{O}}_{\epsilon}$ is supported on the fiber of $e\in{\mathcal{B}}$ in ${\mathcal{V}}$, which is identified with $B\times_{T}T$ in the proof of the lemma above. We denote by $e_{n}$ the $n$-th infinitesimal neighborhood of 00 in ${\mathfrak{b}}$. We denote by $e^{\lambda}_{n}$ the $n$-th infinitesimal neighborhood of $(1,\zeta^{2\lambda})$ in $B\times_{T}T$, and view it as a closed subscheme in ${\mathcal{V}}$. Consider the subsheaf of algebras ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}$ in $\widetilde{{\mathcal{D}}}_{\zeta}$. The ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$-module ${\mathcal{O}}_{\epsilon}$ and the ${\mathcal{O}}_{\mathcal{V}}$-module ${\mathcal{O}}_{e^{\lambda}_{n}}$ define a sheaf of module ${\mathcal{O}}_{(\epsilon,e^{\lambda}_{n})}$ of ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}$. We abbreviate $\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{\lambda}_{n})}=\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{\lambda}_{n})}$. By Lemma 4.8 and formula (4.12), we compute that
(4.13) $R\Gamma(\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{\lambda}_{n})})={\mathbb{M}}/{\mathcal{I}}_{e^{\lambda}_{n}}{\mathbb{M}}=M(-\lambda-2\rho)_{S}/{\mathfrak{m}}_{1}^{n}M(-\lambda-2\rho)_{S},$
as ${\mathfrak{U}}_{\zeta}$-modules, where ${\mathcal{I}}_{e^{\lambda}_{n}}$ is the ideal in ${\mathbb{C}}[B\times_{T}T]$ associated to $e^{\lambda}_{n}$, and ${\mathfrak{m}}_{1}$ is the maximal ideal of $1\in B$ in ${\mathbb{C}}[B]$.
In particular we have $R\Gamma(\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{0}_{n})})=M(-2\rho)_{S}/{\mathfrak{m}}_{1}^{n}M(-2\rho)_{S}$ as ${\mathfrak{U}}_{\zeta}$-modules. Consider the equivalence
$R\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{0},-):D^{b}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod}_{(1,1)}){\ \xrightarrow{\sim}\ }D^{b}\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}).$
We claim that it sends $\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{0}_{n})}$ to ${\mathcal{O}}_{e_{n}}$. Indeed, we have
$R\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{0},\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{0}_{n})})$ $={\mathfrak{M}}_{0}^{\vee}\otimes_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{0}_{n})}$
$={\mathfrak{M}}_{0}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{0}_{n})}.$
By construction that ${\mathfrak{M}}_{0}={\mathcal{O}}^{\rho}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathfrak{M}}_{{-\rho}}$ (see §3.2.3), we obtain that
${\mathfrak{M}}_{0}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{0}_{n})}={\mathfrak{M}}_{{-\rho}}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{-\rho}_{n})}$
as sheaves on ${\mathcal{B}}$. Similarly we have
${\mathfrak{M}}_{{-\rho}}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{-\rho}_{n})}$ $=R\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{{-\rho}},\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{-\rho}_{n})})$
$=\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{{-\rho}}|_{e^{-\rho}_{n}},\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{-\rho}_{n})}).$
By (4.13), we have $R\Gamma(\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{-\rho}_{n})})=M(-\rho)_{S}/{\mathfrak{m}}_{1}^{n}M(-\rho)_{S}$. Using the isomorphisms (3.5) and (4.7), we have
(4.14) $\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{{-\rho}}|_{e^{-\rho}_{n}},\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{-\rho}_{n})})=\mathrm{End}_{({\mathfrak{U}}_{\zeta})_{\widehat{1,\chi_{-\rho}}}}\big{(}M(-\rho)_{S}/{\mathfrak{m}}_{1}^{n}M(-\rho)_{S}\big{)}.$
By the equivalence (4.6), the RHS of (4.14) is equal to ${\mathcal{O}}_{e_{n}}$ (note that we identify the formal neighborhood of 00 in ${\mathfrak{b}}_{S}$ with the formal neighborhood of $(1,1)$ in $B\times_{T}T$).
Finally, under the equivalence (3.1), the sheaf ${\mathcal{O}}_{e_{n}}$ corresponds to the ${\mathrm{A}}$-module ${\mathcal{E}}^{\vee}\otimes_{{\mathcal{O}}_{\tilde{{\mathfrak{g}}}}}{\mathcal{O}}_{e_{n}}={\mathcal{E}}^{\vee,b}_{S}/{\mathfrak{m}}^{n}_{0}{\mathcal{E}}^{\vee,b}_{S}$, where ${\mathfrak{m}}_{0}$ is the maximal ideal of $0\in{\mathfrak{b}}$ in ${\mathbb{C}}[{\mathfrak{b}}]$. Hence we have
${\mathfrak{F}}^{b}({\mathcal{E}}^{\vee,b}_{S}/{\mathfrak{m}}_{0}^{n}{\mathcal{E}}^{\vee,b}_{S})=M(-2\rho)_{S}/{\mathfrak{m}}_{1}^{n}M(-2\rho)_{S},\quad\forall n\geq 0.$
Taking the limits, we obtain that ${\mathfrak{F}}^{b}({\mathcal{E}}^{\vee,b}_{S})=M(-2\rho)_{S}$. ∎
Definition 4.10.
We define the Verma objects in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ to be
${\mathcal{M}}_{x}=({\mathfrak{F}}^{b})^{-1}(M(x\bullet_{l}0)_{S}),\quad x\in W_{\mathrm{ex}}.$
An object in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{Mod}^{T}$ admits Verma flags if it admits a (decreasing, separated, ${\mathbb{Z}}_{\leq 0}$-) filtration with composition factors given by Verma objects.
Corollary 4.11.
Any Verma object in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ admits a natural $B$-equivariant structure.
Proof.
By Proposition 4.9, ${\mathcal{M}}_{w_{0}}={\mathcal{E}}^{\vee,b}_{S}$ is naturally a $B$-equivariant ${\mathrm{A}}^{b}_{S}$-module. We show that in general
(4.15) ${\mathcal{M}}_{t(\nu)w_{0}w^{-1}}={\mathbb{T}}^{c}_{w}({\mathcal{M}}_{w_{0}})\otimes{\mathbb{C}}_{\nu},\quad\forall w\in W,\ \forall\nu\in\Lambda,$
where ${\mathbb{T}}^{c}_{w}={\mathbb{T}}^{c}_{s_{1}}\cdots{\mathbb{T}}^{c}_{s_{n}}$ for a reduced expression $w=s_{1}\cdots s_{n}$. Then since ${\mathbb{T}}^{c}_{w}$ is $B$-equivariant, any Verma object is endowed with a $B$-equivariant structure.
Indeed, for any $w\in W$ and $s\in{\mathrm{I}}$ such that $ws<w$ in the Bruhat order, there is a short exact sequence (see e.g. [23, Thm 7.14(a)])
[…]
which shows that $M(ws\bullet 0)_{S}={\mathbb{T}}^{r}_{s}(M(w\bullet 0)_{S})$. By Proposition 4.1 that the equivalence ${\mathfrak{F}}^{b}$ intertwines the reflection functors on both sides, we have
(4.16) ${\mathcal{M}}_{ws}={\mathbb{T}}^{c}_{s}({\mathcal{M}}_{w}),\quad\text{if}\quad ws<w.$
Apply (4.16) several times on ${\mathcal{M}}_{w_{0}}$ and use the $\Lambda$-translations, we obtain (4.15). ∎
4.4. Fully-faithfulness
For any Noetherian commutative $S$-algebra $R$, one can define the functor
${\mathbb{V}}^{b}_{r,R}:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{R}\rightarrow{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{R\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$
as in §4.2.2. Let ${\mathbb{K}}={\mathrm{Frac}}(S)$ be the fraction field of $S$.
Lemma 4.12.
  1. (1)
    The functor ${\mathbb{V}}^{b}_{r,R}$ is faithful on the objects admitting Verma flags.
  2. (2)
    The functor ${\mathbb{V}}^{b}_{r,{\mathbb{K}}}$ is an equivalence of categories.
Proof.
(1) By definition of ${\mathbb{V}}^{b}_{r,R}$ we have an inclusion
$\mathrm{Hom}({\mathbb{V}}^{b}_{r,R}M_{1},{\mathbb{V}}^{b}_{r,R}M_{2})\hookrightarrow\mathrm{Hom}(T^{-\rho}_{0}M_{1},T^{-\rho}_{0}M_{2}),\quad\forall M_{i}.$
Hence the faithfulness of ${\mathbb{V}}^{b}_{r,R}$ (on objects admitting Verma flags) follows from the faithfulness of $T^{-\rho}_{0}$, see [36, Lem 4.9(2)].
(2) There is an isomorphism
(4.17) ${\mathbb{K}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]\simeq\prod_{w\in W}{\mathbb{K}}.$
It induces a decomposition
(4.18) ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{{\mathbb{K}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}=\bigoplus_{w\in W}{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{\mathbb{K}}.$
Moreover, the isomorphism (4.17) together with the natural map induced from (4.9)
${\mathbb{K}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]\rightarrow Z({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{K}})$
yields a block decomposition
(4.19) ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{K}}=\bigoplus_{w\in W}{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0,w}_{\mathbb{K}},$
such that a Verma module $M(\lambda)_{\mathbb{K}}$ is contained in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0,w}_{\mathbb{K}}$ if and only if $\lambda-w\bullet 0\in l\Lambda$. Under the decompositions (4.18) and (4.19), we can write
${\mathbb{V}}^{b}_{r,{\mathbb{K}}}=\bigoplus_{w\in W}T^{-\rho}_{0}\circ\iota_{w},$
where $\iota_{w}$ is the inclusion from the block labelled by $w$ in (4.19). The functor $T^{-\rho}_{0}\circ\iota_{w}$ is an equivalence with quasi-inverse given by $\mathrm{pr}_{w}\circ T^{0}_{-\rho}$, where $\mathrm{pr}_{w}$ is the corresponding block projection. Hence ${\mathbb{V}}^{b}_{r,{\mathbb{K}}}$ is an equivalence of categories. ∎
Proposition 4.13.
The functors ${\mathbb{V}}^{b}_{r}$ and ${\mathbb{V}}^{b}_{c}$ are fully-faithful on the objects admitting Verma flags.
Proof.
By Corollary 4.7, we only have to prove the assertion for ${\mathbb{V}}^{b}_{r}$. It follows standardly from Lemma 4.12 by the general theory of highest weight categories, see e.g. [35, Prop 2.18]. For readers’ convenience, we present the proof as follows.
We abbreviate ${\mathbb{V}}={\mathbb{V}}^{b}_{r,R}$ and $M_{R}=M\otimes_{S}R$ for any $S$-module $M$ and any $S$-algebra $R$. Let $M_{1},M_{2}$ be objects in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ admitting Verma flags. In particular they are free as $S$-modules. Hence the base change yields an embedding $\mathrm{Hom}(M_{1},M_{2})\subseteq\mathrm{Hom}(M_{1,{\mathbb{K}}},M_{2,{\mathbb{K}}})$. By [32, Thm 38], we have $M_{i}=\bigcap\limits_{\mathrm{ht}(\mathfrak{p})=1}M_{i,S_{\mathfrak{p}}}$ as subspaces in $M_{i,{\mathbb{K}}}$ (the intersection is taken for height 1 prime ideals $\mathfrak{p}$ of $S$). Hence we can identify
(4.20) $\mathrm{Hom}(M_{1},M_{2})=\bigcap_{\mathrm{ht}(\mathfrak{p})=1}\mathrm{Hom}(M_{1,S_{\mathfrak{p}}},M_{2,S_{\mathfrak{p}}})$
as subspaces in $\mathrm{Hom}(M_{1,{\mathbb{K}}},M_{2,{\mathbb{K}}})$. Similarly, since ${\mathbb{V}}M_{i}=T^{-\rho}_{0}M_{i}$ are free $S$-modules, we have
(4.21) $\mathrm{Hom}({\mathbb{V}}M_{1},{\mathbb{V}}M_{2})=\bigcap_{\mathrm{ht}(\mathfrak{p})=1}\mathrm{Hom}({\mathbb{V}}M_{1,S_{\mathfrak{p}}},{\mathbb{V}}M_{2,S_{\mathfrak{p}}})$
as subspaces in $\mathrm{Hom}({\mathbb{V}}M_{1,{\mathbb{K}}},{\mathbb{V}}M_{2,{\mathbb{K}}})$.
By (4.20) and (4.21), it suffices to prove the isomorphism
$\varphi:\mathrm{Hom}(M_{1,S_{\mathfrak{p}}},M_{2,S_{\mathfrak{p}}}){\ \xrightarrow{\sim}\ }\mathrm{Hom}({\mathbb{V}}M_{1,S_{\mathfrak{p}}},{\mathbb{V}}M_{2,S_{\mathfrak{p}}}),$
[…]
Lemma 4.14.
Let $R\rightarrow R^{\prime}$ be a homomorphism of Noetherian commutative $S$-algebras. Suppose that $R^{\prime}$ is flat over $R$. For any $M,M^{\prime}\in{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{R}$, there is a natural isomorphism
$\mathrm{Hom}(M,M^{\prime})\otimes_{R}R^{\prime}\simeq\mathrm{Hom}(M\otimes_{R}R^{\prime},M^{\prime}\otimes_{R}R^{\prime}).$
Proof of Lemma 4.14.
Choose a projective resolution $P_{1}\rightarrow P_{2}\rightarrow M\rightarrow 0$, we have a commutative diagram with exact rows
00{0}${\mathrm{Hom}(M,M^{\prime})_{R^{\prime}}}$${\mathrm{Hom}(P_{2},M^{\prime})_{R^{\prime}}}$${\mathrm{Hom}(P_{1},M^{\prime})_{R^{\prime}}}$00{0}${\mathrm{Hom}(M_{R^{\prime}},M^{\prime}_{R^{\prime}})}$${\mathrm{Hom}(P_{2,R^{\prime}},M^{\prime}_{R^{\prime}})}$${\mathrm{Hom}(P_{1,R^{\prime}},M^{\prime}_{R^{\prime}}),}$
where the subscript $R^{\prime}$ means $-\otimes_{R}R^{\prime}$. By [20, Prop 2.4], the right two vertical maps are isomorphisms, so is the left one. ∎
Consider the commutative diagram
[…]
where the vertical isomorphisms are by Lemma 4.14 applied on the tuple $(M,M^{\prime},R,R^{\prime})$ $=(M_{1,S_{\mathfrak{p}}},M_{2,S_{\mathfrak{p}}},S_{\mathfrak{p}},{\mathbb{K}})$ or $({\mathbb{V}}M_{1,S_{\mathfrak{p}}}$, ${\mathbb{V}}M_{2,S_{\mathfrak{p}}}$, $S_{\mathfrak{p}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$, ${\mathbb{K}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}])$, and the lower horizontal isomorphism is by Lemma 4.12(2). Hence $\varphi\otimes_{S_{\mathfrak{p}}}{\mathbb{K}}$ is an isomorphism. We conclude that $\varphi$ is an isomorphism. ∎
Remark 4.15.
For the non-deformed category ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}$, the functor ${\mathbb{V}}^{b}_{r,{\mathbb{C}}}$ fails to be full-faithful on objects admitting Verma flags.
5. Main result: equivalence ${\mathfrak{F}}$
5.1. Functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$
Recall that in §2.3.3, we construct an algebra homomorphism
${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}\rightarrow Z(U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}).$
As discussions in §4.2.2, the translation functor $T^{-\rho}_{0}:U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,-\rho}_{S}$ can be enhanced to a functor
${\mathbb{V}}_{r}:\ U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
It restricts to a functor
${\mathbb{V}}_{r}:{\mathcal{O}}_{S}^{0}\rightarrow{\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
In fact, ${\mathbb{V}}_{r}$ is equivalent to the quotient functor $\bigoplus\limits_{\lambda\in\Lambda}\mathrm{Hom}(Q(w_{0}\bullet 0+l\lambda)_{S},-)$ (we will not use this below).
Recall that the functor $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$ is represented by the $G$-equivariant $\big{(}{\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}],{\mathrm{A}}\big{)}$-bimodule $\Gamma(\tilde{{\mathfrak{g}}},{\mathcal{E}})$. We can define a functor
${\mathbb{V}}_{c}:({\mathrm{A}}_{S}^{b}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow({\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$
as ${\mathbb{V}}^{b}_{c}$ in §4.2.2. It restricts to a functor
${\mathbb{V}}_{c}:{\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{B}\rightarrow\mathrm{Coh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}).$
For any $s\in{\mathrm{I}}$, one can define the Soergel bimodule
$B_{s}=\big{(}{\mathbb{C}}[{\mathfrak{t}}]\otimes_{{\mathbb{C}}[{\mathfrak{t}}]^{s}}{\mathbb{C}}[{\mathfrak{t}}]\big{)}_{\hat{0}}$
in $\mathrm{Coh}\big{(}({\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})_{\hat{0}}\big{)}$, where $\hat{0}$ represents the completion at $0\in{\mathfrak{t}}/W$. Tensoring with $B_{s}$ defines a functor on $\mathrm{Coh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$ and ${\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$.
It is automatic that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are compatible with $\Lambda$-translations. We now show that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are also compatible with reflection functors associated with $s\in{\mathrm{I}}$.
5.1.1. Compatibility with reflection functors
Let $\omega\in\Xi$. Consider the composition of functors
${\mathcal{O}}^{\omega}_{S}\xrightarrow{T_{\omega}^{0}}{\mathcal{O}}^{0}_{S}\rightarrow{\mathcal{O}}^{0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}\xrightarrow{T^{\omega}_{0}}{\mathcal{O}}^{\omega}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]},$
where the middle one is induced by the homomorphism $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]\rightarrow Z({\mathcal{O}}^{0}_{S})$.
Lemma 5.1 ([37, Prop B.9]).
There is a natural isomorphism
$T_{0}^{\omega}\circ T_{\omega}^{0}\simeq-\otimes_{S^{W_{\omega}}}S$
of functors from ${\mathcal{O}}^{\omega}_{S}$ to ${\mathcal{O}}^{\omega}_{S\otimes_{S^{W}}S}$.
Proposition 5.2.
For any $s\in{\mathrm{I}}$, there is a natural isomorphism
${\mathbb{V}}_{r}\circ\Theta^{r}_{s}=(-\otimes_{S}B_{s})\circ{\mathbb{V}}_{r}$
of functors from ${\mathcal{O}}^{0}_{S}$ to ${\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$.
Proof.
From the lemma above we deduce that
$T^{-\rho}_{0}\circ\Theta^{r}_{s}$ $=T^{-\rho}_{s}\circ T^{s}_{0}\circ T^{0}_{s}\circ T_{0}^{s}$
$=T^{-\rho}_{s}\big{(}T_{0}^{s}(-)\otimes_{S^{s}}S\big{)}$
$=\Theta^{r}_{s}(-)\otimes_{S^{s}}S,$
where for the first and the last equality we use a natural isomorphism $T_{0}^{-\rho}=T_{s}^{-\rho}\circ T_{0}^{s}$. ∎
Proposition 5.3.
For any $s\in{\mathrm{I}}$, there is a natural isomorphism
${\mathbb{V}}_{c}\circ\Theta^{c}_{s}=(-\otimes_{S}B_{s})\circ{\mathbb{V}}_{c},$
such that the counit $\Theta^{c}_{s}\rightarrow 1$ corresponds to the bimodule homomorphism $B_{s}\rightarrow S$ by multiplication.
Proof.
Let $P$ be any (standard) parabolic subgroup of $G$, and let $W_{P}$ be the corresponding parabolic subgroup in $W$. Denote by $\pi^{P}:\tilde{{\mathfrak{g}}}_{P}\rightarrow{\mathfrak{g}}$ the natural projection. There is a ($G$-equivariant) natural isomorphism
$\pi_{P,*}\pi_{P}^{*}\simeq-\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W_{P}]}{\mathbb{C}}[{\mathfrak{t}}]\qquad\text{on}\quad D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{P}).$
Hence we have the following ($G$-equivariant) natural isomorphisms
$\pi_{*}\circ\Theta^{c}_{P}$ $=\pi^{P}_{*}(\pi_{P,*}\pi_{P}^{*})\pi_{P,*}$
$\simeq\pi^{P}_{*}\circ(-\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W_{P}]}{\mathbb{C}}[{\mathfrak{t}}])\circ\pi_{P,*}$
$=\pi_{*}(-)\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W_{P}]}{\mathbb{C}}[{\mathfrak{t}}]$
of functors from $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}})$ to $D^{b}\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$. Consider the special case when $P=P_{s}$, and then it implies our desired isomorphism by base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$. The second assertion is easy to verify. ∎
Denote by ${}^{w}S=S$ the $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-module given by
$S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]\rightarrow S,\quad f\otimes g\mapsto fw(g),\quad\forall f\in S,\ \forall g\in{\mathbb{C}}[{\mathfrak{t}}].$
Recall that the Soergel bimodule $B_{s}$ admits a short exact sequence
(5.1) $0\rightarrow{}^{s}S\rightarrow B_{s}\rightarrow S\rightarrow 0,$
where the first map is induced by $1\mapsto\alpha_{s}\otimes 1-1\otimes\alpha_{s}$ and the second map is by multiplication.
Lemma 5.4.
For $x=t(\lambda)w_{0}w\in W_{\mathrm{ex}}$ with $w\in W$ and $\lambda\in\Lambda$, we have an isomorphism
${\mathbb{V}}_{c}({\mathcal{M}}_{x})={\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w}S\otimes{\mathbb{C}}_{\lambda}.$
Proof.
Firstly we have
${\mathbb{V}}_{c}({\mathcal{M}}_{w_{0}})={\mathbb{V}}_{c}({\mathcal{E}}^{\vee,b}_{S})=\pi_{*}{\mathcal{E}}^{\vee,b}_{S}={\mathcal{O}}_{{\mathfrak{b}}_{S}},$
on which the $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-action factors through the multiplication $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]\rightarrow S$. By the short exact sequence (5.1) and Proposition 5.3, we have a natural isomorphism ${\mathbb{V}}_{c}\circ{\mathbb{T}}^{c}_{s}={\mathbb{V}}_{c}(-)\otimes_{S}{}^{s}S$ for any $s\in{\mathrm{I}}$. It follows that
${\mathbb{V}}_{c}\circ{\mathbb{T}}^{c}_{w}={\mathbb{V}}_{c}(-)\otimes_{S}{}^{w^{-1}}S,\quad\forall w\in W.$
Now we use (4.15) to obtain our desired formula. ∎
5.2. Truncated categories
In this section, we define truncated subcategories in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ which will be shown to correspond to (certain) truncations in ${\mathcal{O}}^{0}_{S}$ defined in §2.3.1 under the equivalence (1.4), and we will study the projective objects in them.
We firstly define the truncation of ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ by $\nu\in\Lambda$ as
${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\leq\nu}:=({\mathfrak{F}}^{b})^{-1}({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0,\leq l\nu}_{S}).$
It is a Serre subcategory of ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$, and each object in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admits a finite decomposition with each direct factor contained in the truncated subcategory by some $\nu$. We define
${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu},\quad\nu\in\Lambda$
to be the full subcategory of modules in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ which are contained in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\leq\nu}$, forgetting the $B$-equivariant structure. Then ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ is a Serre subcategory in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$, and each object in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admits a finite decomposition with each direct factor contained in the truncated subcategory by some $\nu$.
Let $\leq^{\frac{\infty}{2}}$ be the semi-infinite order on $W_{\mathrm{af}}$, which is a partial order generated by
$xs\leq^{\frac{\infty}{2}}x\quad\text{if}\quad x\bullet_{l}0\geq xs\bullet_{l}0,\quad x\in W_{\mathrm{af}},\ s\in{\mathrm{I}}_{\mathrm{af}}.$
Note that $\leq^{\frac{\infty}{2}}$ is invariant under $\Lambda$-translation from the left. It extends naively to a partial order $\leq^{\frac{\infty}{2}}$ on $W_{\mathrm{ex}}=W_{\mathrm{af}}\rtimes\Lambda/{\mathrm{Q}}$ such that two elements are comparable only if they are contained in the same $\Lambda/{\mathrm{Q}}$-coset. Note that $\leq^{\frac{\infty}{2}}$ is independent of $l$. We have the following proposition, whose proof is postponed to §5.2.1.
Proposition 5.5.
We have $\mathrm{Ext}^{i}_{{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}}({\mathcal{M}}_{x},{\mathcal{M}}_{x^{\prime}})\neq 0$ only if $x\leq^{\frac{\infty}{2}}x^{\prime}$.
Consider the induction functor
$\mathrm{ind}_{c}:=({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}^{b}_{S}}-:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}\rightarrow({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}.$
Note that $\mathrm{ind}_{c}(M)=U{\mathfrak{n}}\otimes M$ as vector spaces, for any ${\mathrm{A}}^{b}_{S}$-module $M$.
Lemma 5.6.
For any objects $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admitting Verma flags, the induction $\mathrm{ind}_{c}(P)$ also admits Verma flags.
Proof.
It is enough to prove the assertion when $P$ is a Verma object. By Corollary 4.11, $P$ is a module of ${\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}}$. Consider the homomorphism of $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-modules
(5.2) $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}^{b}_{S}}P=U{\mathfrak{n}}\otimes P\rightarrow P\otimes U{\mathfrak{n}},$
$u\otimes m\mapsto u^{(1)}.m\otimes u^{(2)},\quad\forall u\in U{\mathfrak{n}},\ \forall m\in P,$
where $P\otimes U{\mathfrak{n}}$ is viewed as an $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-module by ${\mathrm{A}}^{b}_{S}$ acting on $P$ and $U{\mathfrak{n}}$ acting on $P\otimes U{\mathfrak{n}}$ diagonally. The homomorphism (5.2) is an isomorphism, whose inverse is given by $m\otimes u\mapsto u^{(2)}\otimes(\mathrm{S}u^{(1)}).m$ ($\mathrm{S}$ is the antipode for $U{\mathfrak{n}}$). We obtain an isomorphism $\mathrm{ind}_{c}(P)\simeq P\otimes U{\mathfrak{n}}$ of $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-modules. It is clear that $P\otimes U{\mathfrak{n}}$ admits a Verma flag with Verma factors $P\otimes{\mathbb{C}}_{\eta}$ for $\eta\geq 0$. ∎
Note the subset $\{t(\lambda)w\}_{w\in W,\lambda\leq\nu}$ in an ideal in $(W_{\mathrm{ex}},\leq^{\frac{\infty}{2}})$. By the lemma above and Proposition 5.5, for any object $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admitting Verma flags, one can define the quotients
$\mathrm{ind}_{c}(P)^{\leq\nu},\quad\nu\in\Lambda$
of $\mathrm{ind}_{c}(P)$ modulo the submodule with Verma factors in $\{{\mathcal{M}}_{t(\lambda)w}\}_{w\in W,\ \lambda\nleq\nu}$. Then the $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-modules $\mathrm{ind}_{c}(P)^{\leq\nu}$ are contained in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$.
Lemma 5.7.
The family $\{\mathrm{ind}_{c}(P)^{\leq\nu}\}$ with $P$ running over a generating family of projective objects in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ forms a generating family of projective objects in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$.
Proof.
Let $x=t(\lambda)w$ where $\lambda\in\Lambda$ and $w\in W$, then the truncation $M(x\bullet_{l}0)^{\leq l\nu}_{S}\neq 0$ only if $\lambda\leq\nu$. Hence the functor $\mathrm{Hom}({\mathcal{M}}_{x},-)$ is nonzero on ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\leq\nu}$ only if $\lambda\leq\nu$. It shows that any homomorphism from $\mathrm{ind}_{c}(P)$ to an object in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ factors through the quotient $\mathrm{ind}_{c}(P)^{\leq\nu}$. It follows natural isomorphisms
(5.3) $\mathrm{Hom}_{{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}}(\mathrm{ind}_{c}(P)^{\leq\nu},-)$ $=\mathrm{Hom}_{({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}}(\mathrm{ind}_{c}(P),-)$
$=\mathrm{Hom}_{{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}}(P,-)$
of functors on ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$. Hence $\mathrm{ind}_{c}(P)^{\leq\nu}$ is projective in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$. On the other hand, for any module $M$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$, there exists a surjection of $\Lambda$-graded ${\mathrm{A}}^{b}_{S}$-modules $P\twoheadrightarrow M$ for some projective $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. Then the induced homomorphism $\mathrm{ind}_{c}(P)\twoheadrightarrow\mathrm{ind}_{c}(P)^{\leq\nu}\rightarrow M$ is a surjection. So we obtain a generating family of projective objects in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$. ∎
Lemma 5.8.
Any $M\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admits a maximal quotient $M^{\leq\nu}$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$. It defines a functor
$(-)^{\leq\nu}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}\rightarrow{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$
left adjoint to the inclusion functor. We have a natural isomorphism
${\mathbb{V}}_{c}\circ(-)^{\leq\nu}=(-)^{\leq\nu}\circ{\mathbb{V}}_{c}.$
Proof.
We may assume $M\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu^{\prime}}$ for some $\nu^{\prime}>>\nu$. By Lemma 5.7, we can choose a presentation $\mathrm{ind}_{c}(P_{1})^{\leq\nu^{\prime}}\rightarrow\mathrm{ind}_{c}(P_{2})^{\leq\nu^{\prime}}\rightarrow M\rightarrow 0$, for some projective object $P_{i}$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. It induces a homomorphism $\mathrm{ind}_{c}(P_{1})^{\leq\nu}\rightarrow\mathrm{ind}_{c}(P_{2})^{\leq\nu}$, and we set $M^{\leq\nu}$ to be its cokernel. Then $M^{\leq\nu}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$. For any $M^{\prime}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$, we have a commutative diagram with exact rows
00{0}${\mathrm{Hom}(M,M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{2})^{\leq\nu^{\prime}},M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{1})^{\leq\nu^{\prime}},M^{\prime})}$00{0}${\mathrm{Hom}(M^{\leq\nu},M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{2})^{\leq\nu},M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{1})^{\leq\nu},M^{\prime}),}$
where the two vertical equalities follow from the first equality in (5.3). Hence the left vertical map is an isomorphism. It shows that $M^{\leq\nu}$ is the maximal quotient of $M$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$, and therefore $(-)^{\leq\nu}$ defines a functor left adjoint to the inclusion.
For the last assertion, by the right exactness of ${\mathbb{V}}_{c}\circ(-)^{\leq\nu}$ and $(-)^{\leq\nu}\circ{\mathbb{V}}_{c}$, it is enough to construct natural isomorphism ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P)^{\leq\nu})={\mathbb{V}}_{c}(\mathrm{ind}_{c}(P))^{\leq\nu}$ for any projective module $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. By Lemma 5.4, the module ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P))$ admits a filtration with composition factors given by ${\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w}S\otimes{\mathbb{C}}_{\lambda}$ for some $\lambda\in\Lambda$ and $w\in W$. By weight consideration, we have
$\mathrm{Ext}^{1}({\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w}S\otimes{\mathbb{C}}_{\lambda},{\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w^{\prime}}S\otimes{\mathbb{C}}_{\mu})\neq 0\quad\text{only if}\quad\lambda<\mu,\quad\forall w,w^{\prime}\in W.$
Note that ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P)^{\leq\nu})$ is the quotient by the submodule extended by the factors ${\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w}S\otimes{\mathbb{C}}_{\lambda}$ with $\lambda\nleq\nu$, which is exactly the truncation ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P))^{\leq\nu}$. ∎
We have a parallel result of Lemma 5.8.
Lemma 5.9.
There is a natural isomorphism
${\mathbb{V}}_{r}\circ(-)^{\leq l\nu}=(-)^{\leq l\nu}\circ{\mathbb{V}}_{r}$
of functors from ${\mathcal{O}}^{0}_{S}$ to ${\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$.
Proof.
By [36, Lem 4.3(3)] we have a natural isomorphism $T^{-\rho}_{0}\circ(-)^{\leq l\nu}=(-)^{\leq l\nu}\circ T^{-\rho}_{0}$, which upgrades to the desired natural isomorphism. ∎
5.2.1. Proof of Proposition 5.5
We translate our problem to a problem of computing the extensions of some constructible sheaves on the affine flag variety, using Bezrukavnikov’s equivalence [8].
Let ${\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ be the completion of ${\mathrm{A}}$ at $0\in{\mathfrak{t}}/W$. Since the restriction of $\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ on $\{e\}\times{\mathfrak{b}}\hookrightarrow\tilde{{\mathfrak{g}}}$ is ${\mathrm{A}}^{b}_{S}$, we can identify
(5.4) ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}=\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}})$
by induction from $B$ to $G$. The ${\mathrm{A}}^{b}_{S}$-module ${\mathcal{M}}_{w_{0}}={\mathcal{E}}^{\vee,b}_{S}$ corresponds to ${\mathcal{E}}^{\vee}_{\hat{0}}:={\mathcal{E}}^{\vee}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}/W]_{\hat{0}}$ under (5.4). Let $\mathrm{St}=\tilde{{\mathfrak{g}}}\times_{{\mathfrak{g}}}\tilde{{\mathfrak{g}}}$ be the Steinberg variety, and let $\widehat{\mathrm{St}}$ be its completion at $0\in{\mathfrak{t}}/W$. Let $\pi_{2}:\widehat{\mathrm{St}}\rightarrow\tilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}$ be the projection to the second factor. There is an equivalence
$\pi_{2}^{*}{\mathcal{E}}_{{\hat{0}}}\otimes^{L}_{\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}}-:\ D^{b}\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}){\ \xrightarrow{\sim}\ }D^{b}\mathrm{Coh}^{G}(\widehat{\mathrm{St}}),$
sending ${\mathcal{E}}^{\vee}_{\hat{0}}$ to the structure sheaf ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{\Delta}}$ of the diagonal $\tilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}$ in $\widehat{\mathrm{St}}$.
Let $\check{G}$ be the Langlands dual group of $G$. Let $L\check{G}$ be the loop group of $\check{G}$. Let $\check{I}$ be the Iwahori subgroup and let $\check{I}_{0}$ be its pro-unipotent radical. The enhanced affine flag variety is $\widetilde{\mathcal{F}{l}}=L\check{G}/\check{I}_{0}$. By [8], there is an equivalence of triangulated monoidal categories
(5.5) $D^{b}\mathrm{Coh}^{G}(\widehat{\mathrm{St}})\simeq\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}}),$
where $\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}})$ is a completion of the category of $\check{I}_{0}$-equivariant monodromic complexes on $\widetilde{\mathcal{F}{l}}$. There are standard and costandard sheaves $j_{x!}$, $j_{x*}$ in $\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}})$ associated to the embedding $j_{x}$ from the Schubert cell labelled by $x\in W_{\mathrm{ex}}$ to $\widetilde{\mathcal{F}{l}}$. The equivalence (5.5) sends ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{\Delta}}$ to $j_{e!}$, and in general ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{\Delta}}(w_{0}\lambda)$ to the Wakimoto sheaf ${\mathcal{J}}_{\lambda}$ (see [8, §3.3]), for any $\lambda\in\Lambda$ (In loc. cit. the line bundle labelled by $\lambda$ is sent to ${\mathcal{J}}_{\lambda}$, and here our correspondence is due to the different convention that ${\mathcal{O}}_{{\mathcal{B}}}(\lambda)$ is semi-ample if and only if $\lambda$ is anti-dominant).
We view $\widehat{\mathrm{St}}$ as the base change of $\pi:\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ (as the second factor) by $\tilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}\rightarrow{\mathfrak{g}}$ (as the first factor), then by Theorem 3.1 there are functors ${\mathbb{T}}^{c}_{s}$, $s\in{\mathrm{I}}$ on $D^{b}\mathrm{Coh}^{G}(\widehat{\mathrm{St}})$. By [8] the equivalence (5.5) intertwines the functor ${\mathbb{T}}^{c}_{s}$ with the convolution with $j_{s!}$ from the right. Therefore, by (4.15) the composed equivalence
$D^{b}({\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B})\simeq\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}})$
sends
${\mathcal{M}}_{t(\nu)w}\mapsto{\mathcal{J}}_{w_{0}\nu}*j_{w_{0}w!},\quad\forall w\in W,\ \forall\nu\in\Lambda.$
Proof of Proposition 5.5.
We suppose $w_{0}x=t(\lambda)w$ and $w_{0}x^{\prime}=t(\lambda^{\prime})w^{\prime}$, where $\lambda,\lambda^{\prime}\in\Lambda$ and $w,w^{\prime}\in W$. Then we have
$\mathrm{Ext}^{i}({\mathcal{M}}_{x},{\mathcal{M}}_{x^{\prime}})$ $=\mathrm{Ext}^{i}({\mathcal{J}}_{\lambda}*j_{w!},{\mathcal{J}}_{\lambda^{\prime}}*j_{w^{\prime}!})$
$=\mathrm{Ext}^{i}({\mathcal{J}}_{\lambda+\mu}*j_{w!},{\mathcal{J}}_{\lambda^{\prime}+\mu}*j_{w^{\prime}!}),\quad\forall\mu\in\Lambda.$
Let $\mu$ be dominant enough, then we have ${\mathcal{J}}_{\lambda+\mu}=j_{t(\lambda+\mu)*}$ (see [8, §3.3]). Since $t(\lambda)$ is maximal in the Bruhat order on $t(\lambda)W$ if $\lambda$ is dominant, we have ${\mathcal{J}}_{\lambda+\mu}*j_{w!}=j_{t(\lambda+\mu)*}*j_{w!}=j_{t(\lambda+\mu)w*}$. Note that
$\mathrm{Ext}^{i}(j_{t(\lambda+\mu)w*},j_{t(\lambda^{\prime}+\mu)w^{\prime}*})=j^{*}_{t(\lambda^{\prime}+\mu)w^{\prime}}j_{t(\lambda+\mu)w*}$
which is nonzero only if $t(\lambda^{\prime}+\mu)w^{\prime}\leq t(\lambda+\mu)w$ in the Bruhat order. Since $t(\lambda^{\prime}+\mu)w^{\prime}$ and $t(\lambda+\mu)w$ translate the fundamental alcove to the dominant alcoves, the Bruhat order on them coincides with the semi-infinite order (see e.g. [39, Claim 4.14]). We deduce that
$\mathrm{Ext}^{i}({\mathcal{M}}_{x},{\mathcal{M}}_{x^{\prime}})\neq 0$
only if $t(\lambda^{\prime}+\mu)w^{\prime}\leq^{\frac{\infty}{2}}t(\lambda+\mu)w$, equivalently $t(\lambda^{\prime})w^{\prime}\leq^{\frac{\infty}{2}}t(\lambda)w$ and $x\leq^{\frac{\infty}{2}}x^{\prime}$. ∎
5.3. Proof of the main result
Recall the induction functor
$\mathrm{ind}_{r}=U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}-:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}.$
Lemma 5.10.
There are natural isomorphisms
${\mathbb{V}}_{r}\circ\mathrm{ind}_{r}\simeq\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}\ ,\quad\text{and}\quad{\mathbb{V}}_{c}\circ\mathrm{ind}_{c}\simeq\mathrm{ind}_{c}\circ{\mathbb{V}}^{b}_{c}.$
Proof.
For any $M$ in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$, we have the following natural isomorphisms of $U^{\mathrm{hb}}_{\zeta}$-modules
${\mathbb{V}}_{r}\circ\mathrm{ind}_{r}(M)$ $=\mathrm{pr}_{-\rho}\big{(}(U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}M)\otimes V(\rho)^{*}\big{)}$
$\simeq\mathrm{pr}_{-\rho}\big{(}U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}(M\otimes V(\rho)^{*})\big{)}$
$=U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}\big{(}\mathrm{pr}_{-\rho}(M\otimes V(\rho)^{*})\big{)}=\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}(M),$
where the second isomorphism is by tensor formula. Since the action of $Z_{\mathrm{HC}}$ on $U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}M$ coincides with the one induced by the $Z_{\mathrm{HC}}$-action on $M$, the equality ${\mathbb{V}}_{r}\circ\mathrm{ind}_{r}(M)=\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}(M)$ is also an isomorphism of $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-modules. It implies the first natural isomorphism. The second isomorphism is straight forward. ∎
Proposition 5.11.
The functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are fully-faithful on the objects admitting Verma flags.
Proof.
We firstly prove the assertion for ${\mathbb{V}}_{c}$ in a special case: let $M\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ admitting Verma flags, and let $Q=\mathrm{ind}_{c}(M^{\prime})^{\leq\nu}$ with $M^{\prime}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admitting Verma flags. We show that the natural map
$\mathrm{Hom}(Q,M)\rightarrow\mathrm{Hom}({\mathbb{V}}_{c}Q,{\mathbb{V}}_{c}M)$
is an isomorphism. To that end, consider the following commutative diagram
[…]
where the horizontal isomorphisms follow from the adjunctions for $(-)^{\leq\nu}$ and $\mathrm{ind}_{c}$, and the equalities
${\mathbb{V}}_{c}(\mathrm{ind}_{c}(M^{\prime})^{\leq\nu})={\mathbb{V}}_{c}(\mathrm{ind}_{c}(M^{\prime}))^{\leq\nu}\quad\text{and}\quad{\mathbb{V}}_{c}(\mathrm{ind}_{c}(M^{\prime}))=\mathrm{ind}_{c}({\mathbb{V}}^{b}_{c}(M^{\prime}))$
by Lemma 5.8 and Lemma 5.10. It is now equivalent to show that the right vertical map is an isomorphism, which is true by Proposition 4.13.
In general, let $M_{1},M_{2}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ admitting Verma flags. By Lemma 5.7, there is a resolution $Q_{1}\rightarrow Q_{2}\rightarrow M_{1}\rightarrow 0$ with $Q_{i}=\mathrm{ind}_{c}(P_{i})^{\leq\nu}$ for some projective objects $P_{i}$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. We have a commutative diagram with exact rows
00{0}${\mathrm{Hom}(M_{1},M_{2})}$${\mathrm{Hom}(Q_{2},M_{2})}$${\mathrm{Hom}(Q_{1},M_{2})}$00{0}${\mathrm{Hom}({\mathbb{V}}_{c}M_{1},{\mathbb{V}}_{c}M_{2})}$${\mathrm{Hom}({\mathbb{V}}_{c}Q_{2},{\mathbb{V}}_{c}M_{2})}$${\mathrm{Hom}({\mathbb{V}}_{c}Q_{1},{\mathbb{V}}_{c}M_{2}),}$
where the vertical isomorphisms are by our previous result in the special case. It shows that ${\mathbb{V}}_{c}$ is fully-faithful on $M_{1}$ and $M_{2}$.
The assertion for ${\mathbb{V}}_{r}$ can be proved similarly, or more directly as in Proposition 4.13. ∎
Recall the following result
Theorem 5.12 ([36, Thm 3.6]).
The functor
${\mathrm{I}}^{\mathrm{hb}}_{1}:\ ({\mathbb{C}}[{\mathfrak{b}}_{S}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{S}$
induces an equivalence of abelian categories
$\mathfrak{G}:\ \mathrm{Coh}^{B}({\mathfrak{b}}_{S}){\ \xrightarrow{\sim}\ }{\mathcal{O}}^{-\rho}_{S},$
sending ${\mathbb{C}}[{\mathfrak{b}}_{S}]\otimes{\mathbb{C}}_{\lambda}$ to $M(-\rho+l\lambda)_{S}$, for any $\lambda\in\Lambda$.
For any commutative Noetherian $S$-algebra $R$, the equivalence $\mathfrak{G}$ induces an equivalence of the categories on both sides after base changed to $R$. In particular, we have
$\mathfrak{G}:\ \mathrm{Coh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}){\ \xrightarrow{\sim}\ }{\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
The following statement is our main result.
Theorem 5.13.
There is an equivalence of $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-linear abelian categories
${\mathfrak{F}}:\ \mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}){\ \xrightarrow{\sim}\ }{\mathcal{O}}_{S}^{0},$
such that
${\mathfrak{F}}\big{(}{\mathbb{T}}^{c}_{w}({\mathcal{E}}^{\vee}_{\hat{0}})\otimes{\mathcal{O}}_{\tilde{{\mathfrak{g}}}}(\nu)\big{)}=M(w_{0}w^{-1}\bullet 0+l\nu)_{S},\quad\forall w\in W,\ \forall\nu\in\Lambda.$
Proof.
Let $P$ be a projective module in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$, and set ${\mathcal{P}}=({\mathfrak{F}}^{b})^{-1}(P)$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$. Set $Q=\mathrm{ind}_{r}(P)$ and ${\mathcal{Q}}=\mathrm{ind}_{c}({\mathcal{P}})$. We have the following isomorphisms by Lemma 5.9 and Lemma 5.10,
${\mathbb{V}}_{r}(Q^{\leq l\nu})=({\mathbb{V}}_{r}Q)^{\leq l\nu}=({\mathbb{V}}_{r}\circ\mathrm{ind}_{r}P)^{\leq l\nu}=(\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}P)^{\leq l\nu},$
and similarly ${\mathbb{V}}_{c}({\mathcal{Q}}^{\leq\nu})=(\mathrm{ind}_{c}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq\nu}$. By Proposition 4.6 and Corollary 4.7, we have
$\mathfrak{G}(\mathrm{ind}_{c}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq\nu}=(\mathfrak{G}\circ\mathrm{ind}_{c}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq l\nu}=(\mathrm{ind}_{r}\circ\mathfrak{G}^{b}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq l\nu}=(\mathrm{ind}_{r}\circ{\mathbb{V}}_{r}^{b}P)^{\leq l\nu}.$
It follows that $\mathfrak{G}\big{(}{\mathbb{V}}_{c}({\mathcal{Q}}^{\leq\nu})\big{)}={\mathbb{V}}_{r}(Q^{\leq l\nu})$. Hence we obtain identifications
${\mathfrak{G}}\circ{\mathbb{V}}_{c}\big{(}\mathrm{Proj}({\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu})\big{)}={\mathbb{V}}_{r}\big{(}\mathrm{Proj}({\mathcal{O}}^{0,\leq l\nu}_{S})\big{)},\quad\nu\in\Lambda$
of subcategories in ${\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$. By Proposition 5.11, it induces an equivalence of additive categories
$\mathrm{Proj}({\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}){\ \xrightarrow{\sim}\ }\mathrm{Proj}({\mathcal{O}}^{0,\leq l\nu}_{S}),\quad\forall\nu\in\Lambda,$
which extends to an equivalence of abelian categories
${\mathfrak{F}}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}^{0}_{S}.$
Then we identify ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}=\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}})$ by induction from $B$ to $G$.
For the second assertion, we consider the following commutative diagram by construction
[…]
Hence ${\mathfrak{F}}({\mathcal{M}}_{x})$ ($x\in W_{\mathrm{ex}}$) is isomorphic to $M(x\bullet_{l}0)_{S}$ as a $\Lambda$-graded ${\mathfrak{U}}^{b}_{\zeta}\otimes S$-module. It forces that ${\mathfrak{F}}({\mathcal{M}}_{x})\cong M(x\bullet_{l}0)_{S}$ as $\Lambda$-graded $U^{\mathrm{hb}}_{\zeta}\otimes S$-modules. Now the second assertion follows from (4.15). ∎
Remark 5.14.
The assumption that $l$ is a power of prime is used in the results in §3.2 concerning the localization theorem for ${\mathfrak{U}}_{\zeta}$. However, it is expected that this assumption is not necessary (see a proof for type $A$ in [41]). For Theorem 5.13, one may remove this assumption by showing that the equivalence class of the category ${\mathcal{O}}^{0}_{S}$ is independent of $\zeta$ (note that similar independence for small quantum groups appeared in [1].)
Let $\pi_{{}_{\widetilde{{\mathcal{N}}}}}:\widetilde{{\mathcal{N}}}=G\times^{B}{\mathfrak{n}}\rightarrow{\mathfrak{g}}$ be the Springer resolution. We set ${\mathrm{A}}^{\prime}$ be the specialization of ${\mathrm{A}}$ at $0\in{\mathfrak{t}}$, which satisfies an equivalence [12]
$D^{b}\mathrm{Coh}(\widetilde{{\mathcal{N}}})\ \simeq\ D^{b}({\mathrm{A}}^{\prime}\text{-}\mathrm{mod}),$
and therefore is called the non-commutative Springer resolution. Recall that ${\mathrm{A}}_{0}$ is the specialization of ${\mathrm{A}}$ at $0\in{\mathfrak{g}}$. We abbreviate ${\mathcal{O}}^{\chi_{0}}_{\mathbb{C}}$ as the full subcategory of ${\mathcal{O}}_{\mathbb{C}}$ of modules vanished by $\ker\chi_{0}\subset Z_{\mathrm{HC}}$.
Corollary 5.15.
There is an equivalences of abelian categories
(5.6) $\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})={\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{0},$
which by restrictions on $0\in{\mathfrak{t}}$ and $0\in{\mathfrak{g}}$ induces equivalences
$\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathrm{A}}^{\prime}){\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{{\chi_{0}}},\quad\text{and}\quad{\mathrm{A}}_{0}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta},0}_{\mathbb{C}}.$
6. ${\mathfrak{F}}$ as equivariantization of ${\mathfrak{F}}^{b}$
In this section we will use Theorem 5.13 to show that the equivalence ${\mathfrak{F}}^{b}$ intertwines rational $B$-actions on both categories, so that ${\mathfrak{F}}$ is the equivalence of their equivariantizations. As an application, we show that ${\mathfrak{F}}$ intertwines reflection functors $\Theta^{r}_{s}$ and $\Theta^{c}_{s}$, for all $s\in{\mathrm{I}}_{\mathrm{af}}$. For simplicity, we only consider the non-deformed versions of ${\mathfrak{F}}^{b}$ and ${\mathfrak{F}}$.
6.1. Rational group action
In this section, we recall some notions about rational action of a linear algebraic group on a category, following [33].
Let $K$ be a linear algebraic group over ${\mathbb{C}}$. Let ${\mathcal{D}}$ be a cocomplete ${\mathbb{C}}$-linear abelian category. For any commutative ${\mathbb{C}}$-algebra $R$, we denote by ${\mathcal{D}}_{R}$ the category of $R$-linear objects in ${\mathcal{D}}$, namely the category of pairs $(M,R\rightarrow\mathrm{End}_{\mathcal{D}}(M))$ with $M\in{\mathcal{D}}$. For any commutative ${\mathbb{C}}$-algebra $R^{\prime}$ with a ring homomorphism $\iota:R\rightarrow R^{\prime}$, we have a functor $\iota_{*}:=R^{\prime}\otimes_{R}-$ from ${\mathcal{D}}_{R}$ to ${\mathcal{D}}_{R^{\prime}}$ (it is well-defined by the cocompleteness). Note that $\iota\otimes_{R}-$ gives a natural morphism $\mathrm{id}\rightarrow\iota_{*}$ of endo-functors on ${\mathcal{D}}_{R}$.
Consider the coordinate ring ${\mathbb{C}}[K]$ of $K$ with comultiplication $\Delta$ and counit $\varepsilon$. A rational action of $K$ on ${\mathcal{D}}$ is the data of
  • a functor $\psi:{\mathcal{D}}\rightarrow{\mathcal{D}}_{{\mathbb{C}}[K]}$;
  • a coassociative natural isomorphism $\sigma:\Delta_{*}\circ\psi{\ \xrightarrow{\sim}\ }\psi^{2}$;
  • a natural isomorphism $\eta:\varepsilon_{*}\circ\psi{\ \xrightarrow{\sim}\ }\mathrm{id}_{\mathcal{D}}$.
Any $R$-point $\iota:{\mathbb{C}}[K]\rightarrow R$ of $K$ associates a functor $\iota_{*}\circ\psi:{\mathcal{D}}\rightarrow{\mathcal{D}}_{R}$, which defines an action of the abstract group $K(R)$ on ${\mathcal{D}}$.
Given a rational $K$-action on ${\mathcal{D}}$, we define the equivariantization ${\mathcal{D}}^{K}$ as the category of $K$-equivariant objects in ${\mathcal{D}}$, namely the category of pairs $(M,M\xrightarrow{\rho_{M}}\psi(M))$ with $M\in{\mathcal{D}}$ and $\rho_{M}$ coassociative and counital, i.e. $\rho_{M}$ satisfies
$\psi(\rho_{M})\circ\rho_{M}=\sigma_{M}\circ(\Delta\otimes_{{\mathbb{C}}[K]}-)\circ\rho_{M},\quad\mathrm{id}_{M}=\eta_{M}\circ(\varepsilon\otimes_{{\mathbb{C}}[K]}-)\circ\rho_{M}.$
Now we consider the action of the monoidal category $\mathrm{Rep}(K)$ on an abelian category ${\mathcal{C}}$. By definition, it is an exact bifunctor
$-\otimes-:\ {\mathcal{C}}\times\mathrm{Rep}(K)\rightarrow{\mathcal{C}}$
with compatibility of the monoidal structure on $\mathrm{Rep}(K)$. Given an action of $\mathrm{Rep}(K)$ on ${\mathcal{C}}$, the de-equivariantization ${\mathcal{C}}_{K}$ is defined as the category of ${\mathbb{C}}[K]$-modules in ${\mathcal{C}}$, namely the category of pairs $(M,\ M\otimes{\mathbb{C}}[K]\xrightarrow{a_{M}}M)$ with $M\in{\mathcal{C}}$ and $a_{M}$ associative and unital. Here the $K$-module structure on ${\mathbb{C}}[K]$ is induced from the right multiplication of $K$.
Suppose ${\mathcal{C}}$ is cocomplete. Then there is a canonical rational $K$-action on ${\mathcal{C}}_{K}$, via the functor
$\psi:{\mathcal{C}}_{K}\rightarrow({\mathcal{C}}_{K})_{\underline{{\mathbb{C}}[K]}},\quad M\mapsto\underline{{\mathbb{C}}[K]}\otimes M,$
where $\underline{{\mathbb{C}}[K]}$ is the algebra ${\mathbb{C}}[K]$ with trivial $K$-action, and ${\mathbb{C}}[K]$ acts on $\underline{{\mathbb{C}}[K]}\otimes M$ via comultiplication. We have the following lemma
Lemma 6.1 ([33, Prop A.2]).
Let ${\mathcal{C}}$ be a cocomplete abelian category, equipped with an action of $\mathrm{Rep}(K)$. There is an equivalence
${\mathcal{C}}{\ \xrightarrow{\sim}\ }({\mathcal{C}}_{K})^{K},\quad M\mapsto M\otimes{\mathbb{C}}[K],$
where $M\otimes{\mathbb{C}}[K]$ is an object in $({\mathcal{C}}_{K})^{K}$ via the ${\mathbb{C}}[K]$-action and the $\underline{{\mathbb{C}}[K]}$-coaction on ${\mathbb{C}}[K]$.
6.2. Coinduction functor
We define ${\mathbf{U}}_{\zeta}$, $\mathbf{u}_{\zeta}$ as the quotients of $U_{\zeta}$, $u_{\zeta}$ by the relations $K_{\lambda}^{l}-1$, $\lambda\in\Lambda$. In this section, we will work with the quantum groups ${\mathbf{U}}^{\mathrm{hb}}_{\zeta}$ and ${\mathfrak{U}}^{n}_{\zeta}$ with triangular decompositions
(6.1) ${\mathbf{U}}^{\mathrm{hb}}_{\zeta}={\mathfrak{U}}^{-}_{\zeta}\otimes{\mathbf{U}}^{0}_{\zeta}\otimes U^{+}_{\zeta},\quad{\mathfrak{U}}^{n}_{\zeta}={\mathfrak{U}}^{-}_{\zeta}\otimes{\mathbf{u}}^{0}_{\zeta}\otimes u^{+}_{\zeta}.$
Note that ${\mathfrak{U}}^{n}_{\zeta}$ is a subalgebra of ${\mathbf{U}}^{\mathrm{hb}}_{\zeta}$.
The category ${\mathcal{O}}_{\mathbb{C}}$ can be interpreted as the category of finitely generated ${\mathbf{U}}^{\mathrm{hb}}_{\zeta}$-modules integrable over ${\mathbf{U}}^{\geq}_{\zeta}$. We define $\widehat{{\mathcal{O}}}_{\mathbb{C}}$ in the same way as ${\mathcal{O}}_{\mathbb{C}}$ but removing the condition of being finitely generated. Then $\widehat{{\mathcal{O}}}_{\mathbb{C}}$ is the ind-completion of ${\mathcal{O}}_{\mathbb{C}}$, and the equivalence (5.6) induces an equivalence
$\widehat{{\mathfrak{F}}}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}{\ \xrightarrow{\sim}\ }\widehat{{\mathcal{O}}}^{0}_{\mathbb{C}}.$
Similarly we have ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}={\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$. The equivalence (4.5) induces an equivalence
${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{mod}_{\chi_{0}}$
by forgetting the $\Lambda$-grading, which further induces an equivalence
$\widehat{{\mathfrak{F}}}^{n}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}.$
Note that $\widehat{{\mathfrak{F}}}$ and $\widehat{{\mathfrak{F}}}^{n}$ are compatible with the forgetful functor.
For a ${\mathfrak{U}}^{n}_{\zeta}$-module $M$, we endow the space $\mathrm{Hom}_{{\mathfrak{U}}^{n}_{\zeta}}({\mathbf{U}}^{\mathrm{hb}}_{\zeta},M)$ with a natural ${\mathbf{U}}^{\mathrm{hb}}_{\zeta}$-module structure induced by the right multiplication of ${\mathbf{U}}^{\mathrm{hb}}_{\zeta}$. We denote by $\mathrm{coind}_{r}(M)$ the maximal submodule of $\mathrm{Hom}_{{\mathfrak{U}}^{n}_{\zeta}}({\mathbf{U}}^{\mathrm{hb}}_{\zeta},M)$ that is integrable over ${\mathbf{U}}^{\geq}_{\zeta}$. It defines a functor
$\mathrm{coind}_{r}:\ {\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}\rightarrow\widehat{{\mathcal{O}}}_{\mathbb{C}}.$
The functor $\mathrm{coind}_{r}$ is compatible with block decompositions, namely we have
$\mathrm{coind}_{r}:\ {\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{\omega}}\rightarrow\widehat{{\mathcal{O}}}^{\omega}_{\mathbb{C}},\quad\forall\omega\in\Xi.$
The following lemma is standard.
Lemma 6.2.
The functor $\mathrm{coind}_{r}$ is right adjoint to the forgetful functor $\widehat{{\mathcal{O}}}_{\mathbb{C}}\rightarrow{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}$. More precisely, for any $M_{1}\in\widehat{{\mathcal{O}}}_{\mathbb{C}}$ and any ${\mathfrak{U}}^{n}_{\zeta}$-module $M_{2}$, there is a natural isomorphism
$\mathrm{Hom}_{{\mathbf{U}}^{\mathrm{hb}}_{\zeta}}(M_{1},\mathrm{coind}_{r}(M_{2}))\simeq\mathrm{Hom}_{{\mathfrak{U}}^{n}_{\zeta}}(M_{1},M_{2}),$
which maps $\varphi$ on the LHS to the homomorphism $m_{1}\in M_{1}\mapsto\varphi(m_{1})(1)$, and $\phi$ on the RHS to the homomorphism $m_{1}\in M_{1}\mapsto\big{(}u\in{\mathbf{U}}^{\mathrm{hb}}_{\zeta}\mapsto\phi(um_{1})\big{)}$.
We define ${\mathcal{C}}({\mathbf{U}}^{\geq}_{\zeta})$ as the category of integrable ${\mathbf{U}}^{\geq}_{\zeta}$-modules. Similarly, we have a functor
$\mathrm{coind}^{b}_{r}:\ {\mathbf{u}}_{\zeta}\text{-}\mathrm{Mod}\rightarrow{\mathcal{C}}({\mathbf{U}}^{\geq}_{\zeta})$
right adjoint to the forgetful functor ${\mathcal{C}}({\mathbf{U}}^{\geq}_{\zeta})\rightarrow{\mathbf{u}}_{\zeta}\text{-}\mathrm{Mod}$. By (6.1), we have isomorphisms
$\mathrm{Hom}_{{\mathfrak{U}}^{n}_{\zeta}}({\mathbf{U}}^{\mathrm{hb}}_{\zeta},M)=\mathrm{Hom}_{{\mathbf{u}}^{\geq}_{\zeta}}({\mathbf{U}}^{\geq}_{\zeta},M),\quad\forall M\in{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod},$
of ${\mathbf{U}}^{\geq}_{\zeta}$-modules. Hence $\mathrm{coind}_{r}(M)=\mathrm{coind}^{b}_{r}(M)$ as ${\mathbf{U}}^{\geq}_{\zeta}$-modules.
There is a Hopf algebra structure on $\mathrm{Hom}_{{\mathfrak{U}}^{n}_{\zeta}}({\mathbf{U}}^{\mathrm{hb}}_{\zeta},{\mathbb{C}})$ induced by the Hopf algebra structure on ${\mathbf{U}}^{\mathrm{hb}}_{\zeta}$, making $\mathrm{coind}_{r}({\mathbb{C}})$ as a Hopf subalgebra. Moreover $\mathrm{coind}_{r}({\mathbb{C}})$ is an algebra in the monoidal category $\widehat{{\mathcal{O}}}_{\mathbb{C}}$. We have
$\mathrm{coind}_{r}({\mathbb{C}})=\mathrm{coind}^{b}_{r}({\mathbb{C}})\subset\mathrm{Hom}_{{\mathbf{u}}^{\geq}_{\zeta}}({\mathbf{U}}^{\geq}_{\zeta},{\mathbb{C}})=(U{\mathfrak{b}})^{*}.$
The natural pairing between ${\mathbb{C}}[B]$ and $U{\mathfrak{b}}$ identifies $\mathrm{coind}_{r}({\mathbb{C}})={\mathbb{C}}[B]$ as Hopf algebras. Therefore, any module $\mathrm{coind}_{r}(M)$ is equipped with a (right) action from ${\mathbb{C}}[B]=\mathrm{coind}_{r}({\mathbb{C}})$.
In the notations in §6.1, $(\widehat{{\mathcal{O}}}_{{\mathbb{C}}})_{B}$, ${\mathcal{C}}({\mathbf{U}}^{\geq}_{\zeta})_{B}$ are the categories of ${\mathbb{C}}[B]$-modules in $\widehat{{\mathcal{O}}}_{\mathbb{C}}$, ${\mathcal{C}}({\mathbf{U}}^{\geq}_{\zeta})$, respectively. Since ${\mathbb{C}}[B]$ is trivial as a ${\mathfrak{U}}^{n}_{\zeta}$-module, for any $M\in(\widehat{{\mathcal{O}}}_{{\mathbb{C}}})_{B}$, the quotient $M\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}$ admits an action from ${\mathfrak{U}}^{n}_{\zeta}$.
Proposition 6.3.
There is an equivalence
$\mathrm{coind}_{r}:\ {\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }(\widehat{{\mathcal{O}}}_{{\mathbb{C}}})_{B},$
with inverse functor $-\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}$.
Proof.
One checks that the adjunction in Lemma 6.2 induces an adjunction
$-\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}:\ (\widehat{{\mathcal{O}}}_{{\mathbb{C}}})_{B}\ \rightleftarrows\ {\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}\ :\mathrm{coind}_{r}.$
We have natural morphisms
(6.2) $\mathrm{coind}_{r}(-)\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}\rightarrow\mathrm{id},\quad\mathrm{id}\rightarrow\mathrm{coind}_{r}(-)\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}.$
Recall that for any ${\mathfrak{U}}^{n}_{\zeta}$-module $M$, $\mathrm{coind}_{r}(M)=\mathrm{coind}^{b}_{r}(M)$ as ${\mathbf{U}}^{\geq}_{\zeta}$-modules. Therefore, to show that (6.2) are isomorphisms, it is enough to show that $\mathrm{coind}^{b}_{r}$ induces an equivalence
$\mathrm{coind}^{b}_{r}:\ {\mathbf{u}}^{\geq}_{\zeta}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathcal{C}}({\mathbf{U}}^{\geq}_{\zeta})_{B}$
with inverse functor $-\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}$. This follows from [4, Thm 2.8] applied to the sequence of Hopf algebras ${\mathbf{u}}^{\geq}_{\zeta}\hookrightarrow{\mathbf{U}}^{\geq}_{\zeta}\twoheadrightarrow U{\mathfrak{b}}$. ∎
By Lemma 6.2, we have
Corollary 6.4.
The category ${\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}$ admits a natural rational $B$-action with an equivalence
$\widehat{{\mathcal{O}}}_{{\mathbb{C}}}{\ \xrightarrow{\sim}\ }({\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod})^{B}.$
Remark 6.5.
Intuitively, the $B$-action on ${\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}$ is induced from the adjoint action of ${\mathbf{U}}^{\geq}_{\zeta}$ on ${\mathfrak{U}}^{n}_{\zeta}$. Indeed, ${\mathbf{u}}^{\geq}_{\zeta}$ acts via inner automorphisms in ${\mathfrak{U}}^{n}_{\zeta}$, and therefore acts trivially on the category ${\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}$. Hence the action of ${\mathbf{U}}^{\geq}_{\zeta}$ on ${\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}$ factors through the quantum Frobenius map.
Similarly, we consider the functor
$\mathrm{coind}_{c}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}\rightarrow{\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B},\quad M\mapsto M\otimes{\mathbb{C}}[B].$
It is right adjoint to the forgetful functor ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}\rightarrow{\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}$, and induces an equivalence
(6.3) $\mathrm{coind}_{c}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B})_{B}.$
6.3. $\widehat{{\mathfrak{F}}}$ as equivariantization of $\widehat{{\mathfrak{F}}}^{n}$
Lemma 6.6.
The equivalence ${\mathfrak{F}}$ (resp. $\widehat{{\mathfrak{F}}}$) is compatible with actions from $\mathrm{rep}(B)$ (resp. $\mathrm{Rep}(B)$).
Proof.
We prove the assertion for ${\mathfrak{F}}$, then the assertion for $\widehat{{\mathfrak{F}}}$ follows. We have a commutative diagram
[…]
Note that ${\mathbb{V}}_{c}$, ${\mathbb{V}}_{r}$ and ${\mathfrak{G}}$ are compatible with $\mathrm{rep}(B)$-action. Note also that the $\mathrm{rep}(B)$-action preserves the subcategory of objects admitting Verma flags. Hence by Proposition 5.11, ${\mathfrak{F}}$ is compatible with $\mathrm{rep}(B)$-action on such subcategory in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$. Since any object can be resolved by objects admitting Verma flags, such compatibility extends to the whole category. ∎
Corollary 6.7.
The equivalence $\widehat{{\mathfrak{F}}}^{n}:{\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}$ is compatible with the natural rational $B$-action on both sides. Therefore, $\widehat{{\mathfrak{F}}}$ can be obtained as the $B$-equivariantization of $\widehat{{\mathfrak{F}}}^{n}$.
Proof.
Since $\mathrm{coind}_{r}$ and $\mathrm{coind}_{c}$ are right adjoint to the forgetful functors, we have a commutative diagram
[…]
By Proposition 6.3 and (6.3), the coinduction functors identify the source categories with the de-equivariantizations of the target categories, and moreover the rational $B$-action on the sources come from the $\mathrm{Rep}(B)$-action on the targets. Hence the compatibility between $\widehat{{\mathfrak{F}}}^{n}$ and $B$-action follows from the compatibility between $\widehat{{\mathfrak{F}}}$ and $\mathrm{Rep}(B)$-action, which is by Lemma 6.6.
The second assertion follows from the commutative diagram
[…]
where the left vertical equivalence is tautological, the right one is by Corollary 6.4. ∎
As an application, we have the following
Corollary 6.8.
The equivalence
${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{0}$
intertwines the reflection functors $\Theta^{c}_{s}$ and $\Theta^{r}_{s}$, for any $s\in{\mathrm{I}}_{\mathrm{af}}$.
Proof.
It is enough to show the compatibility between $\widehat{{\mathfrak{F}}}$ and reflection functors. The reflection functors are compatible with the actions of $B$ and $\mathrm{Rep}(B)$, and with $\widehat{{\mathfrak{F}}}^{n}$ by Proposition 4.1. Hence they are compatible with $(\widehat{{\mathfrak{F}}}^{n})^{B}$, and therefore with $\widehat{{\mathfrak{F}}}$ by the commutative diagram (6.4). ∎
7. Graded multiplicities
Recall that the tilting bundle ${\mathcal{E}}$ on $\tilde{{\mathfrak{g}}}$ can be equipped with a ${\mathbb{C}}^{\times}$-equivariant structure, which makes ${\mathrm{A}}$ a ${\mathbb{C}}^{\times}$-equivariant algebra. Using the equivalence (5.6), we can define a graded version of ${\mathcal{O}}^{0}_{\mathbb{C}}$ via
${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\ :=\ \mathrm{Coh}^{G\times{\mathbb{C}}^{\times}}(\pi_{\tilde{{\mathcal{N}}}}^{*}{\mathrm{A}})\ \simeq\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}},$
with a degrading functor ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\rightarrow{\mathcal{O}}^{0}_{\mathbb{C}}$. In this section, we compute the graded multiplicities of simple module in Verma module (resp. of Verma module in projective module) in ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}$.
7.1. Periodic and generic polynomials
We recall the periodic and generic polynomials introduced by Lusztig [29] and Kato [25], following Soergel’s notations and normalizations [39].
Let $l(-):W_{\mathrm{ex}}\rightarrow{\mathbb{Z}}$ be the length function. The extended affine Hecke algebra ${\mathbb{H}}_{\mathrm{ex}}$ is the ${\mathbb{Z}}[v^{\pm 1}]$-algebra generated by $\{H_{x}\}_{x\in W_{\mathrm{ex}}}$ modulo the relations
$\begin{cases}(H_{s}+v)(H_{s}-v^{-1})=0&\text{for }s\in{\mathrm{I}}_{\mathrm{af}},\\ H_{x}H_{y}=H_{xy}&\text{if }l(x)+l(y)=l(xy).\end{cases}$
There is a ring homomorphism (called the bar-involution) ${\mathbb{H}}_{\mathrm{ex}}\rightarrow{\mathbb{H}}_{\mathrm{ex}}$, $H\mapsto\overline{H}$ given by $\overline{v}=v^{-1}$, $\overline{H_{x}}=(H_{x^{-1}})^{-1}$.
Let ${\mathbb{P}}$ be the free ${\mathbb{Z}}[v^{\pm 1}]$-module with basis $\{H^{\frac{\infty}{2}}_{x}\}_{x\in W_{\mathrm{ex}}}$. It is equipped with a right ${\mathbb{H}}_{\mathrm{ex}}$-module structure (called the periodic Hecke module) via
$H^{\frac{\infty}{2}}_{x}\cdot H_{s}=\begin{cases}H^{\frac{\infty}{2}}_{xs}&\text{if }x\leq^{\frac{\infty}{2}}xs,\\ H^{\frac{\infty}{2}}_{xs}+(v^{-1}-v)H^{\frac{\infty}{2}}_{x}&\text{if }xs\leq^{\frac{\infty}{2}}x,\end{cases}$
for any $x\in W_{\mathrm{ex}}$ and $s\in{\mathrm{I}}_{\mathrm{af}}$. For $\lambda\in\Lambda$ we set $e(\lambda):=\sum_{w\in W}v^{l(w)}\cdot H^{\frac{\infty}{2}}_{t(\lambda)w}$. Let ${\mathbb{P}}^{\circ}$ be the ${\mathbb{H}}_{\mathrm{ex}}$-submodule of ${\mathbb{P}}$ generated by $e(\lambda)$ with $\lambda\in\Lambda$.
Theorem 7.1 ([29, 39]).
  1. (1)
    On ${\mathbb{P}}^{\circ}$ there is a unique involution ${\mathbb{P}}^{\circ}\rightarrow{\mathbb{P}}^{\circ}$, $P\mapsto\overline{P}$ that is ${\mathbb{H}}_{\mathrm{ex}}$-skew-linear (i.e. compatible with the bar-involution) and satisfies $\overline{e(\lambda)}=e(\lambda)$ for each $\lambda\in\Lambda$.
  2. (2)
    For any $x\in W_{\mathrm{ex}}$, there is a unique element $\underline{H}^{\frac{\infty}{2}}_{x}\in{\mathbb{P}}^{\circ}$ such that $\underline{H}^{\frac{\infty}{2}}_{x}=\overline{\underline{H}^{\frac{\infty}{2}}_{x}}$ and $\underline{H}^{\frac{\infty}{2}}_{x}\in H^{\frac{\infty}{2}}_{x}+\sum_{y\leq^{\frac{\infty}{2}}x}v{\mathbb{Z}}[v]\cdot H^{\frac{\infty}{2}}_{y}$.
We express $\underline{H}^{\frac{\infty}{2}}_{x}=\sum_{y}p_{y,x}\cdot H^{\frac{\infty}{2}}_{y}$, then $p_{y,x}\in{\mathbb{Z}}[v]$ is called the periodic Kazhdan–Lusztig polynomial.
Let $\hat{{\mathbb{P}}}$ be the completion of ${\mathbb{P}}$ consisting of the formal series in $\{H^{\frac{\infty}{2}}_{x}\}_{x\in W_{\mathrm{ex}}}$ that are bounded above (with respect to the order $\leq^{\frac{\infty}{2}}$). For $\eta\in\Lambda$, consider the ${\mathbb{Z}}[v^{\pm 1}]$-linear operator $\langle\eta\rangle$ on ${\mathbb{P}}$ by $\langle\eta\rangle(H^{\frac{\infty}{2}}_{x})=H^{\frac{\infty}{2}}_{t(\eta)x}$, for any $x\in W_{\mathrm{ex}}$. We define two families of polynomials $\{q_{y,x}\}_{x,y\in W_{\mathrm{ex}}}$ and $\{q^{\prime}_{y,x}\}_{x,y\in W_{\mathrm{ex}}}$ in ${\mathbb{Z}}[v]$ via the following equations in $\hat{{\mathbb{P}}}$ (for any $x\in W_{\mathrm{ex}}$)
(7.1) $\big{(}\prod_{\alpha\in\Phi^{+}}(1+v^{2}\langle-\alpha\rangle+v^{4}\langle-2\alpha\rangle+\cdots)\big{)}(\underline{H}^{\frac{\infty}{2}}_{x})=\sum_{y}q_{y,x}\cdot H^{\frac{\infty}{2}}_{y},$
(7.2) $\big{(}\prod_{\alpha\in\Phi^{+}}(1+\langle-\alpha\rangle+\langle-2\alpha\rangle+\cdots)\big{)}(\underline{H}^{\frac{\infty}{2}}_{x})=\sum_{y}q^{\prime}_{y,x}\cdot H^{\frac{\infty}{2}}_{y}.$
Here $q_{y,x}$ is called the generic Kazhdan–Lusztig polynomial.
Remark 7.2.
The generic polynomials $\{q_{x,y}\}_{x,y}$ were originally defined as the unique family of polynomials such that the element $\sum_{y}(-1)^{l(x)+l(y)}\overline{q}_{y,x}\cdot H^{\frac{\infty}{2}}_{y}\in\hat{{\mathbb{P}}}$ is contained in $H^{\frac{\infty}{2}}_{x}+\sum_{y}v^{-1}{\mathbb{Z}}[v^{-1}]\cdot H^{\frac{\infty}{2}}_{y}$ and is invariant under the involution, see [25]. Under this definition, the equality (7.1) is proved in [25, Thm 3.5], see also [39, Thm 6.3].
7.2. Graded multiplicities
7.2.1. Grading on a category
A grading on a category ${\mathcal{C}}$ is the data of a triple $({\mathcal{C}}^{\mathrm{gr}},\langle 1\rangle,\upsilon)$, where
  • $\langle 1\rangle$ is an auto-functor on the category ${\mathcal{C}}^{\mathrm{gr}}$ (called the grading shift),
  • $\upsilon$ is a functor $\upsilon:{\mathcal{C}}^{\mathrm{gr}}\rightarrow{\mathcal{C}}$ (called the degrading functor),
together with the data of a natural isomorphism $\upsilon\circ\langle 1\rangle{\ \xrightarrow{\sim}\ }\upsilon$ satisfying a natural isomorphism
(7.3) $\bigoplus_{d}\mathrm{Hom}_{{\mathcal{C}}^{\mathrm{gr}}}(M_{1},M_{2}\langle d\rangle){\ \xrightarrow{\sim}\ }\mathrm{Hom}_{{\mathcal{C}}}(\upsilon M_{1},\upsilon M_{2}),$
for any $M_{1}$, $M_{2}$ in ${\mathcal{C}}^{\mathrm{gr}}$ (Here we abbreviate $\langle d\rangle=\langle 1\rangle^{\circ d}$ for any $d\in{\mathbb{Z}}$). For $M\in{\mathcal{C}}$, an object $\widetilde{M}\in{\mathcal{C}}^{\mathrm{gr}}$ with an isomorphism $\upsilon\widetilde{M}\cong M$ is called a lifting of $M$ in ${\mathcal{C}}^{\mathrm{gr}}$.
For a graded algebra $A=A^{\bullet}$, the category of (finitely generated) graded modules $A^{\bullet}\text{-}\mathrm{gmod}$ naturally provides a grading on $A\text{-}\mathrm{mod}$. For a graded $A$-module $M=M^{\bullet}$, we set $M\langle d\rangle$ as the graded $A$-module with $(M\langle d\rangle)^{i}=M^{i-d}$, for any $i\in{\mathbb{Z}}$.
7.2.2. Graded ${\mathrm{A}}$-modules
Let $P(\lambda)_{\mathbb{C}}$ be the projective cover of $L(\lambda)_{\mathbb{C}}$ in $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$. Let $Z(\lambda)_{\mathbb{C}}=u_{\zeta}\otimes_{u_{\zeta}^{\geq 0}}{\mathbb{C}}_{\lambda}$ be the baby Verma module. Recall that $P^{b}(\lambda)_{\mathbb{C}}$ is the project cover of $L(\lambda)_{\mathbb{C}}$ in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$. We have isomorphisms of ${\mathfrak{U}}^{b}_{\zeta}$-modules
(7.4) $P(\lambda)_{\mathbb{C}}\simeq P^{b}(\lambda)_{\mathbb{C}}\otimes_{Z_{\mathrm{Fr}}^{-}}{\mathbb{C}},\quad Z(\lambda)_{\mathbb{C}}\simeq M(\lambda)_{\mathbb{C}}\otimes_{Z_{\mathrm{Fr}}^{-}}{\mathbb{C}},$
where ${\mathbb{C}}$ is the trivial representation of $Z_{\mathrm{Fr}}^{-}$.
Recall that we have a commutative diagram
[…]
where the left horizontal maps are by forgetful functors, and the right inclusions are induced by the quotients ${\mathrm{A}}_{\mathfrak{n}}\twoheadrightarrow{\mathrm{A}}_{0}$ and ${\mathfrak{U}}^{b}_{\zeta}\twoheadrightarrow u_{\zeta}$. For $x\in W_{\mathrm{ex}}$, we denote by ${\mathcal{L}}_{x}$, ${\mathcal{Z}}_{x}$, ${\mathcal{P}}_{x}$, ${\mathcal{P}}^{b}_{x}$, ${{\mathcal{Q}}}_{x}^{\leq\nu}$ the preimages of $L(x\bullet_{l}0)_{\mathbb{C}}$, $Z(x\bullet_{l}0)_{\mathbb{C}}$, $P(x\bullet_{l}0)_{\mathbb{C}}$, $P^{b}(x\bullet_{l}0)_{\mathbb{C}}$, $Q(x\bullet_{l}0)^{\leq l\nu}_{\mathbb{C}}$ under the vertical equivalences above. By Lemma 2.2, the ${\mathrm{A}}_{\mathfrak{n}}$-module ${\mathcal{L}}_{x}$ can be endowed with a $B$-equivariant structure, which makes ${\mathcal{L}}_{x}$ a simple object in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$. Recall the Verma objects ${\mathcal{M}}_{x}$ ($x\in W_{\mathrm{ex}}$) in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$. Set ${\mathcal{M}}^{\mathbb{C}}_{x}={\mathcal{M}}_{x}\otimes_{S}{\mathbb{C}}$. We call ${\mathcal{M}}^{\mathbb{C}}_{x}$ (resp. ${\mathcal{Z}}_{x}$) the Verma object in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ and ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ (resp. in ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$). A Verma flag is a filtration with composition factors given by Verma objects.
Lemma 7.3.
Let $x\in W_{\mathrm{ex}}$.
  1. (1)
    The objects ${\mathcal{Z}}_{x}$, ${\mathcal{P}}_{x}$, ${\mathcal{P}}^{b}_{x}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ admit liftings $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$, $\widetilde{{\mathcal{P}}}^{b}_{x}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$, and the liftings are unique up to grading shifts.
  2. (2)
    The objects ${\mathcal{L}}_{x}$, ${\mathcal{M}}^{\mathbb{C}}_{x}$, ${\mathcal{Q}}_{x}^{\leq\nu}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ admits liftings ~$\widetilde{{\mathcal{L}}}_{x}$, ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$, $\widetilde{{\mathcal{Q}}}_{x}^{\leq\nu}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$, and the liftings are unique up to grading shifts.
  3. (3)
    Let $P$ be an object in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ with a lifting $\widetilde{P}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$. If $P$ admits a Verma flag, then $\widetilde{P}$ also admits a Verma flag (i.e. a filtration with composition factors of the form ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$). Similar result holds for ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ and ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$.
Proof.
By a standard argument (see e.g. [12]), one can show that any $B$-equivariant simple ${\mathrm{A}}_{\mathfrak{n}}$-module (resp. $T$-equivariant projective module of ${\mathrm{A}}_{\mathfrak{n}}$ or ${\mathrm{A}}_{0}$) can be equipped with a ${\mathbb{C}}^{\times}$-equivariant structure, and up to isomorphism any lifting differs by a ${\mathbb{C}}^{\times}$-character.
By its construction (4.15), any Verma object ${\mathcal{M}}^{\mathbb{C}}_{x}$ can be naturally equipped with a ${\mathbb{C}}^{\times}$-equivariant structure. Let ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$ be a lifting of ${\mathcal{M}}^{\mathbb{C}}_{x}$, then $\widetilde{{\mathcal{Z}}}_{x}=\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}$ provides a lifting of ${\mathcal{Z}}_{x}$ by (7.4). Suppose ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},1}_{x}$, ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},2}_{x}$ are two liftings of ${\mathcal{M}}^{\mathbb{C}}_{x}$. Since $\mathrm{End}({\mathcal{M}}^{\mathbb{C}}_{x})={\mathbb{C}}$, by (7.3) we have $\mathrm{Hom}(\widetilde{{\mathcal{M}}}^{{\mathbb{C}},1}_{x},\widetilde{{\mathcal{M}}}^{{\mathbb{C}},2}_{x}\langle i\rangle)={\mathbb{C}}\delta_{i,d}$ for some $d\in{\mathbb{Z}}$, and any nonzero map from ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},1}_{x}$ to ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},2}_{x}\langle d\rangle$ is an isomorphism. The uniqueness of $\widetilde{{\mathcal{Z}}}_{x}$ follows similarly. This proves (1) as well as the assertion for ${\mathcal{L}}_{x}$ and ${\mathcal{M}}^{\mathbb{C}}_{x}$ in (2).
Now we show (3). We only show the assertion for ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$. We prove by increasing induction on the length of Verma flag $\sum_{x}(P:{\mathcal{M}}^{\mathbb{C}}_{x})$. Let $x$ be maximal (with respect to the order $\leq^{\frac{\infty}{2}}$) among the elements $y\in W_{\mathrm{ex}}$ such that ${\mathcal{M}}^{\mathbb{C}}_{y}$ is a Verma factor of $P$. Then we have $\mathrm{Hom}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle,\widetilde{P})\neq 0$ for some $d$, and any nonzero element gives an inclusion ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle\hookrightarrow\widetilde{P}$ (since any nonzero map ${\mathcal{M}}^{\mathbb{C}}_{x}\rightarrow P$ is an inclusion). Now $\widetilde{P}/\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$ is a lifting of $P/{\mathcal{M}}^{\mathbb{C}}_{x}$, which by induction hypothesis admits a Verma flag. Hence $\widetilde{P}$ admits a Verma flag.
Finally, we show the assertion for ${\mathcal{Q}}_{x}^{\leq\nu}$ in (2). Consider the ${\mathbb{C}}^{\times}$-equivariant $({\mathrm{A}}_{n}\rtimes U{\mathfrak{n}})$-module $\widetilde{{\mathcal{Q}}}_{x}=({\mathrm{A}}_{n}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}_{\mathfrak{n}}}\widetilde{{\mathcal{P}}}^{b}_{x}$. By (3), $\widetilde{{\mathcal{P}}}^{b}_{x}$ admits a Verma flag. Hence by Lemma 5.6, $\widetilde{{\mathcal{Q}}}_{x}$ also admits a Verma flag. By (the non-defomed version of) Proposition 5.5, we have $\mathrm{Ext}^{1}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y},\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y^{\prime}}\langle d\rangle)\neq 0$ only if $y\leq^{\frac{\infty}{2}}y^{\prime}$. We define $\widetilde{{\mathcal{Q}}}_{x}^{\leq\nu}$ as the quotient of $\widetilde{{\mathcal{Q}}}_{x}$ by the submodule composed by ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\mu)y}$ with $\mu\nleq\nu$ and $y\in W$. As in Lemma 2.3(2), $\widetilde{{\mathcal{Q}}}_{x}^{\leq\nu}$ is a lifting of ${\mathcal{Q}}_{x}^{\leq\nu}$. For the uniqueness of the lifting, we set ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times},\leq\nu}$ as the Serre subcategory of ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$ generated by ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\mu)y}$ with $\mu\leq\nu$ and $y\in W$. Then any lifting of ${\mathcal{Q}}_{x}^{\leq\nu}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$ is a projective cover of ~$\widetilde{{\mathcal{L}}}_{x}\langle d\rangle$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times},\leq\nu}$ for some $d$, which thus is unique up to a grading shift. ∎
For an object $P$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$ admitting Verma flags, we set $(P:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle)$ as the multiplicity of the Verma factor ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$, which does not depend on the choice of Verma flag. We define the graded multiplicity as
(7.5) $(P:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x})_{{\mathrm{gr}}}=\sum_{d}(P:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle)\cdot v^{d}\ \in{\mathbb{Z}}[v^{\pm 1}].$
For any object $M$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$, it admits a separated decreasing ${\mathbb{Z}}_{\leq 0}$-filtration (usually of infinite length) with simple composition factors. We set $[M:\widetilde{{\mathcal{L}}}_{x}\langle d\rangle]$ as the multiplicity (which is finite) of ~$\widetilde{{\mathcal{L}}}_{x}\langle d\rangle$ in the factors, which dose not depend on the choice of such filtration, and we define the graded multiplicity
(7.6) $[M:\widetilde{{\mathcal{L}}}_{x}]_{{\mathrm{gr}}}=\sum_{d}[M:\widetilde{{\mathcal{L}}}_{x}\langle d\rangle]\cdot v^{d}\ \in{\mathbb{Z}}[v^{\pm 1}].$
We also introduce similar notations for ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$ and ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
Let $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$ be the space of ${\mathbb{Z}}[v^{\pm 1}]$-linear formal series in {[~$\{[\widetilde{{\mathcal{L}}}_{x}]\}_{x}$ that are bounded above (with respect to the order $\leq^{\frac{\infty}{2}}$). We have an inclusion
$K_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})\rightarrow\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}),\quad[M]\mapsto\sum_{x}[M:\widetilde{{\mathcal{L}}}_{x}]_{{\mathrm{gr}}}\cdot[\widetilde{{\mathcal{L}}}_{x}].$
Note that (the image of) {[~$\{[\widetilde{{\mathcal{M}}}_{x}]\}_{x}$, $\{[\widetilde{{\mathcal{Z}}}_{x}]\}_{x}$ give other topological basis on $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$.
7.2.3. Graded multiplicities
Here is the main result of this section.
Theorem 7.4.
Under the degrading functor ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\rightarrow{\mathcal{O}}_{\mathbb{C}}^{0}$, the modules $L(x\bullet_{l}0)_{\mathbb{C}}$, $M(x\bullet_{l}0)_{\mathbb{C}}$ and $Q(x\bullet_{l}0)_{\mathbb{C}}^{\leq\nu}$ ($x\in W_{\mathrm{ex}}$) admit liftings $\widetilde{L}(x\bullet_{l}0)_{\mathbb{C}}$, $\widetilde{M}(x\bullet_{l}0)_{\mathbb{C}}$ and $\widetilde{Q}(x\bullet_{l}0)_{\mathbb{C}}^{\leq\nu}$ in ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}$, such that the graded multiplicities are given by
(7.7) $(\widetilde{Q}(x\bullet_{l}0)_{\mathbb{C}}^{\leq\nu}:\widetilde{M}(y\bullet_{l}0)_{\mathbb{C}})_{\mathrm{gr}}=\begin{cases}q^{\prime}_{w_{0}y,w_{0}x}(v)&\text{if }y\bullet_{l}0\leq l\nu,\\ 0&\text{if else},\end{cases}$
and
(7.8) $[\widetilde{M}(x\bullet_{l}0)_{\mathbb{C}}:\widetilde{L}(y\bullet_{l}0)_{\mathbb{C}}]_{\mathrm{gr}}=q_{w_{0}x,w_{0}y}(v),$
for any $x,y\in W_{\mathrm{ex}}$.
Before proving Theorem 7.4, we firstly consider the graded multiplicities in ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$. Set ${\mathrm{B}}$ as the $\Lambda$-graded (alternatively, $T$-equivariant) algebra with an isomorphism
$({\mathrm{B}})^{\mathrm{op}}=\bigoplus_{\lambda\in\Lambda}\mathrm{Hom}\big{(}\bigoplus_{w\in W}P(w\bullet 0)_{\mathbb{C}},\bigoplus_{w\in W}P(w\bullet 0-l\lambda)_{\mathbb{C}}\big{)}.$
Note that ${\mathrm{B}}$ is finite dimensional, and there is an equivalence
$\bigoplus_{\lambda}\mathrm{Hom}(\bigoplus_{w}P(w\bullet 0+l\lambda)_{\mathbb{C}},-):\ u_{\zeta}\text{-}\mathrm{mod}_{\mathbb{C}}^{\Lambda,0}{\ \xrightarrow{\sim}\ }{\mathrm{B}}\text{-}\mathrm{mod}^{T}.$
We denote by $L_{x}$, $Z_{x}$, $P_{x}$ the image of $L(x\bullet_{l}0)_{\mathbb{C}}$, $Z(x\bullet_{l}0)_{\mathbb{C}}$, $P(x\bullet_{l}0)_{\mathbb{C}}$ ($x\in W_{\mathrm{ex}}$) under this equivalence. The following theorem is a recollection of results in [1, §18] (with conventions adopted from [39]).
Theorem 7.5 ([1]).
The algebra ${\mathrm{B}}$ admits a $T$-stable Koszul grading ${\mathrm{B}}^{\bullet}$. Under the degrading functor ${\mathrm{B}}^{\bullet}\text{-}\mathrm{gmod}^{T}\rightarrow{\mathrm{B}}\text{-}\mathrm{mod}^{T}$, the modules $L_{x}$, $Z_{x}$, $P_{x}$ ($x\in W_{\mathrm{ex}}$) admit liftings $\widetilde{L}_{x}$, $\widetilde{Z}_{x}$, $\widetilde{P}_{x}$ in ${\mathrm{B}}^{\bullet}\text{-}\mathrm{gmod}^{T}$. Moreover, these liftings can be chosen such that $\widetilde{L}_{t(\lambda)x}=\widetilde{L}_{x}\otimes{\mathbb{C}}_{\lambda}$ for any $\lambda\in\Lambda$ (and similarly for $\widetilde{Z}_{x}$, $\widetilde{P}_{x}$), and
$(\widetilde{P}_{y}:\widetilde{Z}_{x})_{{\mathrm{gr}}}=[\widetilde{Z}_{x}:\widetilde{L}_{y}]_{{\mathrm{gr}}}=p_{w_{0}x,w_{0}y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
(Here the graded multiplicities are defined as in (7.5) and (7.6).)
We use the theorem above to show the following
Proposition 7.6.
The liftings ~$\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$ ($x\in W_{\mathrm{ex}}$) in ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$ can be chosen such that ~$\widetilde{{\mathcal{L}}}_{t(\lambda)x}=\widetilde{{\mathcal{L}}}_{x}\otimes{\mathbb{C}}_{\lambda}$ for any $\lambda\in\Lambda$ (and similarly for $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$), and
$(\widetilde{{\mathcal{P}}}_{y}:\widetilde{{\mathcal{Z}}}_{x})_{{\mathrm{gr}}}=[\widetilde{{\mathcal{Z}}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{{\mathrm{gr}}}=p_{w_{0}x,w_{0}y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
Proof.
Through the equivalences ${\mathrm{B}}\text{-}\mathrm{mod}^{T}\simeq u_{\zeta}\text{-}\mathrm{mod}_{\mathbb{C}}^{\Lambda,0}\simeq{\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$, the grading ${\mathrm{B}}^{\bullet}$ on ${\mathrm{B}}$ induces a different grading, hence a different graded multiplicity, on ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$. Our main task is to compare this graded multiplicity to the natural one induced by ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
By [12, §5.2.3, §5.5] (see also [11, Thm 4.5]), there is a $T\times{\mathbb{C}}^{\times}$-equivariant vector bundle ${\mathcal{E}}^{new}$ on $\tilde{{\mathfrak{g}}}$ satisfying Theorem 3.2(2), such that the graded algebra $\mathrm{End}^{{\mathrm{op}}}_{\tilde{{\mathfrak{g}}}}({\mathcal{E}}^{new})$ is Koszul (where grading is induced from ${\mathbb{C}}^{\times}$-equivariant structure). We can replace ${\mathcal{E}}$ by ${\mathcal{E}}^{new}$, then ${\mathrm{A}}^{\bullet}$ is a Koszul algebra. In particular, the graded algebra ${\mathrm{A}}_{0}^{\bullet}={\mathrm{A}}^{\bullet}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}$ satisfies: (1) ${\mathrm{A}}_{0}^{\bullet}$ is non-negatively graded; (2) ${\mathrm{A}}^{0}_{0}$ is a semi-simple algebra; (3) ${\mathrm{A}}_{0}^{\bullet}$ is generated by ${\mathrm{A}}^{1}_{0}$ over ${\mathrm{A}}^{0}_{0}$.
We can choose a lifting $\widetilde{{\mathcal{P}}}_{x}$ of ${\mathcal{P}}_{x}$ ($x\in W_{\mathrm{ex}}$) such that it is generated by degree 0 elements as a graded ${\mathrm{A}}_{0}^{\bullet}$-module. Define a graded algebra ${\mathrm{B}}^{\prime\bullet}$ via
$({\mathrm{B}}^{\prime\bullet})^{\mathrm{op}}=\bigoplus_{d\in{\mathbb{Z}},\lambda\in\Lambda}\mathrm{Hom}\big{(}\bigoplus_{w\in W}\widetilde{{\mathcal{P}}}_{w},\bigoplus_{w\in W}\widetilde{{\mathcal{P}}}_{w}\otimes{\mathbb{C}}_{-\lambda}\langle d\rangle)\big{)}.$
Then we have an isomorphism of algebras ${\mathrm{B}}\simeq{\mathrm{B}}^{\prime}$. From the properties of ${\mathrm{A}}_{0}^{\bullet}$ (resp. Koszulity of ${\mathrm{B}}^{\bullet}$), we deduce that (1) ${\mathrm{B}}^{\prime\bullet}$ (resp. ${\mathrm{B}}^{\bullet}$) is non-negatively graded; (2) ${\mathrm{B}}^{\prime 0}$ (resp. ${\mathrm{B}}^{0}$) is a semi-simple algebra; (3) ${\mathrm{B}}^{\prime\bullet}$ (resp. ${\mathrm{B}}^{\bullet}$) is generated by ${\mathrm{B}}^{\prime 1}$ (resp. ${\mathrm{B}}^{1}$) over ${\mathrm{B}}^{\prime 0}$ (resp. ${\mathrm{B}}^{0}$). Hence by [6, Prop 2.4.1], we have $\mathrm{rad}({\mathrm{B}}^{\prime})^{i}=\bigoplus_{j\geq i}{\mathrm{B}}^{\prime j}$ and $\mathrm{rad}({\mathrm{B}})^{i}=\bigoplus_{j\geq i}{\mathrm{B}}^{j}$, and therefore we have a ($T$-equivariant) isomorphism of graded algebras
(7.9) ${\mathrm{B}}^{\bullet}{\ \xrightarrow{\sim}\ }\bigoplus_{i}\mathrm{rad}({\mathrm{B}})^{i}/\mathrm{rad}({\mathrm{B}})^{i+1}\simeq\bigoplus_{i}\mathrm{rad}({\mathrm{B}}^{\prime})^{i}/\mathrm{rad}({\mathrm{B}}^{\prime})^{i+1}\ \xleftarrow{\sim}{\mathrm{B}}^{\prime\bullet}.$
It induces an equivalence ${\mathrm{B}}^{\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathrm{B}}^{\prime\bullet}\text{-}\mathrm{gmod}^{T}$. Note that the isomorphism (7.9) and the earlier isomorphism ${\mathrm{B}}\simeq{\mathrm{B}}^{\prime}$ might be different, but they induce the same isomorphism on the semi-simple quotients ${\mathrm{B}}/\mathrm{rad}({\mathrm{B}})\simeq{\mathrm{B}}^{\prime}/\mathrm{rad}({\mathrm{B}}^{\prime})$. Therefore, under the composed equivalence ${\mathrm{B}}^{\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathrm{B}}^{\prime\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$, the simple module $\widetilde{L}_{x}$ is send to ~$\widetilde{{\mathcal{L}}}_{x}$ ($x\in W_{\mathrm{ex}}$) (up to shifts). As projective cover of $\widetilde{L}_{x}$, $\widetilde{P}_{x}$ is send to $\widetilde{{\mathcal{P}}}_{x}$ (up to shifts). Since $\widetilde{Z}_{x}$ is the projective cover of $\widetilde{L}_{x}$ in the Serre subcategory generated by $\widetilde{L}_{y}\langle d\rangle$ with $y\leq^{\frac{\infty}{2}}x$, it follows that $\widetilde{Z}_{x}$ is send to $\widetilde{{\mathcal{Z}}}_{x}$ (up to shifts).
Now, we can deduce our assertion from Theorem 7.5. ∎
Remark 7.7.
Proposition 7.6 might be proved as the analogue statement for restricted enveloping algebra in [11, Thm 1.2], instead of using Theorem 7.5.
Proof of Theorem 7.4.
We fix the liftings ~$\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$ ($x\in W_{\mathrm{ex}}$) in Proposition 7.6. And choose liftings ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$, $\widetilde{{\mathcal{P}}}^{b}_{x}$ such that ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}=\widetilde{{\mathcal{Z}}}_{x}$ and $\widetilde{{\mathcal{P}}}^{b}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}=\widetilde{{\mathcal{P}}}_{x}$. Then we have ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\lambda)x}=\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes{\mathbb{C}}_{\lambda}$ for any $\lambda\in\Lambda$ (and similarly for $\widetilde{{\mathcal{P}}}^{b}_{x}$). Apply $-\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}$ to any Verma flag of $\widetilde{{\mathcal{P}}}^{b}_{x}$, we obtain a Verma flag of $\widetilde{{\mathcal{P}}}_{x}$, since the corresponding statement in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$ is true by the freeness of Verma modules over $Z_{\mathrm{Fr}}^{-}$. Hence we have
(7.10) $(\widetilde{{\mathcal{P}}}^{b}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}=(\widetilde{{\mathcal{P}}}_{x}:\widetilde{{\mathcal{Z}}}_{y})_{\mathrm{gr}}.$
Recall the $({\mathrm{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})$-module $\widetilde{{\mathcal{Q}}}_{x}=({\mathrm{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}_{\mathfrak{n}}}\widetilde{{\mathcal{P}}}^{b}_{x}=U{\mathfrak{n}}\otimes\widetilde{{\mathcal{P}}}^{b}_{x}$, which admits a Verma flag with factors of the form ${\mathbb{C}}_{\eta}\otimes\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}=\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\eta)y}$, where $\eta\geq 0$ and ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}$ is a Verma factor of $\widetilde{{\mathcal{P}}}^{b}_{x}$. Moreover, the graded multiplicities satisfy the following equality in $\hat{{\mathbb{P}}}$,
$\big{(}\prod_{\alpha\in\Phi^{+}}(1+\langle-\alpha\rangle+\langle-2\alpha\rangle+\cdots)\big{)}(\sum_{y}(\widetilde{{\mathcal{P}}}^{b}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}\cdot H^{\frac{\infty}{2}}_{w_{0}y})=\sum_{y}(\widetilde{{\mathcal{Q}}}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}\cdot H^{\frac{\infty}{2}}_{w_{0}y}.$
By (7.2) and (7.10) and Proposition 7.6, we have
$(\widetilde{{\mathcal{Q}}}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}=q^{\prime}_{w_{0}y,w_{0}x}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
Since $\widetilde{{\mathcal{Q}}}^{\leq\nu}_{x}$ is the truncation of $\widetilde{{\mathcal{Q}}}_{x}$ by $\nu$, (7.7) follows.
Now we study the multiplicity [~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{\mathrm{gr}}$. Consider the following sequence induced by Koszul resolution
(7.11) ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}({\mathbb{C}}[{\mathfrak{n}}]\otimes\bigwedge^{\bullet}{\mathfrak{n}}^{*})\rightarrow\widetilde{{\mathcal{Z}}}_{x}\rightarrow 0,$
which is exact, since the corresponding sequence in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$
$M(x\bullet_{l}0)_{\mathbb{C}}\otimes_{Z_{\mathrm{Fr}}^{-}}(Z_{\mathrm{Fr}}^{-}\otimes\bigwedge^{\bullet}{\mathfrak{n}}^{*})\rightarrow Z(x\bullet_{l}0)_{\mathbb{C}}\rightarrow 0$
is exact by the freeness of the $Z_{\mathrm{Fr}}^{-}$-module $M(x\bullet_{l}0)_{\mathbb{C}}$. By abuse of notation, we view $\langle\eta\rangle$ ($\eta\in\Lambda$) as a continuous ${\mathbb{Z}}[v^{\pm 1}]$-linear operator on $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$ sending [~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]$ to [~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\eta)x}]$, for any $x\in W_{\mathrm{ex}}$. By (7.11), we have the following equality in $K_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$,
(7.12) $[\widetilde{{\mathcal{Z}}}_{x}]=\big{(}\prod_{\alpha\in\Phi^{+}}(1-v^{2}\langle-\alpha\rangle)\big{)}[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
In $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$, we express [~$[\widetilde{{\mathcal{L}}}_{y}]$ into (topological) basis $\{[\widetilde{{\mathcal{Z}}}_{x}]\}_{x}$ and {[~$\{[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]\}_{x}$ as
[~$[\widetilde{{\mathcal{L}}}_{y}]=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{Z}}}_{x}]=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
Then (7.12) shows that
(7.13) $\big{(}\prod_{\alpha\in\Phi^{+}}(1-v^{2}\langle-\alpha\rangle)\big{)}(\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}])=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
By [29, §11] (see also [38, Thm 6.1]), there is an equality
(7.14) $\sum_{x}(-1)^{l(x)+l(y)}q_{x,y}(v)\cdot p_{w_{0}x,w_{0}z}(v)=\delta_{y,z},\quad\forall y,z\in W_{\mathrm{ex}}.$
Hence it follows from Proposition 7.6 that
(7.15) [~$[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}=(-1)^{l(x)+l(y)}q_{x,y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
Combining (7.1), (7.13) with (7.15), we deduce that
[~$[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}=(-1)^{l(x)+l(y)}p_{x,y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
By (7.14) again, we have
[~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{\mathrm{gr}}=q_{w_{0}x,w_{0}y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.\qed$

### v4

@@
 Let $G$ be a complex connected and simply-connected semisimple algebraic group, with a Borel subgroup $B$ and a Cartan subgroup $T\subset B$.
-Let $U_{q}$ be Lusztig’s quantum group and let ${\mathfrak{U}}_{q}$ be De Concini–Kac’s quantum group, associated to $G$.
-The hybrid (or mixed) quantum group $U^{\mathrm{hb}}_{q}$ is an algebra admitting a triangular decomposition  [^1] In the present paper we will use the definition $U^{\mathrm{hb}}_{q}={\mathfrak{U}}^{-}_{q}\otimes{\mathfrak{U}}^{0}_{q}\otimes U^{+}_{q}$, for the convenience of deforming category ${\mathcal{O}}$.
-In fact these two definitions lead to the same category ${\mathcal{O}}$ (see [9, Lem 3.1]).
-$U^{\mathrm{hb}}_{q}={\mathfrak{U}}^{-}_{q}\otimes U^{0}_{q}\otimes U^{+}_{q},$
+Let $U_{q}$ be Lusztig’s quantum group [29] and let ${\mathfrak{U}}_{q}$ be De Concini–Kac’s quantum group [18], associated to $G$.
+The hybrid (or mixed) quantum group $U^{\mathrm{hb}}_{q}$ is an algebra admitting a triangular decomposition
+$U^{\mathrm{hb}}_{q}={\mathfrak{U}}^{-}_{q}\otimes{\mathfrak{U}}^{0}_{q}\otimes U^{+}_{q}$,
 which was firstly introduced by Gaitsgory [22] with the perspective of generalizing the Kazhdan–Lusztig equivalence.
-It also appears naturally when one considers the quantum analogue of the BGG category ${\mathcal{O}}$.
+It also appears naturally when one considers the quantum analogue of the BGG category $\mathcal{O}$} [6].
 Recall that the BGG category $\mathcal{O}$} (with integral weights) for the Lie algebra ${\mathfrak{g}}=\mathrm{Lie}(G)$ can be interpreted as the category
 ${\mathfrak{g}}\text{-}\mathrm{mod}^{B}$
-of finitely generated ${\mathfrak{g}}$-modules whose restriction on ${\mathfrak{b}}=\mathrm{Lie}(B)$ is integrable. Therefore, one can define the quantum category ${\mathcal{O}}$ to be the category
+of $B$-equivariant finitely generated $\mathfrak{g}$}-modules on which the actions of ${\mathfrak{b}}=\mathrm{Lie}(B)$ by restriction of $\mathfrak{g}$}-action and by differential of $B$-action coincide.
+Therefore, one can define the quantum category $\mathcal{O}$} to be the category
 ${\mathcal{O}}_{q}={\mathfrak{U}}_{q}\text{-}\mathrm{mod}^{U^{\geq}_{q}}$
-of finitely generated ${\mathfrak{U}}_{q}$-modules equipped with an integrable $U^{\geq}_{q}$-module structure that is compatible through ${\mathfrak{U}}^{\geq}_{q}\rightarrow U^{\geq}_{q}$.
-It is exactly the category ${\mathcal{O}}$ associated to $U^{\mathrm{hb}}_{q}$ with its triangular decomposition.
+of finitely generated ${\mathfrak{U}}_{q}$-modules equipped with an equivariant and integrable $U^{\geq}_{q}$-module structure that is compatible through ${\mathfrak{U}}^{\geq}_{q}\rightarrow U^{\geq}_{q}$.
+It is exactly the category $\mathcal{O}$} (with integral weights) associated to $U^{\mathrm{hb}}_{q}$ with its triangular decomposition.
 1.2. Main result
@@
 Our main goal is to give a coherent model for ${\mathcal{O}}^{0}_{\zeta}$, using the non-commutative Springer (Grothendieck) resolution defined in [11].
-Let ${\mathrm{A}}$ be the ($G$-equivariant) ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$-algebra defined in loc.
+Let $\mathsf{A}$} be the ($G$-equivariant) ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$-algebra defined in loc.
 cit.
-as the non-commutative counterpart of the Grothendieck resolution $\tilde{{\mathfrak{g}}}=G\times^{B}{\mathfrak{b}}\rightarrow{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}$ (where ${\mathfrak{t}}=\mathrm{Lie}(T)$ and $W$ is the Weyl group).
-Let $\pi_{{}_{\tilde{{\mathcal{N}}}}}:\tilde{{\mathcal{N}}}=G\times^{B}{\mathfrak{n}}\rightarrow{\mathfrak{g}}$ be the Springer resolution (where ${\mathfrak{n}}$ is the nilpotent radical of ${\mathfrak{b}}$).
+as the non-commutative counterpart of the Grothendieck resolution $\widetilde{{\mathfrak{g}}}=G\times^{B}{\mathfrak{b}}\rightarrow{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}$ (where ${\mathfrak{t}}=\mathrm{Lie}(T)$ and $W$ is the Weyl group).
+Let $\pi_{{}_{\widetilde{{\mathcal{N}}}}}:\widetilde{{\mathcal{N}}}=G\times^{B}{\mathfrak{n}}\rightarrow{\mathfrak{g}}$ be the Springer resolution (where $\mathfrak{n}$} is the nilpotent radical of $\mathfrak{b}$}).
 Our main result is
-Theorem A (=Corollaries 5.15, 6.8).
+Theorem A (=Corollaries 5.25&6.8).
 There is an equivalence of abelian categories
-(1.1) ${\mathcal{O}}^{0}_{\zeta}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})$
+(1.1) ${\mathcal{O}}^{0}_{\zeta}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})$
 intertwining the reflection functors on both sides.
-Note that a similar result for modular category ${\mathcal{O}}$ was obtained by Losev [26] recently.
-The category $\mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})$ firstly appeared in [10], which was shown to be derived equivalent to a version of the affine Hecke category and coincide with the heart of the “new t-structure” defined by Frenkel–Gaitsgory [21].
+Note that a similar result for modular analogue of BGG category $\mathcal{O}$} was obtained by Losev [25] recently.
+The category $\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})$ firstly appeared in [8], which was shown to be derived equivalent to a version of the affine Hecke category and coincide with the heart of the “new t-structure” defined by Frenkel–Gaitsgory [21].
 It would be interesting to explore the relationship between (1.1) and a conjectural equivalence of Gaitsgory [22, Conj.
@@
 There is an equivalence of abelian categories
-(1.2) ${\mathcal{O}}^{-\rho}_{\zeta}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\tilde{{\mathcal{N}}}).$
-Based on the equivalence above, we establish functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ on both sides of (1.1) that are fully-faithful on the objects admitting Verma flags (in the deformed setting, see §1.5 below).
+[…]
+Based on the equivalence above, we establish functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ (here “rr” and “cc” represent “representation” and “coherent”, respectively) on both sides of (1.1) that are fully-faithful on the objects admitting Verma flags (in fact our constructions are in the deformed setting, see §1.5 below).
 The functors $\mathbb{V}$} are constructed as follows. Let $C={\mathbb{C}}[0\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ be the coinvariant ring. There is an algebra homomorphism (see [7])
@@
 On the coherent side, we define a functor
-${\mathbb{V}}_{c}:\ \mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})\rightarrow\mathrm{Coh}^{G}(\tilde{{\mathcal{N}}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$
-by base change along $\pi_{{}_{\tilde{{\mathcal{N}}}}}:\tilde{{\mathcal{N}}}\rightarrow{\mathfrak{g}}$ of the non-commutative resolution ${\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$.
+${\mathbb{V}}_{c}:\ \mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})\rightarrow\mathrm{Coh}^{G}(\widetilde{{\mathcal{N}}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$
+by base change along $\pi_{{}_{\widetilde{{\mathcal{N}}}}}:\widetilde{{\mathcal{N}}}\rightarrow{\mathfrak{g}}$ of the non-commutative resolution ${\mathsf{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$.
 To relate the images of two functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$, we use a $C$-linear extension of the equivalence (1.2), i.e.
-${\mathcal{O}}^{-\rho}_{\zeta,C}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\tilde{{\mathcal{N}}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}).$
-Then we compare the images of the projective objects under the functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$. To that end, we firstly establish a “de-equvariantization” of Theorem A.
+${\mathcal{O}}^{-\rho}_{\zeta,C}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\widetilde{{\mathcal{N}}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}).$
+Our next step is to compare the images of the projective objects under the functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$. To that end, we firstly establish a “de-equvariantization” of Theorem A.
 1.4. A de-equvariantization
@@
 Let ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T,0}$ be its principal block.
-Note that the restriction of $\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}}$ on the fiber ${\mathfrak{n}}\hookrightarrow\tilde{{\mathcal{N}}}$ is ${\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]$.
+Note that the restriction of $\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}}$ on the fiber ${\mathfrak{n}}\hookrightarrow\widetilde{{\mathcal{N}}}$ is ${\mathsf{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]$.
 Hence we can identify
-$\mathrm{Coh}^{G}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})=\mathrm{Coh}^{B}({\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}])$
+$\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})=\mathrm{Coh}^{B}({\mathsf{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}])$
 by restriction to $\mathfrak{n}$}. We have the following $T$-equivariant version of Theorem A.
@@
 There is an equivalence of abelian categories
-(1.3) ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T,0}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{T}({\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]),$
+(1.3) ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{T,0}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{T}({\mathsf{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]),$
 intertwining the reflection functors on both sides.
@@
 So that we can identify the images of projective objects under the functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$.
-The proof of Proposition 1.2 is based on results of Tanisaki [42], where the localization theorem for ${\mathfrak{U}}_{\zeta}$ was established and it further related ${\mathfrak{U}}_{\zeta}$-modules to ${\mathrm{A}}$-modules by some splitting bundles for the sheaf of differential operators on the quantized flag variety.
+The proof of Proposition 1.2 is based on results of Tanisaki [39], where the localization theorem for ${\mathfrak{U}}_{\zeta}$ was established and it further related ${\mathfrak{U}}_{\zeta}$-modules to $\mathsf{A}$}-modules by some splitting bundles for the sheaf of differential operators on the quantized flag variety.
 Note that it is not clear from their constructions whether the splitting bundles are $U_{\zeta}$-equivariant or not.
 So à priori, one can not derive (1.1) from (1.3), by adding the $U^{\geq}_{\zeta}$-equivariant and $B$-equivariant conditions on both sides.
-However, after proving our main result, we can show that (1.3) dose respect the $B$-actions on both categories as a consequence (see Corollary 6.7).
+However, after proving our main result, we can show that (1.3) does respect the $B$-actions on both categories as a consequence (see Corollary 6.7).
 Finally, we have to show the fully-faithfulness for ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ on the objects admitting Verma flags, which is true only in the deformed settings (see Remark 4.15).
@@
 The deformation category for ${\mathcal{O}}_{\zeta}$ is an $S$-linear category ${\mathcal{O}}_{\zeta,S}$ whose specialization to the residue field is ${\mathcal{O}}_{\zeta}$.
-Let ${\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ be the completion of ${\mathrm{A}}$ at $0\in{\mathfrak{t}}/W$.
+Let $^{\mathsf{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ be the completion of $\mathsf{A}$} at $0\in{\mathfrak{t}}/W$.
 We have a deformed version of Theorem A
@@
 There is an equivalence of $S$-linear abelian categories
-(1.4) ${\mathcal{O}}^{0}_{\zeta,S}{\ \xrightarrow{\sim}\ }\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}),$
-where ${\mathcal{O}}^{0}_{\zeta,S}$ is the principal block of ${\mathcal{O}}_{\zeta,S}$ and $\pi:\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ is the natural projection.
+[…]
+where ${\mathcal{O}}^{0}_{\zeta,S}$ is the principal block of ${\mathcal{O}}_{\zeta,S}$ and $\pi:\widetilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ is the natural projection.
 We have similar constructions as in §1.3 and §1.4 in the deformed setting, which together with the fact that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are fully-faithful on objects admitting Verma flags (see Proposition 5.21) imply Theorem B.
@@
 1.6. Kazhdan–Lusztig polynomial
-Consider the ${\mathbb{C}}^{\times}$-action on ${\mathfrak{g}}$ by $t.x=t^{-2}x$, for any $t\in{\mathbb{C}}^{\times}$, $x\in{\mathfrak{g}}$.
-The ${\mathbb{C}}[{\mathfrak{g}}]$-algebra ${\mathrm{A}}$ can be equipped with a ${\mathbb{C}}^{\times}$-equivariant structure.
-By the equivalence (1.1), we can define a grading on ${\mathcal{O}}^{0}_{\zeta}$ via
-${\mathcal{O}}^{0,{\mathrm{gr}}}_{\zeta}:=\mathrm{Coh}^{G\times{\mathbb{C}}^{\times}}(\pi_{{}_{\tilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})\ \rightarrow\ {\mathcal{O}}^{0}_{\zeta}.$
+[…]
+${\mathcal{O}}^{0,{\mathrm{gr}}}_{\zeta}:=\mathrm{Coh}^{G\times{\mathbb{C}}^{\times}}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})\ \rightarrow\ {\mathcal{O}}^{0}_{\zeta}.$
 We compute the Kazhdan–Lusztig polynomials in ${\mathcal{O}}^{0,{\mathrm{gr}}}_{\zeta}$:
@@
 In Section 3 we give some reminders on the non-commutative Springer resolutions following [11] and the localization theorem for ${\mathfrak{U}}_{\zeta}$ following [39].
-In Section 4 we prove Proposition 1.2, and construct the functors ${\mathbb{V}}_{r}^{b}$, ${\mathbb{V}}^{b}_{c}$ in §4.2, whose fully-faithfulness will be proved in §4.4.
-In §4.2 we also study the induction functors from the categories in (1.3) to the categories in (1.1).
+In Section 4 we prove the de-equivariantization of main result (i.e.
+Proposition 1.2) in §4.1 and its analogue for Steinberg block in §4.2.
+In §4.3 we study the induction functors from the categories in (1.3) to the categories in (1.1).
+We construct the functors ${\mathbb{V}}_{r}^{b}$, ${\mathbb{V}}^{b}_{c}$ in §4.4, whose fully-faithfulness will be proved in §4.6.
 In §4.5 we investigate the image of Verma modules under the equivalence (1.3).
-In Section 5, we define the functors ${\mathbb{V}}_{r}$, ${\mathbb{V}}_{c}$, and show their compatibility with reflection functors.
-§5.2 is a slightly technical section.
-The projective objects in ${\mathcal{O}}_{\zeta}$ only exist after truncations, hence we have to define corresponding truncations on the coherent side and study projective objects in them, which will be done in §5.2.
+In Section 5, we define the functors ${\mathbb{V}}_{r}$, ${\mathbb{V}}_{c}$, and show their compatibility with reflection functors in §5.1.
+The subsection §5.2 is slightly technical: we consider certain truncated subcategories of ${\mathcal{O}}^{0}_{\zeta}$ and of the coherent side, which admit enough projective objects and are compatible with functors ${\mathbb{V}}_{r}$, ${\mathbb{V}}_{c}$.
 In §5.3 we show the fully-faithfulness of ${\mathbb{V}}_{r}$, ${\mathbb{V}}_{c}$ and then prove the main theorem.
-In Section 6, we firstly recall the formalism of a rational group action on a category, then equip a $B$-action on LHS of (1.3) using coinduction functor.
-We apply our main theorem to show that the equivalence (1.3) respects $B$-actions on both categories, and the equivalence (1.1) can be obtained as the $B$-equivariantization of (1.3).
-In Section 7, we firstly recall the periodic and generic polynomials.
-Then we discuss the liftings of the simple, Verma and projective objects in the graded principal block and compute the related graded multiplicities.
+In Section 6, we firstly recall the formalism of a rational group action on a category in §6.1, then equip a $B$-action on LHS of (1.3) using coinduction functor in §6.2.
+In §6.3, we apply our main theorem to show that equivalence (1.1) is the $B$-equivariantization of (1.3), and obtain the compatibility of (1.1) with reflection functors as an application.
+In Section 7, we firstly recall the periodic and generic polynomials in §7.1.
+Then we discuss the liftings of the simple, Verma and projective objects in the graded principal block and compute the related graded multiplicities in §7.2.
 1.8. Conventions
@@
 By formal neighborhood of $Y=f^{-1}(z)$ in $X$, we mean a scheme
-$\mathrm{FN}_{X}(Y):=X_{\hat{z}}:=X\times_{\mathrm{Spec}R}\mathrm{Spec}(R_{\hat{z}}).$
+[…]
 Let $\mathcal{A}$} be a quasi-coherent sheaf of algebras on a scheme $X$.
@@
 Let $(X^{*}(T),X_{*}(T),\Phi,\check{\Phi})$ be the root datum associated with $G$.
-Let $\langle$-,$\rangle$ - , - ⟩ be the natural pairing of $\Lambda:=X^{*}(T)$ and $X_{*}(T)$.
+Let $\langle-,$-$\rangle$ be the natural pairing of $\Lambda:=X^{*}(T)$ and $X_{*}(T)$.
 Let $\Phi^{+}\subset\Phi$ be the subsets of positive roots.
@@
 Let $W_{\mathrm{af}}=W\ltimes{\mathrm{Q}}$ be the affine Weyl group.
+Let ${\mathrm{I}}_{\mathrm{af}}$ be the subset of simple reflections in $W_{\mathrm{af}}$.
 Let $W_{\mathrm{ex}}=W\ltimes\Lambda$ be the extended affine Weyl group.
-We have $W_{\mathrm{ex}}\simeq W_{\mathrm{af}}\rtimes\Lambda/{\mathrm{Q}}$, where $\Lambda/{\mathrm{Q}}$-acts by automorphisms of the affine root system.
-Let ${\mathrm{I}}_{\mathrm{af}}$ be the subset of simple reflections in $W_{\mathrm{af}}$.
+We have $W_{\mathrm{ex}}\simeq W_{\mathrm{af}}\rtimes\Lambda/{\mathrm{Q}}$, where $\Lambda/{\mathrm{Q}}$ acts by automorphisms of the affine root system.
 Denote by $t(\lambda)$ the translation by $\lambda\in\Lambda$ in $W_{\mathrm{ex}}$.
@@
 For any $\omega\in\Xi$, we will fix a representative (still denoted by $\omega$) in the fundamental alcove $\Xi_{\mathrm{sc}}$ of $W_{\mathrm{af}}$-action on $\Lambda$, where
-[…]
+$\Xi_{\mathrm{sc}}:=\{\lambda\in\Lambda\ |\ 0\leq\langle\lambda+\rho,\check{\alpha}\rangle\leq l,\ \forall\alpha\in\Phi^{+}\}\leftrightarrow\Lambda/(W_{\mathrm{af}},\bullet_{l}).$
 […]
@@
 We set ${\mathfrak{U}}^{t}_{\zeta}$ be the algebra with triangular decomposition ${\mathfrak{U}}^{t}_{\zeta}=u_{\zeta}^{-}\otimes{\mathfrak{U}}_{\zeta}^{0}\otimes u_{\zeta}^{+}$.
+For $A=A_{q}$ or $A_{\zeta}$ above, we abbreviate the subalgebras $A^{\leq}:=A^{-}A^{0}$ and $A^{\geq}:=A^{+}A^{0}$.
 2.2.1. Harish-Chandra center
@@
 Let $U^{0}$ be a Noetherian sub-quotient algebra of ${\mathfrak{U}}_{\zeta}^{0}$ or $U_{\zeta}^{0}$ such that the $\Lambda$-action induces an action on $U^{0}$.
-Let $U=\bigoplus_{\lambda\in{\mathrm{Q}}}U_{\lambda}$ be a ${\mathrm{Q}}$-graded ${\mathbb{C}}$-algebra with triangular decomposition $U=U^{-}\otimes U^{0}\otimes U^{+}$, such that
+Let $U=\bigoplus_{\lambda\in{\mathrm{Q}}}U_{\lambda}$ be a Noetherian $\mathrm{Q}$}-graded $\mathbb{C}$}-algebra with triangular decomposition $U=U^{-}\otimes U^{0}\otimes U^{+}$, such that
 $fm=m\tau_{\lambda}(f),\quad\forall f\in U^{0},\ \forall m\in U_{\lambda},$
-and satisfies further conditions as in [37, §2.3].
+and satisfies further conditions as in [35, §2.3].
 We abbreviate $U^{\leq}=U^{-}U^{0}$ and $U^{\geq}=U^{+}U^{0}$.
@@
 It is an abelian subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$.
+If $U^{+}$ is finite dimensional, then we have ${\mathcal{O}}^{U}_{R}=U\text{-}\mathrm{mod}^{\Lambda}_{R}$.
 Define the Verma module
@@
 We have an isomorphism $M^{U}(\lambda)_{R}=U^{-}\otimes R_{\lambda}$ as $U^{\leq}$-modules.
-A Verma flag of an object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ is a (decreasing, separated, ${\mathbb{Z}}_{\leq 0}$-) filtration with composition factors by Verma modules.
-The composition factors of such an object are called Verma factors, which are independent on the choice of Verma flags.
+A Verma flag of an object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ is a filtration of finite length with composition factors by Verma modules.
+The composition factors of such an object are called Verma factors, which (including multiplicities) are independent on the choice of Verma flags.
 If $R={\mathbb{F}}$ is a field, then $M^{U}(\lambda)_{\mathbb{F}}$ admits a unique simple quotient $L^{U}(\lambda)_{\mathbb{F}}$, and $\{L^{U}(\lambda)_{\mathbb{F}}\}_{\lambda\in\Lambda}$ provides a complete family of pairwise non-isomorphic simple objects in ${\mathcal{O}}^{U}_{R}$.
+The following is standard, see e.g. [23, Prop 3.1].
+Lemma 2.1.
+We have
+(2.3) $\mathrm{Hom}(M^{U}(\lambda)_{R},M^{U}(\mu)_{R})\neq 0\quad\text{only if}\ \lambda\leq\mu,$
+(2.4) $\mathrm{Ext}^{1}(M^{U}(\lambda)_{R},M^{U}(\mu)_{R})\neq 0\quad\text{only if}\ \lambda<\mu.$
 2.3.1. Truncated categories
-For any $\nu\in\Lambda$, there is a truncated subcategory $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ consists of the module $M$ in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ such that $M_{\mu}=0$ unless $\mu\leq\nu$.
-Note that $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ is a Serre subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$.
-The category ${\mathcal{O}}^{U,\leq\nu}_{R}:={\mathcal{O}}^{U}_{R}\cap U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ always admits enough projective objects, in contrast to ${\mathcal{O}}^{U}_{R}$.
-And any object in ${\mathcal{O}}^{U}_{R}$ admits a finite decomposition with each direct factor contained in ${\mathcal{O}}^{U,\leq\nu}_{R}$ for some $\nu\in\Lambda$.
-Any projective module in ${\mathcal{O}}^{U,\leq\nu}_{R}$ admits Verma flags.
-If $R$ is a complete local domain with residue field ${\mathbb{F}}$, we write $Q^{U}(\lambda)_{R}^{\leq\nu}$ as the projective cover of $L^{U}(\lambda)_{\mathbb{F}}$ in ${\mathcal{O}}^{U,\leq\nu}_{R}$.
+For any subset $\Omega\subset\Lambda$, there is a truncated subcategory (by $\Omega$)
+$U\text{-}\mathrm{Mod}_{R}^{\Omega}$
+consisting of the module $M$ in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ such that $M_{\mu}=0$ unless $\mu\in\Omega$.
+It is a Serre subcategory of $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$.
 The truncation functor
-$U\text{-}\mathrm{Mod}_{R}^{\Lambda}\rightarrow U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu},\quad M\mapsto\ M^{\leq\nu}:=M\big{/}U.\big{(}\bigoplus_{\mu\nleq\nu}M_{\mu}\big{)}$
-by taking the maximal quotient in $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ is left adjoint to the natural inclusion $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}\hookrightarrow U\text{-}\mathrm{Mod}_{R}^{\Lambda}$.
+$\tau^{\Omega}:\ U\text{-}\mathrm{Mod}_{R}^{\Lambda}\rightarrow U\text{-}\mathrm{Mod}_{R}^{\Omega},\quad M\mapsto M\big/U.\big(\bigoplus_{\mu\notin\Omega}M_{\mu}\big)$
+by taking the maximal quotient in $U\text{-}\mathrm{Mod}_{R}^{\Omega}$ is left adjoint to the natural inclusion functor $U\text{-}\mathrm{Mod}_{R}^{\Omega}\hookrightarrow U\text{-}\mathrm{Mod}_{R}^{\Lambda}$.
+We abbreviate $U\text{-}\mathrm{mod}_{R}^{\Omega}:=U\text{-}\mathrm{mod}_{R}^{\Lambda}\cap U\text{-}\mathrm{Mod}_{R}^{\Omega}$.
+Let now $\Omega\subset\Lambda$ be a bounded above poset ideal.
+Then any finitely generated $U\otimes R$-module in $U\text{-}\mathrm{Mod}_{R}^{\Omega}$ is contained in ${\mathcal{O}}^{U}_{R}$. We set
+${\mathcal{O}}^{U,\in\Omega}_{R}:={\mathcal{O}}^{U}_{R}\cap U\text{-}\mathrm{Mod}_{R}^{\Omega}.$
+The category ${\mathcal{O}}^{U}_{R}$ is the union of its full subcategories ${\mathcal{O}}^{U,\in\Omega}_{R}$, running over all bounded above poset ideals $\Omega\subset\Lambda$.
+The category ${\mathcal{O}}^{U,\in\Omega}_{R}$ always admits enough projective objects, in contrast to ${\mathcal{O}}^{U}_{R}$.
+Moreover, there is a generating family of projective objects in ${\mathcal{O}}^{U,\in\Omega}_{R}$ admitting Verma flags, with factors of the form $M^{U}(\lambda)_{R}$ with $\lambda\in\Omega$, see e.g.
+[20, Lem 2.3].
+If $R$ is local, then any projective module in ${\mathcal{O}}^{U,\in\Omega}_{R}$ admits Verma flags, see e.g.
+[20, Lem 2.5].
+If $R$ is a complete local domain with residue field $\mathbb{F}$}, we write $Q^{U}(\lambda)_{R}^{\Omega}$ as the projective cover of $L^{U}(\lambda)_{\mathbb{F}}$ in ${\mathcal{O}}^{U,\in\Omega}_{R}$, for any $\lambda\in\Omega$.
+Let $R^{\prime}$ be a commutative Noetherian $R$-algebra. We have a natural functor
+$\otimes_{R}R^{\prime}:\ U\text{-}\mathrm{Mod}^{\Lambda}_{R}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda}_{R^{\prime}}.$
+By [20, Prop 2.4], it yields an equivalence
+(2.5) $\mathrm{Proj}({\mathcal{O}}^{U,\in\Omega}_{R})\otimes_{R}R^{\prime}{\ \xrightarrow{\sim}\ }\mathrm{Proj}({\mathcal{O}}^{U,\in\Omega}_{R^{\prime}}),$
+for any bounded above poset ideal $\Omega$ in $\Lambda$.
 Lemma 2.2.
-  1. (1)
-    We have
-    $\mathrm{Hom}(M^{U}(\lambda)_{R},M^{U}(\mu)_{R})\neq 0\quad\text{only if}\ \lambda\leq\mu,$
-    and
-    $\mathrm{Ext}^{1}(M^{U}(\lambda)_{R},M^{U}(\mu)_{R})\neq 0\quad\text{only if}\ \lambda<\mu.$
-  2. (2)
-    For any object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ admitting Verma flags, its truncation in $U\text{-}\mathrm{Mod}_{R}^{\Lambda,\leq\nu}$ is the quotient by the submodule extended by its Verma factors in $\{M^{U}(\lambda)_{R}\}_{\lambda\nleq\nu}$.
+Let $\Omega$ be a poset ideal in $\Lambda$.
+For any object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ admitting a Verma flag, its truncation in $U\text{-}\mathrm{Mod}_{R}^{\Omega}$ is the quotient by the submodule composed by its Verma factors in $\{M^{U}(\lambda)_{R}\}_{\lambda\notin\Omega}$.
+Proof.
+Let $P$ be an object in $U\text{-}\mathrm{Mod}_{R}^{\Lambda}$ admitting a Verma flag.
+By (2.4), there is a short exact sequence $0\rightarrow P_{1}\rightarrow P\rightarrow P_{2}\rightarrow 0$, where $P_{1}$ is the submodule composed by the Verma factors in $\{M^{U}(\lambda)\}_{\lambda\notin\Omega}$ and $P_{2}$ is the quotient module composed by the Verma factors in $\{M^{U}(\lambda)\}_{\lambda\in\Omega}$.
+It is clear that $P_{2}\in U\text{-}\mathrm{Mod}_{R}^{\Omega}$.
+Note that we have
+$\tau^{\Omega}M^{U}(\lambda)_{R}=\begin{cases}M^{U}(\lambda)_{R},&\lambda\in\Omega\\ 0,&\lambda\notin\Omega\end{cases}.$
+Using the fact that $\tau^{\Omega}$ is right exact, we deduce that $\tau^{\Omega}(P_{1})=0$ and that $\tau^{\Omega}(P)=\tau^{\Omega}(P_{2})=P_{2}$. ∎
+For $\Omega=\{\lambda\in\Lambda|\lambda\leq\nu\}$, we abbreviate $\Omega$” and $\in\Omega$” by $\leq\nu$” in all the superscripts above.
+We will drop the superscripts $U$ from the notations above when $U=U^{\mathrm{hb}}_{\zeta}$.
+For example, if $\Omega=\{\lambda\in\Lambda|\lambda\leq\nu\}$ and $U=U^{\mathrm{hb}}_{\zeta}$, we abbreviate ${\mathcal{O}}^{U,\in\Omega}_{R}$, $Q^{U}(\lambda)_{R}^{\Omega}$ by ${\mathcal{O}}^{\leq\nu}_{R}$, $Q(\lambda)_{R}^{\leq\nu}$.
 2.3.2. Projectives and simples in ${\mathcal{O}}_{R}$
-We will drop the superscripts $U$ from the notations above when $U=U^{\mathrm{hb}}_{\zeta}$. We set
+We set
 $S={\mathbb{C}}[T]_{\hat{1}}={\mathbb{C}}[{\mathfrak{t}}]_{\hat{0}}$.
 Then $S$ is naturally an algebra of ${\mathfrak{U}}^{0}_{\zeta}={\mathbb{C}}\langle K_{\lambda}\rangle_{\lambda\in\Lambda}={\mathbb{C}}[T]$.
-Note that $S$ is a deformation ring for both ${\mathfrak{U}}^{b}_{\zeta}$ and $U^{\mathrm{hb}}_{\zeta}$.
-Lemma 2.2 ([37, Lem 3.1]).
+We thus view $S$ as a deformation ring for both ${\mathfrak{U}}^{b}_{\zeta}$ and $U^{\mathrm{hb}}_{\zeta}$.
+Lemma 2.3 ([35, Lem 3.1]).
 The $U_{\zeta}^{\mathrm{hb}}$-module $L(\lambda)_{\mathbb{C}}$ remains simple as a ${\mathfrak{U}}^{b}_{\zeta}$-module, and the ${\mathfrak{U}}^{b}_{\zeta}$-action on $L(\lambda)_{\mathbb{C}}$ factors through the quotient ${\mathfrak{U}}^{b}_{\zeta}\twoheadrightarrow u_{\zeta}$.
@@
 in $U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{R}$.
-Lemma 2.3 ([9, Lem 3.7] and [37, Lem 3.2]).
+Lemma 2.4 ([7, Lem 3.7] and [35, Lem 3.2]).
   1. (1)
@@
   2. (2)
-    The projective cover for $L(\lambda)_{\mathbb{C}}$ in ${\mathcal{O}}^{\leq\nu}_{S}$ (resp.
-${\mathcal{O}}^{\leq\nu}_{{\mathbb{C}}}$) is $Q(\lambda)_{S}^{\leq\nu}=\tau^{\leq\nu}Q(\lambda)_{S}$ (resp.
-$Q(\lambda)_{\mathbb{C}}^{\leq\nu}=\tau^{\leq\nu}Q(\lambda)_{\mathbb{C}}$).
+    Let $\Omega$ be a poset ideal in $\Lambda$, and let $\lambda\in\Omega$.
+The projective cover for $L(\lambda)_{\mathbb{C}}$ in ${\mathcal{O}}^{\in\Omega}_{S}$ (resp.
+${\mathcal{O}}^{\in\Omega}_{{\mathbb{C}}}$) is $Q(\lambda)_{S}^{\Omega}=\tau^{\Omega}Q(\lambda)_{S}$ (resp.
+$Q(\lambda)_{\mathbb{C}}^{\Omega}=\tau^{\Omega}Q(\lambda)_{\mathbb{C}}$).
 2.3.3. Block decomposition
@@
 Then there are algebra homomorphisms $Z_{\mathrm{HC}}\hookrightarrow{\mathfrak{U}}_{\zeta}\rightarrow U$.
-By the Harish-Chandra isomorphism, we have an algebra homomorphism
-${\mathbb{C}}[T/(W,\bullet)]\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda}_{S}).$
-By the $S$-linearity, it extends to a homomorphism
+Let $R$ be a commutative Noetherian $S$-algebra.
+We have algebra homomorphisms
+$S\otimes Z_{\mathrm{HC}}\rightarrow R\otimes Z_{\mathrm{HC}}\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda}_{R}).$
+By the Harish-Chandra isomorphism, it factors through a homomorphism
 (2.6) ${\mathbb{C}}[T_{\hat{1}}\times_{T/W}T/(W,\bullet)]\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda}_{S}).$
@@
 where $W_{\omega}$ is the stabilizer of the $(W,\bullet)$-action on $\zeta^{\omega}\in T$. The idempotents on the RHS of (2.7) yield a block decomposition
-$U\text{-}\mathrm{Mod}^{\Lambda}_{S}=\bigoplus_{\omega\in\Xi}U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S},$
-such that the Verma module $M^{U}(\lambda)_{S}$ is contained in $U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S}$ if and only if $\omega=W_{\mathrm{ex}}\bullet_{l}\lambda$. We have algebra homomorphisms
-${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}/W_{\omega}]_{\hat{0}}\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S}),\quad\forall\omega\in\Xi.$
-The block decomposition induces a decomposition for the subcategory ${\mathcal{O}}^{U}_{S}$, and extends to a decomposition
-${\mathcal{O}}^{U}_{R}=\bigoplus_{\omega\in\Xi}{\mathcal{O}}^{U,\omega}_{R},$
-for any commutative Noetherian $S$-algebra $R$.
-Note that when $U={\mathfrak{U}}^{t}_{\zeta}$ or ${\mathfrak{U}}^{b}_{\zeta}$, the category ${\mathcal{O}}$ for $U$ is exactly $U\text{-}\mathrm{mod}^{\Lambda}_{R}$.
+$U\text{-}\mathrm{Mod}^{\Lambda}_{R}=\bigoplus_{\omega\in\Xi}U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{R},$
+such that the Verma module $M^{U}(\lambda)_{R}$ is contained in $U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{R}$ if and only if $\omega=W_{\mathrm{ex}}\bullet_{l}\lambda$. We have algebra homomorphisms
+(2.8) ${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}/W_{\omega}]_{\hat{0}}\rightarrow Z(U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S}),\quad\forall\omega\in\Xi.$
+The block decomposition induces a decomposition
+${\mathcal{O}}^{U}_{R}=\bigoplus_{\omega\in\Xi}{\mathcal{O}}^{U,\omega}_{R}.$
+Since $u_{\zeta}$, $u_{\zeta}U^{+}_{\zeta}$ are quotients of ${\mathfrak{U}}_{\zeta}^{t}$, ${\mathfrak{U}}^{t}_{\zeta}U^{+}_{\zeta}$, we obtain block decompositions
+$u_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}=\bigoplus_{\omega\in\Xi}u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,\omega}_{\mathbb{C}}\quad\text{and}\quad{\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta}}_{\mathbb{C}}=\bigoplus_{\omega\in\Xi}{\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta},\omega}_{\mathbb{C}}.$
 Remark 2.5.
 Since $U$ is $\mathrm{Q}$}-graded and $U\text{-}\mathrm{Mod}^{\Lambda}_{S}$ consists of $\Lambda$-graded modules, the blocks above labelled by $\omega\in\Xi$ further decompose into direct sums of $\mathrm{e}$} equivalent sub-blocks.
-Since $u_{\zeta}$, $u_{\zeta}U^{+}_{\zeta}$ are quotients of ${\mathfrak{U}}_{\zeta}^{t}$, ${\mathfrak{U}}^{t}_{\zeta}U^{+}_{\zeta}$, we have block decompositions
-$u_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}=\bigoplus_{\omega\in\Xi}u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,\omega}_{\mathbb{C}}\quad\text{and}\quad{\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta}}_{\mathbb{C}}=\bigoplus_{\omega\in\Xi}{\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta},\omega}_{\mathbb{C}}.$
+For a subset $\Omega\subset\Lambda$, we abbreviate $U\text{-}\mathrm{Mod}^{\Omega,\omega}_{S}:=U\text{-}\mathrm{Mod}^{\Lambda,\omega}_{S}\cap U\text{-}\mathrm{Mod}^{\Omega}_{S}$ and ${\mathcal{O}}^{U,\omega,\in\Omega}_{R}:={\mathcal{O}}^{U,\omega}_{R}\cap U\text{-}\mathrm{Mod}^{\Omega}_{S}$.
+Again, for $\Omega=\{\lambda\in\Lambda|\lambda\leq\nu\}$, we abbreviate $\Omega$” and $\in\Omega$” by $\leq\nu$” in the superscripts.
+We will also drop the superscripts $U$ from the notations above when $U=U^{\mathrm{hb}}_{\zeta}$.
 2.3.4. Translation functors
+Let $R$ be a commutative Noetherian $S$-algebra.
 For any $\omega_{1},\omega_{2}\in\Xi$, there is a unique dominant element $\nu$ in the $W$-orbit of $\omega_{2}-\omega_{1}$.
-Denote by $V(\nu)$ the Weyl module for $U_{\zeta}$ of highest weight $\nu$.
+Denote by $V(\nu)$ the Weyl module for $U_{\zeta}$ of highest weight $\nu$, see e.g.
+[2].
 We view $V(\nu)$ as a $U$-module via the natural map $U\rightarrow U_{\zeta}$.
 We define the translation functors by
-$T_{\omega_{1}}^{\omega_{2}}:U\text{-}\mathrm{Mod}^{\Lambda,\omega_{1}}_{S}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda,\omega_{2}}_{S},\quad M\mapsto\mathrm{pr}_{\omega_{2}}(M\otimes V(\nu)),$
-$T^{\omega_{1}}_{\omega_{2}}:U\text{-}\mathrm{Mod}^{\Lambda,\omega_{2}}_{S}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda,\omega_{1}}_{S},\quad M\mapsto\mathrm{pr}_{\omega_{1}}(M\otimes V(\nu)^{*}),$
-where $\mathrm{pr}_{\omega_{i}}$ is the natural projection to the block $U\text{-}\mathrm{Mod}^{\Lambda,\omega_{i}}_{S}$.
-Note that $T_{\omega_{1}}^{\omega_{2}}$ and $T^{\omega_{1}}_{\omega_{2}}$ are exact and biadjoint to each other.
-The translation functors preserve the subcategory ${\mathcal{O}}^{U}_{S}$ and induce translation functors between blocks in ${\mathcal{O}}^{U}_{R}$, for any commutative Noetherian $S$-algebra $R$.
+${\mathsf{T}}_{\omega_{1}}^{\omega_{2}}:U\text{-}\mathrm{Mod}^{\Lambda,\omega_{1}}_{R}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda,\omega_{2}}_{R},\quad M\mapsto\mathrm{pr}_{\omega_{2}}(M\otimes V(\nu)),$
+${\mathsf{T}}^{\omega_{1}}_{\omega_{2}}:U\text{-}\mathrm{Mod}^{\Lambda,\omega_{2}}_{R}\rightarrow U\text{-}\mathrm{Mod}^{\Lambda,\omega_{1}}_{R},\quad M\mapsto\mathrm{pr}_{\omega_{1}}(M\otimes V(\nu)^{*}),$
+where $\mathrm{pr}_{\omega_{i}}$ is the natural projection to the block $U\text{-}\mathrm{Mod}^{\Lambda,\omega_{i}}_{R}$.
+Note that ${\mathsf{T}}_{\omega_{1}}^{\omega_{2}}$ and ${\mathsf{T}}^{\omega_{1}}_{\omega_{2}}$ are exact and biadjoint to each other.
+The translation functors preserve the subcategory ${\mathcal{O}}^{U}_{R}$ and induce translation functors between blocks in ${\mathcal{O}}^{U}_{R}$.
 For any simple reflection $s\in{\mathrm{I}}_{\mathrm{af}}$, we fix an element $\omega_{s}\in\Xi_{\mathrm{sc}}$ whose stabilizer under the $W_{\mathrm{af}}$-action is $\{1,s\}$.
-We abbreviate $T^{0}_{s}=T^{0}_{\omega_{s}}$ and $T_{0}^{s}=T_{0}^{\omega_{s}}$.
-Define the reflection functor on $U\text{-}\mathrm{Mod}^{\Lambda,0}_{S}$ by
-$\Theta^{r}_{s}:=T^{0}_{s}\circ T_{0}^{s}.$
-[…]
+We abbreviate ${\mathsf{T}}^{0}_{s}={\mathsf{T}}^{0}_{\omega_{s}}$ and ${\mathsf{T}}_{0}^{s}={\mathsf{T}}_{0}^{\omega_{s}}$.
+Define the reflection functor on $U\text{-}\mathrm{Mod}^{\Lambda,0}_{R}$ by
+$\Theta^{r}_{s}:={\mathsf{T}}^{0}_{s}\circ{\mathsf{T}}_{0}^{s}.$
+We also consider the functor ${\mathbb{T}}^{r}_{s}:=\ker(\Theta^{r}_{s}\xrightarrow{\text{counit}}1)$.
 We also obtain translations and reflections functors for the blocks of $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$ and ${\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta}}_{\mathbb{C}}$.
-We similarly define translation functors and reflection functors on the categories ${\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{\chi_{\omega}}$ with $\omega\in\Xi$.
+We can similarly define translation functors and reflection functors on the categories ${\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{\chi_{\omega}}$ with $\omega\in\Xi$.
 2.3.5. $\Lambda$-translations
@@
 For any $\nu\in\Lambda$, there is a trivial $U^{\prime}$-module ${\mathbb{C}}_{l\nu}$ supported on the weight $l\nu$.
-It gives auto-equivalences $-\otimes{\mathbb{C}}_{l\nu}$, $\nu\in\Lambda$ on $U^{\prime}\text{-}\mathrm{Mod}^{\Lambda}_{R}$ preserving each blocks, called the $\Lambda$-translations.
+It gives auto-equivalences $\otimes{\mathbb{C}}_{l\nu}$, $\nu\in\Lambda$ on $U^{\prime}\text{-}\mathrm{Mod}^{\Lambda}_{R}$ preserving each blocks, called the $\Lambda$-translations.
 3. Non-commutative Springer resolutions and localization theorem
@@
 3.1. Non-commutative Springer resolutions
-For any standard parabolic subgroup $P$ in $G$, we set $\tilde{{\mathfrak{g}}}_{P}=G\times^{P}\mathrm{Lie}(P)$ and abbreviate $\tilde{{\mathfrak{g}}}=\tilde{{\mathfrak{g}}}_{B}$.
-There is a natural map $\pi_{P}:\tilde{{\mathfrak{g}}}\rightarrow\tilde{{\mathfrak{g}}}_{P}$.
-The Grothendieck resolution is $\pi:\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$.
-If $P=P_{s}$ is the minimal parabolic subgroup associated to $s\in{\mathrm{I}}$, we will abbreviate $\tilde{{\mathfrak{g}}}_{s}=\tilde{{\mathfrak{g}}}_{P_{s}}$ and $\pi_{s}=\pi_{P_{s}}$.
-Denote by ${\mathcal{O}}_{\mathcal{B}}(\lambda)$ the line bundle $G\times^{B}{\mathbb{C}}_{\lambda}\rightarrow{\mathcal{B}}$, and denote by ${\mathcal{O}}_{Y}(\lambda)$ its pullback along any $Y\rightarrow{\mathcal{B}}$.
-Consider the ${\mathbb{C}}^{\times}$-action on ${\mathfrak{g}}$ via $t.x=t^{-2}x$, for any $t\in{\mathbb{C}}^{\times}$, $x\in{\mathfrak{g}}$.
-It induces ${\mathbb{C}}^{\times}$-actions on $\tilde{{\mathfrak{g}}}_{P}$ for any $P$.
+[…]
 3.1.1. Affine braid group actions
 Let $X$ be a $\mathfrak{g}$}-scheme satisfying
-${{\mathcal{T}}or}_{i}^{{\mathcal{O}}_{\mathfrak{g}}}({\mathcal{O}}_{\tilde{{\mathfrak{g}}}},{\mathcal{O}}_{X})=0,\quad\forall i>0.$
-We set $\tilde{{\mathfrak{g}}}_{P,X}:=\tilde{{\mathfrak{g}}}_{P}\times_{\mathfrak{g}}X$ and write $\pi_{P}=\pi_{P}\times_{\mathfrak{g}}X$ by abuse of notations.
-There is a functor $\Theta_{P}^{c}:=\pi_{P}^{*}\pi_{P*}$ on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$.
+${{\mathcal{T}}or}_{i}^{{\mathcal{O}}_{\mathfrak{g}}}({\mathcal{O}}_{\widetilde{{\mathfrak{g}}}},{\mathcal{O}}_{X})=0,\quad\forall i>0.$
+We set $\widetilde{{\mathfrak{g}}}_{P,X}:=\widetilde{{\mathfrak{g}}}_{P}\times_{\mathfrak{g}}X$ and write $\pi_{P}=\pi_{P}\times_{\mathfrak{g}}X$ by abuse of notations.
+There is a functor $\Theta_{P}^{c}:=\pi_{P}^{*}\pi_{P*}$ on $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$.
 We abbreviate $\Theta^{c}_{s}=\Theta^{c}_{P_{s}}$ for any $s\in{\mathrm{I}}$, which are called the reflection functors.
 Theorem 3.1 ([12]).
-For any $s\in{\mathrm{I}}$, there is a self-equivalence ${\mathbb{T}}^{c}_{s}$ on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$ with natural transformations
-[…]
-which is a distinguished triangle when applying on any object in $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$. The assignments
-$\tilde{s}\mapsto({\mathbb{T}}^{c}_{s})^{-1}\quad\text{and}\quad\theta_{\lambda}\mapsto-\otimes^{L}_{{\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{X}}}{\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{X}}(\lambda),\qquad\forall s\in{\mathrm{I}},\ \forall\lambda\in\Lambda,$
-define a (weak, right) $B_{\mathrm{ex}}$-action on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$.
-Moreover, for any closed subgroup $K\subset G\times{\mathbb{C}}^{\times}$ and $K$-equivariant morphism $X\rightarrow{\mathfrak{g}}$, the constructions above are naturally $K$-equivariant.
+For any $s\in{\mathrm{I}}$, there is a self-equivalence ${\mathbb{T}}^{c}_{s}$ on $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$ with natural transformations
+${\mathbb{T}}^{c}_{s}\rightarrow\Theta^{c}_{s}\xrightarrow{\text{counit}}1\xrightarrow{+1},$
+which is a distinguished triangle when applying on any object in $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$. The assignments
+$\tilde{s}\mapsto({\mathbb{T}}^{c}_{s})^{-1}\quad\text{and}\quad\theta_{\lambda}\mapsto-\otimes^{L}_{{\mathcal{O}}_{\widetilde{{\mathfrak{g}}}_{X}}}{\mathcal{O}}_{\widetilde{{\mathfrak{g}}}_{X}}(\lambda),\qquad\forall s\in{\mathrm{I}},\ \forall\lambda\in\Lambda,$
+define a (weak, right) $\mathfrak{B}_{\mathrm{ex}}$-action on $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$.
+Moreover, if $X\rightarrow{\mathfrak{g}}$ is equivariant for a closed subgroup $K\subset G\times{\mathbb{C}}^{\times}$, the constructions above are naturally $K$-equivariant.
 […]
 3.1.2. Non-commutative Springer resolutions
-Let $B^{+}_{\mathrm{ex}}$ be the semi-group of $B_{\mathrm{ex}}$ generated by the canonical lifts $\tilde{s}$ with $s\in{\mathrm{I}}_{\mathrm{af}}$.
+Let $\mathfrak{B}^{+}_{\mathrm{ex}}$ be the semi-group of $\mathfrak{B}_{\mathrm{ex}}$ generated by the canonical lifts $\tilde{s}$ with $s\in{\mathrm{I}}_{\mathrm{af}}$.
 Theorem 3.2 ([11]).
   1. (1)
-    There is a unique t-structure (called the exotic t-structure) on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$ defined by
-    $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})^{\leq}=\{{\mathcal{F}}|\ \pi_{*}(b{\mathcal{F}})\in D^{\leq}\mathrm{Coh}X,\ \forall b\in B^{+}_{\mathrm{ex}}\};$
-    $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})^{\geq}=\{{\mathcal{F}}|\ \pi_{*}(b^{-1}{\mathcal{F}})\in D^{\geq}\mathrm{Coh}X,\ \forall b\in B^{+}_{\mathrm{ex}}\}.$
+    There is a unique t-structure (called the exotic t-structure) on $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$ defined by
+    $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})^{\leq}=\{{\mathcal{F}}|\ \pi_{*}(b{\mathcal{F}})\in D^{\leq}\mathrm{Coh}X,\ \forall b\in\mathfrak{B}^{+}_{\mathrm{ex}}\}$;
+    $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})^{\geq}=\{{\mathcal{F}}|\ \pi_{*}(b^{-1}{\mathcal{F}})\in D^{\geq}\mathrm{Coh}X,\ \forall b\in\mathfrak{B}^{+}_{\mathrm{ex}}\}$.
   2. (2)
-    There exists a $G\times{\mathbb{C}}^{\times}$-equivariant tilting bundle ${\mathcal{E}}$ on $\tilde{{\mathfrak{g}}}$ such that
-    $R\mathcal{H}om_{\tilde{{\mathfrak{g}}}_{X}}({\mathcal{E}}_{X},-):\ D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})\rightarrow D^{b}({\mathrm{A}}_{X}\text{-}\mathrm{mod})$
-    is an equivalence, where ${\mathcal{E}}_{X}:={\mathcal{E}}\times_{\mathfrak{g}}X$, ${\mathrm{A}}_{X}={\mathrm{A}}\otimes_{{\mathcal{O}}_{\mathfrak{g}}}{\mathcal{O}}_{X}$ and ${\mathrm{A}}=\mathcal{E}nd_{\tilde{{\mathfrak{g}}}}^{\mathrm{op}}({\mathcal{E}})$, and such that this equivalence is t-exact with respect to the exotic t-structure on the LHS and the trivial t-structure on the RHS.
+    There exists a $G\times{\mathbb{C}}^{\times}$-equivariant tilting bundle $\mathcal{E}$} on $\widetilde{{\mathfrak{g}}}$ such that
+    $R\mathcal{H}om_{\widetilde{{\mathfrak{g}}}_{X}}({\mathcal{E}}_{X},-):\ D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})\rightarrow D^{\mathrm{b}}({\mathsf{A}}_{X}\text{-}\mathrm{mod})$
+    is an equivalence, where ${\mathcal{E}}_{X}:={\mathcal{E}}\times_{\mathfrak{g}}X$, ${\mathsf{A}}_{X}={\mathsf{A}}\otimes_{{\mathcal{O}}_{\mathfrak{g}}}{\mathcal{O}}_{X}$ and ${\mathsf{A}}=\mathcal{E}nd_{\widetilde{{\mathfrak{g}}}}^{\mathrm{op}}({\mathcal{E}})$, and such that this equivalence is t-exact with respect to the exotic t-structure on the LHS and the trivial t-structure on the RHS.
   3. (3)
-    Moreover, for any closed subgroup $K\subset G\times{\mathbb{C}}^{\times}$ and $K$-equivariant morphism $X\rightarrow{\mathfrak{g}}$, we have similar statements for $D^{b}\mathrm{Coh}^{K}(\tilde{{\mathfrak{g}}}_{X})$.
-By definition, the functor $\pi_{*}$ is t-exact for the exotic t-structure on $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{X})$.
-It yields a functor $\pi_{*}:{\mathrm{A}}_{X}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}X$.
+    Moreover, if $X\rightarrow{\mathfrak{g}}$ is equivariant for a closed subgroup $K\subset G\times{\mathbb{C}}^{\times}$, we have similar statements for $D^{\mathrm{b}}\mathrm{Coh}^{K}(\widetilde{{\mathfrak{g}}}_{X})$.
+By definition, the functor $\pi_{*}$ is t-exact for the exotic t-structure on $D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{X})$.
+It yields a functor $\pi_{*}:{\mathsf{A}}_{X}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}(X)$.
 By [11], the functors $\Theta^{c}_{P}$, $\Theta^{c}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$) are also exact for the exotic t-structure.
-We thus view them as functors on ${\mathrm{A}}_{X}\text{-}\mathrm{mod}$.
+We thus view them as functors on ${\mathsf{A}}_{X}\text{-}\mathrm{mod}$.
 Consider the completion $X={\mathfrak{g}}_{\hat{0}}$, then there is an equivalence
-$D^{b}({\mathrm{A}}_{{\mathfrak{g}}_{\hat{0}}}\text{-}\mathrm{mod})\simeq D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{\hat{0}}),$
+[…]
 whose restriction on the full subcategories of objects set-theoretically supported at $0\in{\mathfrak{g}}$ induces an equivalence
-(3.1) $D^{b}({\mathrm{A}}\text{-}\mathrm{mod}_{0})\simeq D^{b}\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}).$
+[…]
 3.2. Localization theorem for ${\mathfrak{U}}_{\zeta}$
-The localization theorem for ${\mathfrak{U}}_{\zeta}$ was established by Backelin–Kremnizer [5] and Tanisaki [44], [45], [42].
+The localization theorem for ${\mathfrak{U}}_{\zeta}$ was established by Backelin–Kremnizer [4] and Tanisaki [42, 43, 39].
 In this subsection, we give a reminder on the localization theorem mainly following [39].
 3.2.1. Quantum differential operators
-Lunts—Rosenberg [28] introduced the quantized flag variety ${\mathcal{B}}_{\zeta}$ as a non-commutative scheme over ${\mathcal{B}}$.
-Let ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$ be the pushforward of the structure sheave of ${\mathcal{B}}_{\zeta}$ on ${\mathcal{B}}$, which is a coherent sheaf of (in general non-commutative) algebras on ${\mathcal{B}}$.
-In [43], Tanisaki defined the ring of enhanced differential operators on ${\mathcal{B}}_{\zeta}$ (note that in loc.
-cit.
-the convention for flag variety is $B^{-}\backslash G$, which is different from us).
-Let $\widetilde{{\mathcal{D}}}_{\zeta}$ be its localization on ${\mathcal{B}}$.
-There is an algebra homomorphism
+[…]
 (3.2) ${\mathfrak{U}}_{\zeta}\otimes{\mathbb{C}}[T]\rightarrow\Gamma(\widetilde{{\mathcal{D}}}_{\zeta}).$
@@
 Note that $\mathcal{V}$} is affine over $\mathcal{B}$}.
-By [44, Thm 5.2], there is an isomorphism from the central subsheaf of $\widetilde{{\mathcal{D}}}_{\zeta}$ to the structure sheaf ${\mathcal{O}}_{{\mathcal{V}}}$ of ${\mathcal{V}}$.
+By [42, Thm 5.2], there is an isomorphism from the central subsheaf of $\widetilde{{\mathcal{D}}}_{\zeta}$ to the structure sheaf ${\mathcal{O}}_{{\mathcal{V}}}$ of $\mathcal{V}$}.
 We thus also view $\widetilde{{\mathcal{D}}}_{\zeta}$ as a (coherent) sheaf of algebras on $\mathcal{V}$}.
@@
 3.2.2. Localization theorem
-From now on, we assume moreover that $l$ is a power of prime, following [42].
+From now on, we assume moreover that $l$ is a prime power, following [39].
 The homomorphism ${\mathfrak{U}}_{\zeta}\rightarrow\Gamma(\widetilde{{\mathcal{D}}}_{\zeta})$ yields a functor
-$R\Gamma:D^{b}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod})\rightarrow D^{b}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}).$
+$R\Gamma:D^{\mathrm{b}}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod})\rightarrow D^{\mathrm{b}}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}).$
 Theorem 3.3 ([39]).
@@
 the stabilizer of $(W_{\mathrm{ex}},\bullet_{l})$-action on $\lambda$ is trivial), then there is an equivalence
-$R\Gamma:D^{b}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod}_{(\chi,\zeta^{2\lambda})}){\ \xrightarrow{\sim}\ }D^{b}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(\chi,\chi_{\lambda})}).$
+$R\Gamma:D^{\mathrm{b}}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod}_{(\chi,\zeta^{2\lambda})}){\ \xrightarrow{\sim}\ }D^{\mathrm{b}}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(\chi,\chi_{\lambda})}).$
 3.2.3. Splitting Azumaya algebras
@@
 Let $(\chi,t)=(1,\zeta^{\lambda})\in{G^{*}}\times_{T/W}T$. Then we have
-${\mathcal{V}}_{(1,\zeta^{\lambda})}={\mathcal{B}}\times\{(1,\zeta^{\lambda})\}={\mathcal{B}}\quad\text{in}\quad{\mathcal{V}}.$
+${\mathcal{V}}_{(1,\zeta^{\lambda})}={\mathcal{B}}\times\{(1,\zeta^{\lambda})\}\simeq{\mathcal{B}}\quad\text{in}\quad{\mathcal{V}}.$
 We abbreviate ${\mathfrak{M}}_{\lambda}={\mathfrak{M}}^{1}_{\zeta^{2\lambda}}$.
 The equivalence (3.3) above depends on the choice of splitting bundles.
-We will fix and recall the construction of splitting bundles ${\mathfrak{M}}_{\lambda}$ given in [42, §5.4] as follows.
+We will fix and recall the construction of splitting bundles ${\mathfrak{M}}_{\lambda}$ given in [39, §5.4] as follows.
 Firstly, we consider the completion of ${\mathfrak{U}}_{\zeta}$ at $(1,\chi_{-\rho})\in G^{*}\times_{T/W}T/(W,\bullet)$.
@@
 In general, in [41] the author constructed a locally free ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$-module ${\mathcal{O}}^{\lambda}_{{\mathcal{B}}_{\zeta}}$ of rank 1 for any $\lambda\in\Lambda$.
-By [42, Thm 5.3], one can set ${\mathfrak{M}}_{\lambda}={\mathcal{O}}_{{\mathcal{B}}_{\zeta}}^{\lambda+\rho}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathfrak{M}}_{{-\rho}}$.
+By [39, Thm 5.3], one can set ${\mathfrak{M}}_{\lambda}={\mathcal{O}}_{{\mathcal{B}}_{\zeta}}^{\lambda+\rho}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathfrak{M}}_{{-\rho}}$.
 3.2.4. Formal neighborhood of $\mathcal{B}$}
@@
 Since $\kappa$ is étale, $\tilde{\kappa}$ is étale. It induces isomorphisms of formal neighborhoods
-$\mathrm{FN}_{{\mathcal{V}}}({\mathcal{B}}\times\{(1,\zeta^{\lambda})\})=\mathrm{FN}_{\widetilde{G^{*}}}({\mathcal{B}})=\mathrm{FN}_{\widetilde{G}}({\mathcal{B}}),\quad\forall\lambda\in\Lambda.$
+$\mathrm{FN}_{{\mathcal{V}}}({\mathcal{B}}\times\{(1,\zeta^{\lambda})\})\simeq\mathrm{FN}_{\widetilde{G^{*}}}({\mathcal{B}})\simeq\mathrm{FN}_{\widetilde{G}}({\mathcal{B}}),\quad\forall\lambda\in\Lambda.$
 The exponential map $\exp:{\mathfrak{g}}\rightarrow G$ induces isomorphisms of schemes
-${\mathfrak{g}}_{\hat{0}}{\ \xrightarrow{\sim}\ }G_{\hat{1}}\quad\text{and}\quad\tilde{{\mathfrak{g}}}_{\hat{0}}{\ \xrightarrow{\sim}\ }\widetilde{G}_{\hat{1}}.$
-Hence we can identify $\mathrm{FN}_{\widetilde{G}}({\mathcal{B}})=\mathrm{FN}_{\tilde{{\mathfrak{g}}}}({\mathcal{B}})$, and it follows that
-$\mathrm{Coh}_{{\mathcal{B}}\times\{(1,\zeta^{2\lambda})\}}({\mathcal{V}})=\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}),\quad\forall\lambda\in\Lambda.$
+${\mathfrak{g}}_{\hat{0}}{\ \xrightarrow{\sim}\ }G_{\hat{1}}\quad\text{and}\quad\widetilde{{\mathfrak{g}}}_{\hat{0}}{\ \xrightarrow{\sim}\ }\widetilde{G}_{\hat{1}}$.
+[…]
+$\mathrm{Coh}_{{\mathcal{B}}\times\{(1,\zeta^{2\lambda})\}}({\mathcal{V}})\simeq\mathrm{Coh}_{{\mathcal{B}}}(\widetilde{{\mathfrak{g}}}),\quad\forall\lambda\in\Lambda.$
 Therefore, by Theorems 3.3 and 3.4, we have an equivalence
-(3.6) $D^{b}\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}){\ \xrightarrow{\sim}\ }D^{b}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{0})}).$
-We set ${\mathrm{A}}_{0}$ be the specialization of ${\mathrm{A}}$ at $0\in{\mathfrak{g}}$.
+(3.6) $D^{\mathrm{b}}\mathrm{Coh}_{{\mathcal{B}}}(\widetilde{{\mathfrak{g}}}){\ \xrightarrow{\sim}\ }D^{\mathrm{b}}({\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{0})}).$
+We set ${\mathsf{A}}_{0}$ be the specialization of $\mathsf{A}$} at $0\in{\mathfrak{g}}$.
 Proposition 3.5 ([39]).
-The equivalence (3.6) is exact for the exotic t-structure on the LHS and the trivial t-structure on the RHS. Therefore, there are equivalences
-(3.7) ${\mathrm{A}}\text{-}\mathrm{mod}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{0})}$
+The equivalence (3.6) is exact for the exotic t-structure on the LHS and the trivial t-structure on the RHS. Therefore, there is an equivalence
+(3.7) ${\mathsf{A}}\text{-}\mathrm{mod}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}_{(1,\chi_{0})}.$
 Moreover, this equivalence intertwines the reflection functors $\Theta^{c}_{s}$ and $\Theta^{r}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$), and there is a commutative diagram
@@
 Therefore, (3.7) can be specialized to an equivalence
-${\mathrm{A}}_{0}\text{-}\mathrm{mod}{\ \xrightarrow{\sim}\ }u_{\zeta}\text{-}\mathrm{mod}_{\chi_{0}}.$
+${\mathsf{A}}_{0}\text{-}\mathrm{mod}{\ \xrightarrow{\sim}\ }u_{\zeta}\text{-}\mathrm{mod}_{\chi_{0}}.$
 3.2.5. Graded version
@@
 $(K_{\mu}-\zeta^{(\mu,\nu)})^{n}.m_{\nu}=0.$
-By [42, §7.4] there exists a $\Lambda$-grading (in completion sense) on $M_{\chi_{-\rho}}$. It yields an equivalence
+By [39, §7.4] there exists a $\Lambda$-grading (in completion sense) on $M_{\chi_{-\rho}}$. It yields an equivalence
 $M_{\chi_{-\rho}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}}-:\ \mathrm{Coh}_{0}^{T}({\mathfrak{g}}){\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{-\rho})}.$
-Moreover, the splitting bundles ${\mathfrak{M}}_{\lambda}$, $\lambda\in\Lambda$ are endowed with $T$-equivariant structures (in completion sense) by their constructions in §3.2.3.
+Moreover, the splitting bundles ${\mathfrak{M}}_{\lambda}$, $\lambda\in\Lambda$ are thus endowed with $\Lambda$-grading (in completion sense) by their constructions in §3.2.3.
 We have the following graded version of Proposition 3.5.
 Proposition 3.6 ([39]).
-The $T$-equivariant structure on ${\mathfrak{M}}_{{0}}$ induces equivalences
-(3.9) ${\mathrm{A}}\text{-}\mathrm{mod}^{T}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})},$
+The $\Lambda$-grading structure on ${\mathfrak{M}}_{{0}}$ induces an equivalence
+(3.9) ${\mathsf{A}}\text{-}\mathrm{mod}^{T}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})},$
 intertwining the reflection functors $\Theta^{c}_{s}$ and $\Theta^{r}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$). There is a commutative diagram
@@
 Hence (3.9) can be specialized to an equivalence
-${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{{\mathbb{C}}}.$
+${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{{\mathbb{C}}}.$
 4. Equivalence ${\mathfrak{F}}^{b}$
 We define the following scheme and algebras
-${\mathfrak{b}}_{S}={\mathfrak{b}}\times_{\mathfrak{t}}{\mathfrak{t}}_{\hat{0}},\quad{\mathrm{A}}^{b}={\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{b}}],\quad\text{and}\quad{\mathrm{A}}^{b}_{S}={\mathrm{A}}^{b}\otimes_{{\mathbb{C}}[{\mathfrak{b}}]}{\mathbb{C}}[{\mathfrak{b}}_{S}].$
-The exponential map induces isomorphisms of schemes ${\mathfrak{n}}{\ \xrightarrow{\sim}\ }N$ and ${\mathfrak{b}}_{S}{\ \xrightarrow{\sim}\ }B\times_{T}T_{\hat{1}}$.
-Since the equivalence (3.9) is ${\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}$-linear, it specializes to an equivalence
-(4.1) ${\mathfrak{F}}^{\prime}:\ {\mathrm{A}}^{b}\text{-}\mathrm{mod}^{T}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}.$
+${\mathfrak{b}}_{S}={\mathfrak{b}}\times_{\mathfrak{t}}{\mathfrak{t}}_{\hat{0}},\quad{\mathsf{A}}^{b}={\mathsf{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{b}}],\quad\text{and}\quad{\mathsf{A}}^{b}_{S}={\mathsf{A}}^{b}\otimes_{{\mathbb{C}}[{\mathfrak{b}}]}{\mathbb{C}}[{\mathfrak{b}}_{S}].$
+The exponential map induces isomorphisms of schemes ${\mathfrak{n}}{\ \xrightarrow{\sim}\ }N$ and $^{\mathfrak{b}}_{S}{\ \xrightarrow{\sim}\ }B\times_{T}T_{\hat{1}}$.
+Since the equivalence (3.9) is $^{\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}$-linear, it specializes to an equivalence
+(4.1) ${\mathfrak{F}}^{\prime}:\ {\mathsf{A}}^{b}\text{-}\mathrm{mod}^{T}_{0}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}.$
 In this section, we firstly extend (4.1) to an equivalence
-${\mathfrak{F}}^{b}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$
+${\mathfrak{F}}^{b}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$
 and then investigate some properties of this equivalence.
-4.1. Equivalence
-Note that ${\mathrm{A}}^{b}\text{-}\mathrm{mod}^{T}_{0}$ is the full subcategory of ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ of the modules set-theoretically supported at $0\in{\mathfrak{b}}_{S}$.
+4.1. Equivalence ${\mathfrak{F}}^{b}$
+Note that ${\mathsf{A}}^{b}\text{-}\mathrm{mod}^{T}_{0}$ is the full subcategory of ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ of the modules set-theoretically supported at $0\in{\mathfrak{b}}_{S}$.
 We also view ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{0})}$ as a full subcategory of ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ as follows.
@@
 induces an $S$-module structure on $M=\bigoplus_{\lambda}M_{\lambda}$, making $M$ as a module in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$.
-Since the reflection functors $\Theta^{c}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$) on ${\mathrm{A}}\text{-}\mathrm{mod}$ are $G$-equivariant and ${\mathbb{C}}[{\mathfrak{g}}]$-linear, they induce reflection functors (still denoted by $\Theta^{c}_{s}$) on ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{K}$ by base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$, for any subgroup $K$ in $B$.
+Since the reflection functors $\Theta^{c}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$) on ${\mathsf{A}}\text{-}\mathrm{mod}$ are $G$-equivariant and ${\mathbb{C}}[{\mathfrak{g}}]$-linear, they induce reflection functors (still denoted by $\Theta^{c}_{s}$) on ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{K}$ by base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$, for any subgroup scheme $K$ in $B$.
 Proposition 4.1.
 The equivalence $\mathfrak{F}$$^{\prime}$ in (4.1) extends to an $S$-linear equivalence
-(4.2) ${\mathfrak{F}}^{b}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S},$
+(4.2) ${\mathfrak{F}}^{b}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S},$
 compatible with the reflection functors $\Theta^{c}_{s}$, $\Theta^{r}_{s}$ ($s\in{\mathrm{I}}_{\mathrm{af}}$) and $\Lambda$-translations on both sides.
@@
 Consider the specializations
-$S_{n}=S/({\mathfrak{m}}_{1}\cap{\mathbb{C}}[T])^{n+1}S,\qquad{\mathrm{A}}^{b}_{n}={\mathrm{A}}^{b}_{S}/{\mathfrak{m}}_{0}^{n+1}{\mathrm{A}}^{b}_{S}\quad\text{and}\quad{\mathfrak{U}}^{b}_{\zeta,n}={\mathfrak{U}}^{b}_{\zeta}/{\mathfrak{m}}_{1}^{n+1}{\mathfrak{U}}^{b}_{\zeta}.$
-Since ${\mathfrak{F}}^{\prime}$ is ${\mathbb{C}}[{\mathfrak{b}}]_{\hat{0}}$-linear, it yields equivalences
-${\mathfrak{F}}^{\prime}:\ {\mathrm{A}}^{b}_{n}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}},\quad\forall n\geq 0.$
+$S_{n}=S/({\mathfrak{m}}_{1}\cap{\mathbb{C}}[T])^{n+1}S,\qquad{\mathsf{A}}^{b}_{n}={\mathsf{A}}^{b}_{S}/{\mathfrak{m}}_{0}^{n+1}{\mathsf{A}}^{b}_{S}\quad\text{and}\quad{\mathfrak{U}}^{b}_{\zeta,n}={\mathfrak{U}}^{b}_{\zeta}/{\mathfrak{m}}_{1}^{n+1}{\mathfrak{U}}^{b}_{\zeta}.$
+Since $\mathfrak{F}$$^{\prime}$ is $^{\mathbb{C}}[{\mathfrak{b}}]_{\hat{0}}$-linear, it yields equivalences
+${\mathfrak{F}}^{\prime}:\ {\mathsf{A}}^{b}_{n}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}},\quad\forall n\geq 0.$
 There are natural functors by specializations
-$\tau_{n}:{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}\rightarrow{\mathrm{A}}^{b}_{n}\text{-}\mathrm{mod}^{T},\quad\tau_{n}:{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}\rightarrow{\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}},$
+$\tau_{n}:{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}\rightarrow{\mathsf{A}}^{b}_{n}\text{-}\mathrm{mod}^{T},\quad\tau_{n}:{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}\rightarrow{\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}},$
 and $\mathfrak{F}$$^{\prime}$ is clearly compatible with $\tau_{n}$.
-Note that ${\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{n})$ is projective in ${\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}}$, the family $\{{\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{n})\}_{n}$ admits a lifting $P^{b}$ in $\mathrm{Proj}({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S})$.
-We choose a family of isomorphisms ${\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{n})\cong\tau_{n}(P^{b})$, $n\geq 0$ that are compatible with specializations.
+Note that ${\mathfrak{F}}^{\prime}({\mathsf{A}}^{b}_{n})$ is projective in ${\mathfrak{U}}^{b}_{\zeta,n}\text{-}\mathrm{mod}^{\Lambda,0}_{S_{n}}$, the family $\{{\mathfrak{F}}^{\prime}({\mathsf{A}}^{b}_{n})\}_{n}$ admits a lifting $P^{b}$ in $\mathrm{Proj}({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S})$.
+We choose a family of isomorphisms ${\mathfrak{F}}^{\prime}({\mathsf{A}}^{b}_{n})\cong\tau_{n}(P^{b})$, $n\geq 0$ that are compatible with specializations.
 To proceed, we need the following lemmas.
@@
 Proof of Lemma 4.3.
-(1) Since $\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ is a projective morphism to an affine scheme, ${\mathrm{A}}=\mathrm{End}^{\mathrm{op}}_{\tilde{{\mathfrak{g}}}}({\mathcal{E}})$ is a finitely generated ${\mathbb{C}}[{\mathfrak{g}}]$-module by Serre’s Theorem.
-Hence ${\mathrm{A}}^{b}_{S}$ is a finitely generated ${\mathbb{C}}[{\mathfrak{b}}_{S}]$-module.
-Note that ${\mathrm{A}}^{b}_{S}$ is moreover $\Lambda$-graded and complete over $S$.
+(1) Since $\widetilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ is a projective morphism to an affine scheme, ${\mathsf{A}}=\mathrm{End}^{\mathrm{op}}_{\widetilde{{\mathfrak{g}}}}({\mathcal{E}})$ is a finitely generated ${\mathbb{C}}[{\mathfrak{g}}]$-module by Serre’s Theorem.
+Hence ${\mathsf{A}}^{b}_{S}$ is a finitely generated ${\mathbb{C}}[{\mathfrak{b}}_{S}]$-module.
+Note that ${\mathsf{A}}^{b}_{S}$ is moreover $\Lambda$-graded and complete over $S$.
 We view ${\mathbb{C}}[{\mathfrak{b}}_{S}]$ as a ${\mathbb{Z}}_{\geq 0}$-graded ring via ${\mathbb{C}}[{\mathfrak{b}}_{S}]_{i}=\bigoplus_{\mathrm{ht}(\eta)=i}{\mathbb{C}}[{\mathfrak{b}}_{S}]_{-\eta}$, and consider the graded ideal ${\mathfrak{m}}_{0}{\mathbb{C}}[{\mathfrak{b}}_{S}]$, then our assertion follows from Lemma 4.2.
@@
 where the first and the second last equalities are by Lemma 4.3.
-Note that the simple modules of ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ are contained in $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}={\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{0}\text{-}\mathrm{mod}^{T})$, and therefore they admit a surjection from $\bigoplus_{\nu}{\mathfrak{F}}^{\prime}({\mathrm{A}}^{b}_{1}\otimes{\mathbb{C}}_{\nu})=\bigoplus_{\nu}\tau_{1}(P^{b}\otimes{\mathbb{C}}_{l\nu})$.
+Note that the simple modules of ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ are contained in $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}={\mathfrak{F}}^{\prime}({\mathsf{A}}^{b}_{0}\text{-}\mathrm{mod}^{T})$, and therefore they admit a surjection from $\bigoplus_{\nu}{\mathfrak{F}}^{\prime}({\mathsf{A}}^{b}_{1}\otimes{\mathbb{C}}_{\nu})=\bigoplus_{\nu}\tau_{1}(P^{b}\otimes{\mathbb{C}}_{l\nu})$.
 It implies that $\{P^{b}\otimes{\mathbb{C}}_{l\nu}\}_{\nu}$ forms a generating family of projective modules in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$.
@@
 Now our desired equivalence (4.2) follows from the isomorphism (4.4).
-By Proposition 3.6, the equivalence (4.1) is compatible with reflection functors and $\Lambda$-translations. The compatibility for (4.2) follows. ∎
-We set ${\mathrm{A}}_{\mathfrak{n}}:={\mathrm{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]$.
+Finally, by Proposition 3.6, the equivalence (4.1) is compatible with reflection functors and $\Lambda$-translations. The compatibility for (4.2) follows. ∎
+We set ${\mathsf{A}}_{\mathfrak{n}}:={\mathsf{A}}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{n}}]$.
 Corollary 4.4.
 There is an equivalence of abelian categories
-(4.5) ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}.$
-4.2. Steinberg block and functor ${\mathbb{V}}^{b}$
-4.2.1. Equivalence for Steinberg block
-Recall the equivalence (3.4). By the ${\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}$-linearity it specializes to an equivalence
+(4.5) ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}.$
+4.2. Equivalence ${\mathfrak{G}}^{b}$ for Steinberg block
+Recall the equivalence (3.4). By the $^{\mathbb{C}}[{\mathfrak{g}}]_{\hat{0}}$-linearity it specializes to an equivalence
 $M^{b}_{\chi_{-\rho}}\otimes_{{\mathbb{C}}[{\mathfrak{b}}]_{\hat{0}}}-:\ \mathrm{Coh}_{0}^{T}({\mathfrak{b}}){\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{(1,\chi_{-\rho})},$
@@
 (4.7) $M^{b}_{\chi_{-\rho}}\cong M(-\rho)_{S}\otimes_{{\mathbb{C}}[{\mathfrak{b}}_{S}]}{\mathbb{C}}[{\mathfrak{b}}]_{\hat{0}}$.
-We can adjust the $\Lambda$-grading on $M_{\chi_{-\rho}}$ such that the isomorphism (4.7) respects the $\Lambda$-gradings on both sides. Therefore, the equivalence (4.6) is given by
+We can adjust the $\Lambda$-grading on $M_{\chi_{-\rho}}$ such that the isomorphism (4.7) respects the $\Lambda$-gradings (in completion sense) on both sides.
+Therefore, the equivalence (4.6) is given by
 ${\mathfrak{G}}^{b}=M(-\rho)_{S}\otimes_{{\mathbb{C}}[{\mathfrak{b}}_{S}]}-.$
-For any commutative Noetherian $S$-algebra $R$, the equivalence ${\mathfrak{G}}^{b}$ induces an equivalence
-${\mathfrak{G}}^{b}:\ \mathrm{Coh}^{T}({\mathfrak{b}}_{R}){\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{R},$
-where ${\mathfrak{b}}_{R}:={\mathfrak{b}}_{S}\times_{\mathrm{Spec}S}\mathrm{Spec}R$.
-Since the functor $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}})$ is ${\mathbb{C}}[{\mathfrak{g}}]$-linear and $G$-equivariant, it induces a functor $\pi_{*}:{\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}\rightarrow\mathrm{Coh}^{T}({\mathfrak{b}}_{S})$.
+In particular, we have an isomorphism ${\mathfrak{G}}^{b}({\mathbb{C}}[{\mathfrak{b}}_{S}]\otimes{\mathbb{C}}_{\lambda})=M(-\rho+l\lambda)_{S}$ for any $\lambda\in\Lambda$.
+By (2.5) the equivalence ${\mathfrak{G}}^{b}$ induces an equivalence
+${\mathfrak{G}}^{b}:\ \mathrm{Coh}^{T}({\mathfrak{b}}_{R}){\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{R}$
+for any commutative Noetherian $S$-algebra $R$, where ${\mathfrak{b}}_{R}:={\mathfrak{b}}_{S}\times_{\mathrm{Spec}S}\mathrm{Spec}R$.
+[…]
+$\pi_{*}:{\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{K}\rightarrow\mathrm{Coh}^{K}({\mathfrak{b}}_{S})$
+which is represented by the $({\mathbb{C}}[{\mathfrak{b}}_{S}],{\mathsf{A}}_{S}^{b})$-bimodule $\Gamma(\widetilde{{\mathfrak{g}}},{\mathcal{E}})\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{b}}_{S}]$.
+The functor $\pi_{*}$ admits a right adjunction
+$\pi^{!}:=\mathrm{Hom}_{{\mathbb{C}}[{\mathfrak{b}}_{S}]}(\Gamma(\widetilde{{\mathfrak{g}}},{\mathcal{E}})\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{b}}_{S}],-):\ \mathrm{Coh}^{K}({\mathfrak{b}}_{S})\rightarrow{\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{K}.$
 From Proposition 3.6 we deduce the following
 Proposition 4.5.
-There is a commutative diagram
-[…]
-Recall that by [36, §3.1.2] there is a subalgebra ${\mathfrak{U}}^{\mathrm{hb}}_{\zeta}=Z_{\mathrm{Fr}}^{-}\otimes{\mathfrak{U}}^{0}_{\zeta}\otimes U^{+}_{\zeta}$ of $U^{\mathrm{hb}}_{\zeta}$, equipped with a Frobenius map
+There are commutative diagrams
+(4.8) ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$$\mathrm{Coh}^{T}({\mathfrak{b}}_{S})$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S},$$\scriptstyle{{\mathfrak{F}}^{b}}\scriptstyle{\simeq}$$\scriptstyle{\pi_{*}}$$\scriptstyle{{\mathsf{T}}^{-\rho}_{0}}$$\scriptstyle{{\mathfrak{G}}^{b}}\scriptstyle{\simeq}$
+and
+(4.9) ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$$\mathrm{Coh}^{T}({\mathfrak{b}}_{S})$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S}.$$\scriptstyle{{\mathfrak{F}}^{b}}\scriptstyle{\simeq}$$\scriptstyle{{\mathfrak{G}}^{b}}\scriptstyle{\simeq}$$\scriptstyle{\pi^{!}}$$\scriptstyle{{\mathsf{T}}_{-\rho}^{0}}$
+4.3. Induction functors
+Recall that by [34, §3.1.2] there is a subalgebra ${\mathfrak{U}}^{\mathrm{hb}}_{\zeta}=Z_{\mathrm{Fr}}^{-}\otimes{\mathfrak{U}}^{0}_{\zeta}\otimes U^{+}_{\zeta}$ of $U^{\mathrm{hb}}_{\zeta}$, equipped with a Frobenius map
 $\mathrm{Fr}:\ {\mathfrak{U}}^{\mathrm{hb}}_{\zeta}\rightarrow{\mathbb{C}}[B\times_{T}T]\rtimes U{\mathfrak{n}}$
-compatible with the quantum Frobenius map $U^{+}_{\zeta}\rightarrow U{\mathfrak{n}}$, and the isomorphism $Z^{-}_{\mathrm{Fr}}\otimes{\mathfrak{U}}^{0}_{\zeta}\rightarrow{\mathbb{C}}[B\times_{T}T]$ by (2.1), where $T\rightarrow T$ is by the $l$-th power.
+which is compatible with the quantum Frobenius map $U^{+}_{\zeta}\rightarrow U{\mathfrak{n}}$, and with the isomorphism $Z^{-}_{\mathrm{Fr}}\otimes{\mathfrak{U}}^{0}_{\zeta}{\ \xrightarrow{\sim}\ }{\mathbb{C}}[B\times_{T}T]$ induced by (2.1), where $T\rightarrow T$ is by the $l$-th power.
 The map $\mathrm{Fr}$ is a homomorphism of $\Lambda$-graded algebras, where we view ${\mathbb{C}}[B\times_{T}T]\rtimes U{\mathfrak{n}}$ as an $l\Lambda$-graded ring.
@@
 Define an induction functor
-$\mathrm{ind}_{c}:=({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\otimes_{{\mathbb{C}}[{\mathfrak{b}}_{R}]}-:\ \mathrm{Coh}^{T}({\mathfrak{b}}_{R})\rightarrow({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}.$
+$\mathrm{ind}_{c}:=({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\otimes_{{\mathbb{C}}[{\mathfrak{b}}_{R}]}-:\ \mathrm{QCoh}^{T}({\mathfrak{b}}_{R})\rightarrow({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}.$
 Note that $\mathrm{ind}_{c}M=U{\mathfrak{n}}\otimes M$ as vector spaces for any ${\mathbb{C}}[{\mathfrak{b}}_{R}]$-module $M$. We also consider the induction
 $\mathrm{ind}_{r}:=U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}-:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{R}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{R}.$
-Note that $\mathrm{ind}_{r}$ preserves the block decomposition.
+Note that $\mathrm{ind}_{r}$ preserves the block decomposition in §2.3.3.
 Proposition 4.6.
-There is a commutative diagram
-[…]
+Let $R$ be a commutative Noetherian $S$-algebra. There is a commutative diagram
+(4.10) $\mathrm{Coh}^{T}({\mathfrak{b}}_{R})$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{R}$$({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$$U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{R}.$$\scriptstyle{{\mathfrak{G}}^{b}}\scriptstyle{\simeq}$$\scriptstyle{\mathrm{ind}_{c}}$$\scriptstyle{\mathrm{ind}_{r}}$$\scriptstyle{{\mathrm{I}}^{\mathrm{hb}}_{1}}$
 Proof.
@@
 We thus obtain a natural isomorphism ${\mathrm{I}}^{\mathrm{hb}}_{1}\circ\mathrm{ind}_{c}=\mathrm{ind}_{r}\circ{\mathfrak{G}}^{b}$. ∎
-4.2.2. Functors ${\mathbb{V}}^{b}_{r}$ and ${\mathbb{V}}^{b}_{c}$
-Recall that ${\mathrm{A}}=\mathrm{End}^{\mathrm{op}}_{\tilde{{\mathfrak{g}}}}({\mathcal{E}})$ for some tilting bundle ${\mathcal{E}}$ on $\tilde{{\mathfrak{g}}}$.
-By [12, Lem 2.5.3], ${\mathcal{E}}$ has ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}}$ as a direct summand.
-It follows that ${\mathrm{A}}$ is an algebra over $\Gamma({\mathcal{O}}_{\tilde{{\mathfrak{g}}}})={\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$, and that ${\mathrm{A}}$ has ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ as a direct summand as a ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$-module.
-Moreover, the subalgebra ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ is central in ${\mathrm{A}}$.
-Hence ${\mathrm{A}}_{S}^{b}$ is an algebra over ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ and ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\subset Z({\mathrm{A}}_{S}^{b})$.
-We have ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]^{T}={\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}$.
-Therefore, we have an algebra homomorphism
-(4.8) ${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}\rightarrow Z({\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}).$
-Note that the functor $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}})$ is represented by the $({\mathbb{C}}[{\mathfrak{g}}],{\mathrm{A}})$-bimodule $\Gamma(\tilde{{\mathfrak{g}}},{\mathcal{E}})$.
-Hence $\pi_{*}$ can be enhanced to $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$ and induces a functor
-${\mathbb{V}}^{b}_{c}:\ {\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}\rightarrow\mathrm{Coh}^{T}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$
-by base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$.
+Let $R$ be a commutative Noetherian $S$-algebra.
+Let $\nu\in\Lambda$.
+We denote the truncation functor on $U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{R}$ by the poset ideal $\{\lambda\in\Lambda|\lambda\leq\nu\}$ (see §2.3.1) as
+$\tau_{r}^{\leq\nu}:\ U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{R}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\leq\nu}_{R}.$
+We denote by
+${\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{Mod}^{B,\leq\nu}$
+the category of modules in $({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$ whose $T$-weights are $\leq\nu$, which is a subcategory in ${\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{Mod}^{B}$ (since the $U{\mathfrak{n}}$-action on a module with $T$-weights $\leq\nu$ is locally unipotent).
+By similar argument as in §2.3.1, there is a truncation functor
+$\tau^{\leq\nu}_{c}:\ ({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow{\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{Mod}^{B,\leq\nu}$
+by taking the maximal quotient whose $T$-weights are $\leq\nu$, which is left adjoint to the natural inclusion.
+We abbreviate ${\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{mod}^{B,\leq\nu}:={\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{Mod}^{B,\leq\nu}\cap{\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{mod}^{B}$.
+Lemma 4.7.
+There is a natural isomorphism of functors
+(4.11) ${\mathrm{I}}^{\mathrm{hb}}_{1}\circ\tau^{\leq\nu}_{c}=\tau^{\leq-\rho+l\nu}_{r}\circ{\mathrm{I}}^{\mathrm{hb}}_{1}.$
+from $({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$ to $U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\leq-\rho+l\nu}_{R}$.
+Proof.
+By the right exactness of the functors, it is enough to construct natural isomorphism on the projective generators ${\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}}\otimes{\mathbb{C}}_{\lambda}$, with $\lambda\in\Lambda$.
+We have an isomorphism of $\Lambda$-graded $U^{\mathrm{hb}}_{\zeta}\otimes R$-modules
+${\mathrm{I}}^{\mathrm{hb}}_{1}({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}}\otimes{\mathbb{C}}_{\lambda})=U^{\mathrm{hb}}_{\zeta}\otimes_{U^{\mathrm{hb},\geq}_{\zeta}}(U{\mathfrak{n}}\otimes R_{-\rho+l\lambda}),$
+where $U{\mathfrak{n}}\otimes R_{\mu}$ is viewed as a $U^{\mathrm{hb},\geq}_{\zeta}-$module via the map $U^{+}_{\zeta}\otimes{\mathfrak{U}}^{0}_{\zeta}\xrightarrow{\mathrm{Fr}\otimes\tau_{\mu}}U{\mathfrak{n}}\otimes{\mathfrak{U}}^{0}_{\zeta}\rightarrow U{\mathfrak{n}}\otimes R$, for any $\mu\in\Lambda$.
+Hence we have isomorphisms
+${\mathrm{I}}^{\mathrm{hb}}_{1}\big(\tau^{\leq\nu}_{c}({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}}\otimes{\mathbb{C}}_{\lambda})\big)$ $={\mathrm{I}}^{\mathrm{hb}}_{1}\big({\mathbb{C}}[{\mathfrak{b}}_{R}]\otimes\big(U{\mathfrak{n}}/\bigoplus_{\eta\nleq\nu-\lambda}(U{\mathfrak{n}})_{\eta}\big)\otimes{\mathbb{C}}_{\lambda}\big)$
+$=U^{\mathrm{hb}}_{\zeta}\otimes_{U^{\mathrm{hb},\geq}_{\zeta}}(U{\mathfrak{n}}/\bigoplus_{\eta\nleq\nu-\lambda}(U{\mathfrak{n}})_{\eta}\big)\otimes R_{-\rho+l\lambda})$
+$=\tau^{\leq-\rho+l\nu}_{r}\big({\mathrm{I}}^{\mathrm{hb}}_{1}({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}}\otimes{\mathbb{C}}_{\lambda})\big).\qed$
+4.4. Functors ${\mathbb{V}}^{b}$
 We identify the rings
 ${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}=S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]=S\otimes_{S^{W}}S.$
-Recall that by §2.3.3 there is an algebra homomorphism
+Recall that ${\mathsf{A}}=\mathrm{End}^{\mathrm{op}}_{\widetilde{{\mathfrak{g}}}}({\mathcal{E}})$ for some tilting bundle $\mathcal{E}$} on $\widetilde{{\mathfrak{g}}}$.
+By [11, Lem 2.5.3], $\mathcal{E}$} has ${\mathcal{O}}_{\widetilde{{\mathfrak{g}}}}$ as a direct summand.
+It follows that $\mathsf{A}$} is an algebra over $\Gamma({\mathcal{O}}_{\widetilde{{\mathfrak{g}}}})={\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$, and that $\mathsf{A}$} has ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ as a direct summand as a ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$-module.
+Moreover, the subalgebra ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ is central in $\mathsf{A}$}.
+Hence ${\mathsf{A}}_{S}^{b}$ contains a central subalgebra ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$.
+Note that $^{\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]^{T}={\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}$.
+Therefore, we have an algebra homomorphism
+(4.12) ${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}\rightarrow Z({\mathsf{A}}_{S}^{b}\text{-}\mathrm{Mod}^{T}).$
+The functor $\pi_{*}:{\mathsf{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}})$ factors through a functor (still denoted by $\pi_{*}$) ${\mathsf{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$.
+By base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$, we obtain a functor
+(4.13) ${\mathbb{V}}^{b}_{c}:\ {\mathsf{A}}_{S}^{b}\text{-}\mathrm{Mod}^{T}\rightarrow\mathrm{QCoh}^{T}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$
+that is represented by the $({\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}],{\mathsf{A}}^{b}_{S})$-bimodule $\Gamma(\widetilde{{\mathfrak{g}}},{\mathcal{E}})\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{b}}_{S}]$.
+On the other hand, by (2.8) there is an algebra homomorphism
 (4.14) ${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}\rightarrow Z({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}).$
@@
 Consider the following composition
-${\mathbb{V}}^{b}_{r}:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}\rightarrow{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}\xrightarrow{T^{-\rho}_{0}}{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
+${\mathbb{V}}^{b}_{r}:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}\rightarrow{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}\xrightarrow{{\mathsf{T}}^{-\rho}_{0}}{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
 By the identification (3.8) and the compatibility of $\mathfrak{F}$$^{\prime}$ and ${\mathfrak{F}}^{b}$, the two maps (4.12) and (4.14) coincide under the equivalence ${\mathfrak{F}}^{b}$.
-Therefore, we obtain the following enhancement of Proposition 4.5.
-Corollary 4.7.
-There is a commutative diagram
-[…]
+Therefore, we obtain the following enhancement of (4.8)
+(4.15) ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$$\mathrm{Coh}^{T}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$$\scriptstyle{{\mathfrak{F}}^{b}}\scriptstyle{\simeq}$$\scriptstyle{{\mathbb{V}}^{b}_{c}}$$\scriptstyle{{\mathbb{V}}^{b}_{r}}$$\scriptstyle{{\mathfrak{G}}^{b}}\scriptstyle{\simeq}$
 4.5. Verma modules
-In this section, we investigate the objects in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ corresponding to the Verma modules under the equivalence ${\mathfrak{F}}^{b}$, and then show that they admit natural $B$-equivariant structures.
+In this section, we investigate the objects in ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ corresponding to the Verma modules under the equivalence ${\mathfrak{F}}^{b}$, and then show that they admit natural $B$-equivariant structures.
 We define an algebra structure on $(U^{\leq}_{\zeta})^{*}$ by the comultiplication of $U^{\leq}_{\zeta}$, and define an action of $U^{\leq}_{\zeta}$ on $(U^{\leq}_{\zeta})^{*}$ by the right multiplication of $U^{\leq}_{\zeta}$ on itself.
@@
 Then ${\mathbb{C}}_{\zeta}[N^{-}]$ is an algebra of $U^{\leq}_{\zeta}$-modules.
-Set $e=B/B\in{\mathcal{B}}$ and let $U_{e}=N^{-}B/B$ be the big open cell. By [44, Proof of Prop 3.12], there is an algebra isomorphism
+Set $e=B/B\in{\mathcal{B}}$ and let $U_{e}=N^{-}B/B$ be the big open cell. By [42, Proof of Prop 3.12], there is an algebra isomorphism
 $\Gamma(U_{e},{\mathcal{O}}_{{\mathcal{B}}_{\zeta}})\simeq{\mathbb{C}}_{\zeta}[N^{-}],$
@@
 $\Gamma(U_{e},{\mathcal{O}}_{\mathcal{B}})={\mathbb{C}}[N^{-}]=\bigoplus_{\eta\leq 0}\big((U{\mathfrak{n}}^{-})_{\eta}\big)^{*}.$
-The counit $\epsilon:{\mathbb{C}}_{\zeta}[N^{-}]\rightarrow{\mathbb{C}}$ by evaluating at $1\in U^{\leq}_{\zeta}$ defines an ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$-module ${\mathcal{O}}_{\epsilon}$ supported on $e$.
-As ${\mathcal{O}}_{{\mathcal{B}}}$-modules, ${\mathcal{O}}_{\epsilon}={\mathcal{O}}_{e}$.
+[…]
 Let $({\mathfrak{U}}_{\zeta}^{0})_{-2\rho}$ be the rank one ${\mathfrak{U}}_{\zeta}^{0}$-module given by the automorphism $\tau_{-2\rho}:{\mathfrak{U}}_{\zeta}^{0}\rightarrow{\mathfrak{U}}_{\zeta}^{0}$, and view it as a ${\mathfrak{U}}^{\geq}_{\zeta}$-module via the projection ${\mathfrak{U}}^{\geq}_{\zeta}\rightarrow{\mathfrak{U}}_{\zeta}^{0}$.
@@
 where ${\mathbb{C}}[T]\rightarrow Z^{\leq}_{\mathrm{Fr}}$ is given by $k_{\mu}\mapsto K^{2l}_{\mu}$, and ${\mathbb{C}}[T]\rightarrow{\mathbb{C}}[T]$ is given by $k_{\mu}\mapsto k^{l}_{\mu}$ (here we use $\{k_{\mu}\}_{\mu\in\Lambda}$ as the canonical basis of ${\mathbb{C}}[T]$).
-By [44, Lem 6.3], the map (4.10) induces an isomorphism of ${\mathfrak{U}}^{-}_{\zeta}\otimes Z^{0}_{\mathrm{Fr}}$-modules
+By [42, Lem 6.3], the map (4.16) induces an isomorphism of ${\mathfrak{U}}^{-}_{\zeta}\otimes Z^{0}_{\mathrm{Fr}}$-modules
 (4.17) ${\mathfrak{U}}^{-}_{\zeta}\otimes Z^{0}_{\mathrm{Fr}}\otimes_{{\mathbb{C}}[T]}{\mathbb{C}}[T]={\mathfrak{U}}^{-}_{\zeta}\otimes_{{\mathbb{C}}[N]}{\mathbb{C}}[B\times_{T}T]{\ \xrightarrow{\sim}\ }\Gamma(U_{e},\widetilde{{\mathcal{D}}}_{\zeta})\otimes_{{\mathbb{C}}_{\zeta}[N^{-}]}{\mathbb{C}}.$
@@
 Fix basis $\{x_{r}\}_{r}$ in $U^{+}_{\zeta}$ and $\{y_{r}\}_{r}$ in ${\mathfrak{U}}^{-}_{\zeta}$ which are dual to each other under the canonical pairing between $U^{+}_{\zeta}$ and ${\mathfrak{U}}^{-}_{\zeta}$.
-By [44, Prop 4.13]  [^2] Because of our different conventions on the flag variety and the comultiplications of the quantum group, the formulas in loc.
+By [42, Prop 4.13]  [^1] Because of our different conventions on the flag variety and the comultiplications of the quantum group, the formulas in loc.
 cit.
@@
 (4.18) $K^{-2}_{\lambda}$ $=\sum_{r}y_{r}^{(2)}\cdot\mathrm{S}^{-1}(y_{r}^{(1)}).\varphi_{r}\otimes k_{\lambda}$
-$\equiv\sum_{r}\langle 1^{*}_{\lambda},\mathrm{S}(x_{r})y_{r}^{(1)}.1_{\lambda}\rangle\cdot y_{r}^{(2)}\cdot k_{\lambda}$
-$=\sum_{r}\langle 1^{*}_{\lambda},\mathrm{S}(x_{r})y_{r}.1_{\lambda}\rangle\cdot k_{\lambda}$
+$\equiv\sum_{r}\langle 1^{*}_{\lambda},\mathrm{S}(x_{r})y_{r}^{(1)}1_{\lambda}\rangle\cdot y_{r}^{(2)}\cdot k_{\lambda}$
+$=\sum_{r}\langle 1^{*}_{\lambda},\mathrm{S}(x_{r})y_{r}1_{\lambda}\rangle\cdot k_{\lambda}$
 $=\zeta^{4(\rho,\lambda)}k_{\lambda}\ ,$
-where the third equality is because that $\Delta(y_{r})=y_{r}\otimes 1+\sum_{\text{deg}(y_{r}^{(2)})<0}y_{r}^{(1)}\otimes y_{r}^{(2)}$, and the last equality is because that the quantum Casimir element $\sum_{r}\mathrm{S}(x_{r})y_{r}$ acts on $1_{\lambda}\in{\mathrm{W}}(\lambda)$ by scalar $\zeta^{4(\rho,\lambda)}$.
+where the third equality is because $\Delta(y_{r})=y_{r}\otimes 1+\sum_{\text{deg}(y_{r}^{(2)})<0}y_{r}^{(1)}\otimes y_{r}^{(2)}$, and the last equality is because the quantum Casimir element $\sum_{r}\mathrm{S}(x_{r})y_{r}$ acts on $1_{\lambda}\in{\mathrm{W}}(\lambda)$ by scalar $\zeta^{4(\rho,\lambda)}$.
 Finally, we identify
@@
 Then we have ${\mathbb{M}}=\Gamma(U_{e},\widetilde{{\mathcal{D}}}_{\zeta})\otimes_{{\mathbb{C}}_{\zeta}[N^{-}]}{\mathbb{C}}$ as ${\mathfrak{U}}^{\leq}_{\zeta}$-modules.
-It is in fact an isomorphism of ${\mathfrak{U}}_{\zeta}$-modules, since these two modules are $\Lambda$-graded.
+It is in fact an isomorphism of modules of ${\mathfrak{U}}_{\zeta}$, since the two modules are $\Lambda$-graded.
 ∎
-Recall the tilting bundle ${\mathcal{E}}$ in $\tilde{{\mathfrak{g}}}$.
-Let ${\mathcal{E}}^{\vee,b}$ be the restriction of the dual bundle ${\mathcal{E}}^{\vee}$ on $\{e\}\times{\mathfrak{b}}$ in $\tilde{{\mathfrak{g}}}$, and set ${\mathcal{E}}^{\vee,b}_{S}={\mathcal{E}}^{\vee,b}\otimes_{{\mathbb{C}}[{\mathfrak{t}}]}S$, which is an ${\mathrm{A}}^{b}_{S}$-module.
+Recall the tilting bundle $\mathcal{E}$} in $\widetilde{{\mathfrak{g}}}$.
+Let ${\mathcal{E}}^{\vee,b}$ be the restriction of the dual bundle $\mathcal{E}$$^{\vee}$ on $\{e\}\times{\mathfrak{b}}$ in $\widetilde{{\mathfrak{g}}}$, and set ${\mathcal{E}}^{\vee,b}_{S}={\mathcal{E}}^{\vee,b}\otimes_{{\mathbb{C}}[{\mathfrak{t}}]}S$, which is an ${\mathsf{A}}^{b}_{S}$-module.
 Proposition 4.9.
@@
 As a sheave on $\mathcal{V}$}, $\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathcal{O}}_{\epsilon}$ is supported on the fiber of $e\in{\mathcal{B}}$ in $\mathcal{V}$}, which is identified with $B\times_{T}T$ in the proof of the lemma above.
-We denote by $e_{n}$ the $n$-th infinitesimal neighborhood of 00 in ${\mathfrak{b}}$.
+We denote by $e_{n}$ the $n$-th infinitesimal neighborhood of 0 in $\mathfrak{b}$}.
 We denote by $e^{\lambda}_{n}$ the $n$-th infinitesimal neighborhood of $(1,\zeta^{2\lambda})$ in $B\times_{T}T$, and view it as a closed subscheme in $\mathcal{V}$}.
@@
 The ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}$-module ${\mathcal{O}}_{\epsilon}$ and the ${\mathcal{O}}_{\mathcal{V}}$-module ${\mathcal{O}}_{e^{\lambda}_{n}}$ define a sheaf of module ${\mathcal{O}}_{(\epsilon,e^{\lambda}_{n})}$ of ${\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}$.
-We abbreviate $\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{\lambda}_{n})}=\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{\lambda}_{n})}$.
+We abbreviate $\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{\lambda}_{n})}=\widetilde{{\mathcal{D}}}_{\zeta}\otimes_{({\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}})}{\mathcal{O}}_{(\epsilon,e^{\lambda}_{n})}$.
 By Lemma 4.8 and formula (4.18), we compute that
@@
 Consider the equivalence
-$R\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{0},-):D^{b}(\widetilde{{\mathcal{D}}}_{\zeta}\text{-}\mathrm{mod}_{(1,1)}){\ \xrightarrow{\sim}\ }D^{b}\mathrm{Coh}_{{\mathcal{B}}}(\tilde{{\mathfrak{g}}}).$
+[…]
 We claim that it sends $\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{0}_{n})}$ to ${\mathcal{O}}_{e_{n}}$. Indeed, we have
 $R\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{0},\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{0}_{n})})$ $={\mathfrak{M}}_{0}^{\vee}\otimes_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{0}_{n})}$
-$={\mathfrak{M}}_{0}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{0}_{n})}.$
+$={\mathfrak{M}}_{0}^{\vee}\otimes_{({\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}})}{\mathcal{O}}_{(\epsilon,e^{0}_{n})}.$
 By construction that ${\mathfrak{M}}_{0}={\mathcal{O}}^{\rho}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}}{\mathfrak{M}}_{{-\rho}}$ (see §3.2.3), we obtain that
-${\mathfrak{M}}_{0}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{0}_{n})}={\mathfrak{M}}_{{-\rho}}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{-\rho}_{n})}$
+${\mathfrak{M}}_{0}^{\vee}\otimes_{({\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}})}{\mathcal{O}}_{(\epsilon,e^{0}_{n})}={\mathfrak{M}}_{{-\rho}}^{\vee}\otimes_{({\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}})}{\mathcal{O}}_{(\epsilon,e^{-\rho}_{n})}$
 as sheaves on $\mathcal{B}$}. Similarly we have
-${\mathfrak{M}}_{{-\rho}}^{\vee}\otimes_{{\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}}}{\mathcal{O}}_{(\epsilon,e^{-\rho}_{n})}$ $=R\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{{-\rho}},\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{-\rho}_{n})})$
+${\mathfrak{M}}_{{-\rho}}^{\vee}\otimes_{({\mathcal{O}}_{{\mathcal{B}}_{\zeta}}\otimes_{{\mathcal{O}}_{\mathcal{B}}}{\mathcal{O}}_{\mathcal{V}})}{\mathcal{O}}_{(\epsilon,e^{-\rho}_{n})}$ $=R\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{{-\rho}},\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{-\rho}_{n})})$
 $=\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{{-\rho}}|_{e^{-\rho}_{n}},\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{-\rho}_{n})}).$
@@
 (4.20) $\mathcal{H}om_{\widetilde{{\mathcal{D}}}_{\zeta}|_{\mathrm{FN}({\mathcal{B}})}}({\mathfrak{M}}_{{-\rho}}|_{e^{-\rho}_{n}},\widetilde{{\mathcal{D}}}_{\zeta}|_{(\epsilon,e^{-\rho}_{n})})=\mathrm{End}_{({\mathfrak{U}}_{\zeta})_{\widehat{1,\chi_{-\rho}}}}\big(M(-\rho)_{S}/{\mathfrak{m}}_{1}^{n}M(-\rho)_{S}\big).$
-By the equivalence (4.6), the RHS of (4.14) is equal to ${\mathcal{O}}_{e_{n}}$ (note that we identify the formal neighborhood of 00 in ${\mathfrak{b}}_{S}$ with the formal neighborhood of $(1,1)$ in $B\times_{T}T$).
-Finally, under the equivalence (3.1), the sheaf ${\mathcal{O}}_{e_{n}}$ corresponds to the ${\mathrm{A}}$-module ${\mathcal{E}}^{\vee}\otimes_{{\mathcal{O}}_{\tilde{{\mathfrak{g}}}}}{\mathcal{O}}_{e_{n}}={\mathcal{E}}^{\vee,b}_{S}/{\mathfrak{m}}^{n}_{0}{\mathcal{E}}^{\vee,b}_{S}$, where ${\mathfrak{m}}_{0}$ is the maximal ideal of $0\in{\mathfrak{b}}$ in ${\mathbb{C}}[{\mathfrak{b}}]$.
+By the equivalence (4.6), the RHS of (4.20) is equal to ${\mathcal{O}}_{e_{n}}$ (note that we identify the formal neighborhood of 0 in ${\mathfrak{b}}_{S}$ with the formal neighborhood of $(1,1)$ in $B\times_{T}T$).
+Finally, under the equivalence (3.1), the sheaf ${\mathcal{O}}_{e_{n}}$ corresponds to the $\mathsf{A}$}-module ${\mathcal{E}}^{\vee}\otimes_{{\mathcal{O}}_{\widetilde{{\mathfrak{g}}}}}{\mathcal{O}}_{e_{n}}={\mathcal{E}}^{\vee,b}_{S}/{\mathfrak{m}}^{n}_{0}{\mathcal{E}}^{\vee,b}_{S}$, where ${\mathfrak{m}}_{0}$ is the maximal ideal of $0\in{\mathfrak{b}}$ in ${\mathbb{C}}[{\mathfrak{b}}]$.
 Hence we have
@@
 Definition 4.10.
-We define the Verma objects in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ to be
+We define the Verma objects in ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ to be
 ${\mathcal{M}}_{x}=({\mathfrak{F}}^{b})^{-1}(M(x\bullet_{l}0)_{S}),\quad x\in W_{\mathrm{ex}}.$
-An object in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{Mod}^{T}$ admits Verma flags if it admits a (decreasing, separated, ${\mathbb{Z}}_{\leq 0}$-) filtration with composition factors given by Verma objects.
+An object in ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{Mod}^{T}$ admits Verma flags if it admits a filtration of finite length with composition factors given by Verma objects.
+For any $w\in W$, we choose a reduced expression $w=s_{1}\cdots s_{n}$ and set ${\mathbb{T}}^{c}_{w}={\mathbb{T}}^{c}_{s_{1}}\cdots{\mathbb{T}}^{c}_{s_{n}}$.
 Corollary 4.11.
-Any Verma object in ${\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ admits a natural $B$-equivariant structure.
+For any $w\in W$ and $\nu\in\Lambda$, we have
+(4.21) ${\mathcal{M}}_{t(\nu)w_{0}w^{-1}}={\mathbb{T}}^{c}_{w}({\mathcal{E}}^{\vee,b}_{S})\otimes{\mathbb{C}}_{\nu},\quad\forall w\in W,\ \forall\nu\in\Lambda.$
+As a consequence, any Verma object in ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$ admits a natural $B$-equivariant structure.
 Proof.
-By Proposition 4.9, ${\mathcal{M}}_{w_{0}}={\mathcal{E}}^{\vee,b}_{S}$ is naturally a $B$-equivariant ${\mathrm{A}}^{b}_{S}$-module. We show that in general
-(4.15) ${\mathcal{M}}_{t(\nu)w_{0}w^{-1}}={\mathbb{T}}^{c}_{w}({\mathcal{M}}_{w_{0}})\otimes{\mathbb{C}}_{\nu},\quad\forall w\in W,\ \forall\nu\in\Lambda,$
-where ${\mathbb{T}}^{c}_{w}={\mathbb{T}}^{c}_{s_{1}}\cdots{\mathbb{T}}^{c}_{s_{n}}$ for a reduced expression $w=s_{1}\cdots s_{n}$.
-Then since ${\mathbb{T}}^{c}_{w}$ is $B$-equivariant, any Verma object is endowed with a $B$-equivariant structure.
-Indeed, for any $w\in W$ and $s\in{\mathrm{I}}$ such that $ws<w$ in the Bruhat order, there is a short exact sequence (see e.g. [23, Thm 7.14(a)])
-[…]
+For any $w\in W$ and $s\in{\mathrm{I}}$ such that $ws<w$ in the Bruhat order, there is a short exact sequence (see e.g. [23, Thm 7.14(a)])
+$0\rightarrow M(ws\bullet 0)_{S}\rightarrow\Theta^{r}_{s}(M(w\bullet 0))_{S}\xrightarrow{\text{counit}}M(w\bullet 0)_{S}\rightarrow 0,$
 which shows that $M(ws\bullet 0)_{S}={\mathbb{T}}^{r}_{s}(M(w\bullet 0)_{S})$. By Proposition 4.1 that the equivalence ${\mathfrak{F}}^{b}$ intertwines the reflection functors on both sides, we have
 (4.22) ${\mathcal{M}}_{ws}={\mathbb{T}}^{c}_{s}({\mathcal{M}}_{w}),\quad\text{if}\quad ws<w.$
-Apply (4.16) several times on ${\mathcal{M}}_{w_{0}}$ and use the $\Lambda$-translations, we obtain (4.15). ∎
-4.4. Fully-faithfulness
+Applying (4.22) several times on ${\mathcal{M}}_{w_{0}}={\mathcal{E}}^{\vee,b}_{S}$ (Proposition 4.9) and using the $\Lambda$-translations, we obtain (4.21).
+Finally, since ${\mathcal{E}}^{\vee,b}_{S}$ is naturally endowed with a $B$-equivariant structure and ${\mathbb{T}}^{c}_{w}$ is $B$-equivariant, any Verma object is $B$-equivariant.
+∎
+4.6. Fully-faithfulness for ${\mathbb{V}}^{b}$
 For any Noetherian commutative $S$-algebra $R$, one can define the functor
 ${\mathbb{V}}^{b}_{r,R}:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{R}\rightarrow{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,-\rho}_{R\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$
-as in §4.2.2. Let ${\mathbb{K}}={\mathrm{Frac}}(S)$ be the fraction field of $S$.
+as in §4.4 (note that for $S=R$ we abbreviate ${\mathbb{V}}^{b}_{r}={\mathbb{V}}^{b}_{r,S}$). Let ${\mathbb{K}}={\mathrm{Frac}}(S)$ be the fraction field of $S$.
 Lemma 4.12.
@@
 (1) By definition of ${\mathbb{V}}^{b}_{r,R}$ we have an inclusion
-$\mathrm{Hom}({\mathbb{V}}^{b}_{r,R}M_{1},{\mathbb{V}}^{b}_{r,R}M_{2})\hookrightarrow\mathrm{Hom}(T^{-\rho}_{0}M_{1},T^{-\rho}_{0}M_{2}),\quad\forall M_{i}.$
-Hence the faithfulness of ${\mathbb{V}}^{b}_{r,R}$ (on objects admitting Verma flags) follows from the faithfulness of $T^{-\rho}_{0}$, see [36, Lem 4.9(2)].
+$\mathrm{Hom}({\mathbb{V}}^{b}_{r,R}M_{1},{\mathbb{V}}^{b}_{r,R}M_{2})\hookrightarrow\mathrm{Hom}({\mathsf{T}}^{-\rho}_{0}M_{1},{\mathsf{T}}^{-\rho}_{0}M_{2}),\quad\forall M_{i}.$
+Hence the faithfulness of ${\mathbb{V}}^{b}_{r,R}$ (on objects admitting Verma flags) follows from the faithfulness of ${\mathsf{T}}^{-\rho}_{0}$, see [34, Lem 4.9(2)].
 (2) There is an isomorphism
@@
 Under the decompositions (4.24) and (4.25), we can write
-${\mathbb{V}}^{b}_{r,{\mathbb{K}}}=\bigoplus_{w\in W}T^{-\rho}_{0}\circ\iota_{w},$
+${\mathbb{V}}^{b}_{r,{\mathbb{K}}}=\bigoplus_{w\in W}{\mathsf{T}}^{-\rho}_{0}\circ\iota_{w},$
 where $\iota_{w}$ is the inclusion from the block labelled by $w$ in (4.25).
-The functor $T^{-\rho}_{0}\circ\iota_{w}$ is an equivalence with quasi-inverse given by $\mathrm{pr}_{w}\circ T^{0}_{-\rho}$, where $\mathrm{pr}_{w}$ is the corresponding block projection.
+The functor ${\mathsf{T}}^{-\rho}_{0}\circ\iota_{w}$ is an equivalence with quasi-inverse given by $\mathrm{pr}_{w}\circ{\mathsf{T}}^{0}_{-\rho}$, where $\mathrm{pr}_{w}$ is the corresponding block projection.
 Hence ${\mathbb{V}}^{b}_{r,{\mathbb{K}}}$ is an equivalence of categories.
@@
 Proof.
-By Corollary 4.7, we only have to prove the assertion for ${\mathbb{V}}^{b}_{r}$.
+By (4.15) we only have to prove the assertion for ${\mathbb{V}}^{b}_{r}$.
 It follows standardly from Lemma 4.12 by the general theory of highest weight categories, see e.g.
-[35, Prop 2.18].
+[33, Prop 2.18].
 For readers’ convenience, we present the proof as follows.
@@
 Hence the base change yields an embedding $\mathrm{Hom}(M_{1},M_{2})\subseteq\mathrm{Hom}(M_{1,{\mathbb{K}}},M_{2,{\mathbb{K}}})$.
-By [32, Thm 38], we have $M_{i}=\bigcap\limits_{\mathrm{ht}(\mathfrak{p})=1}M_{i,S_{\mathfrak{p}}}$ as subspaces in $M_{i,{\mathbb{K}}}$ (the intersection is taken for height 1 prime ideals $\mathfrak{p}$ of $S$).
+By [30, Thm 38], we have $M_{i}=\bigcap\limits_{\mathrm{ht}(\mathfrak{p})=1}M_{i,S_{\mathfrak{p}}}$ as subspaces in $M_{i,{\mathbb{K}}}$ (the intersection is taken for height 1 prime ideals $\mathfrak{p}$ of $S$).
 Hence we can identify
@@
 Choose a projective resolution $P_{1}\rightarrow P_{2}\rightarrow M\rightarrow 0$, we have a commutative diagram with exact rows
-00{0}${\mathrm{Hom}(M,M^{\prime})_{R^{\prime}}}$${\mathrm{Hom}(P_{2},M^{\prime})_{R^{\prime}}}$${\mathrm{Hom}(P_{1},M^{\prime})_{R^{\prime}}}$00{0}${\mathrm{Hom}(M_{R^{\prime}},M^{\prime}_{R^{\prime}})}$${\mathrm{Hom}(P_{2,R^{\prime}},M^{\prime}_{R^{\prime}})}$${\mathrm{Hom}(P_{1,R^{\prime}},M^{\prime}_{R^{\prime}}),}$
-where the subscript $R^{\prime}$ means $-\otimes_{R}R^{\prime}$. By [20, Prop 2.4], the right two vertical maps are isomorphisms, so is the left one. ∎
+[…]
+where the subscript $R^{\prime}$ means $\otimes_{R}R^{\prime}$. By [20, Prop 2.4], the right two vertical maps are isomorphisms, so is the left one. ∎
 Consider the commutative diagram
-[…]
-where the vertical isomorphisms are by Lemma 4.14 applied on the tuple $(M,M^{\prime},R,R^{\prime})$ $=(M_{1,S_{\mathfrak{p}}},M_{2,S_{\mathfrak{p}}},S_{\mathfrak{p}},{\mathbb{K}})$ or $({\mathbb{V}}M_{1,S_{\mathfrak{p}}}$, ${\mathbb{V}}M_{2,S_{\mathfrak{p}}}$, $S_{\mathfrak{p}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$, ${\mathbb{K}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}])$, and the lower horizontal isomorphism is by Lemma 4.12(2).
+$\mathrm{Hom}(M_{1,S_{\mathfrak{p}}},M_{2,S_{\mathfrak{p}}})\otimes_{S_{\mathfrak{p}}}{\mathbb{K}}$${\mathrm{Hom}({\mathbb{V}}M_{1,S_{\mathfrak{p}}},{\mathbb{V}}M_{2,S_{\mathfrak{p}}})}\otimes_{S_{\mathfrak{p}}}{\mathbb{K}}$$\mathrm{Hom}(M_{1,{\mathbb{K}}},M_{2,{\mathbb{K}}})$$\mathrm{Hom}({\mathbb{V}}M_{1,{\mathbb{K}}},{\mathbb{V}}M_{2,{\mathbb{K}}}).$$\scriptstyle{\varphi\otimes_{S_{\mathfrak{p}}}{\mathbb{K}}}\scriptstyle{\simeq}\scriptstyle{\simeq}\scriptstyle{\simeq}$
+where the vertical isomorphisms are by Lemma 4.14 applied on the tuple $(M,M^{\prime},R,R^{\prime})$ $(M_{1,S_{\mathfrak{p}}},M_{2,S_{\mathfrak{p}}},S_{\mathfrak{p}},{\mathbb{K}})$ or $({\mathbb{V}}M_{1,S_{\mathfrak{p}}}$, ${\mathbb{V}}M_{2,S_{\mathfrak{p}}}$, $S_{\mathfrak{p}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$, ${\mathbb{K}}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}])$, and the lower horizontal isomorphism is by Lemma 4.12(2).
 Hence $\varphi\otimes_{S_{\mathfrak{p}}}{\mathbb{K}}$ is an isomorphism.
@@
 Remark 4.15.
-For the non-deformed category ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}$, the functor ${\mathbb{V}}^{b}_{r,{\mathbb{C}}}$ fails to be full-faithful on objects admitting Verma flags.
+For the non-deformed category ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{\mathbb{C}}$, the functor ${\mathbb{V}}^{b}_{r,{\mathbb{C}}}$ fails to be fully-faithful on objects admitting Verma flags.
 5. Main result: equivalence $\mathfrak{F}$}
-5.1. Functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$
-Recall that in §2.3.3, we construct an algebra homomorphism
+5.1. Functors $\mathbb{V}$}
+We construct the functors ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$, and discuss their compatibility with reflection functors associated to $\mathrm{I}$.
+5.1.1. Construction of $\mathbb{V}$}
+Recall that in (2.8) we construct an algebra homomorphism
 ${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}\rightarrow Z(U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}).$
-As discussions in §4.2.2, the translation functor $T^{-\rho}_{0}:U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,-\rho}_{S}$ can be enhanced to a functor
+By similar discussions as in §4.4, the translation functor ${\mathsf{T}}^{-\rho}_{0}:U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,-\rho}_{S}$ can be enhanced to a functor
 ${\mathbb{V}}_{r}:\ U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
@@
 ${\mathbb{V}}_{r}:{\mathcal{O}}_{S}^{0}\rightarrow{\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
-In fact, ${\mathbb{V}}_{r}$ is equivalent to the quotient functor $\bigoplus\limits_{\lambda\in\Lambda}\mathrm{Hom}(Q(w_{0}\bullet 0+l\lambda)_{S},-)$ (we will not use this below).
-Recall that the functor $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$ is represented by the $G$-equivariant $\big{(}{\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}],{\mathrm{A}}\big{)}$-bimodule $\Gamma(\tilde{{\mathfrak{g}}},{\mathcal{E}})$.
-We can define a functor
-${\mathbb{V}}_{c}:({\mathrm{A}}_{S}^{b}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow({\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$
-as ${\mathbb{V}}^{b}_{c}$ in §4.2.2. It restricts to a functor
-${\mathbb{V}}_{c}:{\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{B}\rightarrow\mathrm{Coh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}).$
+Remark 5.1.
+In fact, ${\mathbb{V}}_{r}$ is equivalent to the quotient functor $\bigoplus\limits_{\lambda\in\Lambda}\mathrm{Hom}(Q(w_{0}\bullet 0+l\lambda)_{S},-)$. We will not use this fact below.
+Recall the functor ${\mathbb{V}}^{b}_{c}$ defined in (4.13). By $U{\mathfrak{n}}$-equivariance it upgrades to a functor
+${\mathbb{V}}_{c}:\ ({\mathsf{A}}_{S}^{b}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow({\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T},$
+which restricts to a functor
+${\mathbb{V}}_{c}:\ {\mathsf{A}}_{S}^{b}\text{-}\mathrm{Mod}^{B}\rightarrow\mathrm{QCoh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}).$
+It is automatic that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are compatible with $\Lambda$-translations.
+5.1.2. Compatibility with reflection functors
+We show that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are compatible with reflection functors associated with $s\in{\mathrm{I}}$. The results of this subsection will not be used in the sequel.
+Let $\omega\in\Xi$. Consider the composition of functors
+${\mathcal{O}}^{\omega}_{S}\xrightarrow{{\mathsf{T}}_{\omega}^{0}}{\mathcal{O}}^{0}_{S}\rightarrow{\mathcal{O}}^{0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}\xrightarrow{{\mathsf{T}}^{\omega}_{0}}{\mathcal{O}}^{\omega}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]},$
+where the middle one is induced by the homomorphism $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]\rightarrow Z({\mathcal{O}}^{0}_{S})$.
+Lemma 5.2 ([35, Prop B.9]).
+There is a natural isomorphism
+${\mathsf{T}}_{0}^{\omega}\circ{\mathsf{T}}_{\omega}^{0}\simeq-\otimes_{S^{W_{\omega}}}S$
+of functors from ${\mathcal{O}}^{\omega}_{S}$ to ${\mathcal{O}}^{\omega}_{S\otimes_{S^{W}}S}$.
 For any $s\in{\mathrm{I}}$, one can define the Soergel bimodule
 $B_{s}=\big({\mathbb{C}}[{\mathfrak{t}}]\otimes_{{\mathbb{C}}[{\mathfrak{t}}]^{s}}{\mathbb{C}}[{\mathfrak{t}}]\big)_{\hat{0}}$
-in $\mathrm{Coh}\big{(}({\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})_{\hat{0}}\big{)}$, where $\hat{0}$ represents the completion at $0\in{\mathfrak{t}}/W$.
-Tensoring with $B_{s}$ defines a functor on $\mathrm{Coh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$ and ${\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$.
-It is automatic that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are compatible with $\Lambda$-translations.
-We now show that ${\mathbb{V}}_{r}$ and ${\mathbb{V}}_{c}$ are also compatible with reflection functors associated with $s\in{\mathrm{I}}$.
-5.1.1. Compatibility with reflection functors
-Let $\omega\in\Xi$. Consider the composition of functors
-${\mathcal{O}}^{\omega}_{S}\xrightarrow{T_{\omega}^{0}}{\mathcal{O}}^{0}_{S}\rightarrow{\mathcal{O}}^{0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}\xrightarrow{T^{\omega}_{0}}{\mathcal{O}}^{\omega}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]},$
-where the middle one is induced by the homomorphism $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]\rightarrow Z({\mathcal{O}}^{0}_{S})$.
-Lemma 5.1 ([37, Prop B.9]).
-There is a natural isomorphism
-$T_{0}^{\omega}\circ T_{\omega}^{0}\simeq-\otimes_{S^{W_{\omega}}}S$
-of functors from ${\mathcal{O}}^{\omega}_{S}$ to ${\mathcal{O}}^{\omega}_{S\otimes_{S^{W}}S}$.
+[…]
 Proposition 5.3.
@@
 From the lemma above we deduce that
-$T^{-\rho}_{0}\circ\Theta^{r}_{s}$ $=T^{-\rho}_{s}\circ T^{s}_{0}\circ T^{0}_{s}\circ T_{0}^{s}$
-$=T^{-\rho}_{s}\big{(}T_{0}^{s}(-)\otimes_{S^{s}}S\big{)}$
+${\mathsf{T}}^{-\rho}_{0}\circ\Theta^{r}_{s}$ $={\mathsf{T}}^{-\rho}_{s}\circ{\mathsf{T}}^{s}_{0}\circ{\mathsf{T}}^{0}_{s}\circ{\mathsf{T}}_{0}^{s}$
+$={\mathsf{T}}^{-\rho}_{s}\big({\mathsf{T}}_{0}^{s}(-)\otimes_{S^{s}}S\big)$
 $=\Theta^{r}_{s}(-)\otimes_{S^{s}}S,$
-where for the first and the last equality we use a natural isomorphism $T_{0}^{-\rho}=T_{s}^{-\rho}\circ T_{0}^{s}$. ∎
+where for the first and the last equality we use a natural isomorphism ${\mathsf{T}}_{0}^{-\rho}={\mathsf{T}}_{s}^{-\rho}\circ{\mathsf{T}}_{0}^{s}$. ∎
 Proposition 5.4.
@@
 Let $P$ be any (standard) parabolic subgroup of $G$, and let $W_{P}$ be the corresponding parabolic subgroup in $W$.
-Denote by $\pi^{P}:\tilde{{\mathfrak{g}}}_{P}\rightarrow{\mathfrak{g}}$ the natural projection.
+Denote by $\pi^{P}:\widetilde{{\mathfrak{g}}}_{P}\rightarrow{\mathfrak{g}}$ the natural projection.
 There is a ($G$-equivariant) natural isomorphism
-$\pi_{P,*}\pi_{P}^{*}\simeq-\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W_{P}]}{\mathbb{C}}[{\mathfrak{t}}]\qquad\text{on}\quad D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}}_{P}).$
+$\pi_{P,*}\pi_{P}^{*}\simeq-\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W_{P}]}{\mathbb{C}}[{\mathfrak{t}}]\qquad\text{on}\quad D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathfrak{g}}}_{P}).$
 Hence we have the following ($G$-equivariant) natural isomorphisms
@@
 $=\pi_{*}(-)\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W_{P}]}{\mathbb{C}}[{\mathfrak{t}}]$
-of functors from $D^{b}\mathrm{Coh}(\tilde{{\mathfrak{g}}})$ to $D^{b}\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$.
-Consider the special case when $P=P_{s}$, and then it implies our desired isomorphism by base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$.
-The second assertion is easy to verify.
-∎
+[…]
 Denote by $^{w}S=S$ the $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-module given by
@@
 5.2. Truncated categories
-In this section, we define truncated subcategories in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ which will be shown to correspond to (certain) truncations in ${\mathcal{O}}^{0}_{S}$ defined in §2.3.1 under the equivalence (1.4), and we will study the projective objects in them.
-We firstly define the truncation of ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ by $\nu\in\Lambda$ as
-${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\leq\nu}:=({\mathfrak{F}}^{b})^{-1}({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0,\leq l\nu}_{S}).$
-It is a Serre subcategory of ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$, and each object in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admits a finite decomposition with each direct factor contained in the truncated subcategory by some $\nu$.
+In this subsection, we review the truncated subcategories of ${\mathcal{O}}^{0}_{S}$ studied in [34, §4.2], which is a refinement of the truncations discussed in §2.3.1.
+Next, we define the corresponding truncated subcategories in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$.
+The advantage of the new truncations is their compatibility with the functors $\mathbb{V}$}, see Lemmas 5.10 and 5.19.
+5.2.1. Truncation in ${\mathcal{O}}^{0}_{S}$
+Definition 5.6.
+The semi-infinite order $\leq^{\frac{\infty}{2}}$ on $W_{\mathrm{ex}}$ is the partial order generated by
+$x\leq^{\frac{\infty}{2}}tx\quad\text{if}\quad x\bullet_{l}0\leq tx\bullet_{l}0,$
+for $x\in W_{\mathrm{ex}}$ and reflection $t\in\bigcup_{x\in W_{\mathrm{af}}}x{\mathrm{I}}_{\mathrm{af}}x^{-1}$ in $W_{\mathrm{af}}$.
+Remark 5.7.
+  1. (1)
+    The order $\leq^{\frac{\infty}{2}}$ is invariant under $\Lambda$-translation from the left.
+Under the decomposition $W_{\mathrm{ex}}=W_{\mathrm{af}}\rtimes\Lambda/{\mathrm{Q}}$, two elements $(x,\gamma)$ and $(x^{\prime},\gamma^{\prime})$ (with $x,x^{\prime}\in W_{\mathrm{af}}$, $\gamma,\gamma^{\prime}\in\Lambda/{\mathrm{Q}}$) are comparable only if $\gamma=\gamma^{\prime}$.
+And we have $(x,\gamma)\leq^{\frac{\infty}{2}}(x^{\prime},\gamma)$ if and only if $x\leq^{\frac{\infty}{2}}x^{\prime}$.
+  2. (2)
+    The order $\leq^{\frac{\infty}{2}}$ coincides with the order $\prec$” defined in [37].
+More precisely, let $\mathrm{Alc}$ be the set of alcoves associated to the $W_{\mathrm{af}}$-action on ${\mathbb{R}}\otimes_{\mathbb{Z}}\Lambda$, equipped with an order $\prec$ (see [37, §4] for details).
+There is a fundamental alcove $A_{\mathrm{fun}}\in\mathrm{Alc}$ given by
+    $A_{\mathrm{fun}}=\{\lambda\in{\mathbb{R}}\otimes_{\mathbb{Z}}\Lambda\ |\ 0<\langle\lambda,\check{\alpha}\rangle<1,\ \forall\alpha\in\Phi^{+}\}.$
+    We have a bijection $W_{\mathrm{af}}{\ \xrightarrow{\sim}\ }\mathrm{Alc}$, $x\mapsto xA_{\mathrm{fun}}$.
+Then the order $\leq^{\frac{\infty}{2}}$ on $W_{\mathrm{af}}$ corresponds to $\prec$ on $\mathrm{Alc}$.
+In particular, the order $\leq^{\frac{\infty}{2}}$ is independent of $l$.
+Under the identification $W_{\mathrm{ex}}\simeq W_{\mathrm{ex}}\bullet_{l}0$, the order $\leq^{\frac{\infty}{2}}$ on $W_{\mathrm{ex}}$ coincides with the order $\uparrow$ introduced in [34, §2.4.2] restricted on $W_{\mathrm{ex}}\bullet_{l}0$.
+Lemma 5.8 ([34, Lem 4.6(1)]).
+Let $\nu,\mu\in\Lambda$ and $w\in W$.
+We have $t(\mu)w\leq^{\frac{\infty}{2}}t(\nu)$ if and only if $\mu\leq\nu$.
+Namely, we have $\{x\in W_{\mathrm{ex}}|x\leq^{\frac{\infty}{2}}t(\nu)\}=\{t(\lambda)w\}_{\lambda\leq\nu,w\in W}$.
+Let $\nu\in\Lambda$. In [34, §4.2] we introduce a truncated subcategory of ${\mathcal{O}}^{0}_{S}$ (denoted by ${\mathcal{O}}_{S}^{\uparrow l\nu}$ in the loc. cit.)
+${\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S}$,
+which is the the full subcategory of modules $M$ in ${\mathcal{O}}^{0}_{S}$ that admit a surjection $Q\twoheadrightarrow M$ from a module $Q$ admitting a Verma flag with factors $M(x\bullet_{l}0)_{S}$ with $x\leq^{\frac{\infty}{2}}t(\nu)$.
+Since $W_{\mathrm{ex}}$ is covered by the poset ideals of the form $\{x\in W_{\mathrm{ex}}|x\leq^{\frac{\infty}{2}}t(\nu)\}$, any module in ${\mathcal{O}}^{0}_{S}$ is a finite direct sum of modules in ${\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S}$ for some $\nu$.
+Recall the induction functor
+$\mathrm{ind}_{r}=U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}-:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}.$
+Proposition 5.9 ([34]).
+  1. (1)
+    There is a truncation functor
+    $\tau_{r}^{\leq^{\frac{\infty}{2}}t(\nu)}:\ U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{S}\rightarrow{\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S}$
+    by taking the maximal quotient in ${\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S}$, which is left adjoint to the natural inclusion.
+  2. (2)
+    The family $\{\tau_{r}^{\leq^{\frac{\infty}{2}}t(\nu)}(\mathrm{ind}_{r}(P))\}$ with $P$ running over a generating family of projective objects in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ forms a generating family of projective objects in ${\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S}$.
+Moreover, any projective object in ${\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S}$ admits a Verma flag, with factors of the form $M(x\bullet_{l}0)_{S}$ with $x\leq^{\frac{\infty}{2}}t(\nu)$}.
+  3. (3)
+    The category ${\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S}$ is a Serre subcategory in ${\mathcal{O}}_{S}$.
+Proof.
+Part (1) is [34, Lem 4.4], where the truncation functor is denoted by $\tau^{\uparrow l\nu}$. Part (2) and (3) are by [34, Lem 4.5(2)&(3)]. ∎
+Lemma 5.10.
+The functor ${\mathbb{V}}_{r}$ restricts to
+${\mathbb{V}}_{r}:\ {\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S}\rightarrow{\mathcal{O}}^{-\rho,\leq-\rho+l\nu}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
+There is a natural isomorphism of functors from $U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ to ${\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$,
+(5.2) ${\mathbb{V}}_{r}\circ\tau_{r}^{\leq^{\frac{\infty}{2}}t(\nu)}=\tau_{r}^{\leq-\rho+l\nu}\circ{\mathbb{V}}_{r}$
+Proof.
+It is an immediate enhancement of [34, Lem 4.7]. ∎
+5.2.2. Truncation by a subset
+We now consider truncations for the category ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$. We start by studying truncations by a subset $\Omega\subset\Lambda$, as §2.3.1 for the representation side.
+We extends the equivalence ${\mathfrak{F}}^{b}$ to the equivalence of ind-completions of both sides
+${\mathfrak{F}}^{b}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T}\ \simeq\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}.$
+Define the truncation of ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T}$ by a subset $\Omega\subset\Lambda$ as
+${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}:=({\mathfrak{F}}^{b})^{-1}({\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Omega,0}_{S}),$
+which is a Serre subcategory of ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T}$.
+We denote the truncation functor by $\Omega$ on ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T}$ corresponding to the one on ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}$ (introduced in §2.3.1) by
+$\tau^{T,\Omega}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T}\rightarrow{\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}.$
+We abbreviate ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\Omega}:={\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}\cap{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$.
+The category ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ is the union of ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\Omega}$, with $\Omega$ running over all bounded above poset ideals in $\Lambda$.
 We define
-${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu},\quad\nu\in\Lambda$
-to be the full subcategory of modules in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ which are contained in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\leq\nu}$, forgetting the $B$-equivariant structure.
-Then ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ is a Serre subcategory in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$, and each object in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admits a finite decomposition with each direct factor contained in the truncated subcategory by some $\nu$.
-Let $\leq^{\frac{\infty}{2}}$ be the semi-infinite order on $W_{\mathrm{af}}$, which is a partial order generated by
-$xs\leq^{\frac{\infty}{2}}x\quad\text{if}\quad x\bullet_{l}0\geq xs\bullet_{l}0,\quad x\in W_{\mathrm{af}},\ s\in{\mathrm{I}}_{\mathrm{af}}.$
-Note that $\leq^{\frac{\infty}{2}}$ is invariant under $\Lambda$-translation from the left.
-It extends naively to a partial order $\leq^{\frac{\infty}{2}}$ on $W_{\mathrm{ex}}=W_{\mathrm{af}}\rtimes\Lambda/{\mathrm{Q}}$ such that two elements are comparable only if they are contained in the same $\Lambda/{\mathrm{Q}}$-coset.
-Note that $\leq^{\frac{\infty}{2}}$ is independent of $l$.
+$({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T,\Omega}$
+to be the full subcategory of modules in $({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$ which are contained in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}$.
+It is a Serre subcategory in $({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$.
+We abbreviate
+${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B,\Omega}:={\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B}\cap({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T,\Omega}.$
+We define $({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T,\Omega}$ and ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\Omega}$ similarly.
+The category ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ is the union of ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\Omega}$, with $\Omega$ running over all bounded above poset ideals in $\Lambda$.
+Lemma 5.11.
+If $\Omega\subset\Lambda$ is a bounded above poset ideal, then objects in $({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T,\Omega}$ (resp.
+$({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T,\Omega}$) are contained in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B,\Omega}$ (resp.
+${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\Omega}$).
+Proof.
+Since $\Omega$ is bounded above, there is a finite subset $\Omega^{\prime}\subset\Lambda$ such that $\Omega\subset-2\rho+l\Omega^{\prime}-\sum_{s\in{\mathrm{I}}}{\mathbb{N}}\alpha_{s}$.
+For any $\lambda\in\Lambda$ and $w\in W$, if $t(\lambda)w\bullet_{l}0=l\lambda+w\bullet 0\in\Omega$, then $\lambda\in\Omega^{\prime}-\sum_{s\in{\mathrm{I}}}{\mathbb{N}}\alpha_{s}$.
+Any module in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Omega,0}_{S}$ is a quotient by a module composed by $M(x\bullet_{l}0)_{S}$ for some $x\in W_{\mathrm{ex}}$ such that $x\bullet_{l}0\in\Omega$.
+Hence any module in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}$ is a quotient of a module composed by ${\mathcal{M}}_{w}\otimes{\mathbb{C}}_{\lambda}$ with $w\in W$ and $\lambda\in\Omega^{\prime}-\sum_{s\in{\mathrm{I}}}{\mathbb{N}}\alpha_{s}$.
+Note that ${\mathcal{M}}_{w}$ is a finitely generated ${\mathsf{A}}^{b}_{S}$-module and ${\mathsf{A}}^{b}_{S}$ is finite over ${\mathbb{C}}[{\mathfrak{b}}_{S}]$, thus ${\mathcal{M}}_{w}$ is finitely generated over ${\mathbb{C}}[{\mathfrak{b}}_{S}]$, and in particular its $T$-weights are contained in $\Omega(w)-\sum_{s\in{\mathrm{I}}}{\mathbb{N}}\alpha_{s}$, for a finite subset $\Omega(w)\subset\Lambda$.
+Hence the $T$-weights of any module in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}$ are contained in $\Omega^{\prime}+\bigcup_{w\in W}\Omega(w)-\sum_{s\in{\mathrm{I}}}{\mathbb{N}}\alpha_{s}$, which is a bounded above subset in $\Lambda$.
+In particular, the $U{\mathfrak{n}}$-action on any module in $({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T,\Omega}$ is locally unipotent, which induces a $B$-action.
+∎
+Lemma 5.12.
+Let $\Omega$ be a subset in $\Lambda$. There is a truncation functor
+$\tau^{U{\mathfrak{n}},\Omega}:\ ({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T,\Omega}$
+by taking the maximal quotient in $({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T,\Omega}$, which is left adjoint to the natural inclusion.
+Proof.
+For any $M\in({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$, we define $\tau^{U{\mathfrak{n}},\Omega}(M)$ as the quotient of $M$ by the ${\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}}$-submodule generated by $\ker(M\rightarrow\tau^{T,\Omega}(M))$.
+Then $\tau^{U{\mathfrak{n}},\Omega}(M)\in({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T,\Omega}$ since it is a quotient of $\tau^{T,\Omega}(M)$ as a $T$-equivariant ${\mathsf{A}}^{b}_{S}$-module.
+Any morphism from $M$ to a module in $({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T,\Omega}$ necessarily factors through $\tau^{U{\mathfrak{n}},\Omega}(M)$.
+Hence $\tau^{U{\mathfrak{n}},\Omega}(M)$ is the maximal quotient of $M$ in $({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T,\Omega}$.
+It gives a functor that is left adjoint to the inclusion.
+∎
+The functor $\tau^{U{\mathfrak{n}},\Omega}$ restricts to a functor
+$\tau^{B,\Omega}:{\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B}\rightarrow{\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B,\Omega}.$
+Consider the functor $\mathrm{coind}_{T}^{B}:\mathrm{Rep}(T)\rightarrow\mathrm{Rep}(B)$ that is right adjoint to the forgetful functor $\mathrm{Rep}(B)\rightarrow\mathrm{Rep}(T)$.
+Since $B/T$ is affine, the functor $\mathrm{coind}_{T}^{B}$ is exact.
+For any $M\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T}$, the action map ${\mathsf{A}}^{b}_{S}\otimes M\rightarrow M$ induces a homomorphism
+${\mathsf{A}}^{b}_{S}\otimes\mathrm{coind}_{T}^{B}(M)=\mathrm{coind}_{T}^{B}({\mathsf{A}}^{b}_{S}\otimes M)\rightarrow\mathrm{coind}_{T}^{B}(M),$
+where the first equality is by tensor identity (since ${\mathsf{A}}^{b}_{S}$ is a $B$-module).
+It yields a $B$-equivariant ${\mathsf{A}}^{b}_{S}$-module structure on $\mathrm{coind}_{T}^{B}(M)$, which defines a functor
+$\mathrm{coind}_{T}^{B}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T}\rightarrow{\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B},$
+which is exact and is right adjoint to the forgetful functor.
+Lemma 5.13.
+Let $\Omega$ be a poset ideal in $\Lambda$.
+  1. (1)
+    The functor $\mathrm{coind}_{T}^{B}$ sends ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}$ to ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B,\Omega}$.
+  2. (2)
+    There is a commutative diagram
+    (5.3) ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B}$${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B,\Omega}$${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T}$${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}.$$\scriptstyle{\tau^{B,\Omega}}$$\scriptstyle{\mathrm{for}}$$\scriptstyle{\mathrm{for}}$$\scriptstyle{\tau^{T,\Omega}}$
+Proof.
+(1) Since any module in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}$ is a quotient by a module that is composed by ${\mathcal{M}}_{x}$ with $x\in W_{\mathrm{ex}}$ such that $x\bullet_{l}0\in\Omega$, by exactness it is enough to consider $\mathrm{coind}_{T}^{B}({\mathcal{M}}_{x})$.
+Since ${\mathcal{M}}_{x}$ is $B$-equivariant by Corollary 4.11, we have an isomorphism $\mathrm{coind}_{T}^{B}({\mathcal{M}}_{x})={\mathcal{M}}_{x}\otimes{\mathbb{C}}[B/T]$, where $B$ acts on ${\mathcal{M}}_{x}\otimes{\mathbb{C}}[B/T]$ diagonally.
+Hence $\mathrm{coind}_{T}^{B}({\mathcal{M}}_{x})$ is composed by ${\mathcal{M}}_{x}\otimes{\mathbb{C}}_{\eta}$ with $\eta\leq 0$, which shows that $\mathrm{coind}_{T}^{B}({\mathcal{M}}_{x})\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B,\Omega}$.
+(2) By (1) we have a commutative diagram
+${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B}$${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B,\Omega}$${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T}$${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega},$$\scriptstyle{\mathrm{coind}_{T}^{B}}$$\scriptstyle{\mathrm{coind}_{T}^{B}}$
+from which one obtains the diagram (5.3) by taking the left adjoint functors. ∎
+Lemma 5.14.
+For any object in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admitting a Verma flag in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$, it also admits a Verma flag in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$.
+Proof.
+Let $P\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admitting a Verma flag in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$.
+By (2.4) and (4.5), there is a short exact sequence $0\rightarrow M\rightarrow P\rightarrow P^{\prime}\rightarrow 0$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$, where $M\simeq{\mathcal{M}}_{x}^{\oplus n}$ ($n\geq 1$) as $T$-equivariant ${\mathsf{A}}^{b}_{S}$-modules and $P^{\prime}$ is composed by ${\mathcal{M}}_{y}$ for some $y\in W_{\mathrm{ex}}$ such that $y\bullet_{l}0\ngeq x\bullet_{l}0$.
+By adjunction we have
+$\mathrm{Hom}_{{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}}({\mathcal{M}}_{x}^{\oplus n},P)=\mathrm{Hom}_{{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}}({\mathcal{M}}_{x}^{\oplus n},P\otimes{\mathbb{C}}[B/T]),$
+where we use the fact that $\mathrm{coind}_{T}^{B}(P)=P\otimes{\mathbb{C}}[B/T]$ since $P$ is $B$-equivariant. There is a short exact sequence in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$,
+$0\rightarrow P\rightarrow P\otimes{\mathbb{C}}[B/T]\rightarrow P\otimes{\mathbb{C}}[B/T]_{<0}\rightarrow 0,$
+where ${\mathbb{C}}[B/T]_{<0}$ is the quotient module of ${\mathbb{C}}[B/T]$ by the zero weight component.
+Since $P\otimes{\mathbb{C}}[B/T]_{<0}$ is composed by ${\mathcal{M}}_{y}\otimes{\mathbb{C}}_{\eta}$ with $\eta<0$ and $y\bullet_{l}0\ngtr x\bullet_{l}0$, by (2.3) and (4.5) any $T$-equivariant (hence any $B$-equivariant) ${\mathsf{A}}^{b}_{S}$-homomorphism from ${\mathcal{M}}_{x}^{\oplus n}$ to $P\otimes{\mathbb{C}}[B/T]$ factors through $M\subset P\subset P\otimes{\mathbb{C}}[B/T]$.
+Hence any $T$-equivariant ${\mathsf{A}}^{b}_{S}$-homomorphism
+${\mathcal{M}}_{x}^{\oplus n}{\ \xrightarrow{\sim}\ }M\hookrightarrow P$
+is $B$-equivariant.
+It shows that $M$ is a $B$-equivariant submodule of $P$ isomorphic to ${\mathcal{M}}_{x}^{\oplus n}$.
+By induction on the length of Verma flag, $P^{\prime}$ admits a Verma flag in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$.
+It completes the proof.
+∎
+Consider the induction functor (which should not be confused by the induction $\mathrm{ind}_{c}$ for ${\mathbb{C}}[{\mathfrak{b}}]_{S}\text{-}\mathrm{Mod}^{T}$ defined earlier)
+$\mathrm{ind}_{c}:=({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\otimes_{{\mathsf{A}}^{b}_{S}}-:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T}\rightarrow({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T},$
+which is left adjoint to the forgetful functor. Note that $\mathrm{ind}_{c}(M)=U{\mathfrak{n}}\otimes M$ as vector spaces, for any ${\mathsf{A}}^{b}_{S}$-module $M$.
+Lemma 5.15.
+Let $\Omega$ be a bounded above poset ideal in $\Lambda$.
+The family $\{\tau^{U{\mathfrak{n}},\Omega}(\mathrm{ind}_{c}(P))\}$ with $P$ running over a generating family of projective objects in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ forms a generating family of projective objects in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\Omega}$.
+Moreover, any projective object in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\Omega}$ admits a Verma flag.
+Proof.
+The module $\tau^{U{\mathfrak{n}},\Omega}(\mathrm{ind}_{c}(P))$ is contained in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\Omega}$ by Lemma 5.11.
+Since $\tau^{U{\mathfrak{n}},\Omega}\circ\mathrm{ind}_{c}$ is left adjoint to the exact forgetful functor from ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\Omega}$ to ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$, it sends projective objects to projective objects.
+For any $M\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\Omega}$, we choose a surjection $P\twoheadrightarrow M$ from a projective module $P$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$.
+It induces surjections $\mathrm{ind}_{c}(P)\rightarrow\tau^{U{\mathfrak{n}},\Omega}(\mathrm{ind}_{c}(P))\rightarrow M$.
+Hence we obtain a generating family of projective objects in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\Omega}$.
+For the second assertion, by Lemma 5.13(1) the forgetful functor ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B,\Omega}\rightarrow{\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}$ admits an exact right adjunction.
+Hence a projective object in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\Omega}$ is projective in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\Omega}$.
+Since (by the corresponding statement for ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Omega,0}_{S}$) any projective object in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\Omega}$ admits a Verma flag, our assertion follows from Lemma 5.14.
+∎
+Till the end of the subsection, we will only consider the truncation by $\Omega=\{\lambda\in\Lambda|\lambda\leq l\nu\}$, and we will abbreviate $\Omega$” by $\leq l\nu$” in the superscripts above.
+5.2.3. Truncation by $\leq^{\frac{\infty}{2}}$
+We now consider the truncation of ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ by semi-infinite order $\leq^{\frac{\infty}{2}}$.
 We have the following proposition, whose proof is postponed to §5.2.4.
 Proposition 5.16.
-We have $\mathrm{Ext}^{i}_{{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}}({\mathcal{M}}_{x},{\mathcal{M}}_{x^{\prime}})\neq 0$ only if $x\leq^{\frac{\infty}{2}}x^{\prime}$.
-Consider the induction functor
-$\mathrm{ind}_{c}:=({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}^{b}_{S}}-:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}\rightarrow({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}.$
-Note that $\mathrm{ind}_{c}(M)=U{\mathfrak{n}}\otimes M$ as vector spaces, for any ${\mathrm{A}}^{b}_{S}$-module $M$.
+We have
+(5.4) $\mathrm{Ext}^{i}_{{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}}({\mathcal{M}}_{x},{\mathcal{M}}_{x^{\prime}})\neq 0\quad\text{only if}\quad x\leq^{\frac{\infty}{2}}x^{\prime}.$
+For a module $Q\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admitting a Verma flag, by (5.4) we can form the quotient
+$\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q)$
+of $Q$ by the submodule that is composed by the Verma factors ${\mathcal{M}}_{x}$ with $x\nleq^{\frac{\infty}{2}}t(\nu)$. We define
+${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$
+as the the full subcategory of modules in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ that admit a surjection $Q\twoheadrightarrow M$ from a module $Q$ admitting a Verma flag with factors ${\mathcal{M}}_{x}$ with $x\leq^{\frac{\infty}{2}}t(\nu)$.
+Since $W_{\mathrm{ex}}$ is covered by the poset ideals of the form $\{x\in W_{\mathrm{ex}}|x\leq^{\frac{\infty}{2}}t(\nu)\}$, any object in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ is a finite direct sum of objects in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ for some $\nu\in\Lambda$.
 Lemma 5.17.
-For any objects $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admitting Verma flags, the induction $\mathrm{ind}_{c}(P)$ also admits Verma flags.
+There is a truncation functor
+$\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}:\ ({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}\rightarrow{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$
+that is left adjoint to the natural inclusion.
 Proof.
-It is enough to prove the assertion when $P$ is a Verma object.
-By Corollary 4.11, $P$ is a module of ${\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}}$.
-Consider the homomorphism of $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-modules
-(5.2) $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}^{b}_{S}}P=U{\mathfrak{n}}\otimes P\rightarrow P\otimes U{\mathfrak{n}},$
-$u\otimes m\mapsto u^{(1)}.m\otimes u^{(2)},\quad\forall u\in U{\mathfrak{n}},\ \forall m\in P,$
-where $P\otimes U{\mathfrak{n}}$ is viewed as an $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-module by ${\mathrm{A}}^{b}_{S}$ acting on $P$ and $U{\mathfrak{n}}$ acting on $P\otimes U{\mathfrak{n}}$ diagonally.
-The homomorphism (5.2) is an isomorphism, whose inverse is given by $m\otimes u\mapsto u^{(2)}\otimes(\mathrm{S}u^{(1)}).m$ ($\mathrm{S}$ is the antipode for $U{\mathfrak{n}}$).
-We obtain an isomorphism $\mathrm{ind}_{c}(P)\simeq P\otimes U{\mathfrak{n}}$ of $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-modules.
-It is clear that $P\otimes U{\mathfrak{n}}$ admits a Verma flag with Verma factors $P\otimes{\mathbb{C}}_{\eta}$ for $\eta\geq 0$.
-∎
-Note the subset $\{t(\lambda)w\}_{w\in W,\lambda\leq\nu}$ in an ideal in $(W_{\mathrm{ex}},\leq^{\frac{\infty}{2}})$.
-By the lemma above and Proposition 5.5, for any object $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admitting Verma flags, one can define the quotients
-$\mathrm{ind}_{c}(P)^{\leq\nu},\quad\nu\in\Lambda$
-of $\mathrm{ind}_{c}(P)$ modulo the submodule with Verma factors in $\{{\mathcal{M}}_{t(\lambda)w}\}_{w\in W,\ \lambda\nleq\nu}$.
-Then the $({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})$-modules $\mathrm{ind}_{c}(P)^{\leq\nu}$ are contained in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$.
-Lemma 5.7.
-The family $\{\mathrm{ind}_{c}(P)^{\leq\nu}\}$ with $P$ running over a generating family of projective objects in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ forms a generating family of projective objects in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$.
-Proof.
-Let $x=t(\lambda)w$ where $\lambda\in\Lambda$ and $w\in W$, then the truncation $M(x\bullet_{l}0)^{\leq l\nu}_{S}\neq 0$ only if $\lambda\leq\nu$.
-Hence the functor $\mathrm{Hom}({\mathcal{M}}_{x},-)$ is nonzero on ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T,\leq\nu}$ only if $\lambda\leq\nu$.
-It shows that any homomorphism from $\mathrm{ind}_{c}(P)$ to an object in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ factors through the quotient $\mathrm{ind}_{c}(P)^{\leq\nu}$.
-It follows natural isomorphisms
-(5.3) $\mathrm{Hom}_{{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}}(\mathrm{ind}_{c}(P)^{\leq\nu},-)$ $=\mathrm{Hom}_{({\mathrm{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}}(\mathrm{ind}_{c}(P),-)$
-$=\mathrm{Hom}_{{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}}(P,-)$
-of functors on ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$.
-Hence $\mathrm{ind}_{c}(P)^{\leq\nu}$ is projective in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$.
-On the other hand, for any module $M$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$, there exists a surjection of $\Lambda$-graded ${\mathrm{A}}^{b}_{S}$-modules $P\twoheadrightarrow M$ for some projective $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$.
-Then the induced homomorphism $\mathrm{ind}_{c}(P)\twoheadrightarrow\mathrm{ind}_{c}(P)^{\leq\nu}\rightarrow M$ is a surjection.
-So we obtain a generating family of projective objects in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$.
+Since ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is contained in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$, any morphism from $M\in({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}$ to an object in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ factors through $\tau^{B,\leq l\nu}M$.
+Hence it is enough to define the functor
+(5.5) $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}\rightarrow{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}.$
+Let $Q^{\prime}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admitting a Verma flag with factors in $\{M(x\bullet_{l}0)_{S}\}_{x\leq^{\frac{\infty}{2}}t(\nu)}$, and let $Q^{\prime}\twoheadrightarrow M^{\prime}$ be a surjection.
+Let $Q\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$ be a projective object.
+Then any morphism from $Q$ to $M^{\prime}$ can be lifted to $Q^{\prime}$.
+By (5.4) any morphism from $Q$ to $Q^{\prime}$ factors through $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q)$.
+Hence any morphism from $Q$ to $M^{\prime}$ factors through $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q)$, which then is the maximal quotient of $Q$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$.
+In general, let $M\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$ and choose a projective resolution $Q_{2}\rightarrow Q_{1}\rightarrow M\rightarrow 0$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$.
+Then we set
+$\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(M):=\mathrm{coker}\big(\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q_{2})\rightarrow\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q_{1})\big).$
+Then $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(M)$ is contained in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$.
+For any $M^{\prime}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$, we have a commutative diagram with exact rows
+[…]
+Hence the left vertical map is an isomorphism, which shows that $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(M)$ is the maximal quotient of $M$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$.
+It gives the desired functor.
 ∎
 Lemma 5.18.
-Any $M\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ admits a maximal quotient $M^{\leq\nu}$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$. It defines a functor
-$(-)^{\leq\nu}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}\rightarrow{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$
-left adjoint to the inclusion functor. We have a natural isomorphism
-${\mathbb{V}}_{c}\circ(-)^{\leq\nu}=(-)^{\leq\nu}\circ{\mathbb{V}}_{c}.$
+  1. (1)
+    The family $\{\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(\mathrm{ind}_{c}(P))\}$ with $P$ running over a generating family of projective objects in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ forms a generating family of projective objects in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$.
+Moreover, any projective object in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ admits a Verma flag (with factors in $\{{\mathcal{M}}_{x}\}_{x\leq^{\frac{\infty}{2}}t(\nu)}$).
+  2. (2)
+    The category ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is a Serre subcategory of ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$.
 Proof.
-We may assume $M\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu^{\prime}}$ for some $\nu^{\prime}>>\nu$.
-By Lemma 5.7, we can choose a presentation $\mathrm{ind}_{c}(P_{1})^{\leq\nu^{\prime}}\rightarrow\mathrm{ind}_{c}(P_{2})^{\leq\nu^{\prime}}\rightarrow M\rightarrow 0$, for some projective object $P_{i}$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$.
-It induces a homomorphism $\mathrm{ind}_{c}(P_{1})^{\leq\nu}\rightarrow\mathrm{ind}_{c}(P_{2})^{\leq\nu}$, and we set $M^{\leq\nu}$ to be its cokernel.
-Then $M^{\leq\nu}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$.
-For any $M^{\prime}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$, we have a commutative diagram with exact rows
-00{0}${\mathrm{Hom}(M,M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{2})^{\leq\nu^{\prime}},M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{1})^{\leq\nu^{\prime}},M^{\prime})}$00{0}${\mathrm{Hom}(M^{\leq\nu},M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{2})^{\leq\nu},M^{\prime})}$${\mathrm{Hom}(\mathrm{ind}_{c}(P_{1})^{\leq\nu},M^{\prime}),}$
-where the two vertical equalities follow from the first equality in (5.3).
-Hence the left vertical map is an isomorphism.
-It shows that $M^{\leq\nu}$ is the maximal quotient of $M$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$, and therefore $(-)^{\leq\nu}$ defines a functor left adjoint to the inclusion.
-For the last assertion, by the right exactness of ${\mathbb{V}}_{c}\circ(-)^{\leq\nu}$ and $(-)^{\leq\nu}\circ{\mathbb{V}}_{c}$, it is enough to construct natural isomorphism ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P)^{\leq\nu})={\mathbb{V}}_{c}(\mathrm{ind}_{c}(P))^{\leq\nu}$ for any projective module $P$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$.
-By Lemma 5.4, the module ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P))$ admits a filtration with composition factors given by ${\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w}S\otimes{\mathbb{C}}_{\lambda}$ for some $\lambda\in\Lambda$ and $w\in W$.
-By weight consideration, we have
-$\mathrm{Ext}^{1}({\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w}S\otimes{\mathbb{C}}_{\lambda},{\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w^{\prime}}S\otimes{\mathbb{C}}_{\mu})\neq 0\quad\text{only if}\quad\lambda<\mu,\quad\forall w,w^{\prime}\in W.$
-Note that ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P)^{\leq\nu})$ is the quotient by the submodule extended by the factors ${\mathcal{O}}_{{\mathfrak{b}}_{S}}\otimes_{S}{}^{w}S\otimes{\mathbb{C}}_{\lambda}$ with $\lambda\nleq\nu$, which is exactly the truncation ${\mathbb{V}}_{c}(\mathrm{ind}_{c}(P))^{\leq\nu}$.
+(1) The proof of the first assertion is similar to the one of Lemma 5.15.
+For the second assertion, for any projective object $P\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$, by construction $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(\mathrm{ind}_{c}(P))=\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}\circ\tau^{B,\leq l\nu}(\mathrm{ind}_{c}(P))$ admits a Verma flag (with factors in $\{{\mathcal{M}}_{x}\}_{x\leq^{\frac{\infty}{2}}t(\nu)}$).
+Hence any projective object $Q\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is a direct summand of an object admitting a Verma flag.
+Under the equivalence (4.2), the image of $Q$ is a direct summand of a module admitting a Verma flag in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$, which thus admits a Verma flag by a standard argument (see e.g.
+[20, Lem 2.5]), and so does $Q$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$.
+Now the assertion follows from Lemma 5.14.
+(2) By definition ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is closed under taking quotient objects.
+Let $M$ be a subobject of $M^{\prime}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$.
+Then the inclusion $M\hookrightarrow M^{\prime}$ factors through the quotient $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(M)$, hence $M=\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(M)$.
+So ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is also closed under taking subobjects.
+Now we show that ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ is closed under extension.
+Let $0\rightarrow M_{1}\rightarrow M\rightarrow M_{2}\rightarrow 0$ be a short exact sequence in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$ with $M_{1},M_{2}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$.
+Then $M\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$, and we can choose a surjection $Q\twoheadrightarrow M$ from a projective object $Q\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq l\nu}$.
+We have short exact sequence $0\rightarrow Q^{\prime}\rightarrow Q\rightarrow\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q)\rightarrow 0$, where $Q^{\prime}$ is the subobject of $Q$ composed by Verma factors ${\mathcal{M}}_{x}$ with $x\nleq^{\frac{\infty}{2}}t(\nu)$}.
+Then $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q^{\prime})=0$, so $\mathrm{Hom}(Q^{\prime},M_{i})=0$ ($i=1,2$).
+It follows that $\mathrm{Hom}(Q^{\prime},M)=0$.
+Hence the surjection $Q\twoheadrightarrow M$ factors through $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(Q)\twoheadrightarrow M$, which implies that $M\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$.
 ∎
-We have a parallel result of Lemma 5.8.
+Recall the truncation functor by $\nu\in\Lambda$ (for any commutative Noetherian $S$-algebra $R$)
+$\tau^{\leq\nu}_{c}:\ ({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow{\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{Mod}^{B,\leq\nu}.$
 Lemma 5.19.
-There is a natural isomorphism
-${\mathbb{V}}_{r}\circ(-)^{\leq l\nu}=(-)^{\leq l\nu}\circ{\mathbb{V}}_{r}$
-of functors from ${\mathcal{O}}^{0}_{S}$ to ${\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$.
+The functor ${\mathbb{V}}_{c}$ restricts to
+${\mathbb{V}}_{c}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}\rightarrow{\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\text{-}\mathrm{mod}^{B,\leq\nu}.$
+We have a natural isomorphism of functors from $({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}$ to ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\text{-}\mathrm{mod}^{B,\leq\nu}$,
+(5.6) ${\mathbb{V}}_{c}\circ\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}=\tau_{c}^{\leq\nu}\circ{\mathbb{V}}_{c}.$
 Proof.
-By [36, Lem 4.3(3)] we have a natural isomorphism $T^{-\rho}_{0}\circ(-)^{\leq l\nu}=(-)^{\leq l\nu}\circ T^{-\rho}_{0}$, which upgrades to the desired natural isomorphism. ∎
+Recall the functor $\pi_{*}:{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}\rightarrow{\mathbb{C}}[{\mathfrak{b}}_{S}]\text{-}\mathrm{mod}^{T}$ and its right adjunction $\pi^{!}$.
+They induces functors (still denoted by $\pi_{*}$ and $\pi^{!}$) for $U{\mathfrak{n}}$-equivariant or $B$-equivariant modules.
+We show the first assertion.
+By the exactness of ${\mathbb{V}}_{r}$, it only needs to show that ${\mathbb{V}}_{r}{\mathcal{M}}_{x}$ is contained in ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\text{-}\mathrm{mod}^{B,\leq\nu}$, for $x\leq^{\frac{\infty}{2}}t(\nu)$.
+By forgetting $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-action to $S$-action and forgetting the $B$-equivariant structure, it is equivalent to show that $\pi_{*}{\mathcal{M}}_{x}$ is contained in ${\mathbb{C}}[{\mathfrak{b}}_{S}]\text{-}\mathrm{mod}^{T,\leq\nu}$.
+Indeed, by [35, Lem 3.8(2)], for any $w\in W$ and $\lambda\in\Lambda$, in ${\mathcal{O}}_{S}$ we have ${\mathsf{T}}^{-\rho}_{0}M(w\bullet 0+l\lambda)_{S}=M(-\rho+l\lambda)_{S}$.
+By (4.8) we deduce that $\pi_{*}{\mathcal{M}}_{t(\lambda)w}={\mathbb{C}}[{\mathfrak{b}}_{S}]\otimes{\mathbb{C}}_{\lambda}$ (as $T$-equivariant ${\mathbb{C}}[{\mathfrak{b}}_{S}]$-modules), which proves the claim.
+We then obtain a natural transformation $\tau_{c}^{\leq\nu}\circ{\mathbb{V}}_{c}\rightarrow{\mathbb{V}}_{c}\circ\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}$.
+To show that it is an isomorphism, by forgetting $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-action to $S$-action, it is enough to show the natural isomorphism
+(5.7) $\tau_{c}^{\leq\nu}\circ\pi_{*}\rightarrow\pi_{*}\circ\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}$
+By [35, Lem 3.8(2)] again, the module ${\mathsf{T}}_{-\rho}^{0}M(-\rho+l\lambda)_{S}$ is composed by $M(w\bullet 0+l\lambda)_{S}$ with $w\in W$.
+By (4.9) and Lemma 5.14, the object $\pi^{!}{\mathbb{C}}[{\mathfrak{b}}_{S}]\otimes{\mathbb{C}}_{\lambda}$ is composed by ${\mathcal{M}}_{t(\lambda)w}$ with $w\in W$, as $T$-equivariant ${\mathsf{A}}^{b}_{S}$-module.
+Any object in ${\mathbb{C}}[{\mathfrak{b}}_{S}]\text{-}\mathrm{mod}^{B,\leq\nu}$ is a quotient of an object composed by ${\mathbb{C}}[{\mathfrak{b}}_{S}]\otimes{\mathbb{C}}_{\lambda}$ with $\lambda\leq\nu$.
+By the exactness of $\pi^{!}$ (it is exact as a functor for $T$-equivariant modules, by (4.9) again), we deduce that $\pi^{!}$ sends ${\mathbb{C}}[{\mathfrak{b}}_{S}]\text{-}\mathrm{mod}^{B,\leq\nu}$ to ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$.
+Hence we have a commutative diagram
+$({\mathsf{A}}^{b}_{S}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}$${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$$({\mathbb{C}}[{\mathfrak{b}}_{S}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}$${\mathbb{C}}[{\mathfrak{b}}_{S}]\text{-}\mathrm{mod}^{B,\leq\nu}.$$\scriptstyle{\pi^{!}}$$\scriptstyle{\pi^{!}}$
+Now (5.7) follows by taking the left adjunctions. ∎
 5.2.4. Proof of Proposition 5.16
 We translate our problem to a problem of computing the extensions of some constructible sheaves on the affine flag variety, using Bezrukavnikov’s equivalence [14].
-Let ${\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ be the completion of ${\mathrm{A}}$ at $0\in{\mathfrak{t}}/W$.
-Since the restriction of $\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ on $\{e\}\times{\mathfrak{b}}\hookrightarrow\tilde{{\mathfrak{g}}}$ is ${\mathrm{A}}^{b}_{S}$, we can identify
-(5.4) ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}=\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}})$
+Let $^{\mathsf{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ be the completion of $\mathsf{A}$} at $0\in{\mathfrak{t}}/W$.
+Since the restriction of $\pi^{*}{\mathsf{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}$ on $\{e\}\times{\mathfrak{b}}\hookrightarrow\widetilde{{\mathfrak{g}}}$ is ${\mathsf{A}}^{b}_{S}$, we can identify
+[…]
 by induction from $B$ to $G$.
-The ${\mathrm{A}}^{b}_{S}$-module ${\mathcal{M}}_{w_{0}}={\mathcal{E}}^{\vee,b}_{S}$ corresponds to ${\mathcal{E}}^{\vee}_{\hat{0}}:={\mathcal{E}}^{\vee}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}/W]_{\hat{0}}$ under (5.4).
-Let $\mathrm{St}=\tilde{{\mathfrak{g}}}\times_{{\mathfrak{g}}}\tilde{{\mathfrak{g}}}$ be the Steinberg variety, and let $\widehat{\mathrm{St}}$ be its completion at $0\in{\mathfrak{t}}/W$.
-Let $\pi_{2}:\widehat{\mathrm{St}}\rightarrow\tilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}$ be the projection to the second factor.
+The ${\mathsf{A}}^{b}_{S}$-module ${\mathcal{M}}_{w_{0}}={\mathcal{E}}^{\vee,b}_{S}$ corresponds to $^{\mathcal{E}}^{\vee}_{\hat{0}}:={\mathcal{E}}^{\vee}\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}/W]_{\hat{0}}$ under (5.8).
+Let $\mathrm{St}=\widetilde{{\mathfrak{g}}}\times_{{\mathfrak{g}}}\widetilde{{\mathfrak{g}}}$ be the Steinberg variety, and let $\widehat{\mathrm{St}}$ be its completion at $0\in{\mathfrak{t}}/W$.
+Let $\pi_{2}:\widehat{\mathrm{St}}\rightarrow\widetilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}$ be the projection to the second factor.
 There is an equivalence
-$\pi_{2}^{*}{\mathcal{E}}_{{\hat{0}}}\otimes^{L}_{\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}}-:\ D^{b}\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}){\ \xrightarrow{\sim}\ }D^{b}\mathrm{Coh}^{G}(\widehat{\mathrm{St}}),$
-sending ${\mathcal{E}}^{\vee}_{\hat{0}}$ to the structure sheaf ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{\Delta}}$ of the diagonal $\tilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}$ in $\widehat{\mathrm{St}}$.
+[…]
+sending $^{\mathcal{E}}^{\vee}_{\hat{0}}$ to the structure sheaf ${\mathcal{O}}_{\widetilde{{\mathfrak{g}}}_{\Delta}}$ of the diagonal $\widetilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}$ in $\widehat{\mathrm{St}}$.
 Let $\check{G}$ be the Langlands dual group of $G$.
@@
 By [14], there is an equivalence of triangulated monoidal categories
-(5.5) $D^{b}\mathrm{Coh}^{G}(\widehat{\mathrm{St}})\simeq\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}}),$
-where $\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}})$ is a completion of the category of $\check{I}_{0}$-equivariant monodromic complexes on $\widetilde{\mathcal{F}{l}}$.
-There are standard and costandard sheaves $j_{x!}$, $j_{x*}$ in $\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}})$ associated to the embedding $j_{x}$ from the Schubert cell labelled by $x\in W_{\mathrm{ex}}$ to $\widetilde{\mathcal{F}{l}}$.
-The equivalence (5.5) sends ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{\Delta}}$ to $j_{e!}$, and in general ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}_{\Delta}}(w_{0}\lambda)$ to the Wakimoto sheaf ${\mathcal{J}}_{\lambda}$ (see [8, §3.3]), for any $\lambda\in\Lambda$ (In loc.
-cit.
-the line bundle labelled by $\lambda$ is sent to ${\mathcal{J}}_{\lambda}$, and here our correspondence is due to the different convention that ${\mathcal{O}}_{{\mathcal{B}}}(\lambda)$ is semi-ample if and only if $\lambda$ is anti-dominant).
-We view $\widehat{\mathrm{St}}$ as the base change of $\pi:\tilde{{\mathfrak{g}}}\rightarrow{\mathfrak{g}}$ (as the second factor) by $\tilde{{\mathfrak{g}}}\times_{{\mathfrak{t}}/W}({\mathfrak{t}}/W)_{\hat{0}}\rightarrow{\mathfrak{g}}$ (as the first factor), then by Theorem 3.1 there are functors ${\mathbb{T}}^{c}_{s}$, $s\in{\mathrm{I}}$ on $D^{b}\mathrm{Coh}^{G}(\widehat{\mathrm{St}})$.
-By [8] the equivalence (5.5) intertwines the functor ${\mathbb{T}}^{c}_{s}$ with the convolution with $j_{s!}$ from the right.
-Therefore, by (4.15) the composed equivalence
-$D^{b}({\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B})\simeq\widehat{D}_{\check{I}_{0}}(\widetilde{\mathcal{F}{l}})$
+[…]
+[…]
+[…]
+[…]
 sends
@@
 $=\mathrm{Ext}^{i}({\mathcal{J}}_{\lambda+\mu}*j_{w!},{\mathcal{J}}_{\lambda^{\prime}+\mu}*j_{w^{\prime}!}),\quad\forall\mu\in\Lambda.$
-Let $\mu$ be dominant enough, then we have ${\mathcal{J}}_{\lambda+\mu}=j_{t(\lambda+\mu)*}$ (see [8, §3.3]).
+Let $\mu$ be dominant enough, then we have ${\mathcal{J}}_{\lambda+\mu}=j_{t(\lambda+\mu)*}$ (see [14, §3.3]).
 Since $t(\lambda)$ is maximal in the Bruhat order on $t(\lambda)W$ if $\lambda$ is dominant, we have ${\mathcal{J}}_{\lambda+\mu}*j_{w!}=j_{t(\lambda+\mu)*}*j_{w!}=j_{t(\lambda+\mu)w*}$.
@@
 Since $t(\lambda^{\prime}+\mu)w^{\prime}$ and $t(\lambda+\mu)w$ translate the fundamental alcove to the dominant alcoves, the Bruhat order on them coincides with the semi-infinite order (see e.g.
-[39, Claim 4.14]).
+[37, Claim 4.14]).
 We deduce that
@@
 5.3. Proof of the main result
-Recall the induction functor
-$\mathrm{ind}_{r}=U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}-:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}.$
+In this section, we prove our main result.
 Lemma 5.20.
-There are natural isomorphisms
-${\mathbb{V}}_{r}\circ\mathrm{ind}_{r}\simeq\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}\ ,\quad\text{and}\quad{\mathbb{V}}_{c}\circ\mathrm{ind}_{c}\simeq\mathrm{ind}_{c}\circ{\mathbb{V}}^{b}_{c}.$
+There is a natural isomorphism
+(5.10) ${\mathbb{V}}_{r}\circ\mathrm{ind}_{r}\simeq\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}$
+of functor from ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$ to $U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$, and a natural isomorphism
+(5.11) ${\mathbb{V}}_{c}\circ\mathrm{ind}_{c}\simeq\mathrm{ind}_{c}\circ{\mathbb{V}}^{b}_{c}$
+of functors from ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ to $({\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T}$.
 Proof.
@@
 Since the action of $Z_{\mathrm{HC}}$ on $U^{\mathrm{hb}}_{\zeta}\otimes_{{\mathfrak{U}}^{b}_{\zeta}}M$ coincides with the one induced by the $Z_{\mathrm{HC}}$-action on $M$, the equality ${\mathbb{V}}_{r}\circ\mathrm{ind}_{r}(M)=\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}(M)$ is also an isomorphism of $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-modules.
-It implies the first natural isomorphism.
-The second isomorphism is straight forward.
-∎
+It implies (5.10).
+The isomorphism (5.11) is straight forward to check. ∎
 Proposition 5.21.
@@
 Proof.
-We firstly prove the assertion for ${\mathbb{V}}_{c}$ in a special case: let $M\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ admitting Verma flags, and let $Q=\mathrm{ind}_{c}(M^{\prime})^{\leq\nu}$ with $M^{\prime}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admitting Verma flags.
+We firstly prove the assertion for ${\mathbb{V}}_{c}$ in a special case: let $M\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ admitting Verma flags, and let $Q=\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(\mathrm{ind}_{c}(M^{\prime}))$ with $M^{\prime}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$ admitting Verma flags.
 We show that the natural map
@@
 is an isomorphism. To that end, consider the following commutative diagram
+$\mathrm{Hom}(Q,M)$$\mathrm{Hom}(\mathrm{ind}_{c}(M^{\prime}),M)$$\mathrm{Hom}(M^{\prime},M)$$\mathrm{Hom}({\mathbb{V}}_{c}Q,{\mathbb{V}}_{c}M)$$\mathrm{Hom}({\mathbb{V}}_{c}(\mathrm{ind}_{c}(M^{\prime})),{\mathbb{V}}_{c}M)$$\mathrm{Hom}({\mathbb{V}}_{c}^{b}M^{\prime},{\mathbb{V}}_{c}^{b}M),$$\scriptstyle{{\mathbb{V}}_{c}}\scriptstyle{\simeq}$$\scriptstyle{{\mathbb{V}}_{c}}\scriptstyle{\simeq}$$\scriptstyle{{\mathbb{V}}_{c}^{b}}\scriptstyle{\simeq}\scriptstyle{\simeq}$
+where the horizontal isomorphisms follow from the adjunctions for $\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}$, $\tau_{c}^{\leq\nu}$ and $\mathrm{ind}_{c}$, and the natural isomorphisms (5.6) and (5.11).
+It is now equivalent to show that the right vertical map is an isomorphism, which is true by Proposition 4.13.
+In general, let $M_{1},M_{2}\in{\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}$ admitting Verma flags.
+By Lemma 5.18, there is a resolution $Q_{2}\rightarrow Q_{1}\rightarrow M_{1}\rightarrow 0$ with $Q_{i}=\tau_{c}^{\leq^{\frac{\infty}{2}}t(\nu)}(\mathrm{ind}_{c}(P_{i}))$ for some projective objects $P_{i}$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$.
+We have a commutative diagram with exact rows
 […]
-where the horizontal isomorphisms follow from the adjunctions for $(-)^{\leq\nu}$ and $\mathrm{ind}_{c}$, and the equalities
-${\mathbb{V}}_{c}(\mathrm{ind}_{c}(M^{\prime})^{\leq\nu})={\mathbb{V}}_{c}(\mathrm{ind}_{c}(M^{\prime}))^{\leq\nu}\quad\text{and}\quad{\mathbb{V}}_{c}(\mathrm{ind}_{c}(M^{\prime}))=\mathrm{ind}_{c}({\mathbb{V}}^{b}_{c}(M^{\prime}))$
-by Lemma 5.8 and Lemma 5.10. It is now equivalent to show that the right vertical map is an isomorphism, which is true by Proposition 4.13.
-In general, let $M_{1},M_{2}\in{\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}$ admitting Verma flags.
-By Lemma 5.7, there is a resolution $Q_{1}\rightarrow Q_{2}\rightarrow M_{1}\rightarrow 0$ with $Q_{i}=\mathrm{ind}_{c}(P_{i})^{\leq\nu}$ for some projective objects $P_{i}$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$.
-We have a commutative diagram with exact rows
-00{0}${\mathrm{Hom}(M_{1},M_{2})}$${\mathrm{Hom}(Q_{2},M_{2})}$${\mathrm{Hom}(Q_{1},M_{2})}$00{0}${\mathrm{Hom}({\mathbb{V}}_{c}M_{1},{\mathbb{V}}_{c}M_{2})}$${\mathrm{Hom}({\mathbb{V}}_{c}Q_{2},{\mathbb{V}}_{c}M_{2})}$${\mathrm{Hom}({\mathbb{V}}_{c}Q_{1},{\mathbb{V}}_{c}M_{2}),}$
 where the vertical isomorphisms are by our previous result in the special case. It shows that ${\mathbb{V}}_{c}$ is fully-faithful on $M_{1}$ and $M_{2}$.
@@
 Recall the following result
-Theorem 5.12 ([36, Thm 3.6]).
+Theorem 5.22 ([34, Thm 3.6]).
 The functor
@@
 sending ${\mathbb{C}}[{\mathfrak{b}}_{S}]\otimes{\mathbb{C}}_{\lambda}$ to $M(-\rho+l\lambda)_{S}$, for any $\lambda\in\Lambda$.
-For any commutative Noetherian $S$-algebra $R$, the equivalence $\mathfrak{G}$ induces an equivalence of the categories on both sides after base changed to $R$. In particular, we have
+By (2.5), for any commutative Noetherian $S$-algebra $R$, the equivalence $\mathfrak{G}$ induces an equivalence of the categories on both sides after base changed to $R$. In particular, we have
 $\mathfrak{G}:\ \mathrm{Coh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}){\ \xrightarrow{\sim}\ }{\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
@@
 There is an equivalence of $S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]$-linear abelian categories
-${\mathfrak{F}}:\ \mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}){\ \xrightarrow{\sim}\ }{\mathcal{O}}_{S}^{0},$
+${\mathfrak{F}}:\ \mathrm{Coh}^{G}(\pi^{*}{\mathsf{A}}_{({\mathfrak{t}}/W)_{\hat{0}}}){\ \xrightarrow{\sim}\ }{\mathcal{O}}_{S}^{0},$
 such that
-${\mathfrak{F}}\big{(}{\mathbb{T}}^{c}_{w}({\mathcal{E}}^{\vee}_{\hat{0}})\otimes{\mathcal{O}}_{\tilde{{\mathfrak{g}}}}(\nu)\big{)}=M(w_{0}w^{-1}\bullet 0+l\nu)_{S},\quad\forall w\in W,\ \forall\nu\in\Lambda.$
+${\mathfrak{F}}\big({\mathbb{T}}^{c}_{w}({\mathcal{E}}^{\vee}_{\hat{0}})\otimes{\mathcal{O}}_{\widetilde{{\mathfrak{g}}}}(\nu)\big)=M(w_{0}w^{-1}\bullet 0+l\nu)_{S},\quad\forall w\in W,\ \forall\nu\in\Lambda.$
 Proof.
-Let $P$ be a projective module in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$, and set ${\mathcal{P}}=({\mathfrak{F}}^{b})^{-1}(P)$ in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$.
+Let $P$ be a projective module in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$, and set ${\mathcal{P}}=({\mathfrak{F}}^{b})^{-1}(P)$ in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$.
 Set $Q=\mathrm{ind}_{r}(P)$ and ${\mathcal{Q}}=\mathrm{ind}_{c}({\mathcal{P}})$.
-We have the following isomorphisms by Lemma 5.9 and Lemma 5.10,
-${\mathbb{V}}_{r}(Q^{\leq l\nu})=({\mathbb{V}}_{r}Q)^{\leq l\nu}=({\mathbb{V}}_{r}\circ\mathrm{ind}_{r}P)^{\leq l\nu}=(\mathrm{ind}_{r}\circ{\mathbb{V}}^{b}_{r}P)^{\leq l\nu},$
-and similarly ${\mathbb{V}}_{c}({\mathcal{Q}}^{\leq\nu})=(\mathrm{ind}_{c}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq\nu}$. By Proposition 4.6 and Corollary 4.7, we have
-$\mathfrak{G}(\mathrm{ind}_{c}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq\nu}=(\mathfrak{G}\circ\mathrm{ind}_{c}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq l\nu}=(\mathrm{ind}_{r}\circ\mathfrak{G}^{b}\circ{\mathbb{V}}_{c}^{b}{\mathcal{P}})^{\leq l\nu}=(\mathrm{ind}_{r}\circ{\mathbb{V}}_{r}^{b}P)^{\leq l\nu}.$
-It follows that $\mathfrak{G}\big{(}{\mathbb{V}}_{c}({\mathcal{Q}}^{\leq\nu})\big{)}={\mathbb{V}}_{r}(Q^{\leq l\nu})$. Hence we obtain identifications
-${\mathfrak{G}}\circ{\mathbb{V}}_{c}\big{(}\mathrm{Proj}({\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu})\big{)}={\mathbb{V}}_{r}\big{(}\mathrm{Proj}({\mathcal{O}}^{0,\leq l\nu}_{S})\big{)},\quad\nu\in\Lambda$
+We have the following isomorphisms by (5.2) and (5.10),
+${\mathbb{V}}_{r}(\tau^{\leq^{\frac{\infty}{2}}t(\nu)}_{r}Q)=\tau^{\leq-\rho+l\nu}_{r}({\mathbb{V}}_{r}Q)=\tau^{\leq-\rho+l\nu}_{r}(\mathrm{ind}_{r}{\mathbb{V}}^{b}_{r}(P)),$
+and similarly by (5.6) and (5.11) we have
+${\mathbb{V}}_{c}(\tau^{\leq^{\frac{\infty}{2}}t(\nu)}_{c}{\mathcal{Q}})=\tau^{\leq\nu}_{c}({\mathbb{V}}_{c}{\mathcal{Q}})=\tau^{\leq\nu}_{c}(\mathrm{ind}_{c}{\mathbb{V}}^{b}_{c}({\mathcal{P}})).$
+By (4.11), (4.10) and (4.15), we have isomorphisms
+$\mathfrak{G}\big(\tau^{\leq\nu}_{c}(\mathrm{ind}_{c}{\mathbb{V}}^{b}_{c}({\mathcal{P}}))\big)$ $=\tau^{\leq-\rho+l\nu}_{r}\circ{\mathrm{I}}^{\mathrm{hb}}_{1}\circ\mathrm{ind}_{c}({\mathbb{V}}^{b}_{c}({\mathcal{P}}))$
+$=\tau^{\leq-\rho+l\nu}_{r}\circ\mathrm{ind}_{r}\big(\mathfrak{G}^{b}({\mathbb{V}}^{b}_{c}({\mathcal{P}}))\big)$
+$=\tau^{\leq-\rho+l\nu}_{r}(\mathrm{ind}_{r}{\mathbb{V}}^{b}_{r}(P)).$
+It follows that $\mathfrak{G}\big({\mathbb{V}}_{c}(\tau^{\leq^{\frac{\infty}{2}}t(\nu)}_{c}{\mathcal{Q}})\big)={\mathbb{V}}_{r}(\tau^{\leq^{\frac{\infty}{2}}t(\nu)}_{r}Q)$.
+By Lemmas 5.9(1) and 5.18(1), we obtain identifications
+${\mathfrak{G}}\circ{\mathbb{V}}_{c}\big(\mathrm{Proj}({\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)})\big)={\mathbb{V}}_{r}\big(\mathrm{Proj}({\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S})\big),\quad\forall\nu\in\Lambda,$
 of subcategories in ${\mathcal{O}}^{-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}$. By Proposition 5.21, it induces an equivalence of additive categories
-$\mathrm{Proj}({\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq\nu}){\ \xrightarrow{\sim}\ }\mathrm{Proj}({\mathcal{O}}^{0,\leq l\nu}_{S}),\quad\forall\nu\in\Lambda,$
+$\mathrm{Proj}({\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B,\leq^{\frac{\infty}{2}}t(\nu)}){\ \xrightarrow{\sim}\ }\mathrm{Proj}({\mathcal{O}}^{0,\leq^{\frac{\infty}{2}}t(\nu)}_{S}),\quad\forall\nu\in\Lambda,$
 which extends to an equivalence of abelian categories
-${\mathfrak{F}}:\ {\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}^{0}_{S}.$
-Then we identify ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}=\mathrm{Coh}^{G}(\pi^{*}{\mathrm{A}}_{({\mathfrak{t}}/W)_{\hat{0}}})$ by induction from $B$ to $G$.
+${\mathfrak{F}}:\ {\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}^{0}_{S}.$
+[…]
 For the second assertion, we consider the following commutative diagram by construction
-[…]
+${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$${\mathcal{O}}^{0}_{S}$${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{T}$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}.$$\scriptstyle{{\mathfrak{F}}}\scriptstyle{\simeq}$$\scriptstyle{\mathrm{for}}$$\scriptstyle{\mathrm{for}}$$\scriptstyle{{\mathfrak{F}}^{b}}\scriptstyle{\simeq}$
 Hence ${\mathfrak{F}}({\mathcal{M}}_{x})$ ($x\in W_{\mathrm{ex}}$) is isomorphic to $M(x\bullet_{l}0)_{S}$ as a $\Lambda$-graded ${\mathfrak{U}}^{b}_{\zeta}\otimes S$-module.
-It forces that ${\mathfrak{F}}({\mathcal{M}}_{x})\cong M(x\bullet_{l}0)_{S}$ as $\Lambda$-graded $U^{\mathrm{hb}}_{\zeta}\otimes S$-modules.
-Now the second assertion follows from (4.15).
+It forces that ${\mathfrak{F}}({\mathcal{M}}_{x})=M(x\bullet_{l}0)_{S}$ as $\Lambda$-graded $U^{\mathrm{hb}}_{\zeta}\otimes S$-modules.
+Now the assertion follows from (4.21).
 ∎
 Remark 5.24.
-The assumption that $l$ is a power of prime is used in the results in §3.2 concerning the localization theorem for ${\mathfrak{U}}_{\zeta}$.
+The assumption that $l$ is a prime power is used in the results in §3.2 concerning the localization theorem for ${\mathfrak{U}}_{\zeta}$.
 However, it is expected that this assumption is not necessary (see a proof for type $A$ in [40]).
@@
 Let $\pi_{{}_{\widetilde{{\mathcal{N}}}}}:\widetilde{{\mathcal{N}}}=G\times^{B}{\mathfrak{n}}\rightarrow{\mathfrak{g}}$ be the Springer resolution.
-We set ${\mathrm{A}}^{\prime}$ be the specialization of ${\mathrm{A}}$ at $0\in{\mathfrak{t}}$, which satisfies an equivalence [12]
-$D^{b}\mathrm{Coh}(\widetilde{{\mathcal{N}}})\ \simeq\ D^{b}({\mathrm{A}}^{\prime}\text{-}\mathrm{mod}),$
+We set $\mathsf{A}$$^{\prime}$ be the specialization of $\mathsf{A}$} at $0\in{\mathfrak{t}}$, which satisfies an equivalence [11]
+$D^{\mathrm{b}}\mathrm{Coh}(\widetilde{{\mathcal{N}}})\ \simeq\ D^{\mathrm{b}}({\mathsf{A}}^{\prime}\text{-}\mathrm{mod}),$
 and therefore is called the non-commutative Springer resolution.
-Recall that ${\mathrm{A}}_{0}$ is the specialization of ${\mathrm{A}}$ at $0\in{\mathfrak{g}}$.
-We abbreviate ${\mathcal{O}}^{\chi_{0}}_{\mathbb{C}}$ as the full subcategory of ${\mathcal{O}}_{\mathbb{C}}$ of modules vanished by $\ker\chi_{0}\subset Z_{\mathrm{HC}}$.
+Recall that ${\mathsf{A}}_{0}$ is the specialization of $\mathsf{A}$} at $0\in{\mathfrak{g}}$.
+We abbreviate ${\mathcal{O}}^{\chi_{0}}_{\mathbb{C}}$ as the full subcategory of modules in $\mathcal{O}$$_{\mathbb{C}}$ on which the action of $Z_{\mathrm{HC}}$ factors through $\chi_{0}:Z_{\mathrm{HC}}\rightarrow{\mathbb{C}}$.
 Corollary 5.25.
 There is an equivalences of abelian categories
-(5.6) $\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathrm{A}})={\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{0},$
+(5.12) $\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}})={\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{0},$
 which by restrictions on $0\in{\mathfrak{t}}$ and $0\in{\mathfrak{g}}$ induces equivalences
-$\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathrm{A}}^{\prime}){\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{{\chi_{0}}},\quad\text{and}\quad{\mathrm{A}}_{0}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta},0}_{\mathbb{C}}.$
+$\mathrm{Coh}^{G}(\pi_{{}_{\widetilde{{\mathcal{N}}}}}^{*}{\mathsf{A}}^{\prime}){\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{{\chi_{0}}},\quad\text{and}\quad{\mathsf{A}}_{0}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}^{u_{\zeta}U^{+}_{\zeta},0}_{\mathbb{C}}.$
+Remark 5.26.
+Our construction can be applied to obtain a new approach to a similar equivalence for modular analogue of BGG category $\mathcal{O}$}, which is previously established by Losev [25, Equ (1.3) in Thm 1.7].
+To be more precise, assume moreover that $l$ is prime and set $\mathbb{F}$} be an algebraically closed field of characteristic $l$.
+Let $G_{\mathbb{F}}$ be a connected semisimple algebraic group over $\mathbb{F}$}, with a Borel subgroup $B_{\mathbb{F}}$ and a Cartan subgroup $T_{\mathbb{F}}$.
+Let $N_{\mathbb{F}}$ be the unipotent radical of $B_{\mathbb{F}}$, and let ${\mathfrak{g}}_{\mathbb{F}}=\mathrm{Lie}(G_{\mathbb{F}})={\mathfrak{n}}^{-}_{\mathbb{F}}\oplus{\mathfrak{t}}_{\mathbb{F}}\oplus{\mathfrak{n}}_{\mathbb{F}}$ be the corresponding triangular decomposition with ${\mathfrak{t}}_{\mathbb{F}}=\mathrm{Lie}(T_{\mathbb{F}})$ and ${\mathfrak{n}}_{\mathbb{F}}=\mathrm{Lie}(N_{\mathbb{F}})$.
+The modular analogue of BGG category $\mathcal{O}$} is the category
+${\mathcal{O}}_{\mathbb{F}}:={\mathfrak{g}}_{\mathbb{F}}\text{-}\mathrm{mod}^{B_{\mathbb{F}}}$
+of $B_{\mathbb{F}}$-equivariant finitely generated ${\mathfrak{g}}_{\mathbb{F}}$-modules on which the actions of ${\mathfrak{b}}_{\mathbb{F}}$ by restriction of ${\mathfrak{g}}_{\mathbb{F}}$-action and by differential of $B_{\mathbb{F}}$-action coincide.
+The category ${\mathcal{O}}_{\mathbb{F}}$ can be interpreted as the category $\mathcal{O}$} (with integral weights) associated to the modular analogue $U^{\mathrm{hb}}_{\mathbb{F}}$ of $U^{\mathrm{hb}}_{\zeta}$, which is an algebra with a triangular decomposition
+$U^{\mathrm{hb}}_{\mathbb{F}}=U{\mathfrak{n}}^{-}_{\mathbb{F}}\otimes U{\mathfrak{t}}_{\mathbb{F}}\otimes\mathrm{Dist}(N_{\mathbb{F}}),$
+where $\mathrm{Dist}(N_{\mathbb{F}})$ is the distribution algebra of $N_{\mathbb{F}}$.
+Let ${\mathsf{A}}_{\mathbb{F}}$ be the non-commutative Grothendieck resolution of ${\mathfrak{g}}^{(1)}_{\mathbb{F}}$ (the Frobenius twist of ${\mathfrak{g}}_{\mathbb{F}}$).
+Let $\pi_{\widetilde{{\mathcal{N}}}^{(1)}_{\mathbb{F}}}:\widetilde{{\mathcal{N}}}^{(1)}_{\mathbb{F}}\rightarrow{\mathfrak{g}}^{(1)}_{\mathbb{F}}$ be the Springer resolution.
+Then there is an equivalence of categories
+$\mathrm{Coh}^{G_{\mathbb{F}}^{(1)}}(\pi^{*}_{\widetilde{{\mathcal{N}}}^{(1)}_{\mathbb{F}}}{\mathsf{A}}_{\mathbb{F}})\ \simeq\ {\mathcal{O}}^{0}_{\mathbb{F}},$
+where ${\mathcal{O}}^{0}_{\mathbb{F}}$ is the principal block of ${\mathcal{O}}_{\mathbb{F}}$.
+To adapt our construction to obtain the equivalence above, one may replace Tanisaki’s localization theorem in §3.2 by the result for $U{\mathfrak{g}}_{\mathbb{F}}$ by Bezrukavnikov–Mirković–Rumynin [10] and replace Bezrukavnikov’s equivalence (used in §5.2.4) by its modular analogue recently proved by Bezrukavnikov–Riche [13].
 6. $\mathfrak{F}$} as equivariantization of ${\mathfrak{F}}^{b}$
 In this section we will use Theorem 5.23 to show that the equivalence ${\mathfrak{F}}^{b}$ intertwines rational $B$-actions on both categories, so that $\mathfrak{F}$} is the equivalence of their equivariantizations.
+This essentially follows the idea in the work of Arkhipov–Gaitsgory [3].
 As an application, we show that $\mathfrak{F}$} intertwines reflection functors $\Theta^{r}_{s}$ and $\Theta^{c}_{s}$, for all $s\in{\mathrm{I}}_{\mathrm{af}}$.
@@
 Now we consider the action of the monoidal category $\mathrm{Rep}(K)$ on an abelian category $\mathcal{C}$}. By definition, it is an exact bifunctor
-$-\otimes-:\ {\mathcal{C}}\times\mathrm{Rep}(K)\rightarrow{\mathcal{C}}$
+$\otimes-:\ {\mathcal{C}}\times\mathrm{Rep}(K)\rightarrow{\mathcal{C}}$
 with compatibility of the monoidal structure on $\mathrm{Rep}(K)$.
@@
 We have the following lemma
-Lemma 6.1 ([33, Prop A.2]).
+Lemma 6.1 ([31, Prop A.2]).
 Let $\mathcal{C}$} be a cocomplete abelian category, equipped with an action of $\mathrm{Rep}(K)$. There is an equivalence
@@
 Then $\widehat{{\mathcal{O}}}_{\mathbb{C}}$ is the ind-completion of $\mathcal{O}$$_{\mathbb{C}}$, and the equivalence (5.12) induces an equivalence
-$\widehat{{\mathfrak{F}}}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}{\ \xrightarrow{\sim}\ }\widehat{{\mathcal{O}}}^{0}_{\mathbb{C}}.$
+$\widehat{{\mathfrak{F}}}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}{\ \xrightarrow{\sim}\ }\widehat{{\mathcal{O}}}^{0}_{\mathbb{C}}.$
 Similarly we have ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}={\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$.
 The equivalence (4.5) induces an equivalence
-${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{mod}_{\chi_{0}}$
+${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{mod}_{\chi_{0}}$
 by forgetting the $\Lambda$-grading, which further induces an equivalence
-$\widehat{{\mathfrak{F}}}^{n}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}.$
+$\widehat{{\mathfrak{F}}}^{n}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}.$
 Note that $\widehat{{\mathfrak{F}}}$ and $\widehat{{\mathfrak{F}}}^{n}$ are compatible with the forgetful functor.
@@
 $\mathrm{Hom}_{{\mathbf{U}}^{\mathrm{hb}}_{\zeta}}(M_{1},\mathrm{coind}_{r}(M_{2}))\simeq\mathrm{Hom}_{{\mathfrak{U}}^{n}_{\zeta}}(M_{1},M_{2}),$
-which maps $\varphi$ on the LHS to the homomorphism $m_{1}\in M_{1}\mapsto\varphi(m_{1})(1)$, and $\phi$ on the RHS to the homomorphism $m_{1}\in M_{1}\mapsto\big{(}u\in{\mathbf{U}}^{\mathrm{hb}}_{\zeta}\mapsto\phi(um_{1})\big{)}$.
+which maps $\varphi$ on the LHS to the homomorphism $m_{1}\in M_{1}\mapsto\varphi(m_{1})(1)$, and maps $\phi$ on the RHS to the homomorphism $m_{1}\in M_{1}\mapsto\big(u\in{\mathbf{U}}^{\mathrm{hb}}_{\zeta}\mapsto\phi(um_{1})\big)$.
 We define ${\mathcal{C}}({\mathbf{U}}^{\geq}_{\zeta})$ as the category of integrable ${\mathbf{U}}^{\geq}_{\zeta}$-modules. Similarly, we have a functor
@@
 $\mathrm{coind}_{r}:\ {\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }(\widehat{{\mathcal{O}}}_{{\mathbb{C}}})_{B},$
-with inverse functor $-\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}$.
+with inverse functor $\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}$.
 Proof.
 One checks that the adjunction in Lemma 6.2 induces an adjunction
-$-\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}:\ (\widehat{{\mathcal{O}}}_{{\mathbb{C}}})_{B}\ \rightleftarrows\ {\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}\ :\mathrm{coind}_{r}.$
+$\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}:\ (\widehat{{\mathcal{O}}}_{{\mathbb{C}}})_{B}\ \rightleftarrows\ {\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}\ :\mathrm{coind}_{r}.$
 We have natural morphisms
@@
 $\mathrm{coind}^{b}_{r}:\ {\mathbf{u}}^{\geq}_{\zeta}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathcal{C}}({\mathbf{U}}^{\geq}_{\zeta})_{B}$
-with inverse functor $-\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}$.
-This follows from [4, Thm 2.8] applied to the sequence of Hopf algebras ${\mathbf{u}}^{\geq}_{\zeta}\hookrightarrow{\mathbf{U}}^{\geq}_{\zeta}\twoheadrightarrow U{\mathfrak{b}}$.
+with inverse functor $\otimes_{{\mathbb{C}}[B]}{\mathbb{C}}$.
+This follows from [3, Thm 2.8] applied to the sequence of Hopf algebras ${\mathbf{u}}^{\geq}_{\zeta}\hookrightarrow{\mathbf{U}}^{\geq}_{\zeta}\twoheadrightarrow U{\mathfrak{b}}$.
 ∎
@@
 Similarly, we consider the functor
-$\mathrm{coind}_{c}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}\rightarrow{\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B},\quad M\mapsto M\otimes{\mathbb{C}}[B].$
-It is right adjoint to the forgetful functor ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}\rightarrow{\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}$, and induces an equivalence
-(6.3) $\mathrm{coind}_{c}:\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B})_{B}.$
+$\mathrm{coind}_{c}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B},\quad M\mapsto M\otimes{\mathbb{C}}[B].$
+It is right adjoint to the forgetful functor ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}$, and induces an equivalence
+(6.3) $\mathrm{coind}_{c}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B})_{B}.$
 6.3. $\widehat{{\mathfrak{F}}}$ as equivariantization of $\widehat{{\mathfrak{F}}}^{n}$
@@
 We prove the assertion for $\mathfrak{F}$}, then the assertion for $\widehat{{\mathfrak{F}}}$ follows. We have a commutative diagram
-[…]
+${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$${\mathcal{O}}^{0}_{S}$$\mathrm{Coh}^{B}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$${\mathcal{O}}^{0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$$\scriptstyle{{\mathfrak{F}}}\scriptstyle{\simeq}$$\scriptstyle{{\mathbb{V}}_{c}}$$\scriptstyle{{\mathbb{V}}_{r}}$$\scriptstyle{{\mathfrak{G}}}\scriptstyle{\simeq}$
 Note that ${\mathbb{V}}_{c}$, ${\mathbb{V}}_{r}$ and $\mathfrak{G}$} are compatible with $\mathrm{rep}(B)$-action.
 Note also that the $\mathrm{rep}(B)$-action preserves the subcategory of objects admitting Verma flags.
-Hence by Proposition 5.11, ${\mathfrak{F}}$ is compatible with $\mathrm{rep}(B)$-action on such subcategory in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$.
+Hence by Proposition 5.21, $\mathfrak{F}$} is compatible with $\mathrm{rep}(B)$-action on such subcategory in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$.
 Since any object can be resolved by objects admitting Verma flags, such compatibility extends to the whole category.
@@
 Corollary 6.7.
-The equivalence $\widehat{{\mathfrak{F}}}^{n}:{\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}$ is compatible with the natural rational $B$-action on both sides.
+The equivalence $\widehat{{\mathfrak{F}}}^{n}:{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}{\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}$ is compatible with the natural rational $B$-action on both sides.
 Therefore, $\widehat{{\mathfrak{F}}}$ can be obtained as the $B$-equivariantization of $\widehat{{\mathfrak{F}}}^{n}$.
@@
 Since $\mathrm{coind}_{r}$ and $\mathrm{coind}_{c}$ are right adjoint to the forgetful functors, we have a commutative diagram
-[…]
+${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}$${\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}}$${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}$$\widehat{{\mathcal{O}}}_{\mathbb{C}}^{0}.$$\scriptstyle{\widehat{{\mathfrak{F}}}^{n}}\scriptstyle{\simeq}$$\scriptstyle{\mathrm{coind}_{c}}$$\scriptstyle{\mathrm{coind}_{r}}$$\scriptstyle{\widehat{{\mathfrak{F}}}}\scriptstyle{\simeq}$
 By Proposition 6.3 and (6.3), the coinduction functors identify the source categories with the de-equivariantizations of the target categories, and moreover the rational $B$-action on the sources come from the $\mathrm{Rep}(B)$-action on the targets.
@@
 The second assertion follows from the commutative diagram
-[…]
+(6.4) ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B}$$\widehat{{\mathcal{O}}}_{\mathbb{C}}^{0}$$({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod})^{B}$$({\mathfrak{U}}^{n}_{\zeta}\text{-}\mathrm{Mod}_{\chi_{0}})^{B},$$\scriptstyle{\widehat{{\mathfrak{F}}}}\scriptstyle{\simeq}\scriptstyle{\simeq}\scriptstyle{\simeq}$$\scriptstyle{(\widehat{{\mathfrak{F}}}^{n})^{B}}\scriptstyle{\simeq}$
 where the left vertical equivalence is tautological, the right one is by Corollary 6.4. ∎
@@
 The equivalence
-${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{0}$
+${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}{\ \xrightarrow{\sim}\ }{\mathcal{O}}_{\mathbb{C}}^{0}$
 intertwines the reflection functors $\Theta^{c}_{s}$ and $\Theta^{r}_{s}$, for any $s\in{\mathrm{I}}_{\mathrm{af}}$.
@@
 7. Graded multiplicities
-Recall that the tilting bundle ${\mathcal{E}}$ on $\tilde{{\mathfrak{g}}}$ can be equipped with a ${\mathbb{C}}^{\times}$-equivariant structure, which makes ${\mathrm{A}}$ a ${\mathbb{C}}^{\times}$-equivariant algebra.
-Using the equivalence (5.6), we can define a graded version of ${\mathcal{O}}^{0}_{\mathbb{C}}$ via
-${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\ :=\ \mathrm{Coh}^{G\times{\mathbb{C}}^{\times}}(\pi_{\tilde{{\mathcal{N}}}}^{*}{\mathrm{A}})\ \simeq\ {\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}},$
+[…]
+${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\ :=\ \mathrm{Coh}^{G\times{\mathbb{C}}^{\times}}(\pi_{\widetilde{{\mathcal{N}}}}^{*}{\mathsf{A}})\ \simeq\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}},$
 with a degrading functor ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\rightarrow{\mathcal{O}}^{0}_{\mathbb{C}}$.
-In this section, we compute the graded multiplicities of simple module in Verma module (resp.
-of Verma module in projective module) in ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}$.
+In this section, we compute the graded multiplicity of simple modules in Verma modules in ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}$, which turns out to be given by the generic Kazhdan–Lusztig polynomials defined in [24] (see also [28, 37]).
+We also obtain formulas for graded multiplicity of Verma module in projective module.
 7.1. Periodic and generic polynomials
 We recall the periodic and generic polynomials introduced by Lusztig [28] and Kato [24], following Soergel’s notations and normalizations [37].
-Let $l(-):W_{\mathrm{ex}}\rightarrow{\mathbb{Z}}$ be the length function.
+Let $\ell(-):W_{\mathrm{ex}}\rightarrow{\mathbb{Z}}$ be the length function.
 The extended affine Hecke algebra ${\mathbb{H}}_{\mathrm{ex}}$ is the ${\mathbb{Z}}[v^{\pm 1}]$-algebra generated by $\{H_{x}\}_{x\in W_{\mathrm{ex}}}$ modulo the relations
-$\begin{cases}(H_{s}+v)(H_{s}-v^{-1})=0&\text{for }s\in{\mathrm{I}}_{\mathrm{af}},\\ H_{x}H_{y}=H_{xy}&\text{if }l(x)+l(y)=l(xy).\end{cases}$
+$\begin{cases}(H_{s}+v)(H_{s}-v^{-1})=0&\text{for }s\in{\mathrm{I}}_{\mathrm{af}},\\ H_{x}H_{y}=H_{xy}&\text{if }\ell(x)+\ell(y)=\ell(xy).\end{cases}$
 There is a ring homomorphism (called the bar-involution) ${\mathbb{H}}_{\mathrm{ex}}\rightarrow{\mathbb{H}}_{\mathrm{ex}}$, $H\mapsto\overline{H}$ given by $\overline{v}=v^{-1}$, $\overline{H_{x}}=(H_{x^{-1}})^{-1}$.
+Recall the semi-infinite order $\leq^{\frac{\infty}{2}}$ in Definition 5.6.
 Let $\mathbb{P}$} be the free ${\mathbb{Z}}[v^{\pm 1}]$-module with basis $\{H^{\frac{\infty}{2}}_{x}\}_{x\in W_{\mathrm{ex}}}$.
-It is equipped with a right ${\mathbb{H}}_{\mathrm{ex}}$-module structure (called the periodic Hecke module) via
-$H^{\frac{\infty}{2}}_{x}\cdot H_{s}=\begin{cases}H^{\frac{\infty}{2}}_{xs}&\text{if }x\leq^{\frac{\infty}{2}}xs,\\ H^{\frac{\infty}{2}}_{xs}+(v^{-1}-v)H^{\frac{\infty}{2}}_{x}&\text{if }xs\leq^{\frac{\infty}{2}}x,\end{cases}$
-for any $x\in W_{\mathrm{ex}}$ and $s\in{\mathrm{I}}_{\mathrm{af}}$.
-For $\lambda\in\Lambda$ we set $e(\lambda):=\sum_{w\in W}v^{l(w)}\cdot H^{\frac{\infty}{2}}_{t(\lambda)w}$.
+It is equipped with a right ${\mathbb{H}}_{\mathrm{ex}}$-module structure (called the periodic Hecke module) via the following formulas for any $x\in W_{\mathrm{ex}}$, $s\in{\mathrm{I}}_{\mathrm{af}}$ and $\gamma\in\Lambda/{\mathrm{Q}}$,
+$H^{\frac{\infty}{2}}_{x}\cdot H_{s}$ $=\begin{cases}H^{\frac{\infty}{2}}_{xs}&\text{if }x\leq^{\frac{\infty}{2}}xs,\\ H^{\frac{\infty}{2}}_{xs}+(v^{-1}-v)H^{\frac{\infty}{2}}_{x}&\text{if }xs\leq^{\frac{\infty}{2}}x;\end{cases}$
+$H^{\frac{\infty}{2}}_{x}\cdot H_{\gamma}$ $=H^{\frac{\infty}{2}}_{x\gamma}.$
+For $\lambda\in\Lambda$ we set $e(\lambda):=\sum_{w\in W}v^{\ell(w)}\cdot H^{\frac{\infty}{2}}_{t(\lambda)w}$.
 Let $\mathbb{P}$$^{\circ}$ be the ${\mathbb{H}}_{\mathrm{ex}}$-submodule of $\mathbb{P}$} generated by $e(\lambda)$ with $\lambda\in\Lambda$.
@@
     On $\mathbb{P}$$^{\circ}$ there is a unique involution $\mathbb{P}$$^{\circ}\rightarrow{\mathbb{P}}^{\circ}$, $P\mapsto\overline{P}$ that is ${\mathbb{H}}_{\mathrm{ex}}$-skew-linear (i.e.
-compatible with the bar-involution) and satisfies $\overline{e(\lambda)}=e(\lambda)$ for each $\lambda\in\Lambda$.
+compatible with the bar-involution of ${\mathbb{H}}_{\mathrm{ex}}$) and satisfies $\overline{e(\lambda)}=e(\lambda)$ for each $\lambda\in\Lambda$.
   2. (2)
     For any $x\in W_{\mathrm{ex}}$, there is a unique element $\underline{H}^{\frac{\infty}{2}}_{x}\in{\mathbb{P}}^{\circ}$ such that $\underline{H}^{\frac{\infty}{2}}_{x}=\overline{\underline{H}^{\frac{\infty}{2}}_{x}}$ and $\underline{H}^{\frac{\infty}{2}}_{x}\in H^{\frac{\infty}{2}}_{x}+\sum_{y\leq^{\frac{\infty}{2}}x}v{\mathbb{Z}}[v]\cdot H^{\frac{\infty}{2}}_{y}$.
-We express $\underline{H}^{\frac{\infty}{2}}_{x}=\sum_{y}p_{y,x}\cdot H^{\frac{\infty}{2}}_{y}$, then $p_{y,x}\in{\mathbb{Z}}[v]$ is called the periodic Kazhdan–Lusztig polynomial.
+We express $\underline{H}^{\frac{\infty}{2}}_{x}=\sum_{y}p_{y,x}(v)\cdot H^{\frac{\infty}{2}}_{y}$, then $p_{y,x}(v)\in{\mathbb{Z}}[v]$ is called the periodic Kazhdan–Lusztig polynomial.
 Let $\hat{{\mathbb{P}}}$ be the completion of $\mathbb{P}$} consisting of the formal series in $\{H^{\frac{\infty}{2}}_{x}\}_{x\in W_{\mathrm{ex}}}$ that are bounded above (with respect to the order $\leq^{\frac{\infty}{2}}$).
 For $\eta\in\Lambda$, consider the ${\mathbb{Z}}[v^{\pm 1}]$-linear operator $\langle\eta\rangle$ on $\mathbb{P}$} by $\langle\eta\rangle(H^{\frac{\infty}{2}}_{x})=H^{\frac{\infty}{2}}_{t(\eta)x}$, for any $x\in W_{\mathrm{ex}}$.
-We define two families of polynomials $\{q_{y,x}\}_{x,y\in W_{\mathrm{ex}}}$ and $\{q^{\prime}_{y,x}\}_{x,y\in W_{\mathrm{ex}}}$ in ${\mathbb{Z}}[v]$ via the following equations in $\hat{{\mathbb{P}}}$ (for any $x\in W_{\mathrm{ex}}$)
-(7.1) $\big{(}\prod_{\alpha\in\Phi^{+}}(1+v^{2}\langle-\alpha\rangle+v^{4}\langle-2\alpha\rangle+\cdots)\big{)}(\underline{H}^{\frac{\infty}{2}}_{x})=\sum_{y}q_{y,x}\cdot H^{\frac{\infty}{2}}_{y},$
-(7.2) $\big{(}\prod_{\alpha\in\Phi^{+}}(1+\langle-\alpha\rangle+\langle-2\alpha\rangle+\cdots)\big{)}(\underline{H}^{\frac{\infty}{2}}_{x})=\sum_{y}q^{\prime}_{y,x}\cdot H^{\frac{\infty}{2}}_{y}.$
-Here $q_{y,x}$ is called the generic Kazhdan–Lusztig polynomial.
+We define two families of polynomials $\{q_{y,x}(v)\}_{x,y\in W_{\mathrm{ex}}}$ and $\{q^{\prime}_{y,x}(v)\}_{x,y\in W_{\mathrm{ex}}}$ in ${\mathbb{Z}}[v]$ via the following equations in $\hat{{\mathbb{P}}}$ (for any $x\in W_{\mathrm{ex}}$)
+(7.1) $\big(\prod_{\alpha\in\Phi^{+}}(1+v^{2}\langle-\alpha\rangle+v^{4}\langle-2\alpha\rangle+\cdots)\big)(\underline{H}^{\frac{\infty}{2}}_{x})=\sum_{y}q_{y,x}(v)\cdot H^{\frac{\infty}{2}}_{y},$
+(7.2) $\big(\prod_{\alpha\in\Phi^{+}}(1+\langle-\alpha\rangle+\langle-2\alpha\rangle+\cdots)\big)(\underline{H}^{\frac{\infty}{2}}_{x})=\sum_{y}q^{\prime}_{y,x}(v)\cdot H^{\frac{\infty}{2}}_{y}.$
+Here $q_{y,x}(v)$ is called the generic Kazhdan–Lusztig polynomial. Note that $q_{y,x}(1)=q_{y,x}^{\prime}(1)$.
 Remark 7.2.
-The generic polynomials $\{q_{x,y}\}_{x,y}$ were originally defined as the unique family of polynomials such that the element $\sum_{y}(-1)^{l(x)+l(y)}\overline{q}_{y,x}\cdot H^{\frac{\infty}{2}}_{y}\in\hat{{\mathbb{P}}}$ is contained in $H^{\frac{\infty}{2}}_{x}+\sum_{y}v^{-1}{\mathbb{Z}}[v^{-1}]\cdot H^{\frac{\infty}{2}}_{y}$ and is invariant under the involution, see [25].
-Under this definition, the equality (7.1) is proved in [25, Thm 3.5], see also [39, Thm 6.3].
+  1. (1)
+    In [37] the periodic (resp.
+generic) polynomials are labeled by alcoves: $p_{y,x}(v)$ (resp.
+$q_{y,x}(v)$) is denoted by $p_{yA_{\mathrm{fun}},xA_{\mathrm{fun}}}(v)$ (resp.
+by $q_{yA_{\mathrm{fun}},xA_{\mathrm{fun}}}(v)$) in the loc.
+cit..
+  2. (2)
+    The generic polynomials $\{q_{x,y}(v)\}_{x,y}$ are originally defined as the unique family of polynomials such that the element $\sum_{y}(-1)^{\ell(x)+\ell(y)}{q}_{y,x}(v^{-1})\cdot H^{\frac{\infty}{2}}_{y}\in\hat{{\mathbb{P}}}$ is contained in $H^{\frac{\infty}{2}}_{x}+\sum_{y}v^{-1}{\mathbb{Z}}[v^{-1}]\cdot H^{\frac{\infty}{2}}_{y}$ and is invariant under the involution on $\hat{{\mathbb{P}}}$ (obtained by extending the involution on $\mathbb{P}$$^{\circ}$), see [24].
+Under this definition, the equality (7.1) is proved in [24, Thm 3.5], see also [37, Thm 6.3].
 7.2. Graded multiplicities
@@
 For a graded $A$-module $M=M^{\bullet}$, we set $M\langle d\rangle$ as the graded $A$-module with $(M\langle d\rangle)^{i}=M^{i-d}$, for any $i\in{\mathbb{Z}}$.
-7.2.2. Graded ${\mathrm{A}}$-modules
+7.2.2. Graded $\mathsf{A}$}-modules
 Let $P(\lambda)_{\mathbb{C}}$ be the projective cover of $L(\lambda)_{\mathbb{C}}$ in $u_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$.
 Let $Z(\lambda)_{\mathbb{C}}=u_{\zeta}\otimes_{u_{\zeta}^{\geq 0}}{\mathbb{C}}_{\lambda}$ be the baby Verma module.
-Recall that $P^{b}(\lambda)_{\mathbb{C}}$ is the project cover of $L(\lambda)_{\mathbb{C}}$ in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$.
+Recall that $P^{b}(\lambda)_{\mathbb{C}}$ is the projective cover of $L(\lambda)_{\mathbb{C}}$ in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$.
 We have isomorphisms of ${\mathfrak{U}}^{b}_{\zeta}$-modules
@@
 Recall that we have a commutative diagram
+(7.5) ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}$${\mathcal{O}}^{0}_{\mathbb{C}}$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}}$$u_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{\mathbb{C}},$$\scriptstyle{\simeq}\scriptstyle{\simeq}\scriptstyle{\simeq}$
+where the left horizontal maps are by forgetful functors, and the right inclusions are induced by the quotients ${\mathsf{A}}_{\mathfrak{n}}\twoheadrightarrow{\mathsf{A}}_{0}$ and ${\mathfrak{U}}^{b}_{\zeta}\twoheadrightarrow u_{\zeta}$.
+For any $x\in W_{\mathrm{ex}}$ and bounded above poset ideal $\Omega\subset\Lambda$, we denote by
+${\mathcal{L}}_{x},\quad{\mathcal{Z}}_{x},\quad{\mathcal{P}}_{x},\quad{\mathcal{P}}^{b}_{x},\quad{{\mathcal{Q}}}_{x}^{\Omega}$
+the preimages of $L(x\bullet_{l}0)_{\mathbb{C}}$, $Z(x\bullet_{l}0)_{\mathbb{C}}$, $P(x\bullet_{l}0)_{\mathbb{C}}$, $P^{b}(x\bullet_{l}0)_{\mathbb{C}}$, $Q(x\bullet_{l}0)^{\Omega}_{\mathbb{C}}$ under the equivalences above.
+By Lemma 2.3, the ${\mathsf{A}}_{\mathfrak{n}}$-module ${\mathcal{L}}_{x}$ can be endowed with a $B$-equivariant structure, which makes ${\mathcal{L}}_{x}$ a simple object in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$.
+Recall the Verma objects ${\mathcal{M}}_{x}$ ($x\in W_{\mathrm{ex}}$) in ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$.
+We set
+${\mathcal{M}}^{\mathbb{C}}_{x}={\mathcal{M}}_{x}\otimes_{S}{\mathbb{C}}$
+and call ${\mathcal{M}}^{\mathbb{C}}_{x}$ (resp.
+${\mathcal{Z}}_{x}$) the Verma object in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ and ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ (resp.
+in ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}$).
+A Verma flag is a finite filtration with composition factors given by Verma objects.
+Lemma 7.3.
+Let $x\in W_{\mathrm{ex}}$, and let $\Omega\subset\Lambda$ be a bounded above poset ideal.
+  1. (1)
+    The objects ${\mathcal{Z}}_{x}$, ${\mathcal{P}}_{x}$, ${\mathcal{P}}^{b}_{x}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ admit liftings $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$, $\widetilde{{\mathcal{P}}}^{b}_{x}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$, and the liftings are unique up to grading shifts.
+  2. (2)
+    The objects ${\mathcal{L}}_{x}$, ${\mathcal{M}}^{\mathbb{C}}_{x}$, ${\mathcal{Q}}_{x}^{\Omega}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ admits liftings $\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$, $\widetilde{{\mathcal{Q}}}_{x}^{\Omega}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$, and the liftings are unique up to grading shifts.
+  3. (3)
+    Let $P$ be an object in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ with a lifting $\widetilde{P}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$.
+If $P$ admits a Verma flag, then $\widetilde{P}$ also admits a Verma flag (i.e.
+a filtration with composition factors of the form $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$).
+Similar result holds for ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ and ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}$.
+Proof.
 […]
-where the left horizontal maps are by forgetful functors, and the right inclusions are induced by the quotients ${\mathrm{A}}_{\mathfrak{n}}\twoheadrightarrow{\mathrm{A}}_{0}$ and ${\mathfrak{U}}^{b}_{\zeta}\twoheadrightarrow u_{\zeta}$.
-For $x\in W_{\mathrm{ex}}$, we denote by ${\mathcal{L}}_{x}$, ${\mathcal{Z}}_{x}$, ${\mathcal{P}}_{x}$, ${\mathcal{P}}^{b}_{x}$, ${{\mathcal{Q}}}_{x}^{\leq\nu}$ the preimages of $L(x\bullet_{l}0)_{\mathbb{C}}$, $Z(x\bullet_{l}0)_{\mathbb{C}}$, $P(x\bullet_{l}0)_{\mathbb{C}}$, $P^{b}(x\bullet_{l}0)_{\mathbb{C}}$, $Q(x\bullet_{l}0)^{\leq l\nu}_{\mathbb{C}}$ under the vertical equivalences above.
-By Lemma 2.2, the ${\mathrm{A}}_{\mathfrak{n}}$-module ${\mathcal{L}}_{x}$ can be endowed with a $B$-equivariant structure, which makes ${\mathcal{L}}_{x}$ a simple object in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$.
-Recall the Verma objects ${\mathcal{M}}_{x}$ ($x\in W_{\mathrm{ex}}$) in ${\mathrm{A}}^{b}_{S}\text{-}\mathrm{mod}^{B}$.
-Set ${\mathcal{M}}^{\mathbb{C}}_{x}={\mathcal{M}}_{x}\otimes_{S}{\mathbb{C}}$.
-We call ${\mathcal{M}}^{\mathbb{C}}_{x}$ (resp.
-${\mathcal{Z}}_{x}$) the Verma object in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ and ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ (resp.
-in ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$).
-A Verma flag is a filtration with composition factors given by Verma objects.
-Lemma 7.3.
-Let $x\in W_{\mathrm{ex}}$.
-  1. (1)
-    The objects ${\mathcal{Z}}_{x}$, ${\mathcal{P}}_{x}$, ${\mathcal{P}}^{b}_{x}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ admit liftings $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$, $\widetilde{{\mathcal{P}}}^{b}_{x}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$, and the liftings are unique up to grading shifts.
-  2. (2)
-    The objects ${\mathcal{L}}_{x}$, ${\mathcal{M}}^{\mathbb{C}}_{x}$, ${\mathcal{Q}}_{x}^{\leq\nu}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ admits liftings ~$\widetilde{{\mathcal{L}}}_{x}$, ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$, $\widetilde{{\mathcal{Q}}}_{x}^{\leq\nu}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$, and the liftings are unique up to grading shifts.
-  3. (3)
-    Let $P$ be an object in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$ with a lifting $\widetilde{P}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$.
-If $P$ admits a Verma flag, then $\widetilde{P}$ also admits a Verma flag (i.e.
-a filtration with composition factors of the form ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$).
-Similar result holds for ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T}$ and ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$.
-Proof.
-By a standard argument (see e.g.
-[12]), one can show that any $B$-equivariant simple ${\mathrm{A}}_{\mathfrak{n}}$-module (resp.
-$T$-equivariant projective module of ${\mathrm{A}}_{\mathfrak{n}}$ or ${\mathrm{A}}_{0}$) can be equipped with a ${\mathbb{C}}^{\times}$-equivariant structure, and up to isomorphism any lifting differs by a ${\mathbb{C}}^{\times}$-character.
-By its construction (4.15), any Verma object ${\mathcal{M}}^{\mathbb{C}}_{x}$ can be naturally equipped with a ${\mathbb{C}}^{\times}$-equivariant structure.
-Let ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$ be a lifting of ${\mathcal{M}}^{\mathbb{C}}_{x}$, then $\widetilde{{\mathcal{Z}}}_{x}=\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}$ provides a lifting of ${\mathcal{Z}}_{x}$ by (7.4).
-Suppose ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},1}_{x}$, ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},2}_{x}$ are two liftings of ${\mathcal{M}}^{\mathbb{C}}_{x}$.
-Since $\mathrm{End}({\mathcal{M}}^{\mathbb{C}}_{x})={\mathbb{C}}$, by (7.3) we have $\mathrm{Hom}(\widetilde{{\mathcal{M}}}^{{\mathbb{C}},1}_{x},\widetilde{{\mathcal{M}}}^{{\mathbb{C}},2}_{x}\langle i\rangle)={\mathbb{C}}\delta_{i,d}$ for some $d\in{\mathbb{Z}}$, and any nonzero map from ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},1}_{x}$ to ~$\widetilde{{\mathcal{M}}}^{{\mathbb{C}},2}_{x}\langle d\rangle$ is an isomorphism.
-The uniqueness of $\widetilde{{\mathcal{Z}}}_{x}$ follows similarly.
-This proves (1) as well as the assertion for ${\mathcal{L}}_{x}$ and ${\mathcal{M}}^{\mathbb{C}}_{x}$ in (2).
+[…]
+Step 2.
 Now we show (3).
-We only show the assertion for ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$.
+We only show the assertion for ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B}$.
 We prove by increasing induction on the length of Verma flag $\sum_{x}(P:{\mathcal{M}}^{\mathbb{C}}_{x})$.
-Let $x$ be maximal (with respect to the order $\leq^{\frac{\infty}{2}}$) among the elements $y\in W_{\mathrm{ex}}$ such that ${\mathcal{M}}^{\mathbb{C}}_{y}$ is a Verma factor of $P$.
-Then we have $\mathrm{Hom}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle,\widetilde{P})\neq 0$ for some $d$, and any nonzero element gives an inclusion ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle\hookrightarrow\widetilde{P}$ (since any nonzero map ${\mathcal{M}}^{\mathbb{C}}_{x}\rightarrow P$ is an inclusion).
+Let $x$ be such that $x\bullet_{l}0$ is maximal among the weights $y\bullet_{l}0$ with ${\mathcal{M}}^{\mathbb{C}}_{y}$ appearing as a Verma factor of $P$.
+Then we have $\mathrm{Hom}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle,\widetilde{P})\neq 0$ for some $d$, and any nonzero element gives an inclusion $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle\hookrightarrow\widetilde{P}$ (since any nonzero map ${\mathcal{M}}^{\mathbb{C}}_{x}\rightarrow P$ is an inclusion).
 Now $\widetilde{P}/\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$ is a lifting of $P/{\mathcal{M}}^{\mathbb{C}}_{x}$, which by induction hypothesis admits a Verma flag.
 Hence $\widetilde{P}$ admits a Verma flag.
-Finally, we show the assertion for ${\mathcal{Q}}_{x}^{\leq\nu}$ in (2).
-Consider the ${\mathbb{C}}^{\times}$-equivariant $({\mathrm{A}}_{n}\rtimes U{\mathfrak{n}})$-module $\widetilde{{\mathcal{Q}}}_{x}=({\mathrm{A}}_{n}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}_{\mathfrak{n}}}\widetilde{{\mathcal{P}}}^{b}_{x}$.
-By (3), $\widetilde{{\mathcal{P}}}^{b}_{x}$ admits a Verma flag.
-Hence by Lemma 5.6, $\widetilde{{\mathcal{Q}}}_{x}$ also admits a Verma flag.
-By (the non-defomed version of) Proposition 5.5, we have $\mathrm{Ext}^{1}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y},\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y^{\prime}}\langle d\rangle)\neq 0$ only if $y\leq^{\frac{\infty}{2}}y^{\prime}$.
-We define $\widetilde{{\mathcal{Q}}}_{x}^{\leq\nu}$ as the quotient of $\widetilde{{\mathcal{Q}}}_{x}$ by the submodule composed by ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\mu)y}$ with $\mu\nleq\nu$ and $y\in W$.
-As in Lemma 2.3(2), $\widetilde{{\mathcal{Q}}}_{x}^{\leq\nu}$ is a lifting of ${\mathcal{Q}}_{x}^{\leq\nu}$.
-For the uniqueness of the lifting, we set ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times},\leq\nu}$ as the Serre subcategory of ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$ generated by ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\mu)y}$ with $\mu\leq\nu$ and $y\in W$.
-Then any lifting of ${\mathcal{Q}}_{x}^{\leq\nu}$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$ is a projective cover of ~$\widetilde{{\mathcal{L}}}_{x}\langle d\rangle$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times},\leq\nu}$ for some $d$, which thus is unique up to a grading shift.
+Step 3.
+We show the assertion for ${\mathcal{Q}}_{x}^{\Omega}$ in (2).
+Recall the truncated categories ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}$, ${\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{B,\Omega}$, and the truncation functors $\tau^{T,\Omega}$, $\tau^{U{\mathfrak{n}},\Omega}$ introduced in §5.2.2.
+We set ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T}={\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T}\cap{\mathsf{A}}^{b}_{S}\text{-}\mathrm{Mod}^{T,\Omega}$ and define ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B,\Omega}$ similarly.
+The truncation functors restrict to
+$\tau^{T,\Omega}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T,\Omega}\quad\text{and}\quad\tau^{U{\mathfrak{n}},\Omega}:\ ({\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B,\Omega}.$
+We set ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times},\Omega}$ as the preimage of ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T,\Omega}$ under the degrading functor ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times}}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T}$, and define ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B\times{\mathbb{C}}^{\times},\Omega}$ similarly.
+Consider the truncation functor
+$\tau^{T\times{\mathbb{C}}^{\times},\Omega}:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times}}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times},\Omega},$
+by sending $M$ to its quotient by the minimal graded submodule containing $\ker(M\rightarrow\tau^{T,\Omega}(M))$.
+The functor $\tau^{T\times{\mathbb{C}}^{\times},\Omega}$ is left adjoint to the natural inclusion.
+We define
+$\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}:\ ({\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times}}\rightarrow{\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{B\times{\mathbb{C}}^{\times},\Omega},$
+by sending $M$ to its quotient by the (graded) ${\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}}$-module generated by $\ker(M\rightarrow\tau^{T\times{\mathbb{C}}^{\times},\Omega}(M))$.
+Then $\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}$ is left adjoint to the natural inclusion.
+We claim that the following natural transformations are isomorphisms
+(7.6) $\tau^{T,\Omega}\circ\upsilon{\ \xrightarrow{\sim}\ }\upsilon\circ\tau^{T\times{\mathbb{C}}^{\times},\Omega}\quad\text{and}\quad\tau^{U{\mathfrak{n}},\Omega}\circ\upsilon{\ \xrightarrow{\sim}\ }\upsilon\circ\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}$
+where $\upsilon$ is the degrading functor.
+Note that the second isomorphism follows from the first one, hence we only need to show the first isomorphism.
+By the right exactness of truncation functors and degrading functors, it is enough to construct isomorphism $\upsilon\circ\tau^{T\times{\mathbb{C}}^{\times},\Omega}(P^{\mathrm{gr}})\simeq\tau^{T,\Omega}(P)$ for any $T$-equivariant projective ${\mathsf{A}}_{\mathfrak{n}}$-module $P$ and its lifting $P^{\mathrm{gr}}$ (which exists by (1)).
+By (3), $P^{\mathrm{gr}}$ admits a Verma filtration.
+By (2.4) and (7.3) applied to the degrading $D^{\mathrm{b}}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}})\rightarrow D^{\mathrm{b}}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T})$, we have $\mathrm{Ext}^{1}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x},\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle)=0$ if $x\bullet_{l}0\nless y\bullet_{l}0$.
+Hence $\tau^{T\times{\mathbb{C}}^{\times},\Omega}(P^{\mathrm{gr}})$ is the quotient of $P^{\mathrm{gr}}$ by the submodule composed by Verma factors $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle$ with $y\bullet_{l}0\notin\Omega$, which shows that $\tau^{T\times{\mathbb{C}}^{\times},\Omega}(P^{\mathrm{gr}})=\tau^{T,\Omega}(P)$ after forgetting the grading structure.
+It proves the claim.
+Finally, consider the induction functor
+$\mathrm{ind}_{c}:=({\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\otimes_{{\mathsf{A}}_{\mathfrak{n}}}-:\ {\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}\rightarrow({\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}.$
+Let $x\in W_{\mathrm{ex}}$ such that $x\bullet_{l}0\in\Omega$.
+The module $\widetilde{{\mathcal{Q}}}_{x}^{\Omega}=\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}(\mathrm{ind}_{c}\widetilde{{\mathcal{P}}}^{b}_{x})$ is the projective cover of $\widetilde{{\mathcal{L}}}_{x}\langle d\rangle$ (for some $d\in{\mathbb{Z}}$) in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times},\Omega}$, which provide a lifting of ${\mathcal{Q}}^{\Omega}_{x}=\tau^{U{\mathfrak{n}},\Omega}(\mathrm{ind}_{c}{{\mathcal{P}}}^{b}_{x})$.
+The uniqueness of lifting follows from the uniqueness of projective cover.
 ∎
-For an object $P$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$ admitting Verma flags, we set $(P:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle)$ as the multiplicity of the Verma factor ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$, which does not depend on the choice of Verma flag.
+For an object $P$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$ admitting Verma flags, we set $(P:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle)$ as the multiplicity of the Verma factor $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle$, which does not depend on the choice of Verma flag.
 We define the graded multiplicity as
 (7.7) $(P:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x})_{{\mathrm{gr}}}=\sum_{d}(P:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\langle d\rangle)\cdot v^{d}\ \in{\mathbb{Z}}[v^{\pm 1}].$
-For any object $M$ in ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$, it admits a separated decreasing ${\mathbb{Z}}_{\leq 0}$-filtration (usually of infinite length) with simple composition factors.
-We set $[M:\widetilde{{\mathcal{L}}}_{x}\langle d\rangle]$ as the multiplicity (which is finite) of ~$\widetilde{{\mathcal{L}}}_{x}\langle d\rangle$ in the factors, which dose not depend on the choice of such filtration, and we define the graded multiplicity
-(7.6) $[M:\widetilde{{\mathcal{L}}}_{x}]_{{\mathrm{gr}}}=\sum_{d}[M:\widetilde{{\mathcal{L}}}_{x}\langle d\rangle]\cdot v^{d}\ \in{\mathbb{Z}}[v^{\pm 1}].$
-We also introduce similar notations for ${\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$ and ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
-Let $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$ be the space of ${\mathbb{Z}}[v^{\pm 1}]$-linear formal series in {[~$\{[\widetilde{{\mathcal{L}}}_{x}]\}_{x}$ that are bounded above (with respect to the order $\leq^{\frac{\infty}{2}}$).
+For any object $M$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}$, it admits a separated decreasing ${\mathbb{Z}}_{\leq 0}$-filtration (may of infinite length) with simple composition factors.
+We set $M:\widetilde{{\mathcal{L}}}_{x}\langle d\rangle]$ as the multiplicity (which is finite) of $\widetilde{{\mathcal{L}}}_{x}\langle d\rangle$ in the factors, which dose not depend on the choice of such filtration, and we define the graded multiplicity
+(7.8) $M:\widetilde{{\mathcal{L}}}_{x}]_{{\mathrm{gr}}}=\sum_{d}[M:\widetilde{{\mathcal{L}}}_{x}\langle d\rangle]\cdot v^{d}\ \in{\mathbb{Z}}[v^{\pm 1}].$
+We also introduce similar notations for ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$ and ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
+Let $\hat{K}_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$ be the space of ${\mathbb{Z}}[v^{\pm 1}]$-linear formal series in $\{[\widetilde{{\mathcal{L}}}_{x}]\}_{x}$ that are bounded above (with respect to the order $\leq^{\frac{\infty}{2}}$).
 We have an inclusion
-$K_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})\rightarrow\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}),\quad[M]\mapsto\sum_{x}[M:\widetilde{{\mathcal{L}}}_{x}]_{{\mathrm{gr}}}\cdot[\widetilde{{\mathcal{L}}}_{x}].$
-Note that (the image of) {[~$\{[\widetilde{{\mathcal{M}}}_{x}]\}_{x}$, $\{[\widetilde{{\mathcal{Z}}}_{x}]\}_{x}$ give other topological basis on $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$.
+$K_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})\rightarrow\hat{K}_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}}),\quad[M]\mapsto\sum_{x}[M:\widetilde{{\mathcal{L}}}_{x}]_{{\mathrm{gr}}}\cdot[\widetilde{{\mathcal{L}}}_{x}].$
+Note that (the images of) $\{[\widetilde{{\mathcal{M}}}_{x}]\}_{x}$ and $\{[\widetilde{{\mathcal{Z}}}_{x}]\}_{x}$ give other topological basis on $\hat{K}_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$.
 7.2.3. Graded multiplicities
@@
 Theorem 7.4.
-Under the degrading functor ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\rightarrow{\mathcal{O}}_{\mathbb{C}}^{0}$, the modules $L(x\bullet_{l}0)_{\mathbb{C}}$, $M(x\bullet_{l}0)_{\mathbb{C}}$ and $Q(x\bullet_{l}0)_{\mathbb{C}}^{\leq\nu}$ ($x\in W_{\mathrm{ex}}$) admit liftings $\widetilde{L}(x\bullet_{l}0)_{\mathbb{C}}$, $\widetilde{M}(x\bullet_{l}0)_{\mathbb{C}}$ and $\widetilde{Q}(x\bullet_{l}0)_{\mathbb{C}}^{\leq\nu}$ in ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}$, such that the graded multiplicities are given by
-(7.7) $(\widetilde{Q}(x\bullet_{l}0)_{\mathbb{C}}^{\leq\nu}:\widetilde{M}(y\bullet_{l}0)_{\mathbb{C}})_{\mathrm{gr}}=\begin{cases}q^{\prime}_{w_{0}y,w_{0}x}(v)&\text{if }y\bullet_{l}0\leq l\nu,\\ 0&\text{if else},\end{cases}$
+Let $\Omega$ be a bounded above poset ideal of $\Lambda$.
+Under the degrading functor ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}\rightarrow{\mathcal{O}}_{\mathbb{C}}^{0}$, the modules $L(x\bullet_{l}0)_{\mathbb{C}}$, $M(x\bullet_{l}0)_{\mathbb{C}}$ and $Q(x\bullet_{l}0)_{\mathbb{C}}^{\Omega}$ ($x\in W_{\mathrm{ex}}$) admit liftings $\widetilde{L}(x\bullet_{l}0)_{\mathbb{C}}$, $\widetilde{M}(x\bullet_{l}0)_{\mathbb{C}}$ and $\widetilde{Q}(x\bullet_{l}0)_{\mathbb{C}}^{\Omega}$ in ${\mathcal{O}}_{\mathbb{C}}^{0,{\mathrm{gr}}}$, such that the graded multiplicities are given by
+[…]
 and
-(7.8) $[\widetilde{M}(x\bullet_{l}0)_{\mathbb{C}}:\widetilde{L}(y\bullet_{l}0)_{\mathbb{C}}]_{\mathrm{gr}}=q_{w_{0}x,w_{0}y}(v),$
+(7.10) $\widetilde{M}(x\bullet_{l}0)_{\mathbb{C}}:\widetilde{L}(y\bullet_{l}0)_{\mathbb{C}}]_{\mathrm{gr}}=q_{w_{0}x,w_{0}y}(v),$
 for any $x,y\in W_{\mathrm{ex}}$.
-Before proving Theorem 7.4, we firstly consider the graded multiplicities in ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
+Before proving Theorem 7.4, we firstly consider the graded multiplicities in ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
 Set $\mathrm{B}$} as the $\Lambda$-graded (alternatively, $T$-equivariant) algebra with an isomorphism
-$({\mathrm{B}})^{\mathrm{op}}=\bigoplus_{\lambda\in\Lambda}\mathrm{Hom}\big{(}\bigoplus_{w\in W}P(w\bullet 0)_{\mathbb{C}},\bigoplus_{w\in W}P(w\bullet 0-l\lambda)_{\mathbb{C}}\big{)}.$
+${\mathrm{B}}^{\mathrm{op}}=\bigoplus_{\lambda\in\Lambda}\mathrm{Hom}\big(\bigoplus_{w\in W}P(w\bullet 0)_{\mathbb{C}},\bigoplus_{w\in W}P(w\bullet 0-l\lambda)_{\mathbb{C}}\big).$
 Note that $\mathrm{B}$} is finite dimensional, and there is an equivalence
@@
 We denote by $L_{x}$, $Z_{x}$, $P_{x}$ the image of $L(x\bullet_{l}0)_{\mathbb{C}}$, $Z(x\bullet_{l}0)_{\mathbb{C}}$, $P(x\bullet_{l}0)_{\mathbb{C}}$ ($x\in W_{\mathrm{ex}}$) under this equivalence.
-The following theorem is a recollection of results in [1, §18] (with conventions adopted from [39]).
+The following theorem is a recollection of results by Andersen–Jantzen–Soergel in [1, §18] (with conventions adopted from [37]).
 Theorem 7.5 ([1]).
@@
 Proposition 7.6.
-The liftings ~$\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$ ($x\in W_{\mathrm{ex}}$) in ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$ can be chosen such that ~$\widetilde{{\mathcal{L}}}_{t(\lambda)x}=\widetilde{{\mathcal{L}}}_{x}\otimes{\mathbb{C}}_{\lambda}$ for any $\lambda\in\Lambda$ (and similarly for $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$), and
+The liftings $\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$ ($x\in W_{\mathrm{ex}}$) in ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$ can be chosen such that $\widetilde{{\mathcal{L}}}_{t(\lambda)x}=\widetilde{{\mathcal{L}}}_{x}\otimes{\mathbb{C}}_{\lambda}$ for any $\lambda\in\Lambda$ (and similarly for $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$), and
 $(\widetilde{{\mathcal{P}}}_{y}:\widetilde{{\mathcal{Z}}}_{x})_{{\mathrm{gr}}}=[\widetilde{{\mathcal{Z}}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{{\mathrm{gr}}}=p_{w_{0}x,w_{0}y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
 Proof.
-Through the equivalences ${\mathrm{B}}\text{-}\mathrm{mod}^{T}\simeq u_{\zeta}\text{-}\mathrm{mod}_{\mathbb{C}}^{\Lambda,0}\simeq{\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$, the grading ${\mathrm{B}}^{\bullet}$ on ${\mathrm{B}}$ induces a different grading, hence a different graded multiplicity, on ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T}$.
-Our main task is to compare this graded multiplicity to the natural one induced by ${\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
-By [12, §5.2.3, §5.5] (see also [11, Thm 4.5]), there is a $T\times{\mathbb{C}}^{\times}$-equivariant vector bundle ${\mathcal{E}}^{new}$ on $\tilde{{\mathfrak{g}}}$ satisfying Theorem 3.2(2), such that the graded algebra $\mathrm{End}^{{\mathrm{op}}}_{\tilde{{\mathfrak{g}}}}({\mathcal{E}}^{new})$ is Koszul (where grading is induced from ${\mathbb{C}}^{\times}$-equivariant structure).
-We can replace ${\mathcal{E}}$ by ${\mathcal{E}}^{new}$, then ${\mathrm{A}}^{\bullet}$ is a Koszul algebra.
-In particular, the graded algebra ${\mathrm{A}}_{0}^{\bullet}={\mathrm{A}}^{\bullet}\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}$ satisfies: (1) ${\mathrm{A}}_{0}^{\bullet}$ is non-negatively graded; (2) ${\mathrm{A}}^{0}_{0}$ is a semi-simple algebra; (3) ${\mathrm{A}}_{0}^{\bullet}$ is generated by ${\mathrm{A}}^{1}_{0}$ over ${\mathrm{A}}^{0}_{0}$.
-We can choose a lifting $\widetilde{{\mathcal{P}}}_{x}$ of ${\mathcal{P}}_{x}$ ($x\in W_{\mathrm{ex}}$) such that it is generated by degree 0 elements as a graded ${\mathrm{A}}_{0}^{\bullet}$-module.
+Through the equivalences ${\mathrm{B}}\text{-}\mathrm{mod}^{T}\simeq u_{\zeta}\text{-}\mathrm{mod}_{\mathbb{C}}^{\Lambda,0}\simeq{\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}$, the grading $\mathrm{B}$$^{\bullet}$ on $\mathrm{B}$} induces a different grading, hence a different graded multiplicity, on ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T}$.
+Our main task is to compare this graded multiplicity to the natural one induced by ${\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$.
+[…]
+We can choose a lifting $\widetilde{{\mathcal{P}}}_{x}$ of ${\mathcal{P}}_{x}$ ($x\in W_{\mathrm{ex}}$) such that it is generated by degree 0 elements as a graded ${\mathsf{A}}_{0}^{\bullet}$-module.
 Define a graded algebra $\mathrm{B}$$^{\prime\bullet}$ via
@@
 Then we have an isomorphism of algebras ${\mathrm{B}}\simeq{\mathrm{B}}^{\prime}$.
-From the properties of ${\mathrm{A}}_{0}^{\bullet}$ (resp.
+From the properties of ${\mathsf{A}}_{0}^{\bullet}$ (resp.
 Koszulity of $\mathrm{B}$$^{\bullet}$), we deduce that (1) $\mathrm{B}$$^{\prime\bullet}$ (resp.
@@
 ${\mathrm{B}}^{0}$).
-Hence by [6, Prop 2.4.1], we have $\mathrm{rad}({\mathrm{B}}^{\prime})^{i}=\bigoplus_{j\geq i}{\mathrm{B}}^{\prime j}$ and $\mathrm{rad}({\mathrm{B}})^{i}=\bigoplus_{j\geq i}{\mathrm{B}}^{j}$, and therefore we have a ($T$-equivariant) isomorphism of graded algebras
+Hence by [5, Prop 2.4.1], we have $\mathrm{rad}({\mathrm{B}}^{\prime})^{i}=\bigoplus_{j\geq i}{\mathrm{B}}^{\prime j}$ and $\mathrm{rad}({\mathrm{B}})^{i}=\bigoplus_{j\geq i}{\mathrm{B}}^{j}$, and therefore we have a ($T$-equivariant) isomorphism of graded algebras
 (7.11) ${\mathrm{B}}^{\bullet}{\ \xrightarrow{\sim}\ }\bigoplus_{i}\mathrm{rad}({\mathrm{B}})^{i}/\mathrm{rad}({\mathrm{B}})^{i+1}\simeq\bigoplus_{i}\mathrm{rad}({\mathrm{B}}^{\prime})^{i}/\mathrm{rad}({\mathrm{B}}^{\prime})^{i+1}\ \xleftarrow{\sim}{\mathrm{B}}^{\prime\bullet}.$
@@
 Note that the isomorphism (7.11) and the earlier isomorphism ${\mathrm{B}}\simeq{\mathrm{B}}^{\prime}$ might be different, but they induce the same isomorphism on the semi-simple quotients ${\mathrm{B}}/\mathrm{rad}({\mathrm{B}})\simeq{\mathrm{B}}^{\prime}/\mathrm{rad}({\mathrm{B}}^{\prime})$.
-Therefore, under the composed equivalence ${\mathrm{B}}^{\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathrm{B}}^{\prime\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathrm{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$, the simple module $\widetilde{L}_{x}$ is send to ~$\widetilde{{\mathcal{L}}}_{x}$ ($x\in W_{\mathrm{ex}}$) (up to shifts).
+Therefore, under the composed equivalence ${\mathrm{B}}^{\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathrm{B}}^{\prime\bullet}\text{-}\mathrm{gmod}^{T}\simeq{\mathsf{A}}_{0}\text{-}\mathrm{mod}^{T\times{\mathbb{C}}^{\times}}$, the simple module $\widetilde{L}_{x}$ is send to $\widetilde{{\mathcal{L}}}_{x}$ ($x\in W_{\mathrm{ex}}$) (up to shifts).
 As projective cover of $\widetilde{L}_{x}$, $\widetilde{P}_{x}$ is send to $\widetilde{{\mathcal{P}}}_{x}$ (up to shifts).
@@
 Remark 7.7.
-Proposition 7.6 might be proved as the analogue statement for restricted enveloping algebra in [11, Thm 1.2], instead of using Theorem 7.5.
+Proposition 7.6 might be proved as the analogue statement for restricted enveloping algebra in [9, Thm 1.2], instead of using Theorem 7.5.
 Proof of Theorem 7.4.
-We fix the liftings ~$\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$ ($x\in W_{\mathrm{ex}}$) in Proposition 7.6.
-And choose liftings ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$, $\widetilde{{\mathcal{P}}}^{b}_{x}$ such that ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}=\widetilde{{\mathcal{Z}}}_{x}$ and $\widetilde{{\mathcal{P}}}^{b}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}=\widetilde{{\mathcal{P}}}_{x}$.
-Then we have ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\lambda)x}=\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes{\mathbb{C}}_{\lambda}$ for any $\lambda\in\Lambda$ (and similarly for $\widetilde{{\mathcal{P}}}^{b}_{x}$).
-Apply $-\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}$ to any Verma flag of $\widetilde{{\mathcal{P}}}^{b}_{x}$, we obtain a Verma flag of $\widetilde{{\mathcal{P}}}_{x}$, since the corresponding statement in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$ is true by the freeness of Verma modules over $Z_{\mathrm{Fr}}^{-}$.
+We fix the liftings $\widetilde{{\mathcal{L}}}_{x}$, $\widetilde{{\mathcal{Z}}}_{x}$, $\widetilde{{\mathcal{P}}}_{x}$ ($x\in W_{\mathrm{ex}}$) in Proposition 7.6.
+And choose liftings $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}$, $\widetilde{{\mathcal{P}}}^{b}_{x}$ such that $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}=\widetilde{{\mathcal{Z}}}_{x}$ and $\widetilde{{\mathcal{P}}}^{b}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}=\widetilde{{\mathcal{P}}}_{x}$.
+Then we have $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\lambda)x}=\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes{\mathbb{C}}_{\lambda}$ for any $\lambda\in\Lambda$ (and similarly for $\widetilde{{\mathcal{P}}}^{b}_{x}$).
+Step 1. We firstly consider the multiplicity $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{\mathrm{gr}}$. Consider the following sequence induced by Koszul resolution
+(7.12) $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}({\mathbb{C}}[{\mathfrak{n}}]\otimes\bigwedge^{\bullet}{\mathfrak{n}}^{*})\rightarrow\widetilde{{\mathcal{Z}}}_{x}\rightarrow 0,$
+which is exact, since the corresponding sequence in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$
+$M(x\bullet_{l}0)_{\mathbb{C}}\otimes_{Z_{\mathrm{Fr}}^{-}}(Z_{\mathrm{Fr}}^{-}\otimes\bigwedge^{\bullet}{\mathfrak{n}}^{*})\rightarrow Z(x\bullet_{l}0)_{\mathbb{C}}\rightarrow 0$
+is exact by the freeness of the $Z_{\mathrm{Fr}}^{-}$-module $M(x\bullet_{l}0)_{\mathbb{C}}$.
+By abuse of notation, we view $\langle\eta\rangle$ ($\eta\in\Lambda$) as a continuous ${\mathbb{Z}}[v^{\pm 1}]$-linear operator on $\hat{K}_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$ sending $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]$ to $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\eta)x}]$, for any $x\in W_{\mathrm{ex}}$.
+By (7.12), we have the following equality in $K_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$,
+(7.13) $\widetilde{{\mathcal{Z}}}_{x}]=\big(\prod_{\alpha\in\Phi^{+}}(1-v^{2}\langle-\alpha\rangle)\big)[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
+In $\hat{K}_{0}({\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$, we express $\widetilde{{\mathcal{L}}}_{y}]$ into (topological) basis $\{[\widetilde{{\mathcal{Z}}}_{x}]\}_{x}$ and $\{[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]\}_{x}$ as
+$\widetilde{{\mathcal{L}}}_{y}]=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{Z}}}_{x}]=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
+Then (7.13) shows that
+(7.14) $\big(\prod_{\alpha\in\Phi^{+}}(1-v^{2}\langle-\alpha\rangle)\big)(\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}])=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
+By [28, §11] (see also [36, Thm 6.1]), there is an equality
+(7.15) $\sum_{x}(-1)^{\ell(x)+\ell(y)}q_{x,y}(v)\cdot p_{w_{0}x,w_{0}z}(v)=\delta_{y,z},\quad\forall y,z\in W_{\mathrm{ex}}.$
+Hence it follows from Proposition 7.6 that
+(7.16) $\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}=(-1)^{\ell(x)+\ell(y)}q_{x,y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
+Combining (7.1), (7.14) with (7.16), we deduce that
+$\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}=(-1)^{\ell(x)+\ell(y)}p_{x,y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
+By (7.15) again, we have
+(7.17) $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{\mathrm{gr}}=q_{w_{0}x,w_{0}y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
+Step 2.
+Now we study the multiplicity $(\widetilde{{\mathcal{Q}}}^{\Omega}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}$.
+Apply $\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}{\mathbb{C}}$ to any Verma flag of $\widetilde{{\mathcal{P}}}^{b}_{x}$, we obtain a Verma flag of $\widetilde{{\mathcal{P}}}_{x}$, since the corresponding statement in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$ is true by the freeness of Verma modules over $Z_{\mathrm{Fr}}^{-}$.
 Hence we have
 (7.18) $(\widetilde{{\mathcal{P}}}^{b}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}=(\widetilde{{\mathcal{P}}}_{x}:\widetilde{{\mathcal{Z}}}_{y})_{\mathrm{gr}}.$
-Recall the $({\mathrm{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})$-module $\widetilde{{\mathcal{Q}}}_{x}=({\mathrm{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\otimes_{{\mathrm{A}}_{\mathfrak{n}}}\widetilde{{\mathcal{P}}}^{b}_{x}=U{\mathfrak{n}}\otimes\widetilde{{\mathcal{P}}}^{b}_{x}$, which admits a Verma flag with factors of the form ${\mathbb{C}}_{\eta}\otimes\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}=\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\eta)y}$, where $\eta\geq 0$ and ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}$ is a Verma factor of $\widetilde{{\mathcal{P}}}^{b}_{x}$.
-Moreover, the graded multiplicities satisfy the following equality in $\hat{{\mathbb{P}}}$,
+Consider the $(T\times{\mathbb{C}}^{\times})$-equivariant ${\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}}$-module
+$\widetilde{{\mathcal{Q}}}_{x}:=\mathrm{ind}_{c}(\widetilde{{\mathcal{P}}}^{b}_{x})=({\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}})\otimes_{{\mathsf{A}}_{\mathfrak{n}}}\widetilde{{\mathcal{P}}}^{b}_{x},$
+[…]
+(7.19) $0={\mathcal{Q}}_{0}\subset{\mathcal{Q}}_{1}\subset{\mathcal{Q}}_{2}\subset\cdots\subset{\mathcal{Q}}_{n}=\widetilde{{\mathcal{Q}}}_{x},$
+with composition factors ${\mathcal{Q}}_{i+1}/{\mathcal{Q}}_{i}$ of the form $\mathrm{ind}_{c}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle)$, where $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle$ are Verma factors of $\widetilde{{\mathcal{P}}}^{b}_{x}$.
+Since $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}$ is a ${\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}}$-module, there is an isomorphism of ${\mathsf{A}}_{\mathfrak{n}}\rtimes U{\mathfrak{n}}$-modules $\mathrm{ind}_{c}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})\simeq\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes U{\mathfrak{n}}$, where $U{\mathfrak{n}}$ acts on $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes U{\mathfrak{n}}$ diagonally.
+There is a separated decreasing filtration
+$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes U{\mathfrak{n}}={\mathcal{M}}_{0}\supset{\mathcal{M}}_{1}\supset\cdots,$
+with composition factors ${\mathcal{M}}_{i}/{\mathcal{M}}_{i+1}$ of the form $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes{\mathbb{C}}_{\eta}$, $\eta\geq 0$. We formally write
+$(\widetilde{{\mathcal{Q}}}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}:=\sum_{i=0}^{n-1}({\mathcal{Q}}_{i+1}/{\mathcal{Q}}_{i}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}.$
+Then there is an equality in $\hat{{\mathbb{P}}}$,
 $\big(\prod_{\alpha\in\Phi^{+}}(1+\langle-\alpha\rangle+\langle-2\alpha\rangle+\cdots)\big)(\sum_{y}(\widetilde{{\mathcal{P}}}^{b}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}\cdot H^{\frac{\infty}{2}}_{w_{0}y})=\sum_{y}(\widetilde{{\mathcal{Q}}}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}\cdot H^{\frac{\infty}{2}}_{w_{0}y}.$
@@
 $(\widetilde{{\mathcal{Q}}}_{x}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}=q^{\prime}_{w_{0}y,w_{0}x}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
-Since $\widetilde{{\mathcal{Q}}}^{\leq\nu}_{x}$ is the truncation of $\widetilde{{\mathcal{Q}}}_{x}$ by $\nu$, (7.7) follows.
-Now we study the multiplicity [~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{\mathrm{gr}}$. Consider the following sequence induced by Koszul resolution
-(7.11) ~$\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}\otimes_{{\mathbb{C}}[{\mathfrak{n}}]}({\mathbb{C}}[{\mathfrak{n}}]\otimes\bigwedge^{\bullet}{\mathfrak{n}}^{*})\rightarrow\widetilde{{\mathcal{Z}}}_{x}\rightarrow 0,$
-which is exact, since the corresponding sequence in ${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{\mathbb{C}}$
-$M(x\bullet_{l}0)_{\mathbb{C}}\otimes_{Z_{\mathrm{Fr}}^{-}}(Z_{\mathrm{Fr}}^{-}\otimes\bigwedge^{\bullet}{\mathfrak{n}}^{*})\rightarrow Z(x\bullet_{l}0)_{\mathbb{C}}\rightarrow 0$
-is exact by the freeness of the $Z_{\mathrm{Fr}}^{-}$-module $M(x\bullet_{l}0)_{\mathbb{C}}$.
-By abuse of notation, we view $\langle\eta\rangle$ ($\eta\in\Lambda$) as a continuous ${\mathbb{Z}}[v^{\pm 1}]$-linear operator on $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$ sending [~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]$ to [~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{t(\eta)x}]$, for any $x\in W_{\mathrm{ex}}$.
-By (7.11), we have the following equality in $K_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$,
-(7.12) $[\widetilde{{\mathcal{Z}}}_{x}]=\big{(}\prod_{\alpha\in\Phi^{+}}(1-v^{2}\langle-\alpha\rangle)\big{)}[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
-In $\hat{K}_{0}({\mathrm{A}}_{\mathfrak{n}}\text{-}\mathrm{mod}^{B\times{\mathbb{C}}^{\times}})$, we express [~$[\widetilde{{\mathcal{L}}}_{y}]$ into (topological) basis $\{[\widetilde{{\mathcal{Z}}}_{x}]\}_{x}$ and {[~$\{[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]\}_{x}$ as
-[~$[\widetilde{{\mathcal{L}}}_{y}]=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{Z}}}_{x}]=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
-Then (7.12) shows that
-(7.13) $\big{(}\prod_{\alpha\in\Phi^{+}}(1-v^{2}\langle-\alpha\rangle)\big{)}(\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}])=\sum_{x}[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}\cdot[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}].$
-By [29, §11] (see also [38, Thm 6.1]), there is an equality
-(7.14) $\sum_{x}(-1)^{l(x)+l(y)}q_{x,y}(v)\cdot p_{w_{0}x,w_{0}z}(v)=\delta_{y,z},\quad\forall y,z\in W_{\mathrm{ex}}.$
-Hence it follows from Proposition 7.6 that
-(7.15) [~$[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{Z}}}_{x}]_{\mathrm{gr}}=(-1)^{l(x)+l(y)}q_{x,y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
-Combining (7.1), (7.13) with (7.15), we deduce that
-[~$[\widetilde{{\mathcal{L}}}_{y}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}]_{\mathrm{gr}}=(-1)^{l(x)+l(y)}p_{x,y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.$
-By (7.14) again, we have
-[~$[\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x}:\widetilde{{\mathcal{L}}}_{y}]_{\mathrm{gr}}=q_{w_{0}x,w_{0}y}(v),\quad\forall x,y\in W_{\mathrm{ex}}.\qed$
+To complete the proof, we show that as a truncation of $\widetilde{{\mathcal{Q}}}_{x}$, the module $\widetilde{{\mathcal{Q}}}^{\Omega}_{x}$ admits graded multiplicity obtained as truncation from the one for $\widetilde{{\mathcal{Q}}}_{x}$, as (7.9).
+As $(T\times{\mathbb{C}}^{\times})$-equivariant ${\mathsf{A}}_{\mathfrak{n}}$-modules, we have $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes U{\mathfrak{n}}\simeq\bigoplus_{\eta}\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\otimes(U{\mathfrak{n}})_{\eta}$.
+Since $\mathrm{Ext}^{1}(\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{x},\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle)=0$ if $x\bullet_{l}0\nless y\bullet_{l}0$, from the filtration (7.19) one can obtain a separated decreasing filtration of $\widetilde{{\mathcal{Q}}}_{x}$ in ${\mathsf{A}}_{\mathfrak{n}}\text{-}\mathrm{Mod}^{T\times{\mathbb{C}}^{\times}}$, with composition factors by Verma objects.
+Hence $\tau^{T\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x})$ (see again the proof of Lemma 7.3 for the truncation functor $\tau^{T\times{\mathbb{C}}^{\times},\Omega}$) is the quotient of $\widetilde{{\mathcal{Q}}}_{x}$ by the submodule composed by $\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y}\langle d\rangle$ with $y\bullet_{l}0\notin\Omega$, and therefore its graded multiplicity of Verma objects is as the RHS of (7.9).
+Since $\widetilde{{\mathcal{Q}}}_{x}^{\Omega}=\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x})$ is a quotient of $\tau^{T\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x})$, we have
+[…]
+On the other hand, by the BGG reciprocity [35, Prop 3.4] (in loc.
+cit.
+it is stated for $\Omega=\{\lambda\in\Lambda|\lambda\leq\nu\}$ for some $\nu$, but the proof works for general $\Omega$) and (7.17) we have
+[…]
+Hence (7.20) is an equality.
+It follows that $\widetilde{{\mathcal{Q}}}_{x}^{\Omega}=\tau^{U{\mathfrak{n}}\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x})=\tau^{T\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x})$, whose graded multiplicity $(\widetilde{{\mathcal{Q}}}_{x}^{\Omega}:\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}=(\tau^{T\times{\mathbb{C}}^{\times},\Omega}(\widetilde{{\mathcal{Q}}}_{x}):\widetilde{{\mathcal{M}}}^{\mathbb{C}}_{y})_{\mathrm{gr}}$ is equal to the RHS of (7.9).
+∎